A Bibliography of ISSAC (International Symposia on Symbolic and Algebraic Computation) Conference Papers

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

30 October 2017 Version 2.23

Title word cross-reference

(1, 2) [EGL04]. (F_{5}) [Fau02]. (n − 1) [BdC09]. (q) [BSC^{+}12]. (T) [LMPX08]. −3 [KR08]. 0∗ [BFS93]. 1 [BC92, JKR05, KR08]. 2 [FvH11, Hie92, Ser12]. 2^{nd} [Sch95b, Sch95a]. 3 [Bur11b, CM96a]. 32 [Wol13]. 3^{rd} [Sch95b, Sch95a]. 4 [Cor01, DEF13, PO08]. 5 [Cor01]. 6 [HS06]. 6R [RGL94]. ^{2} [GR97]. _{3}

[Mül98]. A [Beu12, CDHJ07]. A^{Hash} [WBD^{+}94]. C(x, e^{∫ f(x) dx}) [BF99]. D [LM08, Sal05, GS05]. exp [RSSV96, Str08]. f(x)0(00 − 010)0/02 mod h(x) [BP94]. f(x, y) [CC89]. F 5 [PHW13]. F_{2} [vG96]. F_{q}(t) [GZ03b]. G [Roa97]. GF (q^{m}) [MvV89]. ∈ [BB07]. l [DDS13]. λ [KMO90, Oka89]. ln [PO08]. log [RSSV96, Str08]. M [Ser12]. Z[x] [ASZ00]. μ [DCS05]. N [FS12, MS98, Pot96]. O(n^{3}) [Sto98]. P [Abr06, DS99, KP10, Abb89a, Abb89b, AIRR10, BL12, LRW94, Nus97, Thu96a, Thü96b, Wan92a]. p^{2} [BvzGZ12]. P^{3} [EGL04]. P^{4} [KS11a]. ∥ [Sie93]. q [Abr95b, Abr10, GK12]. QR [Vil07a]. R [Vil07a]. R^{3} [Bro00]. R^{n} [Ric96]. S [Vol93b]. Sat(T) [LMPX08]. s exp(s) = a [CDHJ07]. W [CJK97, CDHJ07]. Z(x_{1}, , x_{r}) [SL93]. Z_{N}[x] [CL01].

-adic [Abb89a, Abb89b, AIRR10, BL12, DDS13, Wan92a, DS99]. -bases [DCS05]. -closed [BFS93, Ser12]. -confluence [KMO90]. -descent [FvH11]. -difference [Abr95b, Abr10]. -differential [BSC^{+}12]. -dimensional [DEF13]. -finite [KP10]. -finiteness [Sal05]. -gcd [BB07]. -Groups [Thu96a, LRW94, Thü96b]. -holonomic [GK12]. -hypergeometric [Beu12]. -loop [BC92]. -manifold [Bur11b]. -manipulator [RGL94]. -matrix [Jea00]. -module [LM08]. -modules [GS05]. -piercing [Nus97]. -Primes [CM96a]. -recursive [Abr06]. -representations [Ser12]. -space [PO08]. -st [BdC09]. -surfaces [PO08]. -transformation [Vol93b]. -year [Wol13].

’05 [Kau05].

1.1 [BP94]. 1979 [Ng79]. 1981 [Wan81]. 1993 [ACM93, Sin93]. 1994 [Hon94]. 1995 [Ait95].

2 [Kre89]. 2006 [Tra06]. 2009 [May09]. 21st [Jef08]. 2nd [Sch96].

3D [Sti89].

4 [BL98].

5-modular [Mül98].

6th [Mor89a].

7-roots [BF91]. ’76 [Jen76]. ’79 [Ng79].

’81 [Wan81]. ’86 [Cha86b]. ’88 [Gia89]. ’89 [Gon89].

’90 [WN90]. ’91 [Wat91]. ’92 [Wan92b]. ’94 [ACM94, Hon94]. ’95 [Ano95a, BGY96, Lev95]. ’96 [Ano95b, CL96, Lak96].

A1PI [TD89]. AAECC [Mor89a]. AAECC-6 [Mor89a]. Abduction [KP94a]. abductive [KSI97]. abelian [FM06, FFdlAGM00, KR86, MMS07]. Absolute [Chè04, GS10, CGKW02, GW97]. Abstract [SS06b, Bel90, CT98, Che90b, Cod90, CO94, Ebe97, ER94, Fat96, Ger98, HMS01, Iva00, Lam97, Nöc00, PB99, San93, SSW09, Sme90, Ste13, vG96, YNT90b, vH13, vH96, vzGN99, vzG09]. AC [BGK94a, Bün91]. AC-term [Bün91]. Accelerated [Web94b, CZ00]. Acceleration [PW02]. accelerator [AZZ91]. accessibility [HKL^{+}09]. accessible [Wan99]. accuracy [GSW12]. accurate [AP05, AT94]. ACE [GP97, PG97]. ACM [ACM93, Gon89, Jen76, Wan81]. ACM-SIGSAM [Gon89]. action [FS12]. actions [BKLG97, CGL91]. addition [CK00b]. Adherence [BBDP05]. adic [Abb89a, Abb89b, AIRR10, BL12, DDS13, DS99, Dum08, Wan92a]. ADIFOR [BCCG92]. adjacency [BBDP05]. Admissible [Zob05]. ADrien [VM99]. Advanced [Har00]. Advances [Fat90]. Affine [JPR05, ALM09, AC04, Far06, Por89, RD05]. aggregate [EMT10]. aided [KK91]. Albert [Jac94]. ALDES [Kre89]. ALDES/SAC [Kre89]. ALDES/SAC-2 [Kre89]. Algebra [Coo95a, DGS94, Fat81, Gut93, Mor89a, Sch94, SS93, Zim93, ACS86, AGLM99, Alb12, Bar03, BS86, BK86, BLWhW91, BR06, BHP07, BCLA82, CD11, Ces97, Cha95, Cha03, CGG^{+}86, CS05, Coo86, Coo95b, Dat03, DGT91, Dev89, Dew92, DF94, DGP02, DGP04, DKLM98, Fat91, Fau09, GSV91, GKWW91, Ger10, Gol86, Gon86a, HMS01, IA98, Jac94, Jur01, Kaj92, KNS11, Kha01, Kol90, Kre89, LHJ92, Lau03, LS12, LÜ10, Leo86, LS03, Li04, MC92, Mar91, MT12, MBC92, NF97, NT92, OW05, Pou06, RLZ86, RSSRV89, RER09, Rd91, RLBS91, San93, SS86, VT98, Vil07b, Vil11, WG91, Wei94b, Wol13, WR06, Yap09, Zim97, Zol91]. algebra-the [Wei94b]. Algebraic [ACM94, Ano95b, AF86, ACFG05, Bro93, Bro07, Cha86b, DD89b, Doo99, GR94, GLV89, Glo98, Gon89, Gut04, Hub10, Jen76, KSY94, Kau05, Küc97, Lak96, LK90, LKK91, Lev95, MCG13, Mor89a, Mor02, Mou01, NF96, RCK04, RF94b, SF94, SE11, Sen03, SXZ11, Sud08, Sul12, SS91, Tho96, Tra06, Tra00, Wan81, Wan92b, WN90, Wat91, Wat10, vdHvH12, Abb89a, ABD86, Abb89b, AB89, AV12, AMR89, AL08, Arn89, BPR98, BS86, BJ98, BL12, Bod04, BCM94, BCS^{+}07, BDE^{+}13, BL98, Bro86, Bro91, Bro98, BD07a, Bro13, BCLA82, BS07, BCGY08, BD07b, Cha90, CMXY09, CDM^{+}10, CDM^{+}11, CKS12, Che90b, DS03, Du94, EKW07, El 13, Enc94, Enc97b, Fat89b, Fat90, FsG04, FGPT02, GC92, GR97, GKWW91, GGS89, GKZ90, Gia89, GV91b, Gon91b, GS98, GL11]. algebraic [Hon90, Hon92, HMS01, HD00, Ida09, JM07, JM09, JPR05, Joh92, Kal90, KS97a, KK06, KYAH08, KYAH09, Kau08, KMH89, KM95, Kut89, Lan91, Lec00, LÜ10, Lev12, LkGID97, LD11, May09, MP90, MG94, MK99, Ng79, Par13, PDS05, Pfl08, Rio92, Rio02, RV93, RV96, Roy12, Ryb91, San04, Sas86, SLA89, Sch98, Sch91b, Sed01b, SS03, Sho99, Sme89, Sme90, Str10, Str12b, Tao90, Ver04, Vol97, WR06, Yok04, Zha93, kMG97, van94, vHM04]. algebraically [YNT90b]. algebraists [AK91b]. Algebras [CT96, EG96, MZ96, BFS93, BHP07, CIK97, GMT89, dG97, GSZ11, Han07, HS06, HP89, Hub10, Iva00, LS03, LSS13, Mor89b, Nör98b, OS91, RLBW92, Sch91a, Shi90, SWMZ12, Zol97, dGW99]. Algorithm [Cab93, CdR96, CM96a, FM96, Gru93, HW86, Jeb93, KM96b, MF94, Sie94, Tsa96, Val93, Web94b, Abr02, AP05, AM90, ARS97, AL08, Arn13, AH08, AT94, Baj86, Bar89b, Bar95, Bea93, BCLV13, Ber94, BF12, BS00, BDPZ09, Bra90, Bro86, Bro94, CL89, Cha03, Che90a, CG90a, CK92, CFS90, CGKW02, CG86, DRJ97, Ebe10, EM99, Enc94, Enc97a, Fau02, FM11, FGPF89, GGV10, GG12, GM06, GKZ07, GM86, GG89, GLL04, GZ08, GKsL02, GMN^{+}91, GT09, GS98, GL11, vHM02, HD09, HE05, Hub04, JM07, JP99, Jea00, Kai94, KL94, KM99, KsLL00, Kal02, KSW10, KS11b, KSS07, KM96a, KC09, KS86, Sca94a, Lan91, Lap06, Lel12, LN97, LkGID97, jL90, LLMM08, Man94a, Mar95, MW01, MMS07, MW00, Mon04, Mor92, Mor95, MF97]. algorithm [Nab07, Nor90, Oak12, PB99, PW02, PQZ11, PHW13, Pfl97, PR08, Pot96, QV95, RWW94, RY08, Rou10, Ryb91, SY11, Sag12, SLA89, Sca94b, Sch91b, SP91b, SP91a, Sch89, Sed01a, Sed01b, SY12, Sho91, Sme89, Sor95, SZ91, Sto98, SWMZ12, SS06c, Tak90a, Tho01, TD89, Tur06, Vil97, Win89a, Win89b, YZ86, dKMW05, van95, vH93]. Algorithmic [Kem08, LRB95, RLBW92, SDJ07, BBDP04, GR10, HM01, MN93, AH94]. Algorithms [ABRC97, BHC95, BCB09, BL10, BAE^{+}00, CM11, CT96, GKsL02, uGK91, GL95, HK04, LS12, LuS91, MZ96, Mor89a, MM01, Oak94, RD05, Web94a, dKMW05, vHM04, AK91a, BCFS91, BK86, BdC09, BCS05, BCSC06, BS09b, BCdP13, BL98, BCGY08, But92, CT98, CIK97, CKP12, CR03, CG92, DDS13, Dew92, EK97, Ebe04, EP11, FR95, Fuk08, GPT89, GM06, Ger10, Gie94, GSGVPN^{+}09, GRA00, GR92, HK98, HKL99, Iva00, JX07, KKS86, KY89, KM95, LM08, MC91, Mon05, Pan01, PS13, PS07, Sal05, SW04, Sch92, SL93, SR89, Sie93, SKCL12, Sto96, Sud08, SW11, SS91, Tra89, Uma06, Wav99, Zol97, kMG97, vzGG97]. alication [GR97]. Alkahest [FNM86]. Almost [ESY06, Mon04]. along [AMBSW11]. ALPAL [Coo92, Pai92]. alternate [CL06]. alternative [Giu09, Ste08, SS02]. always [CDHJ07]. amidst [GS98]. among [MS97, Ric95, YLN95]. amplitudes [BL91]. analog [Tak90a]. analogue [Giu09]. Analyses [GV93, MV90]. Analysis [Ste96, BCLV13, BFP12, BB07, Bür11a, Cet92, Cha95, Cor94, EKW07, Far06, FNM86, GSV91, GVv92, Kha01, KK91, KF91, Lan91, Lev12, MS92, MBC92, Pol98, Ren91, Rd91, SY98, SW89, SXZ11, Sor95, TC89, TMH94, VF92, Vil97, Vol93a, Vol93b, WX05]. analytic [Cha95, Mez10, RRW99]. analytical [CMS89, LCW05]. Analyzing [DS94, GSGVPN^{+}09, GMS90]. and-or [Bel90]. And-parallelism [GP97]. Andrews [Tra00]. annihilators [AZ97]. Announcement [Ano95a]. annual [Jef08]. ansatz [IK12]. ANTS [AH94]. ANTS-I [AH94]. Applicability [Abr02, DJ94]. Application [BLWhW91, BP06, GMS90, GT90, Gut93, KS06, MW05, Pet91, RLBS91, Tra98, VT98, WR06, kMG97, van96, BP98, BHP07, CC89, DP95, Dye94, FS12, HLY86, Han07, HD09, KR93, LSS13, MW94, MSW07, MM97, Tho01, Tou05, TSD91, Wan91]. Applications [HCGW93, Pur86, BBP08, BLPR13a, BS10, BW98, CC90, Fat91, FES10, HI93, HR10, KMO90, Li98, LH08, Par04, RG91, Sal05, SF90, Sto97, SS91, TD90, Tsa10, tW05, WR13, Zim97, vdH04]. applied [Dal92, Har94, VF92, Mor89a]. Applying [Wis89, Wol95]. appraisement [BCK99]. Approach [BGK96, DGS94, SB94, ACS86, BP07, BBDP04, BFP12, BC89a, CKS09, GVv92, Lau03, LY86, LS90, MGv91, MBC92, NS06, RRM09, Ren91, Rus92, Sha89b, Sha89c, SS02]. Approaches [DGS98, OW05]. approximability [HMS01]. approximants [BL09, Cha86a]. Approximate [CGJ99, GKM^{+}04, JFRS06, KYZ06, KL96b, LSYL11, Sas01, CCC98, CGvH^{+}01, JK97, KM03, KY07, KLYZ08, LYZ10, MSW13, Nag02, Ter09, ZD04, Zen09]. Approximately [BS13, Nag11]. Approximation [Tou05, CC89, GS90, KS11c, Laz06]. approximations [Abb89a, Abb89b, GR10, GT09, KRL86, MH01]. April [CM96b]. APS [GD93, LK90]. arbitrary [AK12, FGT01, Kau04, Oka89, ZR99]. arbitrary-order [AK12]. Architecture [JKR05, RC06, WGK^{+}01]. Architecture-aware [JKR05]. Architectures [INPS93, Bis12]. arise [GKZ90]. arising [BHP07, KR93]. Arithmetic [CJSW94, MP90, vG96, BP06, Fat92, FLMS06, HD00, Joh92, JK97, LMS07, Pan00, Rum10, SY98, Wei97]. arithmetics [FS09]. Artin [FS09, GKZ12]. Asian [BP06]. Asir [NT92, NM97]. Aspect [RV93]. aspects [Beu12, MvV89, Mil11]. assignment [Wu93]. assistant [AK91b]. assisted [EK95, FM01]. association [Miy00]. Associative [EG96, CdG07, GSZ11, OS91]. asteroid [VT98]. Asymptotic [Gru93, RSSV96, SS92, Sha89a, Str12a, CEMP05, Sha99, SXZ11]. Asymptotically [Ger10, Kal94, Rei97, SL96]. Atkin [GM06, MMS07]. Atlanta [Ait95]. atlases [Man08]. attractors [Far06]. August [Bro07, Glo98, Jen76, LGG97, Sen03, Tra00, Wan81, WN90]. Austria [BW98, Hon94, Jef08]. Automata [MT94]. Automated [AGLM99, Cho90, CGZ93, Doy99, KS86]. Automatic [Bac96, BGH94, Cha89, CR94, Päs09, RCK04, SB94, ABRC97, Cha03, Fat03, FNM86, GMR94, LD11, MGv91, SW89, VM99, BCCG92]. automatically [MS03b]. automatics [Oll89]. Automating [Pet94]. automation [LD90]. automorphism [Sha03, Sim97]. average [Enc97a]. aware [JKR05]. AXIOM [JT94, Doy99, Gil92, Rio92, SDJ07].

Babel [SRHW94]. baby [Kal02, RC06]. Bäcklund [GKZ91]. Bags [GK94]. balanced [LH08]. Barnes [Ada01]. barrier [Bro92]. Base [ARS99, BCFS91, CFS90, CF92, HLY86, Mar95, Mor95, SS94, WCD08]. Based [ABP94, Bac96, IEE95, LHD94, BL90, BB07, BJ10, CS05, CK02, CR03, DP07, Ebe04, EP11, GKZ07, GSGVPN^{+}09, Hon92, KSI97, McC99, McC01, MB09, Nor90, RC06, RMC07, SRHW94, Sas01, SWMZ12, VT98, WEM98, Win89a, BGY96]. Bases [MR93, Sti93, Tra98, AH08, BS98, BF03, BW91, BGM91, BW98, CD11, CCT08, CL06, CdG07, Cox07, CG86, Cza91, Dah09, DCS05, EP11, ER94, FR09, Fau02, Fau09, FES10, FM11, FES12, FEV13, FS13, GGV10, GM86, GZ08, KSW11, KFS86, KM96c, KN07, LSS13, Lev07, MMM91, MR10, MMT92, Mor89b, Nor98a, PW06, Pet91, Ryb91, Sat98, SS01, Sch96, Ste97, SWMZ12, SS02, SS06c, Wei92a, Wei02, Win89a, ZW06, ZL09, ZLS12, ZL13, Zob05, dGW99, WR07]. Basis [AGK96, Fau94, KM96b, Sie94, AD11, BSS08, Cab93, Cho90, ER13, FSK86, GPT89, GR97, Hie92, Hon96, KM96a, LM94, Man94a, MR11, MZ01, MT12, Nag11, RS12b, SV05, SW11, Tak90a, Tak90b, TD89, TZ02, Vil92, Vil07a, Vil11, Vol03]. Bath [ABD86]. Baxter [GS06, Hie92, OS91]. BC [Doo99]. be [CDS89, Fit89, Gri90b, MS02]. before [RS12a]. Beijing [Kau05]. Ben [KsLL00]. Ben-Or [KsLL00]. Ben-Or/Tiwari [KsLL00]. Bendix [BD86, Win89b]. Berlekamp [CKP12, GKsL02, KL94]. Berlekamp/Massey [CKP12, GKsL02]. Bernstein [ALM09, BL10]. Bessel [DvHK08, vHY10]. betatron [AZZ91]. Better [BBD03, BBDP05, BD02]. between [BHC95, Fau09, OW05, RER09, SW01, Sme90]. Bézier [GP05]. Bezout [CGR10, Gem97]. Bezout-like [CGR10]. bi [BM05]. bi-equational [BM05]. Biased [KC09]. bicubic [GP05]. bidegree [EGL04]. Bidirectional [KJ94]. Bieberbach [Koe03]. Bignum [CJSW94]. bihomogeneous [BD07b]. Binary [Jeb93, Sch91c]. Binding [Coo95a, Coo95b]. Binomial [KM96b, HL86, JZXM01, KM96a, McC03, Shi90]. bio [BHP07]. bio-operations [BHP07]. biochemical [DJLS07]. biological [Lau03, Tou05, WX05]. bisection [EGT10]. bivariate [Ber98, BLS^{+}04, BCCL10, BLPR13a, BLPR13b, CGK^{+}13, CGL09, CGvH^{+}01, DE01, Dah09, DET07, ES12, GW97, Kal90, KK05, KSG04, LMMRRR13, LMS13, Nag02, PS06, vzG07]. Black [BS10, DK95, Ebe00, Sau01, DK98, DPS09, EGG^{+}07, KL94, KN11, Tur05]. Black-box [BS10, DPS09]. Black-box/white-box [BS10]. Blackboard [DJ94]. BLAS [CS05]. Blind [LYZ10]. block [EGG^{+}07, Ebe10, HE05, Tho01, Tur06, Vil97]. blocks [SS94]. blockwise [vdHL12]. BLT [Bet13]. BLT-sets [Bet13]. Bonn [Wat91]. boolean [BDE^{+}13, PT13, BD08, GZ08, GRA00, Kol90]. Boolean-product-witness-matrix [GRA00]. Border [AD11, MT12]. Borel [Lel12]. Borel-fixed [Lel12]. Boston [MCG13]. bound [Ast91, BG07, EMT10, Kal93]. bounding [KS11b]. Bounds [GV00, BY09, ESY06, EMT10, KS10, Koi12, MR10, MR11, PV01, Sch02b, SY12, SU92, Vin91, YZZ13]. box [BS10, DK98, DPS09, Ebe00, EGG^{+}07, KL94, KN11, Ric96, Sau01, Tur05]. box/white [BS10]. Boxes [DK95]. brackets [LH08, OS91]. Branch [DF94]. branches [GR01]. Branges [Koe03]. breaking [Bro92]. BREEFS [Li07]. Brokers [ABP94]. Brown [MW00]. Buchberger [AT94, Bra90, CdR96, CG86, GM86, GMN^{+}91, KS86, LkGID97, Win89b]. bulk [CG01]. Bylaws [Kal96].

C [CS05, WEM98, vdHL13]. CA [SE11]. CABAL [NF97]. CAD [BBDP05, Bro00, BM05, DSS04, McC99, McC01, MB09]. CAD-based [McC99, McC01, MB09]. cage [BG07]. calculating [AM90, CL89, Ter09]. Calculation [CK93, CK96, Hul13, BL91, CGZ93, Vin91, dJ91]. Calculations [KMG96, NF96, IKRT91, SD86, SS91]. calculus [BGJ13, Cor94, LHJ92, Oka89, Sha89a]. California [ACM93]. call [FT03]. call-by-reference [FT03]. call-by-value [FT03]. CALYPSO [Ces97]. Cambridge [HI93]. camera [RTZ03]. Can [CDS89, Fit89, Ste08, Gri90b, KP10, KMZ99, MS02]. Canada [Doo99, Mou01, Bro07, Lev95]. canal [LSWH00]. cancellation [SY98]. Canonical [Fat91, Nor98a, Wei02, ALP03, ARS97, BL00, Car07, CO94, Gil92, KM02, KM89, Sat98]. Cantabria [Gut04]. Careful [Kut89]. Carleman [CDS91, CD99]. carry [YZ86]. carry-free [YZ86]. Cartan [DP07]. CAS [CA07, Kaj92]. CAS/PI [Kaj92]. CAS1 [RLZ86]. CASA [GKWW91]. Case [AGK96, ZB94, AGLM99, BF03, CA07, Dah09, Ebe10, FH97, FES12, FS13, LMS13, RS12a, Sch12, SD86, Wei94b, dKMW05]. cases [Hub04]. Castelnuovo [Giu89]. Catalan [Fee90]. catastrophe [GSV91]. categorical [DGT91]. cavity [MYT09]. Cayley [BC89b, CK04, WM89]. cell [Bro13, McC03]. center [FGPF89]. Central [DGS94]. centralizers [KC09]. certain [KMO90, vzGG97]. Certificates [GKZ12, KNS11]. Certification [Vil07a, KLYZ08, PS12, Sto09]. Certified [LMMRRR13, GT09, MM11, Nag02]. Certifying [GLS98]. Chains [BWZ94, CM95, Zim95, KMZ98, LMP09, Rol89, Zim01]. challenge [Laz06]. Change [PS06, BSS08, CFS90, CF92, FM11]. Changing [BF03]. Character [Thu96a, BF94, Thü96b]. Characteristic [KP94b, Som94, Alb12, BS09b, BM04, Can89, Cho90, Clu03, DPW05, DPS09, FGT01, KW90, LS92, PS07, Sho91]. characters [Gri89]. Chebyshev [BS09a, BJ10, EF89]. check [RGSGG00]. checking [MW05]. Chemistry [Bar03]. China [Kau05]. Chinese [BW91]. chiral [BL91]. Choice [MT94]. chow [Khe02]. circuit [BRS09]. circular [Bou05]. class [Ada97, BC92, DP95, GS98, RB06, SS09]. classes [CS97, Pet94, RW00, Som94]. classical [JKR05]. Classification [FMRD08, Bet13, FM01, LJM08, WX05]. classroom [CGG^{+}86]. Clause [MT94, Bon94]. Clause-Diffusion [Bon94]. Client [LH99]. Client-server [LH99]. Clifford [Li04, Li07]. cliques [Bet13]. close [Sas03]. Closed [BLvH04, CM95, BFS93, BWZ94, BCEW12, Ser12, YNT90b]. closed-form [BWZ94]. closure [DD89b, GT97, Rio92]. CLP [KNKM94]. Clusters [Ste96, JFRS06, Sas03]. Co [Mül98]. CoCoA [BR06]. Code [Bac96, Coo92, Zim93, BGV94, Gat86, MM97, WW89, vHGv89]. Codes [Mor89a, Kor01]. coding [Sud08]. coefficients [AK91a, Abr95b, AvH99, AK12, BLvH04, CG92, Dah09, GZ04, HL86, KYZ06, ZL12, vHY10]. coercion [DR01, Doy99]. coercions [Web94a]. coexisting [FT03]. coherent [FK91b, FK91a]. cohomology [Sch12]. College [Ait95]. collision [Sti89]. collisions [BvzGZ12, vzGGZ10]. column [BBP10, DPS13, ZL13]. column-reduced [BBP10]. combinations [BDE^{+}13]. combinatorial [LLMM08]. Combinatorics [DEF13]. combined [Min03, Oka89, Rus92]. Committed [MT94]. Committed-Choice [MT94]. Committee [Kal96]. Common [DK95, KL96b, AuK93, BCSL12, Fat03, GPT89, KYZ06, Küc91, LN97]. communication [LH99, OIiM^{+}97]. communicative [Pou06]. commutation [RLBW92, Wis89]. commutative [BS86, BR06, FS13, Lev06, Nor98a, RGSGG00]. Commutativity [Wav99]. Comparative [BGH94]. Comparison [KS95, But92, Fat89b, LM08]. comparisons [BB07]. compatible [CFFL11]. Compile [BGH94, Rup10]. Compile-time [BGH94]. compiler [WBD^{+}94]. Complete [BCGY08, CGY07, CM95, Tsa96, Bou05, LJM08, MT00, RTZ03, AD11]. completely [AZ97, BP98]. completeness [AIRR10]. Completion [AGK96, Bün91, BD86, BGK94b, GPT89, MZ11, MZ12, Nor90, RVWW05, Sch92]. Complex [Ahr96, CK96, LH08, CK92, GK12, PZ10, SY11, TCT97]. complexes [BCGW89, Sca94a, Sca94b]. complexification [RV96]. Complexity [BPR98, BLS^{+}04, BCCL10, BCdP13, Gri89, Gri90a, Mor06, Roy12, Sha07, Szá97, Wei97, BCSC06, BGP12, BD07a, Cuc00, DET07, ES12, EGT10, FH97, Fat89a, FEV13, GJV03, HMS01, Jeb97, KK05, LMS13, PT13, PV01, PR08, Tei89, Wu93, vH13, vdHL12]. complexity/efficient [HMS01]. Component [KMG96, ADG99, DS03]. components [BPR98, IA98, LCJ98, WR13]. composed [Min06]. Composition [vzGGZ10, Bea93, BSS08, Hon96, LHK^{+}10, LM02, MZ01, Sch91c]. Compositions [BvzGZ12]. Compound [Pol98]. comprehensive [KSW10, KSW11, KN07, Nab07, SS01, SS02, SS06c, Wei02]. computability [Ser89]. Computation [ACM94, BdC09, Bro93, Bro07, BHKR95, BP94, Cha86b, CO01, CM95, DG95, Fee90, FGT01, GR94, Gie95, Gil92, Glo98, Gon89, HCGW93, Her97, Hol00, Hon94, INPS93, Jef08, Jen76, KMG96, KM02, KN07, Lak96, Lev95, MPV98, Min95, MCG13, Mor02, NM97, ORY09, Ren06, SS09, SE11, Sen03, SW01, SL95, SL96, Str10, Tho96, Tra06, Tra98, Val93, Vil93, Wan81, WN90, Wat91, Wat10, WB93, ZSST96, vdHvH12, ABM99, Baj86, BK86, BCL97, BL09, BCK99, Ber94, Bod10, BH06, BCSL12, BHPS86, BGY96, BCLA82, Buc05, BC92, Cab93, CL96, CS97, Ces97, Cha90, CG92, Dal92, DMR08, DD89a, Dew89, DKS^{+}91, DS03, DR01, DMMRY01, DPW05, DPS13, Dye94, Ebe10, ER13, EGNS12, For01, FGPF89, Fuk08, GR97, GZ03a, Gia89, GL89, Har94, HS02]. computation [Joh92, Juo89, KZ06, KW90, Kir92, Küc91, KM96c, LCJ98, Lap06, LH99, LY86, Li04, LD11, LHK^{+}10, LS90, MW94, MSW07, May09, MR11, Mez10, Miy07, MMT92, Ng79, Nor00, PG97, Rei97, RS12b, SF90, SS06b, Sho99, Sie93, Sme89, Sti97, Tho01, TZ02, Ver04, VF92, Vil94, VC89, Vin91, Vol03, Wan97, Wan99, Wat89, WR07, Yap09, YZ86, ZL09, vHM04, Ano95b, Doo99, Gut04, Kau05, Küc97, Mou01, Tra00, Wan92b]. Computational [LM08, Mil11, Pou06, Beu12, BS07, But86, Ida09, MvV89, Tei89, Zim01]. Computations [AB89, Ahr96, BSG96, Bin96, BP91, CGL91, DP95, GC91, LMP09, Ait95, BS98, BR06, BP90, BP93, Col97, EGG^{+}07, FK91b, FK91a, GJV03, Hie92, HP97, Kol90, KR86, Leh97, LSS13, MV90, NS06, Pan89, Zol91]. compute [CD11, Cha86a, EP11, GT94b, GV00, KMZ99, Oak12, SS06c]. computed [MS02]. Computer [BCLA82, CGG^{+}86, DGS94, Edn93, Edn97, EK95, FK91a, Fat81, GZ90, Gut93, IEE95, Kau04, RSSRV89, Sch94, SD86, SS93, Wol13, Zim93, AGLM99, AL08, Bar03, BLWhW91, BHP07, Ces97, Cha95, Cha03, Coo86, Dat03, Dev89, Dew92, DF94, DKLM98, Fat91, Fau09, FM01, GSV91, GKWW91, Ger10, GL91, IA98, Jur01, Kaj92, Kha01, KK91, Kre89, LHJ92, Lau03, LS03, Lev12, Mar91, MM97, MBC92, NF97, NT92, OW05, Rab89, RLZ86, RER09, Rd91, RLBS91, San93, SS86, SS91, VT98, Zim97]. computer-aided [KK91]. computer-algebraic [Lev12]. computer-assisted [FM01]. Computer-Based [IEE95]. Computing [AV12, AMR89, BP07, BCEW12, BL00, BM91, Che89, CMXY09, CDM^{+}11, CL01, CFT95, CSU00, DZ05, DCS05, DK95, Din08, FES10, FGPT02, GL95, Gru93, GEZ13, GS11, HR03, Hul98, HKZ10, JM10, KSW11, KMO92, Lec00, LRW94, MZ12, MR93, Man94b, MMS07, Mil04, Miy00, Nag11, PS13, RV96, SSW09, SC10, Sim97, Sin93, SV05, Thu96a, Thü96b, Vil96, WZ08, ZB94, ZLS12, ZL13, van94, vzGN99, Bar89b, BPR98, BS86, Bea93, BBDP05, BF12, BP06, BJ10, Che90a, CG90a, CM11, CL06, DW90, Enc94, Fau02, FEV13, GGV10, GM06, GG89, GZ08, GKsL02, GT09, Gri89, GL11, HI93, HW86, HK98, HKL99, Hit05, Kal92, KSW10, LS94, Lel12, LN97, Li07, LH08, MW01, Mon05, Nab07, Pfl97, RY08, Ric91, RC06, RER09, Ryb91, SP91b]. computing [SP91a, Sto96, SWMZ12, Vol97, Zen03, van95]. concepts [BC00]. concrete [Kau11]. Concurrent [LC94, TL94, OIiM^{+}97]. Condensation [CT96, RMC07]. condition [Bür11a, CR91, CdG07, Hie91]. Conditional [DJ94]. Conditions [CK00a, CK06, BOS13, CMS89, GZ90, GKZ90, Pil13, SU92]. Conference [Ait95, BW98, LGG97, Mor89a, Sin93]. configurations [FK91b, FK91a]. confluence [KMO90]. Conform [Sey86]. conformal [Sey86]. congruence [Sch91b, SZ91]. conic [SW10]. conjecture [AJL01, Bad94, Koe03]. conjugacy [CS97]. conjugated [EPB97]. conjugates [BH06]. connected [BPR98, BM04, DS03, Gri90b]. connections [BCEW12]. connectivity [Roy12]. conquer [But86]. Consistency [GR10, GMS90]. constant [Fee90, GZ04, Hie92, Ric92]. constants [CHT^{+}07, RF94a, Ric95]. Constrained [BR94, KK89, HK98, Hon93, KYZ06]. Constraint [ABP94, LC94, TL94, AB01b, Bas97, DS99, GMB01, McC99, RF94b, SF94]. Constraint-Based [ABP94]. Constraints [MR94, BM05, Coh13, McC01, MB09]. Constructing [Bro13, CFYT94, CO94, dG97, GHS00, KM96b, dGW99, KM96a, Tak90a]. Construction [KRL86, Ser12, YT93, BGL05, Bro99, BM05, BGHS99, DP95, Fat90, Hon92, LuS91, Pet94, SY98, VT98]. constructions [CK04, CFL89]. Constructive [LY86, Fra90, GKWW91, Ren91]. constructs [CCF89]. content [Doo98, Doo02]. context [CO94, Ged86, KMO92]. context-free [CO94, KMO92]. contextual [AB01b]. continuation [Cha95, MZ11, RVWW05]. continued [Sha07]. continuous [JR98, SXZ11]. contour [Kir92]. contraction [Bon94]. contributions [BC92]. Control [DGS94, LHD94, KYAH07, MGv91, RD05, dJ91]. Controlled [BMP98]. convergence [Fat89b]. convergent [KFS86, Oka89]. converging [KSS07]. conversion [LM94]. Converting [FT03, Tra98]. Convex [Par13, Sch09]. convexity [HS08]. Conway [Mül98]. Cook [BZ07, Giu09, Zan10]. cooperation [GMB01]. Cooperative [MR94]. coordinate [Por89]. coordinates [HK04]. Coordinating [Doo98]. Coppersmith [Vil97]. Copying [Kes94]. corner [PR06, ZG00]. corners [Rou10]. correct [CKP12]. Correcting [Mor89a, Jur01]. correspondence [AP06]. Coset [Hav91, CG01, RMM98]. cosets [HL03, KM89]. cosmological [CDS89]. cost [KSI97]. cost-based [KSI97]. Costs [Kes94]. counterparts [OS89]. Counting [vzG07, BS13]. coupled [FM01, GKZ91]. Coupling [WB93]. Covering [Nör98b]. Cray [WW89]. create [Ste08]. Creative [BLS13, BCCL10, BCC^{+}13, CK12, Le02]. criteria [BD86, Zob05]. criterion [SW11]. Critical [BD86, El 13, FES12, GEZ10, Nor90]. Critical-pair [BD86, Nor90]. critical-pair/completion [Nor90]. crossing [OS04]. cryptography [CCT08]. cryptology [Fau09, FES10]. cube [GMN^{+}91, KC07]. cubic [Edn97, KR08, Wei94b]. cubics [KM02]. Cumulating [ZB94]. cumulation [MGv91]. curriculum [Dev89]. curve [ARS97, Bro91, DMR08, GR01, Gri90b, Mor92, Sha03, Tei89, Vol93a, Vol93b, Vol97]. curves [AB89, ARS99, AF86, BM91, BCGY08, CK12, CJKS13, DCS05, Du94, EKW07, Giu89, GV91b, GVT95, Gon91b, KS11a, LSYL11, LM94, Man90, Mil11, Moo11, Por89, RSV06, SS95, SW01, SC10, van94]. customizable [ADG99]. cuts [DF94]. cutting [ZG00]. cyclic [AZZ91, BF91, CF92, Han07, Vin91]. cyclicity [Edn97]. cyclotomic [BD89]. Cylindrical [BDE^{+}13, Bro98, BD07a, Bro13, CMXY09, Hon90, Hon92, PV01, Ric91, SLA89, SS03, Str10, Str12b].

D [Mez10]. D-finite [Mez10]. DAE [QV95]. d’Alembert [Abr93a, Abr93b]. D’Alembertian [AP94, AZ96]. Data [TL94, BSG97, NS06]. Data-Parallel [TL94]. database [BIL91]. Dataflow [Yam90]. decidable [Abr10, CC90]. Deciding [AMW08, AW00, BBR93, GE11, MQR98, FFdlAGM00, Roy12]. Decision [CCF89, CC93, Wei89, Win89a]. decoding [KPR^{+}10, Pan00]. decomposable [vzG09]. Decomposing [EPB97, GZ03b]. Decomposition [BFS93, BS93, CGTW95, EG96, FvzGP10, GZ04, GMT89, AP01, AF86, Arn89, BBCM11, BD07a, Bro13, CMXY09, CDM^{+}10, CDM^{+}11, CGJ99, EM99, FP09, GGSZ03, GS05, GS08, GR92, GRS01, Hon90, Hon92, Hub04, LS12, Lec00, Ley08, Mül98, Pfl08, SLA89, Sch07, SS03, Szá97, Wei92b, Zie91, Zip91, Zol97]. decomposition-lemma [Pfl08]. decompositions [BDE^{+}13, Bro98, CM11, DMS^{+}05, Ebe00, Kir92, Maz06, Mil04, PV01, Ric91]. deconvolution [LYZ10]. Dedekind [BP91]. Deduction [DS94]. deductive [BIL91]. defect [Bra89]. defects [FES10]. defined [Car07, Che90b, LSYL11, SW10]. definite [AGLM99, GKZ12, Le02]. definition [CG01, MS92]. definitions [Ser89]. Deflation [MM11]. Degree [FM96, MR10, Sch02b, Ast91, BvzGZ12, CK12, CJKS13, CR03, Ger98, GKZ12, GRS01, HS06, Hub04, KL94, KK06, KvH13, MR11, PDS05, SW01]. Del [HS06]. delay [Pec91]. delay-differential [Pec91]. Delayed [DMN94]. delineability [MB09]. delivery [Fit90]. denominator [KS10, KS11b]. Dense [Por89, ABM99, Alb12, Pan89, SW04]. depend [AuK93]. dependencies [Sme90]. dependent [GP97]. depending [Bou99, Che89, JP99]. depth [Sch05a]. derivation [Cha03, SW89]. derivations [GL89]. derivatives [BG11, FP09, RR97]. derived [MYT09]. Deriving [TCT97, CG90b]. Descartes [ESY06, JKL^{+}06, Sag12]. descent [FvH11]. Describing [Sch09]. descriptions [BPR98, Bay04]. Design [ADG99, Wan99, ABRC97, ACS86, BS86, BCM94, CGR10, Fat90, HVS97, LHJ92, Leo86, MW00, YZ86, CL96]. Desingularization [AvH99, CJKS13, Bod04]. detect [KP10]. Detecting [Bur11b, Sme90]. detection [AD11, BCR13, EK95]. determinant [Kal02, KK08]. Determinantal [Khe02, CK06]. determinants [ABM99, Kal92, KK08]. Determination [RW00, Sch91a, AL08, GR01, Gon86b, HD00, LRB95, RLBW92, RTZ03]. Determining [Sha03, CR91, LSWZ06, SZ91, YNT90a]. Deterministic [PS12, vdW05, ABM99, BBR93, Cet92, IKS09, Sho91]. development [Kre89, Kry86]. developments [Cox07]. DFTs [JX07]. Diagnosing [Jur01]. diagonalisation [vdW06]. diagram [Sch91a]. diagrams [KR86]. Dialogue [Kry86]. Diego [ACM93]. Difference [GV93, AP94, Abr95b, AZ96, AB98, AvH99, Abr08, Abr10, AK12, Bar89a, Bar99, BOS13, CvH09, DT86, GS90, GSV91, GVv92, GY06, GR10, HKL^{+}09, KS10, KS11b, Lev13, LYG13, LD90, LuS91, MS03a, Man08, MV90, ZW06, van98, vHL10, vzGG97]. difference-differential [Lev13, ZW06]. differences [LH08]. differentiable [LSYL11]. Differential [BCS^{+}07, BS07, DG95, GLL04, Hub96, PS93a, RS96, Tho96, Tsa96, Wei93b, Wei93a, van96, AK91a, AP94, AZ96, AMBSW11, ACFG05, Ast91, Bar89b, BP98, BP07, BCB09, Bar10, BBP10, BR12, BS09a, BCS05, BCR08, BS09b, BCSL12, BSC^{+}12, Bou99, BGU03, BOLP95, BL00, BLRR13, Bro89a, Bro91, Bro92, Bro94, BMW97, BF99, BL02, Che90a, CG90a, CDS91, Clu03, Cod92, CSU00, Cor01, CGL91, CG92, DvHK08, DRJ97, FvH11, Fat89b, FsG04, Fre00, Fre03, GKM^{+}04, GZ04, Gri89, Gri90a, HW86, Hub04, Hub10, Koh90, KF91, KvH13, LÜ10, Lev13, LST02, LGY11, Mac89, Mac99, MS03a, MS02, Oak94, Oak12, Oku90, Pec91, PS93b, Pfl08, Pfl97, RRM09, RB91, RW00, RVWW05, RR08, RRW99, Sch89, Sch92, Sch96, Sch07, Sha89b, Sha89c, Sin89]. differential [SU92, Tsa97, Tsa99, Tsa05, Tsa10, UC90, Wei94a, WR13, ZL12, ZW06, Zob05, vHW05, vH07, vHY10]. Differential-Algebraic [Tho96, Pfl08]. differential-equations [Sha89b, Sha89c]. Differentially [RS96]. differentiation [ABRC97, BCCG92, Fat03, MN93, SDJ07, VM99]. Diffusion [Bon94]. Digital [Man08]. dihedral [Ren06]. Dimension [CT96, Lev07, Lev13, Pot96]. Dimensional [Kha01, AH08, BY09, CDS89, CGY07, CIK97, CGT97, Dah09, DEF13, FM11, GT09, KM95, LST02, LSWZ06, Mai98, MR10, Rd91, Shi90, Tak90a, YZZ13]. dimensions [CCG05, EPW89, Hie92]. Diophantine [Gie97, MS99]. Dirac [IKRT91]. direct [RMC07]. directed [RB91]. directions [Roy12]. DISCO [CL96]. Discovering [CMS89]. Discrete [SS01, AP05, GK05, KN07, LSYL11]. discriminant [FMRD08, KS11a]. discriminants [BRS09]. disequations [Cod90]. disjointness [MQR98]. disk [CR03, RC06, RMC07, SDKC10]. disk-based [CR03, RC06, RMC07]. disks [SKCL12]. dispersion [AB00, MW94]. displacement [BJS07]. display [ADG99]. dissolution [MR89]. Distinct [FM96]. Distributed [BSG96, BHKR95, CM96b, DGS94, DS94, LC94, STK97, Tak94, Wik94, ZB94, Ber97, Bon94, Bon97, BGY96, DKS^{+}91, Maw86, Wan97, HVS97]. Diversification [GR11]. Divide [But86]. Divide-and-conquer [But86]. dividing [BIL91]. Division [EKM96, KJ94, Moo11, GZ08, Jeb97, Rol89]. divisions [Kal92]. divisor [AuK93, GPT89]. Divisors [DK95, KL96b, ARS97, KYZ06, Küc91, LN97]. Dixon [CK00a, CK02, CK04, KSY94, KS97b]. Dixon-based [CK02]. DMM [EMT10]. documents [Doo98, Doo02, YF04]. does [LMPX08]. Domain [Bac96, BHPS86, Sek08]. Domains [OW05, Jef93]. doubles [BGHS99]. drawings [TC89]. Drexel [Sen03]. Drinfel’d [BGHS99, Ruf07]. driven [CT91]. DSC [DKS^{+}91]. DTS [BHKR95]. dual [ASW05, HKZ10, MMM91, Vol97]. Duality [MMM95]. duals [GHS00]. dummy [IK91]. Dwork [BLS13]. Dynamic [ASN89, BGDW95, Cab93, GP97, KK91, KN07, Nor06]. dynamical [CD99, RD05, SXZ11, Tiw10, Tou05]. Dynamics [CM95, BLWhW91, dJ91]. Dynkin [Sch91a].

each [DS03]. Early [Ebe03, KsLL00]. Easy [LHK^{+}10, Mor92, GPT89, GV00]. edit [ADG99]. Editing [Doo02]. editor [Pai92]. education [HLY86, WKZZ03]. educational [Jur86, Jur01]. Edusym [Jur86]. Effective [BD89, Sch12, vdH06, El 13]. efficiency [CK02]. Efficient [Bod04, BGHS99, CT98, DSS04, DPW05, EG96, Gie97, HK98, HKL99, KS11c, MC91, MSW07, SL93, Sho99, SL95, Wei89, ZL09, ALP03, BDPZ09, CHT^{+}07, CK92, Ebe97, EGG^{+}07, Fau02, GKW94, GL89, HMS01, Lel12, MR11, Mil89, MM97, MF97, Roc09, SY11, SKCL12, TCT97]. efficiently [Miy00]. eigen [PZ10]. eigen-solving [PZ10]. eigenpolynomial [Ast91]. eigenvalue [For01, GM06, Jea00, MMS07]. Eigenvalues [YT93]. Eigenvectors [YT93]. Eighth [IEE95]. element [Rup10, SW89, VF92]. elementary [Bad94, Bea93, BBD03, BBDP04, BD02, Bro89b, CCF89, CC90, Cha03, Ric92, RF94a, Ric95, Ric96, Str09]. Elements [Lev12, MZ96]. Elimination [MS03a, Bas97, BD07a, BS10, CJ89b, DGS98, DW00, FH97, GLV89, Hon92, Hon93, HD09, LI00, McC99, McC01, MB09, ST11b, Wei94b, Wei97, Wei99]. Elkies [GM06, MMS07]. elliptic [BM91, BLvH04, CA07, Mil11, Mor92]. embedding [Sch91b]. Employing [MR89]. encoding [Pan00]. encoding/decoding [Pan00]. encyclopedia [MS03b]. endomorphism [GHS00]. energy [BL91, SS91, Vin91]. Engel [CdG07]. engineering [SF90, TD90, TMH94]. England [ACM94]. Enhanced [LSS13]. enhancement [EK95]. Enlarging [BHPS86]. Entries [YT93]. enumerating [KM89]. Enumeration [GS06, Tsa96, Bur11b, CG01, Hav91, RMM98]. Environment [ZB94, BCCG92, Dew89, DW90, FLMS06, Kor01, Pur89]. environments [LCJ98, Pur86]. enzyme [CR94]. equal [LMPX08]. equality [HD00]. Equation [Hub96, van96, Bro89a, Bro91, FS04, Gat90a, Gat90b, Hie92, LCW05, Mac89, OS91, Roa94, TD90, Zha93, vdW05]. equational [BM05, KK89, McC99, McC01, MB09]. equationally [Che90b]. Equations [Abd89, Cod90, PS93a, Tho96, AK91a, AP94, ABP95, Abr95b, AZ96, Av97, Abr08, Abr10, ACFG05, Bar89a, Bar89b, Bar99, BP07, BBP08, Bar10, BS09a, BJ98, BOS13, BCS05, BCS^{+}07, BS09b, BSC^{+}12, Bou99, BGU03, Bro86, Bro92, BF99, BL02, CKY89, CvH09, CDS91, Cod92, Coh13, CGR10, CSU00, Cor01, CG92, DvHK08, DT86, DRJ97, FvH11, Fat89b, FR09, FsG04, FM01, FS04, Fre00, FSK86, GPP91, GVBv94, GC91, GKM^{+}04, GKZ90, GKZ91, Ger98, Gie97, GLV89, Gri89, Gri90a, Hon93, Kal90, KYAH09, KS10, KS11b, KR93, KMH89, Koh90, KF91, KvH13, Mac99, Man94b, MK99, Mor06, Oak94, Oak12, Oku90, Pec91, Pet91, PS93b, Pfl08, RGL94, RB91, RW00, Rei00, Ren91, Ric96, Sch89, Sch92, Sch96, Sch07, Sha89b, Sha89c, Sin89, SU92]. equations [Sti97, Tsa99, Tsa05, UC90, Wei94a, YC92, Yok04, van98, vHW05, vH07, vHY10, vHL10, vzGG97]. equidimensional [Lec00]. Equilibria [Gut93, Dat03, FS12]. equilibrium [KR93]. Equivalence [Gri96, BP98, Bar99, DP07, Gon86b, Sha89b, Sha89c]. equivalences [WJLL93, WL93]. equivariant [JX07]. erasure [Pan00]. erasure-resilient [Pan00]. Erlang [Wik94]. Error [Mor89a, MGv91, MBC92, SY98, YZZ13]. Error-Correcting [Mor89a]. errors [CKP12, KY13]. ESF [MS03b]. estimable [CR01]. estimates [BCdP13, DS04]. estimation [Sha89a]. Euclid [Sed01a]. Euclidean [PW02, Pot96, Sor95]. Euro [LGG97, LGG97]. Euro-Par [LGG97]. Euro-Par’97 [LGG97]. EUROSAM [Ng79]. evaluate [CZ00]. Evaluating [CBdCD97, Gem01, KM95]. Evaluation [BGDW95, HK94, KRS94, KRL93, LF95, AB01b, BWZ94, BCR08, CHT^{+}07, Edn97, JKM13, KN07, Mil89, MH97, MF97, Nor00, Nor06, STK97, YL91]. Evaluations [DK95]. evaluators [Fat96]. even [Alb12]. evolution [GZ90, GKZ90, Ren91]. Exact [EKM96, Fuk08, KLYZ08, KJ94, LCW05, Ste13, Yap09, CR91, CS05, CK00a, CT91, EKW07, HD00, KY07, SY11, SWY11, Vil07b]. exactly [BF91]. Examples [GMR94, Wav99]. EXCALC [CDS89]. exchange [GKW94]. Execution [GT94a, SRHW94, Bel90]. Existence [uGKO92, RRW99, Sch91d]. expansion [CEMP05]. Expansions [RSSV96, Sha93, BS09a, PB90, SS92]. expected [EGT10]. expedite [BWZ94]. experience [DSV00, Kry86]. Experimental [But92, Gla97]. Experimenting [TD89]. Experiments [AW89a, Arn89, Bon97, CJSW94, RSV91b, RSV91a]. explains [CJKS13]. explicit [BLRS13, CG92]. exploited [Vin91]. Exploiting [GZ03a, Bel90]. Exponential [KM96c, Ric93, AMW08, Fre03, Mai98, Pfl97]. exponentiation [Nöc00]. exponents [PW06, Yok04]. Expressing [KK08]. expression [BLRS13, Car04, Sas86]. expressions [ASN89, Cha89, DS97, Gon86b, GKW94, IK91, Jef93, KS07, OW05, PU98, Sch05a, Tsa97, YF04]. Extended [GP97, CT98, CO94, ER94, Fat96, Ger98, uGKO92, Iva00, MH95, MS98, Nöc00, PB99, Sto97, vG96, YZ86, vH96, vzGN99, vzG09]. extensible [Kaj92]. extension [ST11a, ST12]. extensions [BR12, BLW05, DP95, DD89b, Ebe97, Fre00, Fre03, vHM02, KS97a, Lev13, PW02, Sch04, Sho99, Tsa97]. extent [Jef93]. exterior [Gri89]. Extracting [YF04]. Extraneous [KS97b, Leh97]. extrema [Fat92].

facility [Gat86]. factor [Vil07a]. factorable [AZ97]. Factoring [BCGW89, CGK^{+}13, Enc97b, GZ03b, KL94, LST02, Min03, SGL04, vH96, BP98, Bro94, CGvH^{+}01, GS10, HvHN11, IKS09, KK05, Sho91, YNT90b, vH13]. Factorization [ASZ00, Clu03, Fre03, FM96, Sie94, ZL12, van96, Bar99, BLS^{+}04, Chè04, CGT97, CGKW02, DGT91, DCS05, Enc97a, GKM^{+}04, JM09, KS97a, Kal03, KYAH07, KSS07, Li07, Luc86, MSW13, NY02, Pan01, Sas01, Sch89, TW91, Tsa97, Tsa99, Tsa10, vG96, Wan90, WM89, Zen09, vzG06]. Factorizations [Tsa96, CL01, KLYZ08]. Factorizer [GL94]. factorizing [Hie92]. factors [Bea93, DPS09, Enc97a, KK06, KS97b, Kir98, LM02, Sek07]. faithful [dG97]. familiar [RL02]. Families [CR97, CR01, Raa13]. fan [DJLS07]. Fast [ABM99, AK91a, Ahr96, Bin96, BCS05, BS09b, BCSL12, Bro89a, Bro09, DDS13, EKW07, EGNS12, FM11, FS09, GM06, Gie94, Gie95, GRA00, HP97, Iva00, KKS86, KS97a, KM06, Kir92, Kir98, LMS07, MW94, Rák95, Sch91c, Sim98, SDKC10, SL96, Tho01, Vil94, vzGG97, Bur11b, CL89, CF92, FLMS06, GZ03a, Ger10, IKRT91, Kai94, Kal94, KMZ99, KC09, LYZ10, Lic05, Luc86, MF97, PS13, Rei97, Sag12, SW04, Sho91, VZ97, YZ86, Pos11]. Faster [BRS09, EGG^{+}07, PS07, Bro86, CKY89, MT00, Rio02]. fat [CO01]. feasibility [BRS09]. feasible [ZR99]. Features [Vin91]. February [BW98]. FELIX [AK91b]. Fenchel [BH06]. Feynman [KR86]. FFPACK [DGP04]. FFT [CM96a, JX07, vdHLS13]. FIDE [LD90]. Field [DG95, ARS99, BF12, DGP02, DGP04, Ebe10, Lan91, Lev13, Ren06, RY08, Rio92, Ryb91, SZ91, ST11a, Tsa97]. Fields [Bin96, KP94b, KL96a, Sha93, Abe90, Alb12, BGM91, BFP12, BCR13, BM91, DP95, DDS13, Ebe97, Ebe00, Ebe03, Enc94, Enc97b, FS09, FGT01, GRS01, vHM02, HE05, Iva00, JM07, JM09, KS97a, KK06, LL93, MW00, MQR98, NY02, ORY09, Pos11, PR08, Rol89, Sch91b, Sho91, Sho99, ST12, TW91, YNT90b, ZR99, vHM04, vdW05, von90, vzGN99, vzG01]. find [Dat03]. Finding [BGM91, JKM13, KK06, Lin04, MS97, PS93a, PS93b, Sch05a, SS94, WR13, YLN95, vHY10, AH08, DPS09, For01, Gem07, Kau08, jL90, MvV89, Oak94, PZ10]. fine [BGK94b, MH97]. fine-grained [BGK94b]. Finite [DGP02, KL96a, Min95, Wei92a, Bad94, Bay04, BGM91, BFP12, BCR13, BM04, BM91, CIK97, DDS13, DGP04, FM06, FS09, GR10, HE05, Iva00, KS97a, KP10, LL04, LL93, Mez10, NS06, NY02, PB99, PR08, RC06, Rup10, SW89, Shi90, Sho91, Sho99, Sim97, TW91, Wei89, tW05, vdW05, von90, vzGN99, vzG01, LD90]. finite-dimensional [Shi90]. finite-rank [LL04]. finitely [BP91, BOLP95, CdG07, MQR98, RGSGG00, dGW99]. finitely-presented [CdG07]. finiteness [BBR93, Sal05, Zob05]. First [AH94, Ait95, AB98, AMBSW11, CG92, Pfl97, RSV91b, RSV91a, Sch91d, Sin89, WBD^{+}94, Zha93, Hon94]. fitting [Hul13]. fixed [Lel12, MZ11, Wat89]. flash [SDKC10]. flat [AL08]. floating [CK00b, GZ03a, KLYZ08, MBC92, Rum10, SY98]. floating-point [MBC92, Rum10, SY98]. flow [CA07, MYT09]. flows [KR93]. fluctuations [Pol98]. Form [CM95, BWZ94, BCEW12, BLvH04, Car07, CDS91, CD99, DSV00, GPP91, Gil92, Jea02, Khe02, MW01, RB91, SR89, Sto98, Vil93, Vil94, WZ08, van95, Gie95, SW04, SL95]. Formal [AMBSW11, DT86, Mac99, Pec91, Bar89b, BC89a, BHP07, Che90a, CG90a, DKLM98, HW86, vdH02, RRW99, van97]. Forms [Val93, Vil96, Abe90, ALP03, BP07, BLV99, BLPR13b, BD08, BS07, CL89, Che89, EGNS12, GLL04, Gie94, GS11, HC05, HL12, KM02, Man08, Mil11, Mou05, Sch91c, SP91b, SP91a, Sto96, SL96]. formula [AP05, Bro99, CG90b, CK06, Fee90, Hon92, Khe02, Oku90]. formulae [EM09, PDS05, Tao90]. formulas [ADG99, Bro09, Hon93, Str10, Str12b]. Formulation [Bac96, CK00a, KS97b]. Formulations [KS95]. Fortran [BGV94, FT03, WW89]. Foundations [RCK04]. Fourier [KS13, Pos11, Arn13, HR10, Mil89, vdH04]. Fourth [ACM93, BGU03]. FOXBOX [DK98]. Fraction [BCL97, BCL02, BL09, BS93, KK08, Kir92]. Fraction-free [BCL97, BCL02, BL09]. fractional [BG11]. fractions [BLRR13, Kir98, Sha07]. fragment [CC90]. frames [SM08]. Framework [BSG96, LD11]. franca [LHK^{+}10]. France [BGY96, FR95, Ng79, Mor02, vdHvH12]. Fraser [Doo99]. Free [MZ96, BCL97, BCL02, BL09, BD08, CO94, GR97, Giu09, KMO92, LSS13, Ste08, YZ86, dGW99]. Frobenius [Ebe00, Sto98]. Full [BS93]. func [GLL04]. Function [Bin96, Roa96, Abe90, Ada01, BL90, BLvH04, CKS09, CJK97, CDHJ07, GK05, GT90, GR92, JM07, JM09, KN11, KY13, KM06, MS02, Miy06, Ryb91, Sch91b, SZ91, Zie91, Zip91, vHM04, vHY10]. Functional [Bra94, DMN94, Ham94, Sch94, AB01a, BBP08, BLW05, Dal92, GRS01, SS92, Sha89b, Sha89c]. Functions [BS93, CdR96, RSSV96, ArRL89, Abr95a, AM90, BWZ94, Bad94, BBD03, BBDP04, Beu12, BCS^{+}07, BCCL10, BLS13, BCC^{+}13, BD02, Bro89b, BL02, Car07, CFFL11, CKS12, Che90b, CA07, CGL91, DvHK08, Gem01, GVT95, GKZ12, HL86, JLVR97, JR98, KY07, KYAH08, Kau08, Kir92, Kno86, Koe03, Le02, MS03b, Mez10, MG94, NM97, Raa13, RL89, RF94a, Roa97, Sha89a, Str08, Str09, Tak90b, Val89, YL91]. fundamental [SZ91]. Further [Nor00, Vil97]. future [Bar03, Buc05, Roy12].

G [BGL05, BG07]. G-graphs [BGL05, BG07]. Galois [Baj86, Col97, CSU00, FFdlAGM00]. Gamma [GK94]. Gamma-Operation [GK94]. GAP [BL98, Miy06, RER09]. gauge [KR86]. gauges [BC92]. Gaussian [CJ89a, FH97, Kai94]. gcd [BB07, GR97, MH95, CCC98, Gem97, vHM02, JM07, Jeb93, KM99, KSS07, LYZ10, LF95, Lic05, MS98, RW94, RWW94, Sed01a, Sor95, Sto97, Ter09, Web94b, YZ86, ZD04, dKMW05, vHM04]. gcd-free [GR97]. gcds [Sme89, BCLV13, Enc94, CL06]. GCUR [GV91b, Gon91b]. GENCRAY [WW89]. gene [Far06]. genera [SP91b, SP91a]. General [Hub96, ACFG05, Cha90, FsG04, MT12, RGL94]. general-purpose [Cha90]. Generalization [Jeb93, SD86, Sti89]. Generalized [Can89, GS05, Mou05, Tsa05, AW89b, BJ98, BS11a, Gem01, GEZ10, KMO90, MPV98, PO08, SW11]. generated [BP91, BOLP95, Far06, MS03b, MQR98, PW06, RGSGG00]. Generating [CT96, JX07, vHKN11, ArRL89, PB99, RL89]. Generation [Bac96, Edn93, RCK04, BGV94, Coo92, EF89, Gat86, GZ90, Kor01, MM97, SW89]. generator [Mil89, WW89]. generators [CO01, GL11, JM10, Tho01]. Generic [AJL01, EKM96, KMY05, MS89, CGL09, FvzGP10, KS11a, Mou93, SS03]. genericity [KM99]. genetic [kMG97]. Genoa [Tra06]. genre [Gri89]. GENTRAN [BGV94]. genus [Bur11b]. geodesy [TSD91]. geometric [CGKW02, EKW07, EM99, Giu09, KSY94, Lec00, LkGID97, Li04, Li07, LH08, RF94b, SF94, WZ08]. geometric-numeric [CGKW02]. geometrical [Sek08, Win89a]. geometries [Cho90]. Geometry [Kap86, Sti93, BS86, BS07, CG90b, CGZ93, El 13, GKWW91, GMR94, HLY86, HS08, KW90, KS86, Kut89, LY86, LL93, Par13, RF94b, SF94, San04, Ver04, WR06, Yap09]. Germany [CL96, LGG97, Glo98, Wat91, Wat10]. giant [Kal02]. Given [DK95, FIK86, GKZ12, MMM91, Sek08]. Global [GEZ10, Oku90, Din08, Fat92, KLYZ08]. globally [FS12]. gmp [GKZ07]. gmp-based [GKZ07]. Good [PR08, BGM91]. Gosper [AP05]. GRADIENT [MN93]. Graeffe [Gem01]. grain [MH97]. grained [BGK94b]. gram [MZ11, EKM96]. Granularity [LHD94]. Graph [Kes94, BM04, Gla97]. graphs [ASN89, BGL05, BG07, MSW07, OS04, RB91, RV93]. Grassmannian [Ruf07]. Greater [GPT89]. Greatest [DK95, KL96b, AuK93, KYZ06, Küc91, LN97]. green [WL93, WJLL93]. Grenoble [vdHvH12]. Griffiths [BLS13]. Gröbner [BW98, KFS86, MMT92, AGK96, AH08, BS98, BF03, BW91, BD08, CD11, Cab93, CCT08, Cho90, CdG07, Cox07, CG86, Cza91, Dah09, DJLS07, EP11, ER13, FR09, Fau94, Fau02, Fau09, FES10, FM11, FES12, FEV13, FS13, FSK86, GGV10, GZ08, Giu09, GL94, Hie92, Hon96, KSW10, KSW11, KM96a, KM96b, KM96c, KN07, LSS13, Lev07, MR93, Man94a, MMM91, MMM95, MR10, MR11, MZ01, Mor89b, Nab07, Nag11, PW06, Pet91, RS12b, Sat98, SS01, Sie94, Ste97, Sti93, SW11, SWMZ12, SS02, SS06c, Tak90a, Tak90b, Tra98, TD89, Tra89, TZ02, Wei92a, Wei02, Win89a, ZW06, dGW99]. Gröbner-free [BD08]. Group [MR93, Uma06, Bea93, BGU03, But86, Col97, CFL89, CSU00, FS12, FS13, FFdlAGM00, GSGVPN^{+}09, Hul13, KR86, KM89, Mül98, Rei00, RF94b, SF94, Sha03]. Group-theoretic [Uma06]. group-theoretical [KR86]. Groups [CFT95, Min95, Rák95, Thu96a, ASW05, AW89b, BCFS91, BBR93, Bay04, BIL91, CS97, CFS90, CF92, CFYT94, CR03, CO94, EP98, FM06, Hol00, KS13, KMO92, KC09, LRW94, LM02, Mar95, Miy00, Miy06, Miy07, Mor95, NS06, PB99, RW00, RC06, RER09, SS94, Sim97, Sim98, Thü96b, Wei94a, Wis89]. growth [Sha99, Sim98]. Gsolve [Bro86]. Guaranteed [Bro99]. Guarded [DS97].

H [MSV09]. H-LLL [MSV09]. Habicht [GLRR89a, GLRR89b]. Hagenberg [Hon94, Jef08]. Hagenberg/Linz [Hon94]. Half [Lic05]. Half-GCD [Lic05]. Hamel [CA07]. handling [BC00]. handprinted [Mar91]. Hankel [BP93, Hit05, SL93]. Hankel-like [BP93]. hardness [BCR13]. hardware [Bis12, GZ03a, YZ86]. Hardy [Sha93]. harmonic [CEMP05]. having [BLvH04]. Hawaii [Küc97]. heaps [MP09]. Hecke [CT96]. Heights [Jen76, CGK^{+}13]. held [Ait95]. Hensel [Ber98, BL12, SY98, YT93]. Hermite [BCC^{+}13, CL89, GS11, HL12, MW01, SR89, SL95, SL96, Vil96, Vol03]. Heuristic [GGS89, Cza91, Kau08, LF95]. Hierarchical [BGK96, GG89]. High [Ahr96, CK93, CT96, Fat03, FP09, Ger98, JKL^{+}06, KRS94, Sto02, BP06, CDS89, CR03, FLMS06, GZ03a, KL94, KS97a, Päs09, SS91]. high-degree [KL94]. high-dimensional [CDS89]. High-level [Fat03, FLMS06]. High-order [Sto02]. High-performance [JKL^{+}06]. High-Precision [Ahr96, BP06]. higher [BCB09, BBP10, Bod10, Hub04]. higher-order [BCB09, BBP10]. Hilbert [CdR96, DP95, GKZ12, Kap86, Lel12, LLMM08, MS02, Rou10]. Hirota [Hie91]. HJLS [CSV13]. HNF [BF12]. holonomic [CKS09, GK12, GS10, IK12]. Homogeneous [RS96, Wei92b, Bou99, Eri04, FEV13, KR08, Li04, McC12, Miy07, Rab89, UC90]. homology [DSV00, RER09, BdC09]. Homomorphism [Bra94]. homotopies [SVV10]. homotopy [GS98, RVWW05]. Honest [Fat92, ABK95]. Householder [MSV09]. huge [EGL04]. hull [CK04]. humans [BC00]. Hybrid [CCG05, CK93, DE01, DMMRY01, GF92, KZ06, Kor01, Nam95, Sza11, BFP12, KsLL00, NM90, Tiw10]. hyperbolic [Hol00, Tsa05]. hypercircles [RSV04]. hyperelliptic [Ber94, SS95, Sha03, SZ91]. Hyperexponential [LL04, BR12, BCC^{+}13, GLL04, JKM13]. Hypergeometric [AB00, PS93a, Roa96, AP01, Abr02, ALP03, AM90, Beu12, CK12, CHT^{+}07, HL86, KvH13, PS93b, YL91]. hypersurface [Sch12]. hypersurfaces [BD04, LM94]. hypothesis [uGKO92].

IAMC [WGK^{+}01]. IBM [RLZ86]. IBM-PC [RLZ86]. ideal [BOLP95, BM04, CT98, Lap06, MP06, RGSGG00]. Ideals [KM96b, AJL01, BL10, BdC09, BL00, CR97, CO01, Eri04, FS13, FGT01, GY06, GS10, JFRS06, JFSMR08, JZXM01, KFS86, KM96a, LL04, Lev06, LST02, MMM91, MR10, Mil04, MZ01, PW06, Rou10, SS09]. identifiability [Oll89]. identification [CR94]. identities [Bur91, Cha03, CKS09, HP89, IK91, Kau04]. identity [RF94a]. IEEE [IEE95]. if [RGSGG00]. II [BCGY08, DGT91, DD89b, Hon96, Wei89]. III [BC92, FNM86]. image [LYZ10, Lin04, Sti97]. images [von90]. imaginary [Rol89]. Implementation [AP06, BGDW95, DJ94, FLMS06, GK94, MF94, RG91, TL94, Web94b, Ber94, CL96, CG90a, DGW97, Fra90, GKZ07, GS89, Gla97, Gon86a, GR92, IKRT91, KSI97, Lel12, Luc86, MW00, MF97, NY02, Rio92, RMC07, SLA89, VM99]. implementations [JKL^{+}06, LPR99]. implementing [CG86]. implicit [DMR08, MH97, WR07]. Implicitization [BD07b, GVT95, LM94, WCD08, Bus01, RSV06]. impossibility [GKZ12]. imprimitive [BGU03]. imprimitivity [SS94]. Improved [Bro00, KY89, BFP12, BS00, Bro94, KVY89, MGv91]. Improvement [OIiM^{+}97, Hon90, Kir98, Miy06]. Improvements [Hub04]. improves [GR11]. Improving [GSW12, HP97]. in-place [HR10]. inconsistency [GLS98]. incremental [GGV10]. Indefinite [Abr95a, AP01, MW94, Sch05a, Wan91]. Independent [DMN94, DW90]. Index [Tho96, BBDP05, GPP91]. indices [IK91, YL91]. Indicial [KMG96]. inductive [MS92]. inequalities [FMRD08, GK05, GEZ13, HC05, Rab89, Sch09]. inequality [CMS89]. inequations [Mor06]. inexact [Zen03, ZD04]. infallible [CK92]. infeasibility [LLMM08]. inference [Web94a]. infimum [GE11]. infinite [AH08, LRB95, Tak90a]. infinity [Por89]. informal [BC89a]. Inheritance [Har94, DR01]. inhomogeneous [AZ96]. initial [BGP12]. input [CKP12, Mar91]. insimplification [RGL94]. Integer [BZ07, DSV00, Gie95, KJ94, SL96, Sto09, Web94b, CS05, FH97, GKZ07, Gie94, Jeb97, LkGID97, PS13, Rei00, SW04, SR89, Ste13, Sto96, Vol03, Wei99, YZ86, Zan10]. integers [BS13, BF12, CJ89a, MW00]. Integrability [BOS13, Bou05, GZ90, GKZ90]. integrable [BCEW12, FM01, Hie91]. Integral [GT97, YLN95, Bün91, EF95, Küc91, Ryb91, Tak90a, YC92]. integrals [AM90, Ada97, AMBSW11, Ber94, BS11a, CG92, Kir92, MH01, Raa13, Sch91d, Sin89, YL91]. Integrated [GL91, Dew89, DW90]. Integration [Fat81, Jef93, JLVR97, Kau08, Kno86, Raa13, Av97, AGLM99, BSG97, Bad94, BLRR13, CCG05, CA07, CG92, Ged86, GS89, GF92, JR98, NM90, Nor90, RB91, Rup10, Rus92, SW10, Tak90b, Tsa05]. Integro [RR08, RRM09]. Integro-differential [RR08, RRM09]. integrodifferential [Tsa97]. intelligent [ABK95]. intensive [SKCL12]. Interactions [Fau09]. Interactive [Coo95a, Coo95b, Doo98, Doo02]. INTERCOMP [GL91]. interconnection [Pur86]. interesting [Ser89]. Interface [BHC95, Sch94, DW90, Kaj92, LHJ92, Leo86, OW05, SS86]. Interfacing [vdHL13]. International [ACM94, AH94, Ait95, Ano95b, Bro93, Bro07, CL96, Doo99, Gia89, Glo98, Gon89, Gut04, Hon94, Jef08, Kau05, Küc97, Lak96, LGG97, Lev95, MCG13, Mor89a, Mor02, Mou01, Ng79, SE11, Sen03, Tra06, Tra00, Wan92b, WN90, Wat91, Wat10, vdHvH12, FR95, May09]. Internet [Wan99]. Interoperating [RER09]. Interpolation [MF94, BW91, BCR08, BJ10, CKP12, GLsL06, GR11, uGK91, KY89, KLW90, KsLL00, KY07, KN11, Tao90]. interpolations [ZR99]. interpolatory [CGR10]. interpretation [Che90b]. Intersection [Lev06, ALM09, MT00]. Intersections [Nam95]. interval [Fat92, GMB01, Joh92, LJM08, PU98, SS06a, Sek07, Sha89a]. intractability [MR90]. Introduction [Ham94, BL90]. introductory [LHJ92]. invariance [McC03]. Invariant [Val93, BLRS13, Bro98, FS12, FS13, Kem08, PS13]. Invariants [DG95, RCK04, CvHL10, HW86, Hub10, HL12, Lev13, LI00, SM08, YZ05]. inverses [JM10, SS92, SS09]. Inversibility [Oll89]. Inversion [BD04, Ebe97, EGG^{+}07, GT90]. investigate [CDS89]. Investigation [Gut93, AZZ91, RLBS91]. investigations [GVBv94]. Invited [Ham94, Wan94, Ida09]. involutive [WZ08]. involving [CCF89, CGL91, GK05]. IRENA [Dew89, KR93]. Iris [Leo86]. irreducibility [Gri90a, KM03, Nag02]. Irreducible [Min95, vzG01, Ast91, BP07, CvH09, GR01, MS97, Mil04, Zen09, vHL10]. irredundant [Mil04]. IRREGULAR [FR95, Che90a, Som94]. irregularly [FR95]. ISAAC [Gut04]. Isabelle [BHC95]. isolate [Sag12]. isolated [jL90]. Isolating [MS09]. isolation [CGY07, CGL09, CJ89b, CK92, GG12, JK97, MM11, MSW13, SY11, Sha07, SY12, Str08, Str09, ST11a, ST12]. isomorphism [Han07, Shi90]. isomorphisms [Du94, Miy00]. isotopic [BCGY08]. ISSAC [Gia89, Gon89, Jef08, Wan92b, ACM94, Ano95a, Ano95b, Doo99, Glo98, Kal96, Kau05, Küc97, Lak96, Lev95, MCG13, Mor02, Mou01, SE11, Sen03, Tra00, WN90, Wat91, Wat10, vdHvH12]. ISSAC’93 [Bro93]. issues [BLS^{+}04, HS02]. Italy [Gia89, Mor89a, Tra06]. iterated [For01, Raa13]. iteration [Gem01]. Iterative [Zan10, BMP98, Fre00, GSW12, Rua01, Ter09].

Jacobi [Moo11]. Jacobian [HS02]. Jacobians [Pai92]. Jameson [GVBv94]. Janet [Sch96]. Japan [BGY96, HI93, WN90]. Java [BCK99, Nor00]. Jeffery [CA07]. Jordan [Gil92]. Jose [SE11]. July [ACM94, Bro93, Bro07, Cha86b, Doo99, Gia89, Gon89, Gut04, Jef08, Kau05, Küc97, Lak96, Lev95, May09, Mor89a, Mor02, Mou01, Tra06, Wat91, Wat10, vdHvH12]. June [BGY96, IEE95, MCG13, Ng79, SE11].

Kaltofen [Laz06]. Karatsuba [Jeb97]. Karlsruhe [CL96]. kenzo [RER09]. kernel [tW05]. key [CCT08]. KIAS [May09]. Kiev [Bro93]. kinematic [RGL94]. kinetics [CR94]. Kingdom [ACM94]. Klein [vHW05]. knapsack [Chè04]. Knowledge [ABP94, HLY86]. knowledge-base [HLY86]. Knuth [BD86, Win89b]. Korea [May09]. Koszul [BdC09]. Kovacic [vH93]. Krylov [Ebe04]. Krylov-based [Ebe04]. Kumar [AP06]. KUX1 [FSK86].

Labeled [GL89]. Lacunary [KK05, CGK^{+}13, GR08, KK06]. Lagrangian [Ruf07]. Lagrangians [CDS89]. Lambert [CJK97, CDHJ07]. Lanczos [EK97, Ebe10, HE05]. Language [KNKM94, BCM94, BC89b, Dal92, Gil92, GL91, Kor01, OIiM^{+}97, VF92]. Languages [DMN94, KRS94, MT94, BHP07, CCF89, HI93]. Laplace [SM08, Tsa05]. Large [SY09, BJS07, CFYT94, DSV00, GKZ07, KC09, Mil04, RC06, RMC07, SDKC10, Vol03, Zol97, von90]. Lattice [Bin96, BLWhW91, Vil92, Vil07a, Vil11, vdHLS13]. Lattices [Len13, CSV13]. Laurent [LSWZ06]. law [Ruf07]. layer [GVBv94]. Lazy [KRS94, van97]. least [Sau01]. left [BCSL12]. Lehmer [Sed01a, Sor95]. Leibniz [AP05]. lemma [Pfl08, BMMT94]. letterplace [LSS13]. level [Fat03, FLMS06]. lexicographic [Cza91]. lexicographical [Dah09]. Library [Coo95a, Alb12, Ces97, CS05, Coo95b]. Lidskii [Jea00]. Lie [ASW05, BK86, CdG07, GKZ91, dG97, HS06, LÜ10, LRB95, MZ96, RLBW92, Sch91a, VT98, Zol91, Zol97, dGW99]. Lifting [DMS^{+}05, BL12, Sch02b, Sto02, Wan92a]. lifts [MMS07]. Lightweight [DKLM98]. Like [KP94b, Bod10, BJS07, CGR10, BP93, Kal94]. Lille [Mor02]. Limitations [Baj86]. limits [GG89]. linBox [KMY05]. line [FIK86]. Linear [Bac96, Bro92, CD11, Fre00, KP94b, PS93a, Tsa96, AK91a, AP94, ABP95, Abr95b, AZ96, AB98, AvH99, AB01a, Abr08, AK12, Alb12, AW00, ACJ03, BCFS91, Bar89a, Bar89b, BP98, BP07, BBP08, BCB09, Bar10, BBP10, BR12, BS09a, BOS13, BCR13, BAE^{+}00, BCS05, BCSC06, BJS07, BCSL12, Bou99, BGU03, BLPR13b, Bro94, BF99, BL02, BLW05, BLvH04, BD07b, CKY89, Cet92, CvH09, CCT04, CS05, Cod92, CSU00, Cor01, DT86, DRJ97, DGP02, DGP04, EGG^{+}06, FvH11, Fat89b, FES10, Fre03, FSK86, GM86, GR10, Gie97, GLS98, GSW12, Gri90a, GS08, GSZ11, GEZ13, HW86, JKM13, Kal94, KNS11, KS11a, KS10, KS11b, LÜ10, LST02, Mac99, MC92, Mar95, MW01, Mor95, MS99, MS00, MQR98, Oak94, PS93b, Pfl08, Pfl97, QV95, RTZ03]. linear [Rei00, SWY11, SS94, Sch89, Sch07, Sch09, SU92, Sit91, Ste13, Tho01, Tsa97, Tsa05, UC90, Vil97, Vil07b, Vil11, Wei99, WR06, YC92, van98, vHW05, vH07, vHY10, vHL10]. linear-trigonometric [AW00]. Linearizablity [KR08]. linearization [CDS91, CD99]. linearly [KYZ06, Min03]. lingua [LHK^{+}10]. links [Bur11b]. Linz [BW98, Hon94]. Liouvillian [CvH09, CCT04, Cor01, Kno86, Sin89, SU92, UC90, vHL10]. LISA [MS92]. Lisp [Fat91, FT03]. List [Bra94]. List-Homomorphism [Bra94]. Lists [Tak94]. Littlewood [SVV10]. LLL [BF03, MSV09]. Local [DJ94, DW00, GS90, CvHL10, CGL09, EGNS12, HMS01, Sed01b, Vol93a, Vol93b]. locally [Wei89]. Locating [SS06a]. loci [FES10, FN97]. Loewy [GS05, GS08, Sch07]. Loewy-decomposition [GS05]. Log [DJ94, BS11a]. log-sine [BS11a]. logarithm [Oak12]. logarithmic [Ada97, Kau08]. logarithms [Fre00]. Logic [BGH94, DMN94, KP94a, LC94, LHD94, Bel90, CBdCD97, OIiM^{+}97, Yam90, BKLG97]. logically [LuS91]. logics [OS89]. London [Mou01]. long [Li07, Zan10]. Lookup [Fat89a]. Loop [RCK04, BC92]. Low [BCSC06, BL91, Ger98, MZ12, MR10]. low-dimensional [MR10]. low-energy [BL91]. low-rank [MZ12]. Lower [BY09, Koi12, Vin91]. Lubbock [IEE95]. Lüroth [BLRS13]. Lyon [FR95].

M4RIE [Alb12]. MA [HI93, MCG13]. Macaulay [BS86]. Machine [SRHW94, Ber97, RSSRV89, VF92]. Machines [Gri96, BC00, Lam97]. Macintosh [DR86]. MACSYMA [EF89, CG90a, Gol86]. Magma [BCM94, Ste08]. magnetic [Bou05]. maintenance [ASN89]. make [JT94]. mammograms [EK95]. management [KMY05]. manifold [Bur11b]. Manifolds [Val93, FGPF89]. manipulating [DK98, FIK86, Wan91]. Manipulation [Ng79, EF89, Fat90, LMMRRR13, MM97, Oku90, Tao90, Tsa97]. manipulator [Jur86, RGL94, Yam90]. many [Kau04]. MAPLE [GS89, GF92, WEM98, AB01b, BHC95, Ber97, CA07, MN93, SF90, Sie93, Ste08, VM99, YC92]. MAPLE/MATLAB/C [WEM98]. MapleTensor [KMG96]. mapping [MS02, Sey86]. mappings [Oll89]. maps [FN97, SW01]. March [Sin93]. Markovian [MS92]. markup [Doo98, Doo02]. MARS [WEM98]. Marseille [Ng79]. Massey [CKP12, GKsL02]. massive [Hie92]. Massively [DD89a, INPS93, Ber97]. Master [BGK96]. Master-Slave [BGK96]. matcher [Coo86]. Matching [MT94, Fra90]. math [BSG97]. MATHEMAGIX [vdHL13]. Mathematica} [Har00, Soi95, Ste08]. Mathematical [Soi95, ZSST96, BC00, BC89a, Doo98, Doo02, Fat03, GKW94, LH99, Päs09, Pur86, Pur89, SD86, TSD91, Wan97, Wan99, YF04]. Mathematics [WKZZ03, BF94, CGG^{+}86, Dev89, Fat96, Fat97, RG91]. MathScribe [SS86]. MATLAB [WEM98, Ste08]. Matrices [KL96a, SL95, SL96, Vil96, YT93, ABM99, AW89a, ASW05, BLV99, BCL02, BBCM11, BP93, BZ07, Che89, CS05, DE01, DE02, DSV00, DPS09, EGNS12, EP97, FES10, FM11, Gem97, Gie94, GS11, Hit05, HS02, JFSMR08, JP99, KKS86, Kal92, Pan89, PS13, SS11, SW04, SL93, SS06b, Sto96, Tur05, Vil93, Vil94, Vol03, ZL13]. Matrix [CFT95, Gie95, Pan00, BBR93, Bar99, BCL97, BB07, Bod10, BDPZ09, CFYT94, DCS05, Ebe97, Ebe04, EGG^{+}07, Gem07, Gil92, GJV03, GSGVPN^{+}09, dG97, GEZ13, HP97, IKRT91, Iva00, Jea00, Jea02, JM10, KMY05, KC09, MZ11, MZ12, MSW07, Mül98, Pai92, SY09, Sch09, SSW09, SR89, SV05, Sto09, Str12a, Tho01, Uma06, GRA00]. matrix-based [BB07]. Maui [Küc97]. Maximal [Mon04, RW00]. Maximizing [Lam97]. maximum [Jef93]. May [ACM93, Ait95]. MBA [PQZ11]. Means [SS93, Sti93, BD04, GS90, GSV91, KRL86, Lec00]. Mechanical [KRL93, CG90b, VC89]. mechanics [TMH94]. Medical [IEE95]. meeting [Jef08]. meets [Sag12]. meganode [RSV91a, RSV91b]. Meijer [Roa97]. Melnikov [GT94b]. membership [JZXM01]. Memory [BDPZ09, CR03, NF96, RW94, Ber97, KMY05, RWW94, SDKC10, SKCL12, Wan90, Wan92a]. Memory-based [CR03]. memory-intensive [SKCL12]. meromorphic [AK12]. meshes [LuS91]. meshing [BCGY08]. message [OIiM^{+}97]. Meta [Fau90]. metasystem [GT90]. Method [CK93, CK96, EC93, GV93, HK94, Nam95, Av97, Akr86, BL90, BWZ94, Bon94, BLS13, CGL09, Cho90, CJ89b, CGT97, DJLS07, DT86, DP07, ESY06, FFdlAGM00, GS89, HKZ10, JKL^{+}06, KM89, Le02, LÜ10, LSWZ06, LD90, MZ11, MW05, RTZ03, Ric95, Sti89, Ter09, Wat89, Zen03]. Methodology [ZSST96, Fra90]. Methods [BGH94, CG90b, Baj86, Bar10, BB07, BMP98, CZ00, CK02, DKLM98, El 13, EGT10, GSV91, GGS89, Gem07, Ida09, MR90, Pec91, Rum10, Rup10, Sau01, Sza11, VT98, Zan10, SS91]. Metric [LI00, YZ05]. MGTP [HK94]. microcalcifications [EK95]. microcomputer [Jur86]. middle [vdH03]. MIMD [RSSRV89]. Minimal [AZ97, AP01, CO01, Sch05a, Sho99, ZLS12]. minimised [LkGID97]. Minimum [KSS07, MZ11]. minimum-rank [MZ11]. Minion [Pur89]. misconceptions [Jur01]. Mixed [Gat91, VZ97, Wei99, van96]. Mmxlib [vdH06]. Model [GT94a, BL91, Bel90, Che90b, Li04, LCW05, MW05, SKCL12]. Modeling [DR01, DJLS07]. models [Baj86, CDS89, CR94, Eri04, Far06, Vin91]. modern [Bis12, Ger10]. modified [Bon94, MZ11]. Modula-2 [Kre89]. Modular [KLW90, MF94, Nor06, Enc94, Enc97a, FS13, vHM02, JM07, KM99, KN07, LN97, LS90, Mil11, Mül98, RWW94, RY08, Sme89, dKMW05, vH96]. module [ABRC97, BF12, FSK86, LM08, Tak90a]. modules [BP91, CIK97, GS05, LSWZ06, RMC07, ZW06]. Moduli [FN97]. modulo [LMP09, MP06, MS98]. Moment [JFSMR08, BBCM11, MZ12]. Monad [HCB94]. monic [dKMW05]. monodromy [MPV98]. Monoid [MR93, Rei00, RGSGG00]. monoids [WJLL93, WL93]. Monomial [BS98, ER94, BdC09, Fre00, Mil04, Nör98b, Rou10]. monomials [Kal93]. monotonic [BKLG97]. Monster [RC06]. Montréal [Lev95]. Morehouse [Ait95]. Moreno [AJL01]. Moreno-Socías [AJL01]. Moser [Bar95, BP07]. Moser-reduction [BP07]. motion [RLBS91, VT98]. moves [KC07]. Moving [SM08, BGJ13]. MP [GKW94]. MPI [Coo95b, Coo95a]. MPP [BSG96, KNKM94]. MPP-Oriented [KNKM94]. Multi [AGK96, BGK94a, RY08, Kor01, Küc91, YNT90b]. multi-environment [Kor01]. multi-language [Kor01]. Multi-modular [RY08]. Multi-Threaded [AGK96, BGK94a, Küc91]. multi-variate [YNT90b]. Multidimensional [KMZ98, BH06]. Multigraded [CdR96]. Multihomogeneous [BBCM11, DE02, EM09]. multilinear [RG91]. Multipath [GT94a]. Multiple [BCLV13, CCG05, CGT97, CEMP05, vHM02, MK99, Pet94, SS06a, ST12, Zen03]. multiple-point [Pet94]. multiples [BCSL12]. multiplication [Bis12, BZ07, Bod10, BDPZ09, FM11, GKZ07, GSGVPN^{+}09, HR10, vdH03, KMY05, MP09, Roc09, Sch05b, Sim98, SDKC10, Str12a, Uma06, Zan10, van97, vdHL12]. multiplications [HP97]. Multiplicities [MMM95, DPS09, GV00]. multiplicity [DZ05, WZ08]. multiplier [BL10]. multiply [Miy07]. multipoint [MF97]. Multipolynomial [MC92, MC91]. Multiprecision [CK00b]. Multiprocessor [BGK96]. Multiprocessors [RW94, CBdCD97, RWW94, Wan90, Wan92a]. multiterm [Sas86]. Multivariate [GL95, KS95, Lev13, MF94, RW94, Sch05b, Ste96, BW91, Cha86a, CGKW02, Din08, EMT10, FP09, FvzGP10, GKM^{+}04, GLsL06, GV91a, Gon91a, GE11, HWSZ00, HKZ10, JM09, KY89, KLW90, KK06, KY13, KS11b, Lev07, RWW94, Rua01, SY98, Sas01, ZR99, vdHL12]. Munich [Wat10].

Nash [Dat03]. Navier [GVBv94]. NC [GR97, MP90]. Near [SY12, Sto96, BAE^{+}00]. near-rings [BAE^{+}00]. Nearest [KL96b, HK98, HKL99, Hit05, Sek08]. Nearly [BCFS91, Mar95, Mor95, Pan01, SS94]. necessary [GZ90, GKZ90, SU92]. neighborhood [Che90a, CG90a]. Nested [Sha93, Sch04, Sch05a]. Network [Fat97, DJLS07]. networks [Far06, KR93, Wan91]. Neumann [Sat98]. neural [Ait95, Wan91]. Newton [AP05, Bar89b, DRJ97, KSG04, PR08, Sag12]. Nielsen [RMM98]. Nilpotency [Rák95]. Nilpotent [CDS91]. no [Akr86]. NODES [Cod92]. Noether [Vol93a, Vol93b]. Noetherian [GT97, Wei92a]. noise [KY13]. Non [Bac96, Bou05, CCT04, CdG07, ACJ03, BKLG97, BAE^{+}00, CKY89, CKS09, Cod92, DMR08, DRJ97, FS13, GS10, HP89, LMS13, Lev06, LuS91, Nor98a, OS04, dKMW05, KR86, Wei92a]. non-Abelian [KR86]. Non-associative [CdG07]. non-commutative [Lev06, Nor98a]. non-crossing [OS04]. non-holonomic [CKS09, GS10]. non-identities [HP89]. Non-Linear [Bac96, ACJ03, BAE^{+}00, CKY89, DRJ97]. Non-Liouvillian [CCT04]. non-modular [FS13]. non-monic [dKMW05]. non-monotonic [BKLG97]. non-Noetherian [Wei92a]. non-orthogonal [LuS91]. non-reduced [DMR08]. non-singular [LMS13]. nonalgebraic [Ric91]. Nonassociative [BHP07, Jac94]. nonclassical [OS89]. noncommutative [LS03, Mor89b]. noncomputability [Ser89]. nondivisibility [uGKO92]. Nonlinear [Edn93, KRL93, Ren91, AZZ91, CDS89, FNM86, Gat91, GZ90, GKZ90, GKZ91, HR03, KR93, LCW05, Pol98, Roa94, Rus92, Tsa99, dJ91]. nonlinearities [KR08]. nonrelativistic [BC92, VC89]. Norfolk [Sin93]. norm [Pet91]. Normal [Abe90, Gie95, SW04, SL95, SL96, Val93, BLV99, BGM91, BD08, Che89, CDS91, CD99, GLL04, HL12, Hul98, MW01, Mou05, RSV06, SR89, Sto96, Sto98, Vil93, Vil94, van95]. normalizability [Oka89]. Normalization [SS11]. normalizer [Miy06]. normalizers [LRW94, LM02, Miy00]. Normalizing [Edn93]. Norman [GS89]. norms [Enc97b]. Notation [ZSST96]. notations [Har00]. Note [OS89, RMM98, RSV06, Vol03]. November [BKLG97]. NP [AD11, AIRR10]. NP-complete [AD11]. NP-completeness [AIRR10]. null [Ebe04]. nullifying [McC03]. nullspace [SV05, ZLS12]. Nullstellensatz [Kap86, LLMM08]. Number [AH94, ABD86, BF12, BS11b, BM91, CR91, DP95, Enc94, Enc97a, Enc97b, vHM02, JM09, Joh92, Kal93, KK06, Lan91, LS90, MW00, Rol89, SY98, Sme89, Zim97, vzG09]. Numbers [GR94, Abb89a, Abb89b, Bür11a, CZ00, GR97, GGS89, JPR05, Maw86, MP90, MG94, Mor92, RL02, Rio02, vdH06]. Numeric [KM95, SWY11, WB93, Zim93, CCG05, CGKW02, Far06, GVv92, GVBv94, Gat91, GF92, GLsL06, KZ06, NM90, RTZ03, RVWW05, Roa94, RD05, Rus92, Sza11, WR07]. Numeric-symbolic [KM95, SWY11]. Numerical [AC04, Fat81, GV93, Ged86, Ley08, Nam95, SS95, TZ02, Ver04, Cha95, Cha89, CGY07, Dew89, Dew92, GW97, GR01, Gat86, GT09, JKM13, Kir92, Kir98, Kor01, MV90, Mez10, PU98, WR06, Yap09]. numers [GV00].

NumGfun [Mez10].

Object [CM96b, Gla97, ACS86, Sti97, VF92, BGY96]. Object-based [BGY96]. object-oriented [ACS86, VF92]. object/image [Sti97]. objects [DK98]. OBPDC [BGY96]. observability [Sed01b]. observations [RL02]. obstacles [GS98]. obtain [Jef93]. Obtaining [CT91]. October [HI93]. ODE [DP07, Edn93, Rus92, Sch95b]. ODEs [BLvH04, CCT04, FNM86, JKM13, LSYL11, Sch95a]. One [GMN^{+}91, AIRR10, DS03, GRS01, KM95, LSWZ06, Mai98, Por89, Rei00, WCD08]. one-dimensional [KM95, LSWZ06]. one-sided [Rei00]. Ontario [Bro07, Mou01, Cha86b]. open [Bro13, Ste08, YZ05]. OpenMath [DGW97]. Operating [Mar91]. Operation [GK94]. Operations [DJ94, Wan94, BHP07, CT98, GGS89]. Operator [Tsa96, ABP95, Arn89, Ast91, Gol86, Hon90, SS03]. Operators [van96, AvH99, BCR08, BCSL12, Bro94, BMW97, CJKS13, CGL91, Fre03, GS90, Gon86a, GSZ11, LST02, RRM09, RG91, RR08, Tsa97, ZL12]. Optical [BF94]. Optimal [Bay04, KM96b, ARS97, BZ07, CR91, KM96a, LkGID97, Mon04, Pan01, SY12, Sto96]. optimality [CK02, KLYZ08]. Optimization [Bac96, CCC98, TD90, Zim93, Zim95, CMS89, Fuk08, GEZ10, HMS01, KYAH07, KYAH08, LKK91, Par13, tW05, vHGv89]. Optimizing [MT94]. optimum [Din08]. options [BP06]. Or/Tiwari [KsLL00]. orbit [AB00, Bay04, Bou05]. orbital [RLBS91]. Order [CK12, ABRC97, AB98, AK12, BCB09, BBP10, BGU03, Bro92, BIL91, CCT04, CJKS13, Cor01, FvH11, FP09, GS08, KvH13, McC03, PS06, Pfl97, Sch95a, Sch95b, Sch96, SU92, Sto02, Zha93, ZL09, vHW05, vH07, vHL10]. Order-degree [CK12, CJKS13]. order-invariance [McC03]. ordered [Rio92]. ordering [BF03, FM11, HL03]. orderings [Fre00, Lev07, Zob05]. Orders [Wei93a, DSS04, Wei93b]. Ordinary [Hub96, Min95, Tsa96, ACFG05, Bar10, BR12, BCR08, BCSL12, Bro92, Bro94, BF99, BL02, CDS91, Cod92, CG92, DRJ97, FsG04, Gri90a, Hub04, Mac89, Oku90, Sch89, Sch96, Tsa97, Tsa99, ZL12]. ore [CJKS13, Av97, GZ03b, LL04, LN97, Li98, LSWZ06]. Oregon [Gon89]. organising [BL98]. organize [Pur89]. Oriented [CM96b, KNKM94, ACS86, VZ97, VF92]. origami [Ida09]. orthogonal [LL04, LuS91]. Orthogonalization [EKM96]. oscillating [Sha89a]. oscillations [AZZ91]. oscillators [CT91]. other [AZ96, EGG^{+}07, LM08, RC06, Zol97]. outlier [CKP12, KY13]. Output [KPR^{+}10, AL08, Kal02]. Output-sensitive [KPR^{+}10, Kal02]. overdetermined [GS99]. overview [HVS97, Sza11]. Oxford [ACM94]. Oz [HVS97].

p} [Clu03]. PA [Sen03]. PAC [GR94, RSV91b, RSV91a]. Package [Sch94, ABD86, BGV94, DGP04, EF89, FK91a, GKWW91, GV91b, Gon91b, LD90, MS92, Mez10, Zol91, vHGv89]. packages [Pol98]. Padé [BCL97, BL09, CL89, Cha86a]. Painleve [Ren91]. pair [BD86]. pair/completion [Nor90]. papers [BGY96]. Par’97 [LGG97]. Para [Sch94]. Para-Functional [Sch94]. paradigm [LSS13]. Parallel [ACM93, BP90, BP93, Bis12, BGK96, BP94, CM96a, CJSW94, Coo95a, CT96, FM96, GR94, GK94, GT94a, GL94, HCGW93, Ham94, Hon94, INPS93, KP94a, KP94b, KRS94, MP09, MF94, PG97, RW94, SRHW94, Sch94, SR89, Sie94, Sin93, Tak94, TL94, Tra98, Vil92, Wan90, Wan92a, Wan94, Web94b, WB93, ZB94, Ait95, AT94, Ber97, BGY96, BGK94b, Ces97, Cha90, Che90b, Coo95b, CG01, CBdCD97, DD89a, DGS98, Ebe97, Fit89, GR97, Gie97, Gla97, HP97, KKS86, KSI97, Lam97, LGG97, Mou93, MF97, Nöc00, NF97, RC06, RMC07, SLA89, Sed01a, SW89, SKCL12, TCT97, VZ97, Vil94, Wan97, FR95, HI93]. parallel/distributed [Wan97]. Parallelising [Bra94]. Parallelism [DMN94, LHD94, Bel90, GP97, MH97]. Parallelization [BGH94, Fau94, HK94, RWW94, Ber98, Bra90, LD11]. Parallelizing [GD93, Sie93, SB94]. parameter [AuK93, GK05, JP99]. parameterization [AC04]. parameterizations [van94]. parameterized [BOS13, BD04]. parameters [Abr10, Bou99, Che89, YL91]. Parametric [KYAH07, SZ10, ARS99, GC91, GC92, GVT95, LJM08, Mor06, PW06, SS09, Sit91, Yok04]. parametrisation [HS06, vdW06]. parametrization [LSWH00, RSV06, Sch98, Sch00, WCD08]. parametrizations [BD07b, Man90, Sch02a]. Parametrized [EGL04]. PARCS [KNKM94]. PARDI [BLM01]. Parsing [ZSST96]. part [Kau08]. Partial [BS93, FFGMGV02, KS10, PDS05, Kir92, Kir98, LST02, Oak94, RR97, Sch92, SS03, STK97, Tsa05, Tsa10]. partition [Col97, HC05]. parts [AHY05]. PASCO [Hon94]. Paso [BKLG97]. Passau [LGG97]. past [Bar03]. path [ASN89, GS98, MR89]. paths [ASN89]. PC [IKRT91, RLZ86]. PDE [ACJ03, GMS90, GR10, GS08, Hie91, HR03, WR06, WR07, LD90]. PDEs [GV93, LRB95, RLBW92, Wol95]. pencil [Sch00]. Penny [MH97]. perfect [GR08]. Performance [KRS94, JKL^{+}06, Lam97, Päs09]. Performing [KMG96]. periodic [FNM86]. permanent [Koi12]. Permutation [CFT95, Rák95, BCFS91, Bea93, BCF89, CS97, CFS90, CF92, CFYT94, CR03, EPB97, EP98, LRW94, LM02, Miy00, Miy06, Miy07, RMC07]. permutations [SDKC10]. perspective [FMRD08]. perspective-three-point [FMRD08]. perturbation [Jea00]. perturbative [SS91]. perturbed [Jea02, Mac99]. Pezzo [HS06]. Pfaffian [PV01, RB06]. Pfaffians [IK12]. phase [ASZ00]. Philadelphia [Sen03]. phylogenetic [Eri04]. physical [Dye94, SS91]. physicists [GL91]. physics [SD86, SS91]. PI [Kaj92]. Picard [BLW05]. piecewise [Car07, Far06, JLVR97, JR98]. piecewise-continuous [JR98]. piercing [Nus97]. PIMs [LS92]. Pincherle [DT86]. pipe [LSWH00]. place [HR10]. plain [BCLV13]. planar [Edn97, KSG04, OS04, TC89]. Plane [HS08, Bus01, EKW07, FS12, GV91b, Gon91b, GS98, GS08, GS10, Tei89, Vol97]. please [GMN^{+}91]. plotting [ABK95, Fat92]. Plural [LS03]. Poincaré [Rou10]. point [Che90a, CG90a, CK00b, CFL89, DS03, EPW89, El 13, FMRD08, GZ03a, HM01, KLYZ08, LÜ10, MZ11, MBC92, OS04, Pet94, Por89, Rum10, SY98, WCD08, Wat89]. points [BM91, CO01, FES12, Gri90b, Lel12, WR13]. Poisson [OS91]. Polar [DS03]. poly [BBDP04]. poly-algorithmic [BBDP04]. Pólya [PR06]. polycyclic [AW89b]. polyene [Vin91]. polygons [KSG04]. polyhedral [Fuk08]. polylogarithms [Bad94, MPV98]. polymorphic [Dal92]. Polynomial [BSG96, Bin96, CIK97, CK93, CK96, CGTW95, DG95, For01, JK97, Kal03, KL96b, LM02, MMM95, MF94, RCK04, SW10, Ste96, SL95, Vil96, Wan94, YT93, YLN95, AK91a, ABP95, Abr95b, AvH99, AK12, AMW08, ALM09, BFS93, BBR93, BLV99, BG11, BBCM11, BFP12, BF12, BP90, BB07, Bis12, BZ07, BMP98, BLS^{+}04, BCS05, Bou99, BHP07, BJ10, CKY89, CR91, CM11, CL06, CGL09, Chè04, CK04, CK06, CK92, CKP12, CGR10, CGT97, CGJ99, CSU00, CO94, CG92, Cuc00, DZ05, DCS05, Din08, DPW05, DPS09, ES12, ER94, Enc97a, FR09, FMRD08, FS12, FLMS06, FGT01, FN97, FSK86, GY06, Gat90a, Gat90b, Gem07, GZ90, GKZ90, GGSZ03, GR08, GJV03, GLV89, GS99, GE11, GS11, HvHN11, HR10, HS08, HK98, HKL99, HKZ10]. polynomial [IKS09, JZXM01, Kal90, KY89, KLW90, KS97a, KM99, KYAH07, Kau04, KSS07, Kir98, Küc91, LS03, LJM08, LF95, LI00, jL90, Luc86, MZ12, Mai98, MW94, Man94b, MM11, MS02, MR10, Maz06, MS97, MYT09, MP06, MP09, Mor06, Mou05, MF97, Nor98a, NF97, NY02, Oak12, Pan00, PZ10, PS07, RRM09, Ren06, RY08, Roc09, Sag12, Sas01, Sat98, SS09, Sch91d, Sed01b, SS06a, Sek08, Sme89, Ste97, SV05, Str12b, SWMZ12, Sza11, Vil93, Vil94, vG96, Wan90, Wei92a, tW05, WR13, YZZ13, Zen09, ZL12, ZL13, ZR99, vHM04, vdHL12, vzG06]. polynomial-based [BJ10]. polynomial-exponential [AMW08]. polynomial-nonlinear [GKZ90]. Polynomial-time [LM02, GGSZ03]. polynomial-type [tW05]. Polynomials [DK95, GL95, Gri96, KL96b, RW94, RS96, Ric93, von90, AuK93, AIRR10, BCGW89, BCL02, BL10, BCR13, BD89, BD08, Bün91, Can89, CGK^{+}13, CGvH^{+}01, CGKW02, EG07, EGT10, Enc94, Enc97b, FP09, FvzGP10, FIK86, GW97, GKM^{+}04, GZ04, GG12, Gem97, Gem01, GKsL02, GZ03b, GLsL06, GLV89, GV91a, Gon91a, uGKO92, GV00, GEZ10, GKZ12, HWSZ00, JM07, JM09, Jea02, KL94, KM03, KK05, KYZ06, KK06, KSG04, LS94, LMMRRR13, Lev07, Lev13, LN97, Li98, MS03a, MS09, Min03, Min06, Moo11, MS98, MM01, Nag02, ORY09, Pan01, PW06, Par04, RWW94, RR08, SGL04, Sek07, Sho91, Sho99, ST98, Ter09, TW91, VZ97, Wei92b, YNT90a, YNT90b, Zen03, ZD04, kMG97, vH13, vdHLS13, vH96, vzG07, vzG09, vzGGZ10]. Polytime [GS98]. polytope [OS04]. polytopes [DEF13, SGL04, San04]. Pommaret [GZ08]. poor [Pos11]. Popov [Vil96]. portable [Kaj92, WW89]. Portland [Gon89]. pose [RTZ03]. position [CGL09]. positions [Mou93]. Positive [KP94b, AHY05, BS09b, HKZ10, KP10, YZZ13]. positive-dimensional [YZZ13]. positivity [Pil13]. PostScript [YF04]. Power [Abr08, ACJ03, BSS08, BSC^{+}12, Koe03, BBD03, Bod10, vdH02, HR03, KFS86, NF97, PB90, RRW99, Sch05b, Wat89, Wis89, van97]. power-commutation-presentations [Wis89]. PowerMath [DR86]. powers [BMW97, GR08, Leh97, vzGN99]. PPOPP [ACM93]. Practical [EC93, HvHN11, Jeb97, RS12b, TW91, GRA00, GR92, NY02]. Practice [ACM93, BLS03, DS97, LMS07, Maz06, PS13, SW04]. Precessions [SS93]. Precision [Ahr96, CK93, BP06, GZ03a, KSS07]. PRECOMAS [Fra90]. Preconditioners [Tur05, Ebe10]. Preconditioning [SL95, PQZ11]. prefix [KMO90]. preprocessing [PZ10]. Presburger [Wei97]. present [Bar03]. presentation [Doo98, Doo02]. presentations [CO94, KMO92, Wis89]. presented [CdG07, vHM02, KM89, dGW99]. Presenting [DS94]. preserve [MH01]. pricing [BP06]. Primality [Dav92, CJ89a, GW97, KVY89, Mor92]. primary [Ley08]. Primbasissatz [LM94]. prime [Maw86]. Primes [CM96a]. Primitive [MZ96, Kal90]. Principal [ALM09, RGSGG00]. principle [BLS03]. Principles [ACM93]. Probabilistic [Bür11a, GV93, Mon05, Sed01b, BCLV13]. probability [CJ89a]. problem [AB00, AW89b, BL90, BP98, Bar99, BW91, BG07, Bus01, Chè04, CCC98, FMRD08, GSZ11, GRA00, Han07, Jea00, JZXM01, WJLL93, KMO90, LCJ98, Laz06, MH95, MYT09, MM97, MS98, Pur86, Pur89, Ric92, RF94a, SD86, Shi90, Sti89, Sto97, TSD91, WL93, YZ05]. problems [Abr10, AMW08, AW00, BGP12, BCF89, CMS89, FR95, HKL99, Koh90, Nus97, Sau01, SF90, SVV10, TMH94, VC89, tW05]. procedure [BD86, BGK94b, GK05, Win89b]. procedures [CCF89, CC93, Pil13, Wei89]. Proceedings [ACM94, Glo98, Kau05, Lak96, Lev95, MCG13, Mor02, SE11, Sen03, Wat10, vdHvH12, Ait95, Bro93, Doo99, Gut04, Jen76, Küc97, Mou01, Tra00, Wan81, WN90, Wat91, ACM93, AH94, Ait95, Bro07, Cha86b, Gon89, IEE95, Jef08, May09, Mor89a, Sin93, Tra06, Wan92b, CL96, FR95, Gia89, HI93, LGG97]. process [AB01b, BL91, Cha95]. Processes [DS94, BP90, MS92]. Processing [Sin93, SB94, Tak94, WB93, Bas97, LGG97, MF97]. Processor [Ebe97]. Processor-efficient [Ebe97]. product [CCG05, Cha03, EPB97, GRA00, vdH03, Li07, PB99, Sch05a, Zol97]. Products [BCR08, CZ00, KMZ99, Sha99]. profiles [DPS13]. Program [Bra94, KF91, Leo86, Mil89, MM97, Pol98, Yam90]. Programming [ACM93, BCM94, CM96b, Ham94, KP94a, LC94, RS12a, Sch94, TL94, Wik94, BKLG97, CBdCD97, FLMS06, Gil92, GSW12, GMB01, LK90, LkGID97, MS89, SKCL12, Ste13]. Programs [BGH94, GD93, LHD94, SRHW94, SB94, Wol95, Bel90, Fat03, FIK86, LKK91, Päs09, SS91, TCT97]. Progress [Cha90, KMG96, Wan94, Jac94, Vil07b, Vil11, WGK^{+}01]. project [BJK^{+}97]. projection [Arn89, Bro00, DSS04, Hon90, McC99, SS03]. projections [BY09, CSV13, EGG^{+}07, KS11a]. projective [Bus01, MT00, vzGGZ10]. Prolog [GT94a]. proof [Cha03, HC05, Sta90]. proofs [Bon94, Fat03, uGKO92, KW90, Kau04, MR89]. propagation [McC01]. Proper [Sch00, Man90]. Properties [KL96a, WG91, Zim01]. proposal [BSG97]. protocol [GKW94, Wan99]. protocols [BSG97]. prototyping [WGK^{+}01]. proved [BF91]. prover [HLY86]. Proving [Sti93, AGLM99, Bon94, Bon97, CGZ93, GK05, GMR94, Kap86, KS86, LL93, LLMM08, Mor92, OS89, Pil13]. Pseudofactors [HWSZ00]. pseudogroups [LRB95]. PSLQ [CSV13]. public [CCT08]. Puiseux [AV12, PR08]. purpose [Cha90]. puzzles [EP98].

Q [Dum08]. Q-adic [Dum08]. QCD [BL91]. QR [KSS07]. QR-factorization [KSS07]. Quadratic [KNS11, Fat89b, Hon93, Rol89, Sch91c, SP91b, SP91a, WCD08]. Quadratic-time [KNS11]. quadrature [Tao90]. quadtree [AW89a]. Qualitative [Rd91]. quandles [McC12]. quantified [SZ10]. Quantifier [CJ89b, Hon93, Wei94b, Bas97, BD07a, BS10, DGS98, DW00, Hon92, HD09, McC99, McC01, MB09, ST11b, Wei99]. quantitative [PR06]. Quantum [Gri96, Hie92, VC89, KS13]. quartic [Moo11]. quasi [FEV13]. quasi-homogeneous [FEV13]. Quasilinear [Tho96]. quaternary [SP91b, SP91a]. query [Bas97]. questions [Vil07b]. quotient [Mon04, MT12]. Quotients [YLN95].

radical [BOLP95, Dah09, FGT01, JFRS06, JFSMR08, KS07, Lap06]. radicals [FFGMGV02, HM01, Sme90]. radius [HKZ10]. Rainbow [Bet13]. RAM [SDKC10]. Ramanujan [Bur91, Fee90]. Ramis [DT86]. Random [EGT10, CFS90, Gon86b, Pol98]. Randomized [AIRR10, PQZ11, EK97, Iva00]. randomly [Far06]. Rank [KL96a, RB06, BJS07, CCF89, DPS13, Ebe04, FES10, LL04, MZ11, MZ12, MSW07, SY09, SV05, Sto09, Tur06]. Rankings [RR97]. Rational [Abr95b, AB98, ALP03, Bar89b, Bar99, Bin96, BLPR13a, BS93, CD99, FsG04, HS06, HL12, MP06, MS00, Sch98, Zie91, Zip91, van96, van98, AK91a, Abr95a, ASW05, ARS97, AIRR10, BCGW89, Bar95, BCCL10, BLS13, BS11b, Cha89, CFFL11, CKS12, CZ00, DCS05, EGG^{+}06, GLL04, Gem01, GGSZ03, Gie94, uGK91, GEZ13, GR92, HKL^{+}09, KLW90, KY07, KN11, KM06, Kir92, KRL86, Le02, LSYL11, Lic05, MS02, Mon04, Oll89, PDS05, Pfl08, RSV06, SWY11, Sch00, SW01, SC10, van94, vHY10]. rationality [vH93]. rationalizing [KLYZ08]. rationally [Vol93a, Vol93b]. rationals [vH13]. reachability [GE11]. readable [HC05]. Real [CK93, Joh92, PZ10, Rio92, Str08, Str09, AHY05, BG11, BRS09, Bro89b, CGY07, CJ89b, DS03, DET07, EKW07, El 13, EGT10, FMRD08, FGPT02, GV91b, Gon91b, GMR94, HKL99, HD09, HD00, HKZ10, JK97, KS11c, KSS07, Koi12, LS90, MZ12, MS09, MP90, PT13, Rio02, Roy12, Sag12, Sch98, SS06a, Sek07, Sha07, SY12, ST11a, ST12, ST11b, WX05, Wei94b, Wei99, WR13, YZZ13, vdH06]. real-integer [Wei99]. realization [AM90]. Reals [MR94, Din08, Tiw10]. reasoning [BKLG97, BC89a, Cha89, Cho90, KSY94, KSI97, KK89, Rd91]. recipe [Li07]. Recognition [Rák95, BF94]. recognizing [HK04, Ric95]. reconstruction [Bon94, BS11b, KM06, Mon04]. Recording [DS94]. recover [KS11a]. Recovery [Abb89a, Abb89b, KY13, Lic05]. Rectangular [SL95, ZG00, HP97, LuS91]. rectifying [LÜ10]. Rectilinear [Nus97]. recurrence [CvHL10]. Recurrences [Zim95, BWZ94, BCSC06, Fat89a, KMZ98, TCT97, Zim01]. Recurrent [Zim93]. recursion [ESY06]. Recursive [JR98, Abr06, BP90, LSWZ06, SKCL12]. reduce [Ebe10, AM90, BL91, BGV94, BHPS86, BC92, CC89, Fit89, Fit90, GMS90, GS90, Gat86, GV91b, Gon91b, KRL86, Kry86, KR86, LD90, Mac89, Ren91, RG91, TSD91, vHGv89, PB90]. Reduced [KM96b, BBP10, DMR08, GR01, Jea02, KM96a, SS11]. reducedness [Vil07a]. reducible [BP98, Bod04, vzG07]. Reducing [Kes94, FS04]. Reduction [CFL89, Koh90, RB91, Sch92, BL90, Bar89a, BP07, BCL02, BCC^{+}13, Bro89a, Fau02, GLL04, JP99, Jea00, Kai94, PR08, QV95, RMM98, RB06, Sch91c, Vil92, Vil11]. reductions [HMS01]. redundant [KPR^{+}10]. Reexamining [MR90]. reference [FT03]. refined [KS11b]. refinement [GSW12, PT13]. Refutational [KW90]. regions [HS08, SW10]. Regular [BBP08, Man90, SS93, AF86, BCB09, BL00, BM04, CG90a, WJLL93, LMP09, MSW07, Sat98, WL93]. regularity [Giu89]. regulator [SZ91]. related [Abr10, BdC09, CCF89, EP98, HKL99, JLVR97, SW11, Vil07b]. relating [CK04]. Relations [Zim93, CC90, CvHL10, MS97, RLBW92, Sas01]. Relative [Col97, Gut93, CJ89a, DP95]. relatively [ARS97]. Relaxed [BL12, vdH03]. Reliable [Ebe04, CL89, HE05]. reliance [Ebe10]. remainder [BW91, Kal90]. Remarks [MZ01, Nöc00]. reordered [CGT97]. reparametrization [ARS97]. replacement [PB99]. replicable [NM97]. Report [KMG96, Wan94, Cha90, Jac94, WGK^{+}01, WBD^{+}94]. Representation [BOLP95, AW89a, DK98, FM06, Maw86, MT12, Por89, Sha99, VZ97]. Representations [CFT95, Min95, Roa96, ALP03, BS98, BLPR13a, CFYT94, Fat91, dG97, GKZ12, Roa97, Ser12, Zol91, Zol97]. representatives [BL00]. represented [CDM^{+}11, Str10, Str12b]. Representing [BC00]. research [Roy12]. Residual [Bus01]. residue [KPR^{+}10, SS09]. residues [CKS12]. resilient [Pan00]. Resolution [Sta90, BS00, GPP91, MS92, Mor06, OS89, Pet91, Tei89]. resolutions [Lec00]. Resolvent [GY06, Leh97]. Resolvents [Val89, Col97]. resonance [AZZ91]. resonant [KR08]. respect [Lev07, YL91]. responses [CT91]. restricted [LM02]. restriction [ARS99]. result [Ser89, vH93]. Resultant [EC93, KS95, Bus01, CK02, CK04, CK06, DE01, DE02, DEF13, EP97, EM09, Kal93, KS97b, LGY11, LYG13, MC91, WEM98]. resultant-based [WEM98]. Resultants [Min06, CK00a, KSY94, KSG04, Leh97, MC92, Min03, Mon05, ZG00]. results [BCR13, Bra89, El 13, Gon86b, LL93, Roy12, Rum10, Ser89, vzGGZ10]. revisited [Dav92, Dum08, PHW13, Zen09]. rewrite [Har94, Oka89]. Rewriting [BGK94a, BGK96, AB01b, AW89b, BGJ13, BCF89, ER13, Gla97, KMO90, Sas86, Wis89]. Rham [Sch12]. Riccati [KYAH09]. Riccatti [Zha93]. Richardson [SVV10]. Riemann [BBDP05, uGKO92]. right [LN97]. rigorous [BJ10, Rum10]. Rings [MR93, AH08, BAE^{+}00, BP91, CdG07, EGNS12, GZ08, GT97, GHS00, JZXM01, Juo89, LL04, Nor98a, Por89, Rei00, Sat98, SS09, Wei92a]. Riquier [WR07]. Risa [NT92, NM97]. Risa/Asir [NT92, NM97]. Risch [Bro89a, Bro91, GS89]. Ritt [Szá97]. robot [Mou93, Sti89]. Robotic [CM95]. Rogers [Bur91]. Roider [Sti89]. Rome [Gia89, Mor89a]. Root [CGL09, CK96, BCR13, CJ89b, CK92, For01, GG12, Gem07, HKL99, JK97, KS11c, LJM08, MSW13, MvV89, PZ10, PT13, SY11, Sch91a, Sha07, SY12, Str08, Str09, ST11a, ST12]. root-finding [Gem07, PZ10]. rootfinder [Rua01]. rootfinding [Pan01]. Roots [Ahr96, CK93, YLN95, AHY05, AIRR10, BF91, BG11, CR91, CGT97, GK12, GT09, HK98, JFRS06, KYAH07, KYAH09, Koi12, LS90, MS97, MS09, Sag12, Sas03, Zen03]. Rostock [Glo98]. Rota [GS06, GSZ11]. routines [SW89]. row [BBP10, BCL02, DPS13, SS11]. row- [BBP10]. Rubik [KC07]. rule [Har94]. rules [Sas86, Sta90]. run [Fit89]. Runtime [KRS94].

s [Tou05, Vol93a]. s-systems [Tou05]. S-transformation [Vol93a]. SAC-2 [Kre89]. SAGBI [FR09]. SAGBI-Gröbner [FR09]. sampler [Cox07]. sampling [PB99]. San [ACM93, SE11]. Santander [Gut04]. SAT [SZ10]. Satellite [Gut93, SS93, Bou05, RLBS91]. satisfaction [RF94b, SF94]. Satisfiability [ZB94]. Sato [ALM09, BL10]. Scalable [CG01, CBdCD97]. scalar [Mac99]. scalars [KLYZ08]. scale [Mil04]. scaled [EM09]. scales [CR94]. scaling [LÜ10]. scalings [HL12]. scheduling [BDPZ09]. Scheme [MR94, Lel12]. Schemes [GV93, IKS09, CGR10, GSV91, GVv92, GVBv94, LuS91, MV90, Miy00, ORY09]. Schmidt [EKM96]. Schönhage [GKZ07]. Schoof [GM06, MMS07]. Schreier [FS09]. Schrödinger [GKZ91, LCW05]. Schubert [SVV10]. Schur [CGT97, HC05]. sciences [Dye94]. Scientific [Sin93, Ait95, DW90, SF90]. Scotland [Tra00]. Scratchpad [Bur91, DGT91, DD89b, JT94]. Search [ZB94, AK91a, Bon97, OS91, Wav99]. Searching [EF95, Hie91, ASZ00]. Secant [CK96]. Second [KvH13, ABRC97, CCT04, FvH11, FR95, vHW05, vH07, vHL10]. second-order [ABRC97]. Seidenberg [HKZ10]. select [Dew92]. selected [BGY96, Man94b]. selection [Cza91, GMN^{+}91]. Selfintersections [GP05]. semantic [Coo86]. Semantics [DJ94]. Semi [GR01, SB94, BPR98, CDM^{+}10, CDM^{+}11, GS98, LÜ10, RV96, Wei94a]. semi-algebraic [BPR98, CDM^{+}10, CDM^{+}11, GS98, RV96]. Semi-Automatic [SB94]. semi-groups [Wei94a]. Semi-numerical [GR01]. semi-rectifying [LÜ10]. semialgebraic [Gri90b, Sch09, Str10, Str12b]. semidefinite [Par13]. semidefiniteness [HKZ10]. semisimple [Sch91a]. semismooth [BS13]. semisymmetric [BGL05]. sense [Fat03]. sensitive [Kal02, KPR^{+}10]. Seoul [May09]. separated [EG07]. Separating [BLPR13b]. separation [EMT10]. September [CL96, FR95, Hon94]. sequence [CJK97, GLRR89a, GLRR89b, KP10]. sequences [Abr06, GK12, Kal90, KS06, Tho01]. Sequential [CT96]. Series [Fat89b, Gru93, Abr08, AV12, AMR89, ACJ03, BBD03, BSS08, BSC^{+}12, CCG05, CJK97, EF89, GG89, vdH02, HR03, KFS86, Koe03, Mil89, NF97, PB90, PR08, Rou10, RRW99, Sch05b, Wat89, YL91, van97]. server [LH99]. servers [Fat97]. set [BPR98, CCF89, CC93, Cho90, DS03, Gri90b, GL11, KW90, Lin04, OS04, RV96]. set-theoretic [CC93, GL11]. sets [Bod04, CDM^{+}11, DS04, EPW89, KM95, LMS07, PB99, PS06, PV01, Roy12, Sch02b, Sch09, Str10, Str12b, Bet13]. several [KM03, KYZ06, Lev07]. shape [BMMT94, BMMT94]. Shared [RW94, RWW94, Wan90, Wan92a]. shared-memory [Wan90, Wan92a]. Sharp [DS04]. Shift [Gri96, JKR05, VZ97]. Shift-Equivalence [Gri96]. Shifted [BLV99]. Shiftless [GGSZ03]. Shifts [GL95, GKsL02, LS94, vzGG97]. short [uGKO92]. shorten [MR89]. Shortest [Rol89, GS98]. SIAM [Sin93]. sided [Rei00]. sign [CJ89b]. Signature [ASW05, EP11, ER13, MG94, SW11, SWMZ12]. Signature-based [EP11, SWMZ12]. signum [JLVR97]. SIGPLAN [ACM93]. SIGSAM [Gon89]. SIMATH [Zim97]. similarity [KKS86]. Simon [Doo99]. Simple [Hon92, Kau08, Mül98, RC06, SY11, Sag12, SY12, SS06c, WCD08, Zol91, Zol97]. simplicial [LMMRRR13]. Simplification [Bro89b, Bro98, Le02, PU98, Sas86, Sch02a, Zim95, BBD03, BD02, BS10, Car04, Fat03, IK91, MP06, kMG97]. simplifications [Bro09]. simplified [Ric95]. simplifier [Fau90]. simplifies [Vol93a, Vol93b]. simplify [Ebe10]. simplifying [BBDP04]. Simultaneous [DPS13, BL09]. Simultaneously [BBP10, KSW11]. sine [BS11a]. single [Bro13, FS04, Sch04]. single-nested [Sch04]. Singular [CGTW95, KL96b, ST98, BSC^{+}12, BCGY08, Che90a, CG90a, Hit05, Kal94, KYZ06, LMS13, LM08, MM11, MS00, Tur05, vzG07]. singularities [BS00, SC10, Som94, Tei89, Vol93a, Vol93b]. singularly [Mac99]. singularly-perturbed [Mac99]. six [KC07]. Sixth [Sin93]. Size [Dah09, Coh13, SY12]. Skeleton [Bra94]. skew [BCL02, JZXM01, RRM09, Wei92a]. skewly [Min06]. Slave [BGK96]. Slice [Rou10]. Small [KP94b, McC12, Alb12, BCFS91, Bet13, Ebe00, Ebe03, Ebe10, HE05, KK06, Mar95, Mor95, SY09, SS94, Sho91, SV05]. small-base [Mar95, Mor95, SS94]. smallest [Lin04]. Smith [DSV00, EGNS12, Gie95, SW04, Sto96, Vil93, Vil94]. smooth [DS03, Du94]. Snowbird [Wan81]. Socías [AJL01]. Software [IA98, Kre89, LCJ98, Fuk08, LHK^{+}10, Pur86]. solid [LY86]. soliton [Hie91]. Solution [Hub96, KP94b, KYAH09, Ric96, Bro99, CDHJ07, Gat90a, Gat90b, Gat91, Gie97, Hon92, Kal94, MH95, Rab89, Sto97, Sza11, Vil97, WX05, WR13]. Solutions [BL02, MYT09, PS93a, van96, AK91a, AP94, ABP95, Abr95b, AZ96, Av97, AB98, AB01a, AK12, AMBSW11, ACJ03, ACFG05, Bar89b, Bar99, BCB09, BCEW12, BS09a, BCS05, BSC^{+}12, Bou99, BLvH04, CvH09, CCT04, CMS89, Che90a, CG90a, Coh13, Cor01, DT86, Fat89b, FsG04, FSK86, Ger98, GEZ13, HR03, JKM13, KvH13, LL04, LMS13, LCW05, LSYL11, MZ12, Mac99, Man94b, MS00, Oak94, PS93b, Pfl97, RRW99, SU92, UC90, YZZ13, Zha93, van98, vHY10, vHL10]. Solvability [HM01, Baj86, FFGMGV02, YNT90a]. solvable [SWMZ12]. solve [AW89b, Sti89]. solver [Cod92, DP07, Mac89, SWY11, TD90, WEM98]. solvers [GSW12]. Solving [BR12, BJ98, BFP12, BJS07, BF99, BCF89, CKY89, CvHL10, Coh13, CGR10, Cuc00, DvHK08, EGG^{+}06, EP98, FR09, FS12, GC92, GKZ90, GS99, Hie92, Kal90, KMH89, Laz06, LC94, Mai98, MMM95, MR94, MT00, Rei00, SVV10, Str12b, YC92, vHW05, vH07, Bar10, BMP98, BGP12, Bro86, DZ05, DET07, DS99, ES12, ER94, EGT10, FMRD08, Giu09, LCJ98, LMS13, LD90, Mou05, MS99, PZ10, Pur86, Pur89, Roa94, SD86, Ste97, SZ10, TSD91, VC89, Wei94a, vdW05]. Some [Bra89, CJSW94, LL93, MvV89, Nöc00, RL02, SF90, Vil07b, AZ96, Abr10, Ast91, BC89a, CC90, CHT^{+}07, Fat03, GS99, Miy07, Pan89, VC89, tW05]. source [BCCG92, Ste08]. Space [MR11, Roc09, ARS97, CHT^{+}07, DMR08, Ebe04, GR01, KM96c, MW01, PO08, SC10]. Space- [Roc09]. Space-efficient [MR11, CHT^{+}07]. Spaces [Sti93, Bay04, FN97]. Spain [Gut04]. Sparse [CKP12, EC93, GL95, KY13, LGY11, LYG13, MF94, RW94, AIRR10, BCR13, Bis12, DE01, DSV00, EGG^{+}06, EGNS12, EP97, FM11, GG12, Gie97, GLS98, GLsL06, uGK91, uGKO92, JM07, KY89, KLW90, KsLL00, KY07, LS94, Maw86, Min03, MP09, RWW94, Vil97, dKMW05]. sparsest [GKsL02]. Sparsity [HS02]. Special [BS11a, BL02, CKS09, GK05, WJLL93, Kno86, MS03b, Wei94a, WL93, vzGN99]. specified [JM10]. Specifying [LPR99]. spectral [KYAH07]. spectrum [Str12a]. speed [Nab07]. speed-up [Nab07]. splitting [ORY09, Ren06, RY08]. sporadic [Mül98]. Square [Ahr96, Gil92]. squares [GEZ10, KLYZ08, Par04, Sau01]. squaring [Bod10]. St [Tra00, BdC09]. Stability [GSV91, GV93, WX05, Cet92, Cha89, GVv92, GVBv94, HS08, MV90, SXZ11, TZ02]. Stabilization [Ste97, TZ02]. stabilizers [CFL89]. Stack [SRHW94]. Stack-based [SRHW94]. staggered [GM86]. standard [GPT89, RB91, Zob05]. standards [LH99]. Star [Sha99, Coo95b, Coo95a]. STAR/MPI [Coo95b, Coo95a]. state [CT91]. statement [vzGGZ10]. stationary [Pet94]. statistics [CR97, Sul12]. steady [CT91, MYT09]. steady-state [CT91]. Steering [Kal96]. steps [Kal02]. steps/giant [Kal02]. stiff [Rus92]. stochastic [Cet92, Tsa10]. Stokes [GVBv94]. story [Kal03]. straight [FIK86]. straight-line [FIK86]. straightening [Ruf07]. Strassen [Bod10, BDPZ09, GKZ07]. Strassen-like [Bod10]. strata [Bay04]. strategies [CCC98, GMN^{+}91, Hav91, OS89]. strategy [AT94, Cza91]. strategy-accurate [AT94]. stratified [CC93]. Strong [Oka89, BL90]. Strongly [HMS01, MSW07]. Strongly-local [HMS01]. structural [Oll89]. Structure [CEMP05, CM95, Ric93, CFFL11, DZ05, EP97, Iva00, LS92, LRB95, NS06, Oak94, Pan00, PS13, WZ08]. Structured [BB07, Gem07, vdHLS13, FR95, JM10, Nag11, Pan89]. structures [BCM94, BHP07, Har94, HMS01, SSW09]. student [Jur01]. Studies [SS93]. Study [AGK96, BGH94, ZB94, AGLM99, CA07, GS90, MSW07, RS12a, SD86]. Sturm [GLRR89a, GLRR89b]. Sub [BCR13, PV01]. Sub-linear [BCR13]. sub-Pfaffian [PV01]. subalgebraic [Nor98a]. subalgebras [Lev06]. subdivision [BCGY08, CGR10]. subdivisions [Bon97]. subexponential [Gri90b]. Subexpressions [DMN94]. subfields [vHKN11]. subgroups [But92, CO94, Hul98, HL03, Hul13, KMO92, Mar95, Mor95, Sim97]. sublanguages [CCF89]. Submersive [HKL^{+}09]. submodules [LSWZ06]. subresultant [GV91a, Gon91a, Lan91, Li98, Man94a, Sas03]. subresultants [BD04, Rei97]. subroutines [DGP02]. subsets [Cod90]. substitution [Abr93a, Abr93b, Yam90]. success [Kal03]. Succinct [FM06]. suffice [KC07]. Sufficiency [Cor94]. sugar [GMN^{+}91]. Sugarbush [BP94]. SUI [DW90]. suited [Bod10]. Sum [AHY05, KYAH07, KYAH09, Sas01, Sch04, Sch05a]. Summation [HL86, ArRL89, AP05, Abr06, GGSZ03, KS06, KS07, MW94, RL89, Sch04]. summations [Wan91]. Sums [Par04, Abr95a, AP01, CSV13, CZ00, CEMP05, GEZ10, KLYZ08, Le02]. sums-of-squares [KLYZ08]. super [BP07]. super-irreducible [BP07]. Superalgebras [MZ96]. superschemes [Miy07]. Supersparse [KN11, KK05, KK06]. Support [CK04, AGLM99, EM09]. Surface [Nam95, vdW06, BLWhW91, GP05, KS11a, Khe02, Sch02a]. Surface-to-Surface [Nam95]. surfaces [AV12, BGJ13, BD07b, DCS05, EGL04, FGPT02, GVT95, HS06, LSWH00, PDS05, PO08, RV93, Sch98, Sch00, WCD08]. survey [BJK^{+}97]. Switzerland [Lak96]. syllogistics [CC93]. Sylow [But92, Mar95, Mor95]. Sylvester [Kal93, KSG04, ZG00]. Sylvester-resultants [KSG04, ZG00]. Symbolic [ArRL89, ACM94, Ano95b, Bac96, Bar10, BK86, BCK99, BH06, Bro93, Bro07, BHKR95, Cha86b, CM96b, Coo95a, CM95, DG95, Doo99, Far06, Fat96, GV93, GVBv94, Gat90a, Gat90b, GLsL06, Glo98, Gon89, GMB01, Gut04, HCGW93, Her97, Hon94, Ida09, IK91, INPS93, Jef08, Jen76, Juo89, KYAH08, Kau05, KS07, KMG96, KRL93, Küc97, Lak96, LSWH00, Lev95, Li04, May09, MV90, MCG13, Mor02, Mou01, Nam95, Ng79, RL89, RVWW05, Roa94, Rup10, Sch04, SE11, Sen03, SW89, SB94, Sti97, Tak94, Tao90, Tho96, TC89, Tra06, Tra00, Tsa97, Wan81, Wan92b, WN90, Wat91, Wat10, WB93, WR07, dJ91, vdHvH12, AGLM99, BCLA82, Buc05, CL96, CCG05, Ces97, Cha90, Coo92, Coo95b, Dal92, DD89a, Dew89, DKS^{+}91, DK98, DR01, Dye94, Fat97, FGPF89, GVv92, Gat91]. symbolic [Ged86, GG89, GF92, Gon86a, HI93, Har94, Jur86, KZ06, KS06, Kol90, KM95, LCJ98, LMMRRR13, LY86, Leo86, Li07, LH08, LHK^{+}10, MW05, MS92, NM90, Nor00, Pai92, PG97, PU98, RGL94, RTZ03, RD05, Rus92, SWY11, SF90, SS06b, SD86, Sie93, Sza11, TD90, TMH94, Ver04, VF92, VC89, Vin91, Gia89, Ng79]. Symbolic-interval [GMB01]. Symbolic-numeric [GVBv94, GLsL06, RVWW05, Roa94, WR07, CCG05, GVv92, Gat91, GF92, KZ06, RTZ03, Sza11]. Symbolic-Numerical [GV93, Nam95, MV90, PU98]. symbolic/numeric [NM90, RD05, Rus92]. Symmetric [BGL05, SS93, ASW05, AH08, BMW97, FS12, GLV89, GVT95, HC05, JX07, KS13, Rab89, Val89, vdHLS13]. symmetrical [Gat91].

Symmetries [Sch95a, Sch95b, FR09, GKZ91, IK91, LÜ10, LRB95, Wol95]. symmetrizing [Le02]. symmetry [FN97, Gat90a, Gat90b, KF91, Len13, RLBW92, RW00]. Symposium [ACM93, ACM94, AH94, Ano95b, BKLG97, Bro93, Bro07, CL96, Cha86b, Doo99, Gia89, Glo98, Gon89, Gut04, Hon94, IEE95, Jef08, Jen76, Kau05, Küc97, Lak96, Lev95, MCG13, Mor02, Mou01, Ng79, SE11, Sen03, Tra06, Tra00, Wan81, Wan92b, WN90, Wat91, Wat10, vdHvH12, May09]. Symsac [Cha86b, Wan81, Jen76]. Synchronization [BR94, DMN94]. syntactic [BSG97]. synthesis [Päs09, ST11b, Tiw10, Win89b]. System [GR94, KRS94, KMG96, MMM95, TL94, ACS86, AM90, BS86, BR06, BGJ13, BL98, CR91, Cha90, CK04, CK06, DR86, DKS^{+}91, DK98, DW90, DR01, Edn97, ES12, ER94, Fat96, Fit90, FS04, FIK86, FSK86, GMS90, GS90, GT09, Gri89, Gri90a, GL91, HR03, Jac94, Jur01, KR08, Kor01, KM89, LHJ92, LK90, LS03, MH97, Mou05, MS99, NT92, Oak94, Oka89, Pur86, RLZ86, San93, SWY11, Sch91a, Sey86, TMH94, TD89, TSD91, Zim97]. Systematic [KK91, OS91]. Systems [BR94, Coo95a, CGTW95, CM95, DG95, Edn93, EG07, IEE95, KP94b, KRL93, Kol90, Ste96, Zim93, AB98, AB01a, AK12, AMBSW11, Bar89a, BP98, BP07, BBP08, BCB09, Bar10, BBP10, BR12, BCL97, BL12, BFP12, BMP98, BJS07, BLPR13a, BLPR13b, BHP07, Bro86, BLW05, BCF89, CL96, CKY89, Cet92, Che90a, CG90a, CDS91, CD99, CDM^{+}10, CM11, CGY07, CGL09, CKS09, Clu03, Coo95b, CGT97, CBdCD97, Cuc00, DE01, DZ05, DET07, DKLM98, EGG^{+}06, EM09, Fat90, Fat91, FR09, FMRD08, FS12, FEV13, GMS90, GC92, GY06, Gat90a, Gat90b, Gat91, GZ90, GKZ90, GR10, Gie97, GLS98, GS99, Gon86a, HKL^{+}09, HI93, HW86, Hub04, Kaj92, Kal90, Kal94, KSW10, KSW11, KR93, KPR^{+}10, KK91, KMH89, KMO90, Lau03, LMS13, LÜ10, LM08]. systems [LD11, jL90, MZ12, Mai98, MM11, Mar91, Maz06, MYT09, MK99, Mor06, MS00, Nab07, Pet91, Pfl08, Pfl97, RB91, RVWW05, Rei00, Ric96, RER09, RD05, RB06, Rus92, RRW99, SXZ11, Sit91, SS86, Sta90, Ste97, ST98, Str12b, Sza11, Tiw10, Tou05, Tsa05, Tsa10, Vil97, WX05, WR13, YZZ13, Yok04, dJ91]. syzygies [BD07b, CD11, MMT92].

tableau [MR90]. tableaux [MR89]. Tables [Thu96a, Col97, EF95, Fat89a, Thü96b]. tadpoles [KC09]. talk [Ida09]. tame [Str09]. Tangent [CK96]. Tangent-Secant [CK96]. Tarski [Bro09]. Taylor [JKR05, vzGG97]. Technique [CT96, Zim93]. techniques [ARS99, AW89b, BC89a, Coo92, DMS^{+}05, EF95, FLMS06, MC91, MS92, Sch02b, TD90, Wis89]. technology [HLY86]. Telescopers [CKS12, Sch05a]. telescoping [BCCL10, BLS13, BCC^{+}13, CK12]. Tellegen [BLS03]. Tensor [KMG96, Zol97, CCG05, Cha03, EPB97, IK91, kMG97]. tensor-product [CCG05]. Term [BGK94a, BGK96, Wei93a, BGJ13, Lev07, Oka89, Wei93b, Bün91]. Term-Orders [Wei93a, Wei93b]. Termination [Pil13, Ebe03, KsLL00, PHW13]. terms [Abr02, ALP03, Bad94, BL02, CR01, DvHK08, vH07]. ternary [HC05, KM02, SP91b, SP91a]. test [GW97, Gri90b, vdH02, KVY89, KMO90, Sed01b]. Testing [Gri96, Dav92, Gri90a, Hie91, Nag02]. tests [BD89]. tetrahedra [LI00, YZ05]. Tetrahedrizing [EPW89]. tetrahedron [Kau11]. Texas [IEE95]. TFT [vdHLS13]. Their [DK95, Abb89a, Abb89b, CGL91, HKL^{+}09, Hub10, Par04, RB91]. Theorem [Ric93, Sti93, AGLM99, Bon94, Bon97, CGZ93, GMR94, HLY86, Kap86, KS86, LL93, OS89, tW05, vHW05, PR06]. theorema [BJK^{+}97]. Theorems [BHC95, KW90, Kut89, RRW99, Sch91d, Wav99]. theoretic [CC93, GL11, Uma06, Zim97]. theoretical [KR86]. theories [BC92, KR86, Wei89]. theorizing [RS12a]. Theory [AH94, Tiw10, But86, CCF89, CC90, Cuc00, GSV91, GV91a, Gon91a, Kem08, Koh90, LM08, Li98, LMS07, Man94a, MS03a, Maz06, Oku90, RF94b, SF94, Sit91, Sud08, Tou05, VT98]. There [Akr86, BF91]. thesis [Giu09]. thin [GVBv94]. thin-layer [GVBv94]. third [GS08, LGG97, SU92, vH07]. third-order [GS08]. Threaded [AGK96, BGK94a, Küc91]. three [EPW89, EG07, FMRD08, Hie91, KvH13, CCF89]. three-soliton [Hie91]. tight [ESY06]. Time [CHT^{+}07, Kes94, BCFS91, BFS93, BBR93, BF12, BGH94, Cet92, CR94, CIK97, CO94, GGSZ03, Gri90b, HvHN11, KNS11, LM02, MS02, Roc09, Rup10, SS94, Sed01b]. Time-and [CHT^{+}07]. time-efficient [Roc09]. time-extended [CO94]. time-varying [Cet92]. Tiwari [KsLL00]. Toeplitz [BP93, BJS07, Kal94, KP94b, KL96a]. Toeplitz- [BJS07]. Toeplitz-like [BP93, Kal94, KP94b]. Tokyo [BGY96, WN90]. toolbox [MM97]. toolkit [Wan91]. Tools [Wan97, BJ10]. Toom [BZ07, Zan10]. topology [DMR08, FGPT02]. torational [LSYL11]. Toric [Eri04, Khe02]. Total [HL03]. towers [DDS13, FS09]. trace [IKRT91, JFSMR08, Tra89]. Tracing [Nam95, NF96]. Trager [Enc97a]. transcendence [GRS01]. transcendental [Tak90b, Tsa97]. transform [Dum08, HR10, vdH04, Arn13, KS13]. Transformation [Edn93, Tsa10, GT90, Sta90, Vol93b, Vol93a]. Transformations [Zim95, BAE^{+}00, GLV89, JPR05, Tsa05]. transforms [VT98, Pos11]. Transition [BR94, GT90]. transitive [Miy07]. translations [Kut89]. translator [BCCG92]. transputers [RSV91b, RSV91a]. Tree [Sie94, ESY06, SY12]. Trees [DG95, CG92, GL89, Wu93]. trends [El 13, Fat90]. triangle [Rab89, SW10]. Triangular [CDM^{+}10, Maz06, CMXY09, CM11, CDM^{+}11, CGY07, DS04, DMS^{+}05, LMS07, PS06, Sch02b]. triangulation [Hub04]. triangulation-decomposition [Hub04]. Triangulations [San04, Bur11b]. trigonometric [AW00, MM01]. trinomials [vzG01]. trivial [Hul13]. trivial-fitting [Hul13]. truncated [Arn13, HR10, vdH04]. truth [Bro98]. truth-invariant [Bro98]. Tutorial [Ham94, Wan94, Ste13, Sza11, vH13]. Twenty [KC07]. Twenty-six [KC07]. Twisting [GK12]. two [BF03, BCdP13, Chè04, Ebe97, EG07, Gri90b, GR92, HK04, KK08, KS11a, LM08, Mai98, MB09, OW05, RB06]. two-dimensional [Mai98]. TX [BKLG97]. type [AM90, BL98, LÜ10, San93, Sat98, Web94a, tW05, vHY10]. typed [Oka89]. typeset [BF94]. typesetting [Soi95].

UFD [HK04]. Ukraine [Bro93]. unbalanced [Zan10]. uncoupling [BCdP13]. undamped [CT91]. undecidable [Abr10]. undergraduate [CGG^{+}86]. Understanding [Car04, SD86]. Uniform [Bas97, PW06, NS06, Wu93]. uniformity [Wei97]. uniformization [SS95]. unimodularity [PS12]. uniqueness [RRW99]. Unirational [GRS01]. unit [SZ91]. unitary [ASW05]. United [ACM94]. units [RSV04]. Unity [YLN95, GK12]. Univariate [PB90, Pan01, ST11a, ST12, BLPR13a, GG12, LS94, Ter09, TW91, Wan90, Wan92a, Zen09, vH13, vzG09]. universal [LS12]. Université [Mor02]. University [Bro07, Doo99, Glo98, Gut04, Mou01, Sen03, Tra00]. Unleashing [Dev89]. unmixed [FES12]. unnested [Sme90]. unsolvable [Cod90]. unspecified [KS06, Raa13]. update [Wol13]. Upper [Koi12, BG07, Kal93]. US/Japan [HI93]. USA [Ait95, HI93, Küc97, MCG13, SE11, Sen03]. Usage [KR86, BL91]. use [Hie92, VC89, Wei94a]. used [BC89a, CDS89]. user [BC89b, DW90, LHJ92, Leo86, SS86]. Using [AW89b, Bra94, Dat03, Dew92, DG95, Gri96, KS11a, KMG96, AC04, BBDP05, BBCM11, BLS13, BGL05, BCF89, BM05, BP94, CvHL10, Cho90, CK00a, Coo92, CFT95, CT96, CG01, CSU00, CG92, DCS05, EGG^{+}07, Fat03, FR09, FES10, Fee90, For01, GG89, Gil92, GVT95, GEZ10, HS06, HL86, vdH03, IK91, Joh92, JK97, Kal90, KYAH07, KYAH09, Kap86, KSY94, KN07, KS86, LCJ98, LÜ10, Maw86, MMS07, Miy00, MMT92, MP09, MSV09, MBC92, RB91, Rum10, Sha89a, Sha07, Sie93, SKCL12, ST11b, SS06c, Tou05, dGW99]. Uspensky [Akr86]. usually [BC89a]. Utah [Wan81].

VA [Sin93]. valence [DSV00]. valid [Jef93]. valuations [AK12]. Value [CGTW95, BGP12, FT03]. values [BS11a, CKP12, GEZ10, KY13]. Vancouver [Doo99]. Vandermonde [BJS07]. Vandermonde-like [BJS07]. variable [AIRR10]. variables [BF03, Chè04, EG07, HK04, KM03, Kau04, RB06]. Variant [HD09, Kai94, Kal02]. variate [YNT90b]. variation [CJ89b]. variations [Cor94]. varieties [AC04, DS03, MB09]. variety [ALM09, EM99, FMRD08, GL11, Lec00]. Various [KS95]. varying [Cet92]. Vector [BS11b, DG95, Sti93, CGZ93, GT94b]. Vectorisation [HCB94]. vectors [GV00]. Vegas [KVY89]. Verification [HP89, Rum10, ST11b, GR10, Tiw10]. Verified [YZZ13]. version [Bar95, BC89b]. vertex [Bur11b]. very [CR03, Zan10]. Vessiot [BLW05]. via [BP07, BRS09, BD07b, CCG05, Cha03, CDS91, CMXY09, CKS12, DSV00, Enc97b, GKM^{+}04, GKsL02, GHS00, HKZ10, KLYZ08, KW90, LYZ10, LI00, MZ11, MZ12, SGL04]. viable [Ste08]. vicinity [AZZ91]. view [Buc05, CSV13, DGT91, HM01]. viewpoint [Cuc00, Sek08]. Virtual [BG11]. vision [MM97]. vortices [FS12].

Walk [Tra98]. was [vzG06]. Waterloo [Bro07, Cha86b]. weak [Sta90]. weakly [FNM86]. WEB [Kre89, WKZZ03]. WEB/Modula [Kre89]. WEB/Modula-2 [Kre89]. Weierstrass [Rua01, van95]. weight [KR86]. Weinstein [Koe03]. Western [Mou01]. whether [FFdlAGM00, Gri90b]. which [AuK93, GKZ90, GV00, MH01]. white-box [BS10]. Who [vzG06]. Wiedemann [Tho01, Tur06, Vil97]. Winograd [BDPZ09]. without [CGK^{+}13, Fau02, Kal92, Leh97, MR11, vdW06]. witness [GRA00]. Witt [Juo89]. WME [WKZZ03]. word [AW89b, Hol00, KMO90]. word-hyperbolic [Hol00]. words [Abd89, GS06]. Working [GV91b, Gon91b]. Workshop [BGY96, Her97, FR95, HI93]. worst [FH97]. worst-case [FH97]. Wrap [Kes94]. Wu [MW05, Szá97].

XIV [CCF89]. Xne [FSK86].

Y.M. [BC92]. Yang [Hie92, OS91]. year [Wol13]. years [BW98]. York [Jen76]. Yorktown [Jen76].

Zeilberger [Abr02, IK12]. Zero [Ric93, Ste96, BY09, CGY07, CGT97, Dah09, Fau02, FM11, GT09, vdH02, LST02, Ric95, Sas01, Sek08, dJ91]. zero-dimensional [BY09, CGY07, CGT97, Dah09, FM11, GT09, LST02]. zero-sum [Sas01]. zero-test [vdH02]. zeros [CGY07, jL90, MM11, MK99, PR06, SS06a]. Zippel [KsLL00]. Zolotarev [Laz06]. Zurich [Lak96].

References

Abhyankar:1989:CAC

[AB89] S. S. Abhyankar and C. L. Bajaj. Computations with algebraic curves. In Gianni [Gia89], pages 274–284. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Abramov:1998:RSF

[AB98] S. A. Abramov and M. A. Barkatou. Rational solutions of first order linear difference systems. In Gloor [Glo98], pages 124– 131. ISBN 1-58113-002-3. LCCN QA155.7.E4 E88 1998. URL

             http://www.acm.org/pubs/citations/proceedings/issac/
             281508/p124-abramov/.

Abramov:2000:HDO

[AB00] Sergei A. Abramov and Manuel Bronstein. Hypergeometric dispersion and the orbit problem. In Traverso [Tra00], pages 8–13. ISBN 1-58113-218-2. LCCN QA76.95.I59 2000. URL

http://www.acm.org/pubs/articles/proceedings/issac/
345542/p8-abramov/p8-abramov.pdf;   http://www.acm.
org/pubs/citations/proceedings/issac/345542/p8-abramov/

. ACM order number 505000.

Abramov:2001:SLF

[AB01a] Sergei A. Abramov and Manuel Bronstein. On solutions of linear functional systems. In Mourrain [Mou01], pages 1–6. ISBN 1-58113-417-7. LCCN QA76.95.I59 2001. ACM order number 505010.

Armando:2001:MEP

[AB01b] Alessandro Armando and Clemens Ballarin. Maple’s evaluation process as constraint contextual rewriting. In Mourrain [Mou01], pages 32–37. ISBN 1-58113-417-7. LCCN QA76.95.I59 2001. ACM order number 505010.

Abbot:1989:RAN

[Abb89a] J. Abbot. Recovery of algebraic numbers from their p-adic approximations. In Gonnet [Gon89], pages 112–120. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989. US$29.00. ACM order number: 505890. English and French.

Abbott:1989:RAN

[Abb89b] John Abbott. Recovery of algebraic numbers from their p-adic approximations. In Gonnet [Gon89], pages 112– 120. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989.

US$29.00. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/74540/p112-abbott/. ACM order num-

ber: 505890. English and French.

Abbott:1986:BAN

[ABD86] J. A. Abbott, R. J. Bradford, and J. H. Davenport. The Bath algebraic number package. In Char [Cha86b], pages 250– 253. ISBN 0-89791-199-7 (paperback). LCCN QA155.7.E4

A281 1986. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/32439/p250-abbott/. ACM order no.

505860.

Abdulrab:1989:EW

[Abd89] H. Abdulrab. Equations in words. In Gianni [Gia89], pages 508–520. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Aberer:1990:NFF

[Abe90] K. Aberer. Normal forms in function fields. In Watanabe and Nagata [WN90], pages 1–7. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95

.I57 1990. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/96877/p1-aberer/.

Avitzur:1995:HIP

[ABK95] Ron Avitzur, Olaf Bachmann, and Norbert Kajler. From honest to intelligent plotting. In Levelt [Lev95], pages 32–41. ISBN 0-89791-699-9. LCCN QA 76.95 I59 1995. URL http://www.

acm.org:80/pubs/citations/proceedings/issac/220346/

p32-avitzur/. ACM order number: 505950.

Abbott:1999:FDC

[ABM99] John Abbott, Manuel Bronstein, and Thom Mulders. Fast deterministic computation of determinants of dense matrices. In Dooley [Doo99], pages 197–204. ISBN 1-58113-073-2.

LCCN QA76.95 .I57 1999. URL http://www.acm.org/pubs/
contents/proceedings/issac/309831/.

Andreoli:1994:CKB

[ABP94] J.-M. Andreoli, U. M. Borghoff, and R. Pareschi. Constraintbased knowledge brokers. In Hong [Hon94], pages 1–11. ISBN 981-02-2040-5. LCCN QA76.642.I58 1994.

Abramov:1995:PSL

[ABP95] Sergei A. Abramov, Manuel Bronstein, and Marko Petkovšek. On polynomial solutions of linear operator equations. In Levelt [Lev95], pages 290–296. ISBN 0-89791-699-9. LCCN QA 76.95

I59 1995. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/220346/p290-abramov/. ACM order

number: 505950.

Abramov:1993:DS

[Abr93a] S. A. Abramov. On d’Alembert substitution. In Bronstein [Bro93], pages 20–26. ISBN 0-89791-604-2. LCCN QA 76.95

I59 1993. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/164081/p20-abramov/. ACM order num-

ber: 505930.

Abramov:1993:DSP

[Abr93b] S. A. Abramov. On d’Alembert substitution. In Bronstein [Bro93], pages 20–26. ISBN 0-89791-604-2. LCCN QA 76.95 I59 1993. ACM order number: 505930.

Abramov:1995:ISR

[Abr95a] S. A. Abramov. Indefinite sums of rational functions. In Levelt [Lev95], pages 303–308. ISBN 0-89791-699-9. LCCN QA 76.95

I59 1995. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/220346/p303-abramov/. ACM order

number: 505950.

Abramov:1995:RSL

[Abr95b] S. A. Abramov. Rational solutions of linear difference and q-difference equations with polynomial coefficients. In Levelt [Lev95], pages 285–289. ISBN 0-89791-699-9. LCCN QA 76.95

I59 1995. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/220346/p285-abramov/. ACM order

number: 505950.

Abramov:2002:AZA

[Abr02] S. A. Abramov. Applicability of Zeilberger’s algorithm to hypergeometric terms. In Mora [Mor02], pages 1–7. ISBN 1-58113-484-3. LCCN QA76.95. URL http://www.lifl.fr/

ISSAC2002/.

Abramov:2006:SRS

[Abr06] S. A. Abramov. On the summation of P -recursive sequences. In Trager [Tra06], pages 17–22. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

Abramov:2008:PSL

[Abr08] Sergei A. Abramov. Power series and linear difference equations. In Jeffrey [Jef08], pages 1–2. ISBN 1-59593-904-0. LCCN ????

Abramov:2010:SDU

[Abr10] S. A. Abramov. On some decidable and undecidable problems related to q-difference equations with parameters. In Watt [Wat10], pages 311–317. ISBN 1-4503-0150-9. LCCN QA76.95 .I59 2010.

Abate:1997:ADS

[ABRC97] Jason Abate, Christian Bischof, Lucas Roh, and Alan Carle. Algorithms and design for a second-order automatic differentiation module. In Küchlin [Küc97], pages 149–155. ISBN 0-89791-875-4. LCCN QA76.95. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/258726/p149-abate/.

Aruliah:2004:NPA

[AC04] D. A. Aruliah and Robert M. Corless. Numerical parameterization of affine varieties using. In Gutierrez [Gut04], pages 12–18. ISBN 1-58113-827-X. LCCN QA76.95 .I57 2004.

Aroca:2005:AGS

[ACFG05] J. M. Aroca, J. Cano, R. Feng, and X. S. Gao. Algebraic general solutions of algebraic ordinary differential equations. In Kauers [Kau05], pages 29–36. ISBN 1-59593-095-7. LCCN ???? ACM Order Number 505050.

Aroca:2003:PSS

[ACJ03] F. Aroca, J. Cano, and F. Jung. Power series solutions for nonlinear PDE’s. In Senda [Sen03], pages 15–22. ISBN 1-58113-641-2. LCCN QA76.95. ACM order number 505030.

ACM:1993:PFA

[ACM93] ACM, editor. Proceedings of the Fourth ACM SIGPLAN Symposium on Principles and Practice of Parallel Programming, PPOPP: San Diego, California, May 19–22, 1993, volume 28(7) of ACM SIGPLAN Notices. ACM Press, New York, NY 10036, USA, 1993. ISBN 0-89791-589-5. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.642.A27 1993.

ACM:1994:IPI

[ACM94] ACM, editor. ISSAC ’94: Proceedings of the 1994 International Symposium on Symbolic and Algebraic Computation: July 20– 22, 1994, Oxford, England, United Kingdom. ACM Press, New York, NY 10036, USA, 1994. ISBN 0-89791-638-7. LCCN QA76.95.I59 1994.

Abdali:1986:OOA

[ACS86] S. K. Abdali, Guy W. Cherry, and Neil Soiffer. An objectoriented approach to algebra system design. In Char [Cha86b], pages 24–30. ISBN 0-89791-199-7 (paperback). LCCN

QA155.7.E4 A281 1986. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/32439/p24-abdali/. ACM

order no. 505860.

Ananth:2011:BBD

[AD11] Prabhanjan Vijendra Ananth and Ambedkar Dukkipati. Border basis detection is NP-complete. In Schost and Emiris [SE11], pages 11–18. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Adamchik:1997:CLI

[Ada97] Victor Adamchik. A class of logarithmic integrals. In Küchlin [Küc97], pages 1–8. ISBN 0-89791-875-4. LCCN

QA76.95. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/258726/p1-adamchik/.

Adamchik:2001:BF

[Ada01] V. S. Adamchik. On the Barnes function. In Mourrain [Mou01], pages 15–20. ISBN 1-58113-417-7. LCCN QA76.95.I59 2001. ACM order number 505010.

Arsac:1999:DCC

[ADG99] Olivier Arsac, Stéphane Dalmas, and Marc Gaétano. Design of a customizable component to display and edit formulas. In Dooley [Doo99], pages 283–290. ISBN 1-58113-073-2.

LCCN QA76.95 .I57 1999. URL http://www.acm.org/pubs/
contents/proceedings/issac/309831/.

Arnborg:1986:ADR

[AF86] S. Arnborg and H. Feng. Algebraic decomposition of regular curves. In Char [Cha86b], pages 53–55. ISBN 0-89791-199-7 (paperback). LCCN QA155.7.E4 A281 1986. URL http://www.

acm.org:80/pubs/citations/proceedings/issac/32439/p53-

arnborg/. ACM order no. 505860.

Amrhein:1996:CSM

[AGK96] Beatrice Amrhein, Oliver Gloor, and Wolfgang Küchlin. A case study of multi-threaded Gröbner basis completion. In Lakshman Y. N. [Lak96], pages 95–102. ISBN 0-89791-796-0. LCCN

QA 76.95 I59 1996. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/236869/p95-amrhein/.

Adams:1999:ATP

[AGLM99] A. A. Adams, H. Gottliebsen, S. A. Linton, and U. Martin. Automated theorem proving in support of computer algebra: symbolic definite integration as a case study. In Dooley [Doo99], pages 253–260. ISBN 1-58113-073-2.

LCCN QA76.95 .I57 1999. URL http://www.acm.org/pubs/
contents/proceedings/issac/309831/.

Adleman:1994:ANT

[AH94] L. M. Adleman and M.-D. Huang, editors. Algorithmic Number Theory. First International Symposium, ANTS-I. Proceedings. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 0-387-58691-1 (New York), 3-540-58691-1 (Berlin). LCCN QA241.A43 1994.

Aschenbrenner:2008:AFS

[AH08] Matthias Aschenbrenner and Christopher J. Hillar. An algorithm for finding symmetric Gröbner bases in infinite dimensional rings. In Jeffrey [Jef08], pages 117–124. ISBN 1-59593-904-0. LCCN ????

Ahrendt:1996:FHC

[Ahr96] Timm Ahrendt. Fast high-precision computations of complex square roots. In Lakshman Y. N. [Lak96], pages 142–149. ISBN 0-89791-796-0. LCCN QA 76.95 I59 1996. URL http://www.

acm.org:80/pubs/citations/proceedings/issac/236869/
p142-ahrendt/.

Anai:2005:SRP

[AHY05] Hirokazu Anai, Shinji Hara, and Kazuhiro Yokoyama. Sum of roots with positive real parts. In Kauers [Kau05], pages 21–28. ISBN 1-59593-095-7. LCCN ???? ACM Order Number 505050.

Avendano:2010:RNC

[AIRR10] Martín Avendaño, Ashraf Ibrahim, J. Maurice Rojas, and Korben Rusek. Randomized NP-completeness for p-adic rational roots of sparse polynomials in one variable. In Watt [Wat10], pages 331–338. ISBN 1-4503-0150-9. LCCN QA76.95 .I59 2010.

Aityan:1995:PNP

[Ait95] S. K. Aityan, editor. Proceedings of neural, parallel and scientific computations: proceedings of the First International Conference on Neural, Parallel and Scientific Computations held at Morehouse College, Atlanta, USA, May 28–31, 1995, volume 1 of Proceedings of Neural Parallel and Scientific Computations. Dynamic Publishers, Inc, Atlanta, GA, 1995. ISBN 0-9640398-9-3, 0-9640398-8-5. LCCN QA76.87 .I58 1995.

Aguirre:2001:GIM

[AJL01] Edith Aguirre, Abdul Salam Jarrah, and Reinhard Laubenbacher. Generic ideals and Moreno-Socías conjecture. In Mourrain [Mou01], pages 21–23. ISBN 1-58113-417-7. LCCN QA76.95.I59 2001. ACM order number 505010.

Abramov:1991:FAS

[AK91a] S. A. Abramov and K. Yu. Kvashenko. Fast algorithms to search for the rational solutions of linear differential equations with polynomial coefficients. In Watt [Wat91], pages 267–270. ISBN 0-89791-437-6. LCCN QA 76.95

I59 1991. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/120694/p267-abramov/.

Apel:1991:FAA

[AK91b] Joachim Apel and Uwe Klaus. FELIX: an assistant for algebraists. In Watt [Wat91], pages 382–389. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/120694/p382-apel/.

Abramov:2012:VMS

[AK12] S. A. Abramov and D. E. Khmelnov. On valuations of meromorphic solutions of arbitrary-order linear difference systems with polynomial coefficients. In van der Hoeven and van Hoeij [vdHvH12], pages 12–19. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Akritis:1986:TNU

[Akr86] Alkiviadis G. Akritis. There is no “Uspensky’s method”. In Char [Cha86b], pages 88–90. ISBN 0-89791-199-7 (paperback). LCCN QA155.7.E4 A281 1986. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/32439/p88-akritis/.

ACM order no. 505860.

Antritter:2008:TCA

[AL08] Felix Antritter and Jean Lévine. Towards a computer algebraic algorithm for flat output determination. In Jeffrey [Jef08], pages 7–14. ISBN 1-59593-904-0. LCCN ????

Albrecht:2012:MLD

[Alb12] Martin R. Albrecht. The M4RIE library for dense linear algebra over small fields with even characteristic. In van der Hoeven and van Hoeij [vdHvH12], pages 28–34. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Andres:2009:PIB

[ALM09] Daniel Andres, Viktor Levandovskyy, and Jorge Martìn Morales. Principal intersection and Bernstein–Sato polynomial of an affine variety. In May [May09], pages 231–238. ISBN 1-60558-609-9. LCCN ????

Abramov:2003:RCF

[ALP03] S. A. Abramov, H. Q. Le, and M. Petkovšek. Rational canonical forms and efficient representations of hypergeometric terms. In Senda [Sen03], pages 7–14. ISBN 1-58113-641-2. LCCN QA76.95. ACM order number 505030.

Adamchik:1990:ACI

[AM90] V. S. Adamchik and O. I. Marichev. The algorithm for calculating integrals of hypergeometric type functions and its realization in REDUCE system. In Watanabe and Nagata [WN90], pages 212–224. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95 .I57 1990. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/96877/p212-
adamchik/.

Aparicio-Monforte:2011:FFI

[AMBSW11] Ainhoa Aparicio-Monforte, Moulay A. Barkatou, Sergi Simon, and Jacques-Arthur Weil. Formal first integrals along solutions of differential systems I. In Schost and Emiris [SE11], pages 19–26. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Alonso:1989:CAS

[AMR89] M. E. Alonso, T. Mora, and M. Raimondo. Computing with algebraic series. In Gonnet [Gon89], pages 101– 111. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989.

US$29.00. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/74540/p101-alonso/. ACM order num-

ber: 505890. English and French.

Achatz:2008:DPE

[AMW08] Melanie Achatz, Scott McCallum, and Volker Weispfenning. Deciding polynomial-exponential problems. In Jeffrey [Jef08], pages 215–222. ISBN 1-59593-904-0. LCCN ????

Anonymous:1995:IA

[Ano95a] Anonymous. ISSAC ’95: Announcement. ACM SIGNUM Newsletter, 30(2):12–??, 1995. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Anonymous:1995:IIS

[Ano95b] Anonymous. ISSAC ’96: International Symposium on Symbolic and Algebraic Computation. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 29 (3&4):19, December 1995. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Abramov:1994:DSL

[AP94] Sergei A. Abramov and Marko Petkovšek. D’Alembertian solutions of linear differential and difference equations. In ACM [ACM94], pages 169–174. ISBN 0-89791-638-7. LCCN

QA76.95.I59 1994. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/190347/p169-abramov/.

Abramov:2001:MDI

[AP01] Sergei A. Abramov and M. Petkovsek. Minimal decomposition of indefinite hypergeometric sums. In Mourrain [Mou01], pages 7–14. ISBN 1-58113-417-7. LCCN QA76.95.I59 2001. ACM order number 505010.

Abramov:2005:GAA

[AP05] S. A. Abramov and M. Petkovšsek. Gosper’s algorithm, accurate summation, and the discrete Newton–Leibniz formula. In Kauers [Kau05], pages 5–12. ISBN 1-59593-095-7. LCCN ???? ACM Order Number 505050.

Abo:2006:IKC

[AP06] Hirotachi Abo and Chris Peterson. Implementation of Kumar’s correspondence. In Trager [Tra06], pages 9–16. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

Arnborg:1989:EPO

[Arn89] S. Arnborg. Experiments with a projection operator for algebraic decomposition. In Gianni [Gia89], pages 177–182. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Arnold:2013:NTF

[Arn13] Andrew Arnold. A new truncated Fourier Transform algorithm. In Monagan et al. [MCG13], pages 15–22. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

A:1989:SSG

[ArRL89] R. A., J. r. Ravenscroft, and E. A. Lamagna. Symbolic summation with generating functions. In Gonnet [Gon89], pages 228–233. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989.

US$29.00. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/74540/p228-ravenscroft/. ACM order

number: 505890. English and French.

Andradas:1997:ROR

[ARS97] Carlos Andradas, Tomás Recio, and J. Rafael Sendra. A relatively optimal rational space curve reparametrization algorithm through canonical divisors. In Küchlin [Küc97], pages 349–355. ISBN 0-89791-875-4. LCCN QA76.95. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/258726/p349-
andradas/.

Andradas:1999:BFR

[ARS99] Carlos Andradas, Taomás Recio, and J. Refael Sendra. Base field restriction techniques for parametric curves. In Dooley [Doo99], pages 17–22. ISBN 1-58113-073-2. LCCN QA76.95

.I57 1999.  URL http://www.acm.org/pubs/contents/
proceedings/issac/309831/.

Ausiello:1989:DMP

[ASN89] G. Ausiello, A. Marchetti Spaccamela, and U. Nanni. Dynamic maintenance of paths and path expressions on graphs. In Gianni [Gia89], pages 1–12. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Astrelin:1991:BDI

[Ast91] A. V. Astrelin. A bound of degree of irreducible eigenpolynomial of some differential operator. In Watt [Wat91], pages 265–266. ISBN 0-89791-437-6. LCCN QA 76.95

I59 1991. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/120694/p265-astrelin/.

Adams:2005:SSR

[ASW05] Jeffrey Adams, B. David Saunders, and Zhendong Wan. Signature of symmetric rational matrices and the unitary dual of Lie groups. In Kauers [Kau05], pages 13–20. ISBN 1-59593-095-7. LCCN ???? ACM Order Number 505050.

Abbott:2000:FSP

[ASZ00] John Abbott, Victor Shoup, and Paul Zimmermann. Factorization in Z[x]: the searching phase. In Traverso [Tra00], pages 1–7. ISBN 1-58113-218-2. LCCN QA76.95.I59 2000. URL http:/

/www.acm.org/pubs/articles/proceedings/issac/345542/
p1-abbott/p1-abbott.pdf;   http://www.acm.org/pubs/
citations/proceedings/issac/345542/p1-abbott/. ACM

order number 505000.

Attardi:1994:SPB

[AT94] G. Attardi and C. Traverso. A strategy-accurate parallel Buchberger algorithm. In Hong [Hon94], pages 12–21. ISBN 981-02-2040-5. LCCN QA76.642.I58 1994.

Abramov:1993:GCD

[AuK93] S. A. Abramov and K. Y. u. Kvashenko. On the greatest common divisor of polynomials which depend on a parameter. In Bronstein [Bro93], pages 152–156. ISBN 0-89791-604-2. LCCN QA 76.95 I59 1993. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/164081/p152-

abramov/. ACM order number: 505930.

Abramov:1997:MIS

[Av97] Sergei A. Abramov and Mark van Hoeij. A method for the integration of solutions of Øre equations. In Küchlin [Küc97], pages 172–175. ISBN 0-89791-875-4. LCCN

QA76.95. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/258726/p172-abramov/.

Adrovic:2012:CPS

[AV12] Danko Adrovic and Jan Verschelde. Computing Puiseux series for algebraic surfaces. In van der Hoeven and van Hoeij [vdHvH12], pages 20–27. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Abramov:1999:DLD

[AvH99] Sergei A. Abramov and Mark van Hoeij. Desingularization of linear difference operators with polynomial coefficients. In Dooley [Doo99], pages 269–275. ISBN 1-58113-073-2.

LCCN QA76.95 .I57 1999. URL http://www.acm.org/pubs/
contents/proceedings/issac/309831/.

Abdali:1989:EQR

[AW89a] S. K. Abdali and D. S. Wiset. Experiments with quadtree representation of matrices. In Gianni [Gia89], pages 96–108. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Avenhaus:1989:URT

[AW89b] J. Avenhaus and D. Wißmann. Using rewriting techniques to solve the generalized word problem in polycyclic groups. In Gonnet [Gon89], pages 322–337. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989. US$29.00. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/74540/p322-

avenhaus/. ACM order number: 505890. English and French.

Anai:2000:DLT

[AW00] Hirokazu Anai and Volker Weispfenning. Deciding lineartrigonometric problems. In Traverso [Tra00], pages 14– 22. ISBN 1-58113-218-2. LCCN QA76.95.I59 2000. URL

http://www.acm.org/pubs/articles/proceedings/issac/
345542/p14-anai/p14-anai.pdf;   http://www.acm.org/
pubs/citations/proceedings/issac/345542/p14-anai/. ACM

order number 505000.

Abramov:1996:DSI

[AZ96] Sergei A. Abramov and Eugene V. Zima. D’Alembertian solutions of inhomogeneous linear equations (differential, difference, and some other). In Lakshman Y. N. [Lak96], pages 232–240. ISBN 0-89791-796-0. LCCN QA 76.95

I59 1996. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/236869/p232-abramov/.

Abramov:1997:MCF

[AZ97] Sergei A. Abramov and Eugene V. Zima. Minimal completely factorable annihilators. In Küchlin [Küc97], pages 290–297.

ISBN 0-89791-875-4. LCCN QA76.95. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/258726/p290-
abramov/.

Amirkhanov:1991:BOV

[AZZ91] I. V. Amirkhanov, E. P. Zhidkov, and I. E. Zhidkova. The betatron oscillations in the vicinity of nonlinear resonance in cyclic accelerator investigation. In Watt [Wat91], pages 452–453. ISBN 0-89791-437-6. LCCN QA 76.95

I59 1991. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/120694/p452-amirkhanov/.

Bacher:1996:AGO

[Bac96] Rainer Bacher. Automatic generation of optimization code based on symbolic non-linear domain formulation. In Lakshman Y. N. [Lak96], pages 283–291. ISBN 0-89791-796-0. LCCN

QA 76.95 I59 1996. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/236869/p283-bacher/.

Baddoura:1994:CIF

[Bad94] Jamil Baddoura. A conjecture on integration in finite terms with elementary functions and polylogarithms. In ACM [ACM94], pages 158–162. ISBN 0-89791-638-7. LCCN

QA76.95.I59 1994. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/190347/p158-baddoura/.

Binder:2000:ANR

[BAE^{+}00] Franz Binder, Erhard Aichinger, Jürgen Ecker, Christof Nöbauer, and Peter Mayr. Algorithms for near-rings of non-linear transformations. In Traverso [Tra00], pages 23– 29. ISBN 1-58113-218-2. LCCN QA76.95.I59 2000. URL

http://www.acm.org/pubs/articles/proceedings/issac/
345542/p23-binder/p23-binder.pdf;   http://www.acm.
org/pubs/citations/proceedings/issac/345542/p23-binder/

. ACM order number 505000.

Bajaj:1986:LAS

[Baj86] Chanderjit Bajaj. Limitations to algorithm solvability: Galois methods and models of computation. In Char [Cha86b], pages 71–76. ISBN 0-89791-199-7 (paperback). LCCN

QA155.7.E4 A281 1986. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/32439/p71-bajaj/. ACM or-

der no. 505860.

Barkatou:1989:RLS

[Bar89a] M. A. Barkatou. On the reduction of linear systems of difference equations. In Gonnet [Gon89], pages 1–6. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989. US$29.00. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/74540/p1-barkatou/

. ACM order number: 505890. English and French.

Barkatou:1989:RNA

[Bar89b] M. A. Barkatou. Rational Newton algorithm for computing formal solutions of linear differential equations. In Gianni [Gia89], pages 183–195. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Barkatou:1995:RVM

[Bar95] A. Barkatou. A rational version of Moser’s algorithm. In Levelt [Lev95], pages 297–302. ISBN 0-89791-699-9. LCCN QA 76.95

I59 1995. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/220346/p297-barkatou/. ACM order

number: 505950.

Barkatou:1999:RSM

[Bar99] M. A. Barkatou. Rational solutions of matrix difference equations: the problem of equivalence and factorization. In Dooley [Doo99], pages 277–282. ISBN 1-58113-073-2.

LCCN QA76.95 .I57 1999. URL http://www.acm.org/pubs/
contents/proceedings/issac/309831/.

Barnett:2003:CCA

[Bar03] Michael P. Barnett. Chemistry and computer algebra: past, present, future. In Senda [Sen03], pages 1–2. ISBN 1-58113-641-2. LCCN QA76.95. ACM order number 505030.

Barkatou:2010:SMS

[Bar10] Moulay A. Barkatou. Symbolic methods for solving systems of linear ordinary differential equations. In Watt [Wat10], pages 7–8. ISBN 1-4503-0150-9. LCCN QA76.95 .I59 2010.

Basu:1997:UQE

[Bas97] Saugata Basu. Uniform quantifier elimination and constraint query processing. In Küchlin [Küc97], pages 21–27. ISBN 0-89791-875-4. LCCN QA76.95. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/258726/p21-basu/.

Bayer:2004:ODO

[Bay04] Thomas Bayer. Optimal descriptions of orbit spaces and strata of finite groups. In Gutierrez [Gut04], pages 19–26. ISBN 1-58113-827-X. LCCN QA76.95 .I57 2004.

Bini:2007:SMB

[BB07] Dario A. Bini and Paola Boito. Structured matrix-based methods for polynomial ∈-gcd: analysis and comparisons. In Brown [Bro07], pages 9–16. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Bernardi:2011:MPD

[BBCM11] Alessandra Bernardi, Jérôme Brachat, Pierre Comon, and Bernard Mourrain. Multihomogeneous polynomial decomposition using moment matrices. In Schost and Emiris [SE11], pages 35–42. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Beaumont:2003:BSE

[BBD03] James Beaumont, Russell Bradford, and James H. Davenport. Better simplification of elementary functions through power series. In Senda [Sen03], pages 30–36. ISBN 1-58113-641-2. LCCN QA76.95. ACM order number 505030.

Beaumont:2004:PAA

[BBDP04] James C. Beaumont, Russell J. Bradford, James H. Davenport, and Nalina Phisanbut. A poly-algorithmic approach to simplifying elementary functions. In Gutierrez [Gut04], pages 27–34. ISBN 1-58113-827-X. LCCN QA76.95 .I57 2004.

Beaumont:2005:ABT

[BBDP05] James C. Beaumont, Russell J. Bradford, James H. Davenport, and Nalina Phisanbut. Adherence is better than adjacency: computing the Riemann index using CAD. In Kauers [Kau05], pages 37–44. ISBN 1-59593-095-7. LCCN ???? ACM Order Number 505050.

Barkatou:2008:RSL

[BBP08] Moulay A. Barkatou, Gary Broughton, and Eckhard Pflügel. Regular systems of linear functional equations and applications. In Jeffrey [Jef08], pages 15–22. ISBN 1-59593-904-0. LCCN ????

Barkatou:2010:SRC

[BBP10] Moulay A. Barkatou, Carole El Bacha, and Eckhard Pflügel. Simultaneously row- and column-reduced higher-order linear differential systems. In Watt [Wat10], pages 45–52. ISBN 1-4503-0150-9. LCCN QA76.95 .I59 2010.

Babai:1993:DFM

[BBR93] László Babai, Robert Beals, and Daniel Rockmore. Deciding finiteness of matrix groups in deterministic polynomial time. In Bronstein [Bro93], pages 117–126. ISBN 0-89791-604-2. LCCN

QA 76.95 I59 1993. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/164081/p117-babai/. ACM

order number: 505930.

BoydelaTour:1989:FAS

[BC89a] T. Boy de la Tour and R. Caferra. A formal approach to some usually informal techniques used in mathematical reasoning. In Gianni [Gia89], pages 402–406. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Butler:1989:CVU

[BC89b] G. Butler and J. Cannon. Cayley, version 4: the user language. In Gianni [Gia89], pages 456–466. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Burnel:1992:CCY

[BC92] A. Burnel and H. Caprasse. The computation of 1-loop contributions in Y.M. theories with class III nonrelativistic gauges and REDUCE. In Wang [Wan92b], pages 103–107. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN QA76.95.I59 1992. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/143242/p103-burnel/

. ACM order number: 505920.

Barendegt:2000:RHM

[BC00] Henk Barendegt and Arjeh M. Cohen. Representing and handling mathematical concepts by humans and machines. In Traverso [Tra00], pages 6–?? ISBN 1-58113-218-2.

LCCN QA76.95.I59 2000. URL http://www.acm.org/pubs/
citations/proceedings/issac/345542/p6-barendegt/. ACM

order number 505000.

Barkatou:2009:ARS

[BCB09] Moulay A. Barkatou, Thomas Cluzeau, and Carole El Bacha. Algorithms for regular solutions of higher-order linear differential systems. In May [May09], pages 7–14. ISBN 1-60558-609-9. LCCN ????

Bostan:2013:HRC

[BCC^{+}13] Alin Bostan, Shaoshi Chen, Frédéric Chyzak, Ziming Li, and Guoce Xin. Hermite reduction and creative telescoping for hyperexponential functions. In Monagan et al. [MCG13], pages 77–84. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

Bischof:1992:AAD

[BCCG92] Christian Bischof, Alan Carle, George Corliss, and Andreas Griewank. ADIFOR: Automatic differentiation in a source translator environment. In Wang [Wan92b], pages 294– 302. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN QA76.95.I59 1992. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/143242/p294-

bischof/. ACM order number: 505920.

Bostan:2010:CCT

[BCCL10] Alin Bostan, Shaoshi Chen, Frédéric Chyzak, and Ziming Li. Complexity of creative telescoping for bivariate rational functions. In Watt [Wat10], pages 203–210. ISBN 1-4503-0150-9. LCCN QA76.95 .I59 2010.

Bostan:2013:CET

[BCdP13] Alin Bostan, Frédéric Chyzak, and Élie de Panafieu. Complexity estimates for two uncoupling algorithms. In Monagan et al. [MCG13], pages 85–92. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

Barkatou:2012:CCF

[BCEW12] M. A. Barkatou, T. Cluzeau, C. El Bacha, and J.-A. Weil. Computing closed form solutions of integrable connections. In van der Hoeven and van Hoeij [vdHvH12], pages 43–50. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Brown:1989:SPP

[BCF89] C. Brown, G. Cooperman, and L. Finkelstein. Solving permutation problems using rewriting systems. In Gianni [Gia89], pages 364–377. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Babai:1991:NLT

[BCFS91] László Babai, Gene Cooperman, Larry Finkelstein, and Ákos Seress. Nearly linear time algorithms for permutation groups with a small base. In Watt [Wat91], pages 200–209. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/120694/p200-
babai/.

Bajaj:1989:FRP

[BCGW89] C. Bajaj, J. Canny, T. Garrity, and J. Warren. Factoring rational polynomials over the complexes. In Gonnet [Gon89], pages 81–90. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989.

US$29.00. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/74540/p81-bajaj/. ACM order number:

505890. English and French.

Burr:2008:CSA

[BCGY08] Michael Burr, Sung Woo Choi, Benjamin Galehouse, and Chee K. Yap. Complete subdivision algorithms, II: isotopic meshing of singular algebraic curves. In Jeffrey [Jef08], pages 87–94. ISBN 1-59593-904-0. LCCN ????

Bernardin:1999:SCJ

[BCK99] Laurent Bernardin, Bruce Char, and Erich Kaltofen. Symbolic computation in Java: an appraisement. In Dooley [Doo99], pages 237–244. ISBN 1-58113-073-2. LCCN QA76.95

.I57 1999.  URL http://www.acm.org/pubs/contents/
proceedings/issac/309831/.

Beckermann:1997:FCM

[BCL97] Bernhard Beckermann, Stan Cabay, and George Labahn. Fraction-free computation of matrix Padé systems. In Küchlin [Küc97], pages 125–132. ISBN 0-89791-875-4. LCCN

QA76.95. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/258726/p125-beckermann/.

Beckermann:2002:FFR

[BCL02] Bernhard Beckermann, Howard Cheng, and George Labahn. Fraction-free row reduction of matrices of skew polynomials.

In Mora [Mor02], pages 8–15. ISBN 1-58113-484-3. LCCN

QA76.95. URL http://www.lifl.fr/ISSAC2002/.

Buchberger:1982:CAS

[BCLA82] Bruno Buchberger, George Edward Collins, Rudiger Loos, and R. Albrecht. Computer algebra: symbolic and algebraic computation, volume 4 of Computing. Supplementum. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 0-387-81684-4. vi + 283 pp. LCCN QA155.7.E4 C65 1982.

Berthe:2013:MGP

[BCLV13] Valérie Berthé, Jean Creusefond, Loïck Lhote, and Brigitte Vallée. Multiple GCDs. Probabilistic analysis of the plain algorithm. In Monagan et al. [MCG13], pages 37–44. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

Bosma:1994:PAS

[BCM94] Wieb Bosma, John Cannon, and Graham Matthews. Programming with algebraic structures: design of the Magma language. In ACM [ACM94], pages 52–57. ISBN 0-89791-638-7. LCCN QA76.95.I59 1994. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/190347/p52-bosma/.

Bostan:2008:POD

[BCR08] Alin Bostan, Frédéric Chyzak, and Nicolas Le Roux. Products of ordinary differential operators by evaluation and interpolation. In Jeffrey [Jef08], pages 23–30. ISBN 1-59593-904-0. LCCN ????

Bi:2013:SLR

[BCR13] Jingguo Bi, Qi Cheng, and J. Maurice Rojas. Sub-linear root detection, and new hardness results, for sparse polynomials over finite fields. In Monagan et al. [MCG13], pages 61–68. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

Bostan:2005:FAP

[BCS05] Alin Bostan, Thomas Cluzeau, and Bruno Salvy. Fast algorithms for polynomial solutions of linear differential equations. In Kauers [Kau05], pages 45–52. ISBN 1-59593-095-7. LCCN ???? ACM Order Number 505050.

Bostan:2007:DEA

[BCS^{+}07] Alin Bostan, Frédéric Chyzak, Bruno Salvy, Grégoire Lecerf, and Éric Schost. Differential equations for algebraic functions. In Brown [Bro07], pages 25–32. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Bostan:2006:LCA

[BCSC06] A. Bostan, F. Chyzak, B. Salvy, and T. Cluzeau. Low complexity algorithms for linear recurrences. In Trager [Tra06], pages 31–38. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

Bostan:2012:FCC

[BCSL12] Alin Bostan, Frédéric Chyzak, Bruno Salvy, and Ziming Li. Fast computation of common left multiples of linear ordinary differential operators. In van der Hoeven and van Hoeij [vdHvH12], pages 99–106. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Bachmair:1986:CPC

[BD86] Leo Bachmair and Nachum Dershowitz. Critical-pair criteria for the Knuth–Bendix completion procedure. In Char [Cha86b], pages 215–217. ISBN 0-89791-199-7 (paperback). LCCN QA155.7.E4 A281 1986. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/32439/p215-

bachmair/. ACM order no. 505860.

Bradford:1989:ETC

[BD89] R. J. Bradford and J. H. Davenport. Effective tests for cyclotomic polynomials. In Gianni [Gia89], pages 244–251. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Bradford:2002:TBS

[BD02] Russell Bradford and James H. Davenport. Towards better simplification of elementary functions. In Mora [Mor02], pages 16– 22. ISBN 1-58113-484-3. LCCN QA76.95. URL http://www.

lifl.fr/ISSAC2002/.

Buse:2004:IPH

[BD04] Laurent Busé and Carlos D’Andrea. Inversion of parameterized hypersurfaces by means of subresultants. In Gutierrez [Gut04], pages 65–71. ISBN 1-58113-827-X. LCCN QA76.95 .I57 2004.

Brown:2007:CQE

[BD07a] Christopher W. Brown and James H. Davenport. The complexity of quantifier elimination and cylindrical algebraic decomposition. In Brown [Bro07], pages 54–60. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Buse:2007:IBP

[BD07b] Laurent Busé and Marc Dohm. Implicitization of bihomogeneous parametrizations of algebraic surfaces via linear syzygies. In Brown [Bro07], pages 69–76. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Brickenstein:2008:GFN

[BD08] Michael Brickenstein and Alexander Dreyer. Gröbner-free normal forms for Boolean polynomials. In Jeffrey [Jef08], pages 55–62. ISBN 1-59593-904-0. LCCN ????

Bigatti:2009:CSK

[BdC09] Anna M. Bigatti and Eduardo Sáenz de Cabezón. Computation of the (n−1)-st Koszul Homology of monomial ideals and related algorithms. In May [May09], pages 31–38. ISBN 1-60558-609-9. LCCN ????

Bradford:2013:CAD

[BDE^{+}13] Russell Bradford, James H. Davenport, Matthew England, Scott McCallum, and David Wilson. Cylindrical algebraic decompositions for boolean combinations. In Monagan et al. [MCG13], pages 125–132. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

Boyer:2009:MES

[BDPZ09] Brice Boyer, Jean-Guillaume Dumas, Clément Pernet, and Wei Zhou. Memory efficient scheduling of Strassen–Winograd’s matrix multiplication algorithm. In May [May09], pages 55–62. ISBN 1-60558-609-9. LCCN ????

Beals:1993:EAC

[Bea93] Robert Beals. An elementary algorithm for computing the composition factors of a permutation group. In Bronstein [Bro93], pages 127–134. ISBN 0-89791-604-2. LCCN QA 76.95

I59 1993. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/164081/p127-beals/. ACM order num-

ber: 505930.

Belmesk:1990:EME

[Bel90] M. Belmesk. An execution model for exploiting andor parallelism in logic programs (abstract). In Watanabe and Nagata [WN90], page 288. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95

.I57 1990. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/96877/p288-belmesk/.

Bertrand:1994:INA

[Ber94] Laurent Bertrand. On the implementation of a new algorithm for the computation of hyperelliptic integrals. In ACM [ACM94], pages 211–215. ISBN 0-89791-638-7. LCCN

QA76.95.I59 1994. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/190347/p211-bertrand/.

Bernardin:1997:MMP

[Ber97] Laurent Bernardin. Maple on a massively parallel, distributed memory machine. In Küchlin [Küc97], pages 217–222. ISBN 0-89791-875-4. LCCN QA76.95. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/266670/p217-
bernardin/.

Bernardin:1998:BHP

[Ber98] Laurent Bernardin. On bivariate Hensel and its parallelization. In Gloor [Glo98], pages 96–100. ISBN 1-58113-002-3. LCCN

QA155.7.E4 E88 1998. URL http://www.acm.org/pubs/
citations/proceedings/issac/281508/p96-bernardin/.

Betten:2013:RCC

[Bet13] Anton Betten. Rainbow cliques and the classification of small BLT-sets. In Monagan et al. [MCG13], pages 53–60. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

Beukers:2012:HFC

[Beu12] Frits Beukers. A-hypergeometric functions: computational aspects. In van der Hoeven and van Hoeij [vdHvH12], pages 1–2. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Backelin:1991:HWP

[BF91] Jörgen Backelin and Ralf Fröberg. How we proved that there are exactly 924 cyclic 7-roots. In Watt [Wat91], pages 103–111. ISBN 0-89791-437-6. LCCN QA 76.95

I59 1991. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/120694/p103-backelin/.

Berman:1994:OCR

[BF94] Benjamin P. Berman and Richard J. Fateman. Optical character recognition for typeset mathematics. In ACM [ACM94], pages 348–353. ISBN 0-89791-638-7. LCCN

QA76.95.I59 1994. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/190347/p348-berman/.

Bronstein:1999:SLO

[BF99] Manuel Bronstein and Anne Fredet. Solving linear ordi-

nary differential equations over C(x, e^{∫ f(x) dx}). In Dooley [Doo99], pages 173–179. ISBN 1-58113-073-2. LCCN QA76.95

.I57 1999.  URL http://www.acm.org/pubs/contents/
proceedings/issac/309831/.

Basiri:2003:COG

[BF03] Abdolali Basiri and Jean-Charles Faugère. Changing the ordering of Gröbner bases with LLL: case of two variables. In Senda [Sen03], pages 23–29. ISBN 1-58113-641-2. LCCN QA76.95. ACM order number 505030.

Biasse:2012:PTA

[BF12] Jean-François Biasse and Claus Fieker. A polynomial time algorithm for computing the HNF of a module over the integers of a number field. In van der Hoeven and van Hoeij [vdHvH12], pages 75–82. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Bettale:2012:SPS

[BFP12] Luk Bettale, Jean-Charles Faugère, and Ludovic Perret. Solving polynomial systems over finite fields: improved analysis of the hybrid approach. In van der Hoeven and van Hoeij [vdHvH12], pages 67–74. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Babai:1993:DCA

[BFS93] László Babai, Katalin Friedl, and Markus Stricker. Decomposition of 0∗-closed algebras in polynomial time. In Bronstein [Bro93], pages 86–94. ISBN 0-89791-604-2. LCCN QA 76.95

I59 1993. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/164081/p86-babai/. ACM order num-

ber: 505930.

Bretto:2007:GGC

[BG07] Alain Bretto and Luc Gillibert. G-graphs for the cage problem: a new upper bound. In Brown [Bro07], pages 49–53. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Bembe:2011:VRR

[BG11] Daniel Bembé and André Galligo. Virtual roots of a real polynomial and fractional derivatives. In Schost and Emiris [SE11], pages 27–34. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Broadbery:1995:IDE

[BGDW95] P. A. Broadbery, T. Gómez-Díaz, and S. M. Watt. On the implementation of dynamic evaluation. In Levelt [Lev95], pages 77–84. ISBN 0-89791-699-9. LCCN QA 76.95

I59 1995. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/220346/p77-broadbery/. ACM order

number: 505950.

Bueno:1994:CSM

[BGH94] F. Bueno, M. Garcia de la Banda, and M. Hermenegildo. A comparative study of methods for automatic compile-time parallelization of logic programs. In Hong [Hon94], pages 63–73. ISBN 981-02-2040-5. LCCN QA76.642.I58 1994.

Brunick:1999:ECD

[BGHS99] Gerard P. Brunick, Edward L. Green, Lenwood S. Heath, and Craig A. Struble. Efficient construction of Drinfel’d doubles. In Dooley [Doo99], pages 45–52. ISBN 1-58113-073-2.

LCCN QA76.95 .I57 1999. URL http://www.acm.org/pubs/
contents/proceedings/issac/309831/.

Boady:2013:TRS

[BGJ13] Mark Boady, Pavel Grinfeld, and Jeremy Johnson. A term rewriting system for the calculus of moving surfaces. In Monagan et al. [MCG13], pages 69–76. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

Buendgen:1994:MAT

[BGK94a] R. Buendgen, M. Goebel, and W. Kuechlin. Multi-threaded AC term rewriting. In Hong [Hon94], pages 84–93. ISBN 981-02-2040-5. LCCN QA76.642.I58 1994.

Bundgen:1994:FPC

[BGK94b] Reinhard Bündgen, Manfred Göbel, and Wolfgang Küchlin. A fine-grained parallel completion procedure. In ACM [ACM94], pages 269–277. ISBN 0-89791-638-7. LCCN

QA76.95.I59 1994. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/190347/p269-bundgen/.

Buendgen:1996:MAP

[BGK96] R. Buendgen, M. Goebel, and W. Kuechlin. A master-slave approach to parallel term rewriting on a hierarchical multiprocessor. In Calmet and Limongelli [CL96], pages 183–194. ISBN 3-540-61697-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.S88I576 1996.

Bretto:2005:SSG

[BGL05] Alain Bretto, Luc Gillibert, and Bernard Laget. Symmetric and semisymmetric graphs construction using G-graphs. In Kauers [Kau05], pages 61–67. ISBN 1-59593-095-7. LCCN ???? ACM Order Number 505050.

Beth:1991:FGN

[BGM91] T. Beth, W. Geiselmann, and F. Meyer. Finding (good) normal bases in finite fields. In Watt [Wat91], pages 173–178. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/120694/p173-
beth/.

Bournez:2012:CSI

[BGP12] Olivier Bournez, Daniel S. Graça, and Amaury Pouly. On the complexity of solving initial value problems. In van der Hoeven and van Hoeij [vdHvH12], pages 115–121. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Boucher:2003:FOL

[BGU03] Delphine Boucher, Philippe Gaillard, and Felix Ulmer. Fourth order linear differential equations with imprimitive group. In Senda [Sen03], pages 45–49. ISBN 1-58113-641-2. LCCN QA76.95. ACM order number 505030.

Borst:1994:GRP

[BGV94] W. N. Borst, V. V. Goldman, and J. A. Van Hulzen. GEN-TRAN 90: a REDUCE package for the generation of Fortran 90 code. In ACM [ACM94], pages 45–51. ISBN 0-89791-638-7. LCCN QA76.95.I59 1994. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/190347/p45-borst/.

Briot:1996:OBP

[BGY96] Jean-Pierre Briot, Jean-Marc Geib, and Akinori Yonezawa, editors. Object-based parallel and distributed computation: France– Japan Workshop, OBPDC ’95, Tokyo, Japan, June 21–23, 1995: selected papers, volume 1107 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-61487-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.64 .F7 1995.

Borwein:2006:SCM

[BH06] Jonathan M. Borwein and Chris H. Hamiltony. Symbolic computation of multidimensional Fenchel conjugates. In Trager [Tra06], pages 23–30. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

Ballarin:1995:TAI

[BHC95] Clemens Ballarin, Karsten Homann, and Jacques Calmet. Theorems and algorithms: An interface between Isabelle and Maple. In Levelt [Lev95], pages 150–157. ISBN 0-89791-699-9. LCCN QA 76.95 I59 1995. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/220346/p150-

ballarin/. ACM order number: 505950.

Bubeck:1995:DSC

[BHKR95] T. Bubeck, M. Hiller, W. Kuechlin, and W. Rosenstiel. Distributed symbolic computation with DTS. In Ferreira and Rolim [FR95], pages 231–248. CODEN LNCSD9. ISBN 3-540-60321-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.642 .I59 1995.

Bremner:2007:NSP

[BHP07] Murray R. Bremner, Michael J. Hancock, and Yunfeng Piao. Nonassociative structures on polynomial algebras arising from bio-operations on formal languages: an application of computer algebra to nonassociative systems. In Brown [Bro07], pages 41–48. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Bradford:1986:ERD

[BHPS86] R. J. Bradford, A. C. Hearn, J. A. Padget, and E. Schrüfer. Enlarging the REDUCE domain of computation. In Char [Cha86b], pages 100–106. ISBN 0-89791-199-7 (paperback). LCCN QA155.7.E4 A281 1986. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/32439/p100-

bradford/. ACM order no. 505860.

Butler:1991:DDG

[BIL91] Greg Butler, Sridhar S. Iyer, and Susan H. Ley. A deductive database of the groups of order dividing 128. In Watt [Wat91], pages 210–218. ISBN 0-89791-437-6. LCCN QA 76.95

I59 1991. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/120694/p210-butler/.

Binder:1996:FCL

[Bin96] Franz Binder. Fast computations in the lattice of polynomial rational function fields. In Lakshman Y. N. [Lak96], pages 43–48. ISBN 0-89791-796-0. LCCN QA 76.95

I59 1996. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/236869/p43-binder/.

Biscani:2012:PSP

[Bis12] Francesco Biscani. Parallel sparse polynomial multiplication on modern hardware architectures. In van der Hoeven and van Hoeij [vdHvH12], pages 83–90. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Beringer:1998:SSA

[BJ98] Frédéric Beringer and Françoise Jung. Solving “generalized algebraic equations”. In Gloor [Glo98], pages 222–227. ISBN 1-58113-002-3. LCCN QA155.7.E4 E88 1998. URL http://

www.acm.org/pubs/citations/proceedings/issac/281508/
p222-beringer/.

Brisebarre:2010:CIP

[BJ10] Nicolas Brisebarre and Mioara Joldeş. Chebyshev interpolation polynomial-based tools for rigorous computing. In Watt [Wat10], pages 147–154. ISBN 1-4503-0150-9. LCCN QA76.95 .I59 2010.

Buchberger:1997:STP

[BJK^{+}97] Bruno Buchberger, Tudor Jebelean, Franz Kriftner, Mircea Marin, Elena Tomuţa, and Daniela Vāsaru. A survey of the theorema project. In Küchlin [Küc97], pages 384–391. ISBN 0-89791-875-4. LCCN QA76.95. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/258726/p384-
buchberger/.

Bostan:2007:STV

[BJS07] Alin Bostan, Claude-Pierre Jeannerod, and Éric Schost. Solving Toeplitz- and Vandermonde-like linear systems with large displacement rank. In Brown [Bro07], pages 33–40. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Beck:1986:SAL

[BK86] Robert E. Beck and Bernard Kolman. Symbolic algorithms for Lie algebra computation. In Char [Cha86b], pages 85–87. ISBN 0-89791-199-7 (paperback). LCCN

QA155.7.E4 A281 1986. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/32439/p85-beck/. ACM or-

der no. 505860.

Baral:1997:LPN

[BKLG97] C. Baral, V. S. Kreinovich, V. Lifschitz, and M. Gelfond, editors. Logic programming, non-monotonic reasoning and reasoning about actions: Symposium — November 1995, El Paso, TX, volume 21(2) of Annals of Mathematics and Artificial Intelligence. Baltzer Science, Basel, Switzerland, 1997. ISSN 1012-2443.

Baaz:1990:SPR

[BL90] M. Baaz and A. Leitsch. A strong problem reduction method based on function introduction. In Watanabe and Nagata [WN90], pages 30–37. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95

.I57 1990. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/96877/p30-baaz/.

Belkov:1991:RUC

[BL91] Alexander A. Bel’kov and Alexander V. Lanyov. REDUCE usage for calculation of low-energy process amplitudes in chiral QCD model. In Watt [Wat91], pages 454–455. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991. URL http://www.acm.org:

80/pubs/citations/proceedings/issac/120694/p454-bel_
kov/.

Breuer:1998:GTS

[BL98] Thomas Breuer and Steve Linton. The GAP 4 type system organising algebraic algorithms. In Gloor [Glo98], pages 38– 45. ISBN 1-58113-002-3. LCCN QA155.7.E4 E88 1998. URL

http://www.acm.org/pubs/citations/proceedings/issac/
281508/p38-breuer/.

Boulier:2000:CCR

[BL00] François Boulier and François Lemaire. Computing canonical representatives of regular differential ideals. In Traverso [Tra00], pages 38–47. ISBN 1-58113-218-2. LCCN QA76.95.I59 2000.

URL http://www.acm.org/pubs/articles/proceedings/issac/
345542/p38-boulier/p38-boulier.pdf;   http://www.acm.
org/pubs/citations/proceedings/issac/345542/p38-boulier/

. ACM order number 505000.

Bronstein:2002:SLO

[BL02] Manuel Bronstein and Sébastien Lafaille. Solutions of linear ordinary differential equations in terms of special functions. In Mora [Mor02], pages 23–28. ISBN 1-58113-484-3. LCCN

QA76.95. URL http://www.lifl.fr/ISSAC2002/.

Beckermann:2009:FFC

[BL09] Bernhard Beckermann and George Labahn. Fraction-free computation of simultaneous Padé approximants. In May [May09], pages 15–22. ISBN 1-60558-609-9. LCCN ????

Berkesch:2010:ABS

[BL10] Christine Berkesch and Anton Leykin. Algorithms for Bernstein–Sato polynomials and multiplier ideals. In Watt [Wat10], pages 99–106. ISBN 1-4503-0150-9. LCCN QA76.95 .I59 2010.

Berthomieu:2012:RPA

[BL12] Jérémy Berthomieu and Romain Lebreton. Relaxed p-adic Hensel lifting for algebraic systems. In van der Hoeven and van Hoeij [vdHvH12], pages 59–66. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Boulier:2001:P

[BLM01] François Boulier, François Lemaire, and Marc Moreno Maza. PARDI! In Mourrain [Mou01], pages 38–47. ISBN 1-58113-417-7. LCCN QA76.95.I59 2001. ACM order number 505010.

Bouzidi:2013:RUR

[BLPR13a] Yacine Bouzidi, Sylvain Lazard, Marc Pouget, and Fabrice Rouillier. Rational univariate representations of bivariate systems and applications. In Monagan et al. [MCG13], pages 109– 116. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

Bouzidi:2013:SLF

[BLPR13b] Yacine Bouzidi, Sylvain Lazard, Marc Pouget, and Fabrice Rouillier. Separating linear forms for bivariate systems. In Monagan et al. [MCG13], pages 117–124. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

Boulier:2013:IDF

[BLRR13] François Boulier, François Lemaire, Georg Regensburger, and Markus Rosenkranz. On the integration of differential fractions. In Monagan et al. [MCG13], pages 101–108. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

Basson:2013:EEL

[BLRS13] Romain Basson, Reynald Lercier, Christophe Ritzenthaler, and Jeroen Sijsling. An explicit expression of the Lüroth invariant. In Monagan et al. [MCG13], pages 31–36. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

Bostan:2003:TPP

[BLS03] A. Bostan, G. Lecerf, and É. Schost. Tellegen’s principle into practice. In Senda [Sen03], pages 37–44. ISBN 1-58113-641-2. LCCN QA76.95. ACM order number 505030.

Bostan:2004:CIB

[BLS^{+}04] A. Bostan, G. Lecerf, B. Salvy, É. Schost, and B. Wiebelt. Complexity issues in bivariate polynomial factorization. In Gutierrez [Gut04], pages 42–49. ISBN 1-58113-827-X. LCCN QA76.95 .I57 2004.

Bostan:2013:CTR

[BLS13] Alin Bostan, Pierre Lairez, and Bruno Salvy. Creative telescoping for rational functions using the Griffiths–Dwork method. In Monagan et al. [MCG13], pages 93–100. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

Beckermann:1999:SNF

[BLV99] Bernhard Beckermann, George Labahn, and Gilles Villard. Shifted normal forms of polynomial matrices. In Dooley [Doo99], pages 189–196. ISBN 1-58113-073-2. LCCN QA76.95

.I57 1999.  URL http://www.acm.org/pubs/contents/
proceedings/issac/309831/.

Burger:2004:CFS

[BLvH04] Reinhold Burger, George Labahn, and Mark van Hoeij. Closed form solutions of linear ODEs having elliptic function coefficients. In Gutierrez [Gut04], pages 58–64. ISBN 1-58113-827-X. LCCN QA76.95 .I57 2004.

Bronstein:2005:PVE

[BLW05] Manuel Bronstein, Ziming Li, and Min Wu. Picard–vessiot extensions for linear functional systems. In Kauers [Kau05], pages 68–75. ISBN 1-59593-095-7. LCCN ???? ACM Order Number 505050.

Berndt:1991:ACA

[BLWhW91] R. Berndt, A. Lock, G. Witte, and C. h. Wöll. Application of computer algebra to surface lattice dynamics. In Watt [Wat91], pages 433–438. ISBN 0-89791-437-6. LCCN QA 76.95

I59 1991. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/120694/p433-berndt/.

Buchmann:1991:CNP

[BM91] Johannes Buchmann and Volker Müller. Computing the number of points of elliptic curves over finite fields. In Watt [Wat91], pages 179–182. ISBN 0-89791-437-6. LCCN QA 76.95

I59 1991. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/120694/p179-buchmann/.

Brunat:2004:CIF

[BM04] Josep M. Brunat and Antonio Montes. The characteristic ideal of a finite, connected, regular graph. In Gutierrez [Gut04], pages 50–57. ISBN 1-58113-827-X. LCCN QA76.95 .I57 2004.

Brown:2005:UBE

[BM05] Christopher W. Brown and Scott McCallum. On using biequational constraints in CAD construction. In Kauers [Kau05], pages 76–83. ISBN 1-59593-095-7. LCCN ???? ACM Order Number 505050.

Becker:1994:SSL

[BMMT94] Eberhard Becker, Teo Mora, Maria Grazia Marinari, and Carlo Traverso. The shape of the Shape Lemma. In ACM [ACM94], pages 129–133. ISBN 0-89791-638-7. LCCN

QA76.95.I59 1994. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/190347/p129-becker/.

Bondyfalat:1998:CIM

[BMP98] Didier Bondyfalat, Bernard Mourrain, and Victor Y. Pan. Controlled iterative methods for solving polynomial systems. In Gloor [Glo98], pages 252–259. ISBN 1-58113-002-3. LCCN

QA155.7.E4 E88 1998. URL http://www.acm.org/pubs/
citations/proceedings/issac/281508/p252-bondyfalat/.

Bronstein:1997:SPD

[BMW97] Manuel Bronstein, Thom Mulders, and Jacques-Arthur Weil. On symmetric powers of differential operators. In Küchlin [Küc97], pages 156–163. ISBN 0-89791-875-4. LCCN

QA76.95. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/258726/p156-bronstein/.

Bodnar:2004:EDR

[Bod04] Gábor Bodnár. Efficient desingularization of reducible algebraic sets. In Gutierrez [Gut04], pages 35–41. ISBN 1-58113-827-X. LCCN QA76.95 .I57 2004.

Bodrato:2010:SLM

[Bod10] Marco Bodrato. A Strassen-like matrix multiplication suited for squaring and higher power computation. In Watt [Wat10], pages 273–280. ISBN 1-4503-0150-9. LCCN QA76.95 .I59 2010.

Boulier:1995:RRF

[BOLP95] F. Boulier, F. Ollivier, D. Lazard, and M. Petitot. Representation for the radical of a finitely generated differential ideal. In Levelt [Lev95], pages 158–166. ISBN 0-89791-699-9. LCCN QA 76.95 I59 1995. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/220346/p158-

boulier/. ACM order number: 505950.

Bonacina:1994:RPD

[Bon94] M. P. Bonacina. On the reconstruction of proofs in distributed theorem proving with contraction: a modified Clause-Diffusion method. In Hong [Hon94], pages 22–33. ISBN 981-02-2040-5. LCCN QA76.642.I58 1994.

Bonacina:1997:ESS

[Bon97] Maria Paola Bonacina. Experiments with subdivisions of search in distributed theorem proving. In Küchlin [Küc97], pages 88– 100. ISBN 0-89791-875-4. LCCN QA76.95. URL http://www.

acm.org:80/pubs/citations/proceedings/issac/266670/
p88-bonacina/.

Bessonov:2013:ICP

[BOS13] Mariya Bessonov, Alexey Ovchinnikov, and Maxwell Shapiro. Integrability conditions for parameterized linear difference equations. In Monagan et al. [MCG13], pages 45–52. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

Boucher:1999:APS

[Bou99] Delphine Boucher. About the polynomial solutions of homogeneous linear differential equations depending on parameters. In Dooley [Doo99], pages 261–268. ISBN 1-58113-073-2.

LCCN QA76.95 .I57 1999. URL http://www.acm.org/pubs/
contents/proceedings/issac/309831/.

Boucher:2005:NCI

[Bou05] Delphine Boucher. Non complete integrability of a magnetic satellite in circular orbit. In Kauers [Kau05], pages 53–60. ISBN 1-59593-095-7. LCCN ???? ACM Order Number 505050.

Bini:1990:PPC

[BP90] D. Bini and V. Pan. Parallel polynomial computations by recursive processes. In Watanabe and Nagata [WN90], page 294. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95 .I57 1990. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/96877/p294-bini/.

Bosma:1991:CFG

[BP91] Wieb Bosma and Michael Pohst. Computations with finitely generated modules over Dedekind rings. In Watt [Wat91], pages 151–156. ISBN 0-89791-437-6. LCCN QA 76.95

I59 1991. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/120694/p151-bosma/.

Bini:1993:PCT

[BP93] Dario Bini and Victor Pan. Parallel computations with Toeplitz-like and Hankel-like matrices. In Bronstein [Bro93], pages 193–200. ISBN 0-89791-604-2. LCCN QA 76.95

I59 1993. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/164081/p193-bini/. ACM order num-

ber: 505930.

Burke-Perline:1994:PCU

[BP94] T. Burke-Perline. The parallel computation of f(x)0(00 − 010)0/02 mod h(x) using Sugarbush 1.1. In Hong [Hon94], pages 74–83. ISBN 981-02-2040-5. LCCN QA76.642.I58 1994.

Barkatou:1998:EPL

[BP98] M. A. Barkatou and E. Pflügel. On the equivalence problem of linear differential systems and its application for factoring completely reducible systems. In Gloor [Glo98], pages 268– 275. ISBN 1-58113-002-3. LCCN QA155.7.E4 E88 1998. URL

http://www.acm.org/pubs/citations/proceedings/issac/
281508/p268-barkatou/.

Boyle:2006:AHP

[BP06] Phelim Boyle and Alex Potapchik. Application of high-precision computing for pricing arithmetic Asian options. In Trager [Tra06], pages 39–46. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

Barkatou:2007:CSI

[BP07] Moulay A. Barkatou and Eckhard Pflügel. Computing superirreducible forms of systems of linear differential equations via Moser-reduction: a new approach. In Brown [Bro07], pages 1–8. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Basu:1998:CCS

[BPR98] Saugata Basu, Richard Pollack, and Marie-Françoise Roy. Complexity of computing semi-algebraic descriptions of the connected components of a semi-algebraic set. In Gloor [Glo98], pages 25–29. ISBN 1-58113-002-3. LCCN QA155.7.E4

E88 1998. URL http://www.acm.org/pubs/citations/
proceedings/issac/281508/p25-basu/.

Briek:1994:SCT

[BR94] S. Briek and A. Rauzy. Synchronization of constrained transition systems. In Hong [Hon94], pages 54–62. ISBN 981-02-2040-5. LCCN QA76.642.I58 1994.

Bigatti:2006:CSC

[BR06] Anna Bigatti and Lorenzo Robbiano. CocoA: a system for computations in commutative algebra. In Trager [Tra06], page 6. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

Barkatou:2012:SLO

[BR12] Moulay A. Barkatou and Clemens G. Raab. Solving linear ordinary differential systems in hyperexponential extensions. In van der Hoeven and van Hoeij [vdHvH12], pages 51–58. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Bradford:1989:SRD

[Bra89] R. Bradford. Some results on the defect. In Gonnet [Gon89], pages 129–135. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989.

US$29.00. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/74540/p129-bradford/. ACM order

number: 505890. English and French.

Bradford:1990:PBA

[Bra90] R. Bradford. A parallelization of the Buchberger algorithm. In Watanabe and Nagata [WN90], page 296. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95

.I57 1990. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/96877/p296-bradford/.

Bratvold:1994:PFP

[Bra94] T. A. Bratvold. Parallelising a functional program using a listhomomorphism skeleton. In Hong [Hon94], pages 44–53. ISBN 981-02-2040-5. LCCN QA76.642.I58 1994.

Bronstein:1986:GFA

[Bro86] Manuel Bronstein. Gsolve: a faster algorithm for solving systems of algebraic equations. In Char [Cha86b], pages 247– 249. ISBN 0-89791-199-7 (paperback). LCCN QA155.7.E4

A281 1986. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/32439/p247-bronstein/. ACM order no.

505860.

Bronstein:1989:FRR

[Bro89a] M. Bronstein. Fast reduction of the Risch differential equation. In Gianni [Gia89], pages 64–72. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Bronstein:1989:SRE

[Bro89b] M. Bronstein. Simplification of real elementary functions. In Gonnet [Gon89], pages 207–211. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989. US$29.00. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/74540/p207-

bronstein/. ACM order number: 505890. English and French.

Bronstein:1991:RDE

[Bro91] Manuel Bronstein. The Risch differential equation on an algebraic curve. In Watt [Wat91], pages 241–246. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/120694/p241-
bronstein/.

Bronstein:1992:LOD

[Bro92] Manuel Bronstein. Linear ordinary differential equations: breaking through the order 2 barrier. In Wang [Wan92b], pages 42–48. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN QA76.95.I59 1992. URL http://www.

acm.org:80/pubs/citations/proceedings/issac/143242/

p42-bronstein/. ACM order number: 505920.

Bronstein:1993:IPI

[Bro93] Manuel Bronstein, editor. ISSAC’93: proceedings of the 1993 International Symposium on Symbolic and Algebraic Computation, July 6–8, 1993, Kiev, Ukraine. ACM Press, New York, NY 10036, USA, 1993. ISBN 0-89791-604-2. LCCN QA 76.95 I59 1993. ACM order number: 505930.

Bronstein:1994:IAF

[Bro94] Manuel Bronstein. An improved algorithm for factoring linear ordinary differential operators. In ACM [ACM94], pages 336–

340. ISBN 0-89791-638-7. LCCN QA76.95.I59 1994. URL http:

//www.acm.org:80/pubs/citations/proceedings/issac/190347/
p336-bronstein/.

Brown:1998:STI

[Bro98] Christopher W. Brown. Simplification of truth-invariant cylindrical algebraic decompositions. In Gloor [Glo98], pages 295–

301. ISBN 1-58113-002-3. LCCN QA155.7.E4 E88 1998. URL

http://www.acm.org/pubs/citations/proceedings/issac/
281508/p295-brown/.

Brown:1999:GSF

[Bro99] Christopher W. Brown. Guaranteed solution formula construction. In Dooley [Doo99], pages 137–144. ISBN 1-58113-073-2.

LCCN QA76.95 .I57 1999. URL http://www.acm.org/pubs/
contents/proceedings/issac/309831/.

Brown:2000:IPC

[Bro00] Christopher W. Brown. Improved projection for CAD’s of R^{3}. In Traverso [Tra00], pages 48–53. ISBN 1-58113-218-2. LCCN QA76.95.I59 2000. URL http:/

/www.acm.org/pubs/articles/proceedings/issac/345542/
p48-brown/p48-brown.pdf;   http://www.acm.org/pubs/
citations/proceedings/issac/345542/p48-brown/. ACM

order number 505000.

Brown:2007:PIS

[Bro07] C. W. Brown, editor. Proceedings of the 2007 International Symposium on Symbolic and Algebraic Computation, July 29– August 1, 2007, University of Waterloo, Waterloo, Ontario, Canada. ACM Press, New York, NY 10036, USA, 2007. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Brown:2009:FST

[Bro09] Christopher W. Brown. Fast simplifications for Tarski formulas. In May [May09], pages 63–70. ISBN 1-60558-609-9. LCCN ????

Brown:2013:CSO

[Bro13] Christopher W. Brown. Constructing a single open cell in a cylindrical algebraic decomposition. In Monagan et al. [MCG13], pages 133–140. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

Bihan:2009:FRF

[BRS09] Frédéric Bihan, J. Maurice Rojas, and Casey E. Stella. Faster real feasibility via circuit discriminants. In May [May09], pages 39–46. ISBN 1-60558-609-9. LCCN ????

Bayer:1986:DMS

[BS86] D. Bayer and M. Stillman. The design of Macaulay: a system for computing in algebraic geometry and commutative algebra. In Char [Cha86b], pages 157–162. ISBN 0-89791-199-7 (paperback). LCCN QA155.7.E4 A281 1986. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/32439/p157-

bayer/. ACM order no. 505860.

Bronstein:1993:FPF

[BS93] Manuel Bronstein and Bruno Salvy. Full partial fraction decomposition of rational functions. In Bronstein [Bro93], pages 157–160. ISBN 0-89791-604-2. LCCN QA 76.95

I59 1993. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/164081/p157-bronstein/. ACM order

number: 505930.

Bachmann:1998:MRG

[BS98] Olaf Bachmann and Hans Schönemann. Monomial representations for Gröbner bases computations. In Gloor [Glo98], pages 309–316. ISBN 1-58113-002-3. LCCN QA155.7.E4 E88 1998.

URL http://www.acm.org/pubs/citations/proceedings/
issac/281508/p309-bachmann/.

Bodnar:2000:IAR

[BS00] Gábor Bodnár and Josef Schicho. An improved algorithm for the resolution of singularities. In Traverso [Tra00], pages 30– 37. ISBN 1-58113-218-2. LCCN QA76.95.I59 2000. URL

http://www.acm.org/pubs/articles/proceedings/issac/
345542/p30-bodnar/p30-bodnar.pdf;   http://www.acm.
org/pubs/citations/proceedings/issac/345542/p30-bodnar/

. ACM order number 505000.

Burgisser:2007:DFC

[BS07] Peter Bürgisser and Peter Scheiblechner. Differential forms in computational algebraic geometry. In Brown [Bro07], pages 61–68. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Benoit:2009:CES

[BS09a] Alexandre Benoit and Bruno Salvy. Chebyshev expansions for solutions of linear differential equations. In May [May09], pages 23–30. ISBN 1-60558-609-9. LCCN ????

Bostan:2009:FAD

[BS09b] Alin Bostan and Éric Schost. Fast algorithms for differential equations in positive characteristic. In May [May09], pages 47– 54. ISBN 1-60558-609-9. LCCN ????

Brown:2010:BBW

[BS10] Christopher W. Brown and Adam Strzeboński. Black-box/ white-box simplification and applications to quantifier elimination. In Watt [Wat10], pages 69–76. ISBN 1-4503-0150-9. LCCN QA76.95 .I59 2010.

Borwein:2011:SVG

[BS11a] Jonathan M. Borwein and Armin Straub. Special values of generalized log-sine integrals. In Schost and Emiris [SE11], pages 43–50. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Bright:2011:VRN

[BS11b] Curtis Bright and Arne Storjohann. Vector rational number reconstruction. In Schost and Emiris [SE11], pages 51–58. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Bach:2013:ACS

[BS13] Eric Bach and Jonathan P. Sorenson. Approximately counting semismooth integers. In Monagan et al. [MCG13], pages 23–30. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

Bostan:2012:PSS

[BSC^{+}12] Alin Bostan, Bruno Salvy, Muhammad F. I. Chowdhury, Éric Schost, and Romain Lebreton. Power series solutions of singular (q)-differential equations. In van der Hoeven and van Hoeij [vdHvH12], pages 107–114. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Bachmann:1996:MFD

[BSG96] Olaf Bachmann, Hans Schönemann, and Simon Gray. MPP: A framework for distributed polynomial computations. In Lakshman Y. N. [Lak96], pages 103–112. ISBN 0-89791-796-0. LCCN

QA 76.95 I59 1996. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/236869/p103-bachmann/.

Bachmann:1997:PSD

[BSG97] Olaf Bachmann, Hans Schönemann, and Simon Gray. A proposal for syntactic data integration math protocols. In Küchlin [Küc97], pages 165–175. ISBN 0-89791-875-4. LCCN

QA76.95. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/266670/p165-bachmann/.

Bostan:2008:PSC

[BSS08] Alin Bostan, Bruno Salvy, and Éric Schost. Power series composition and change of basis. In Jeffrey [Jef08], pages 269–276. ISBN 1-59593-904-0. LCCN ????

Buchberger:2005:VFS

[Buc05] Bruno Buchberger. A view on the future of symbolic computation. In Kauers [Kau05], page 1. ISBN 1-59593-095-7. LCCN ???? ACM Order Number 505050.

Bundgen:1991:CIP

[Bün91] Reinhard Bündgen. Completion of integral polynomials by ACterm completion. In Watt [Wat91], pages 70–78. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991. URL http://www.

acm.org:80/pubs/citations/proceedings/issac/120694/
p70-bundgen/.

Burge:1991:SRI

[Bur91] William H. Burge. Scratchpad and the Rogers–Ramanujan identities. In Watt [Wat91], pages 189–190. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/120694/p189-burge/.

Burgisser:2011:PAC

[Bür11a] Peter Bürgisser. Probabilistic analysis of condition numbers. In Schost and Emiris [SE11], pages 5–6. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Burton:2011:DGV

[Bur11b] Benjamin A. Burton. Detecting genus in vertex links for the fast enumeration of 3-manifold triangulations. In Schost and Emiris [SE11], pages 59–66. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Buse:2001:RRP

[Bus01] Laurent Busé. Residual resultant over the projective plane and the implicitization problem. In Mourrain [Mou01], pages 48– 55. ISBN 1-58113-417-7. LCCN QA76.95.I59 2001. ACM order number 505010.

Butler:1986:DCC

[But86] Greg Butler. Divide-and-conquer in computational group theory. In Char [Cha86b], pages 59–64. ISBN 0-89791-199-7 (paperback). LCCN QA155.7.E4 A281 1986. URL http://www.

acm.org:80/pubs/citations/proceedings/issac/32439/p59-

butler/. ACM order no. 505860.

Butler:1992:ECA

[But92] Greg Butler. Experimental comparison of algorithms for Sylow subgroups. In Wang [Wan92b], pages 251–262. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN QA76.95.I59 1992. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/143242/p251-butler/

. ACM order number: 505920.

Blankertz:2012:CCD

[BvzGZ12] Raoul Blankertz, Joachim von zur Gathen, and Konstantin Ziegler. Compositions and collisions at degree p^{2}. In van der Hoeven and van Hoeij [vdHvH12], pages 91–98. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Becker:1991:CRP

[BW91] Thomas Becker and Volker Weispfenning. The Chinese remainder problem, multivariate interpolation, and Gröbner bases. In Watt [Wat91], pages 64–69. ISBN 0-89791-437-6. LCCN

QA 76.95 I59 1991. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/120694/p64-becker/.

Buchberger:1998:YGB

[BW98] Bruno Buchberger and Franz Winkler, editors. 33 years of Gröbner bases: Gröbner bases and applications: Conference — February 1998, Linz, Austria, number 251 in London Mathematical Society Lecture Note Series. Cambridge University Press, Cambridge, UK, 1998. ISBN 0-521-63298-6. LCCN QA251.3.G76 1998.

Bachmann:1994:CRM

[BWZ94] Olaf Bachmann, Paul S. Wang, and Eugene V. Zima. Chains of recurrences — a method to expedite the evaluation of closedform functions. In ACM [ACM94], pages 242–249. ISBN 0-89791-638-7. LCCN QA76.95.I59 1994. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/190347/p242-
bachmann/.

Brownawell:2009:LBZ

[BY09] W. Dale Brownawell and Chee K. Yap. Lower bounds for zerodimensional projections. In May [May09], pages 79–86. ISBN 1-60558-609-9. LCCN ????

Bodrato:2007:IPM

[BZ07] Marco Bodrato and Alberto Zanoni. Integer and polynomial multiplication: towards optimal Toom–Cook matrices. In Brown [Bro07], pages 17–24. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Corless:2007:JHF

[CA07] Robert M. Corless and Dawit Assefa. Jeffery–Hamel flow with Maple: a case study of integration of elliptic functions in a CAS. In Brown [Bro07], pages 108–115. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Caboara:1993:DAG

[Cab93] Massimo Caboara. A dynamic algorithm for Gröbner basis computation. In Bronstein [Bro93], pages 275–283. ISBN 0-89791-604-2. LCCN QA 76.95 I59 1993. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/164081/p275-

caboara/. ACM order number: 505930.

Canny:1989:GCP

[Can89] J. Canny. Generalized characteristic polynomials. In Gianni [Gia89], pages 293–299. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Carette:2004:UES

[Car04] Jacques Carette. Understanding expression simplification. In Gutierrez [Gut04], pages 72–79. ISBN 1-58113-827-X. LCCN QA76.95 .I57 2004.

Carette:2007:CFP

[Car07] Jacques Carette. A canonical form for piecewise defined functions. In Brown [Bro07], pages 77–84. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Costa:1997:EPL

[CBdCD97] Vítor Santos Costa, Ricardo Bianchini, and Inês de Castro Dutra. Evaluating parallel logic programming systems on scalable multiprocessors. In Küchlin [Küc97], pages 58–67. ISBN 0-89791-875-4. LCCN QA76.95. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/266670/p58-santos_
costa/.

ChaffyCamus:1989:ARA

[CC89] C. Chaffy-Camus. An application of REDUCE to the approximation of f(x, y). In Gianni [Gia89], pages 73–84. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Cantone:1990:DFE

[CC90] D. Cantone and V. Cutello. A decidable fragment of the elementary theory of relations and some applications. In Watanabe and Nagata [WN90], pages 24–29. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95

.I57 1990. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/96877/p24-cantone/.

Cantone:1993:DPS

[CC93] Domenico Cantone and Vincenzo Cutello. Decision procedures for stratified set-theoretic syllogistics. In Bronstein [Bro93], pages 105–110. ISBN 0-89791-604-2. LCCN QA 76.95

I59 1993. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/164081/p105-cantone/. ACM order

number: 505930.

Chin:1998:OSA

[CCC98] Paulina Chin, Robert M. Corless, and George F. Corliss. Optimization strategies for the approximate GCD problem. In Gloor [Glo98], pages 228–235. ISBN 1-58113-002-3. LCCN

QA155.7.E4 E88 1998. URL http://www.acm.org/pubs/
citations/proceedings/issac/281508/p228-chin/.

Cantone:1989:DPE

[CCF89] D. Cantone, V. Cutello, and A. Ferro. Decision procedures for elementary sublanguages of set theory. XIV. Three languages involving rank related constructs. In Gianni [Gia89], pages 407– 422. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Carvajal:2005:HSN

[CCG05] Orlando A. Carvajal, Frederick W. Chapman, and Keith O. Geddes. Hybrid symbolic-numeric integration in multiple dimensions via tensor-product series. In Kauers [Kau05], pages 84–91. ISBN 1-59593-095-7. LCCN ???? ACM Order Number 505050.

Chan:2004:NLS

[CCT04] L. Chan and E. S. Cheb-Terrab. Non-Liouvillian solutions for second order linear ODEs. In Gutierrez [Gut04], pages 80–86. ISBN 1-58113-827-X. LCCN QA76.95 .I57 2004.

Caboara:2008:GBP

[CCT08] Massimo Caboara, Fabrizio Caruso, and Carlo Traverso. Gröbner bases for public key cryptography. In Jeffrey [Jef08], pages 315–324. ISBN 1-59593-904-0. LCCN ????

Chen:1999:RNF

[CD99] Guoting Chen and Jean Della Dora. Rational normal form for dynamical systems by Carleman linearization. In Dooley [Doo99], pages 165–172. ISBN 1-58113-073-2. LCCN QA76.95

.I57 1999.  URL http://www.acm.org/pubs/contents/
proceedings/issac/309831/.

Cabarcas:2011:LAC

[CD11] Daniel Cabarcas and Jintai Ding. Linear algebra to compute syzygies and Gröbner bases. In Schost and Emiris [SE11], pages 67–74. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Cicalo:2007:NAG

[CdG07] Serena Cicalò and Willem de Graaf. Non-associative Gröbner bases, finitely-presented Lie rings and the Engel condition. In Brown [Bro07], pages 100–107. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Corless:2007:SEA

[CDHJ07] Robert M. Corless, Hui Ding, Nicholas J. Higham, and David J. Jeffrey. The solution of s exp(s) = a is not always the Lambert W function of a. In Brown [Bro07], pages 116–121. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Chen:2010:TDS

[CDM^{+}10] Changbo Chen, James H. Davenport, John P. May, Marc Moreno Maza, Bican Xia, and Rong Xiao. Triangular decomposition of semi-algebraic systems. In Watt [Wat10], pages 187–194. ISBN 1-4503-0150-9. LCCN QA76.95 .I59 2010.

Chen:2011:CSA

[CDM^{+}11] Changbo Chen, James H. Davenport, Marc Moreno Maza, Bican Xia, and Rong Xiao. Computing with semi-algebraic sets represented by triangular decomposition. In Schost and Emiris [SE11], pages 75–82. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Caboara:1996:MHF

[CdR96] Massimo Caboara, Gabriel de Dominicis, and Lorenzo Robbiano. Multigraded Hilbert functions and Buchberger algorithm. In Lakshman Y. N. [Lak96], pages 72–78. ISBN 0-89791-796-0. LCCN QA 76.95 I59 1996. URL http://www.

acm.org:80/pubs/citations/proceedings/issac/236869/
p72-caboara/.

Caprasse:1989:CEB

[CDS89] H. Caprasse, J. Demaret, and E. Schrufer. Can EXCALC be used to investigate high-dimensional cosmological models with nonlinear Lagrangians? In Gianni [Gia89], pages 116–124. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Chen:1991:NNF

[CDS91] Guoting Chen, Jean Della Dora, and Laurent Stolovitch. Nilpotent normal form via Carleman linearization (for systems of ordinary differential equations). In Watt [Wat91], pages 281–288. ISBN 0-89791-437-6. LCCN QA 76.95

I59 1991. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/120694/p281-chen/.

Costermans:2005:SAE

[CEMP05] C. Costermans, J. Y. Enjalbert, Hoang Ngoc Minh, and M. Petitot. Structure and asymptotic expansion of multiple harmonic sums. In Kauers [Kau05], pages 100–107. ISBN 1-59593-095-7. LCCN ???? ACM Order Number 505050.

Cesari:1997:CCA

[Ces97] Giovanni Cesari. CALYPSO: a computer algebra library for parallel symbolic computation. In Küchlin [Küc97], pages 204–216. ISBN 0-89791-875-4. LCCN QA76.95. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/266670/p204-
cesari/.

Cetinkaya:1992:SAL

[Cet92] Cetin Cetinkaya. On stability analysis of linear stochastic and time-varying deterministic systems. In Wang [Wan92b], pages 278–283. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN QA76.95.I59 1992. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/143242/p278-

cetinkaya/. ACM order number: 505920.

Cooperman:1992:FCB

[CF92] Gene Cooperman and Larry Finkelstein. A fast cyclic base change for permutation groups. In Wang [Wan92b], pages 224– 232. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN QA76.95.I59 1992. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/143242/p224-

cooperman/. ACM order number: 505920.

Chen:2011:SCR

[CFFL11] Shaoshi Chen, Ruyong Feng, Guofeng Fu, and Ziming Li. On the structure of compatible rational functions. In Schost and Emiris [SE11], pages 91–98. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Cooperman:1989:RGC

[CFL89] G. Cooperman, L. Finkelstein, and E. Luks. Reduction of group constructions to point stabilizers. In Gonnet [Gon89], pages 351–356. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989.

US$29.00. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/74540/p351-cooperman/. ACM order

number: 505890. English and French.

Cooperman:1990:RBC

[CFS90] G. Cooperman, L. Finkelstein, and N. Sarawagi. A random base change algorithm for permutation groups. In Watanabe and Nagata [WN90], pages 161–168. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95

.I57 1990. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/96877/p161-cooperman/.

Cooperman:1995:CMG

[CFT95] Gene Cooperman, Larry Finkelstein, and Michael Tselman. Computing with matrix groups using permutation representations. In Levelt [Lev95], pages 259–264. ISBN 0-89791-699-9. LCCN QA 76.95 I59 1995. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/220346/p259-

cooperman/. ACM order number: 505950.

Cooperman:1994:CPR

[CFYT94] Gene Cooperman, Larry Finkelstein, Bryant York, and Michael Tselman. Constructing permutation representations for large matrix groups. In ACM [ACM94], pages 134–138. ISBN 0-89791-638-7. LCCN QA76.95.I59 1994. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/190347/p134-
cooperman/.

Czapor:1986:IBA

[CG86] S. R. Czapor and K. O. Geddes. On implementing Buchberger’s algorithm for Gröbner bases. In Char [Cha86b], pages 233– 238. ISBN 0-89791-199-7 (paperback). LCCN QA155.7.E4

A281 1986. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/32439/p233-czapor/. ACM order no.

505860.

Chen:1990:IAM

[CG90a] G. Chen and I. Gil. The implementation of an algorithm in Macsyma: computing the formal solutions of differential systems in the neighborhood of regular singular point. In Watanabe and Nagata [WN90], page 307. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95

.I57 1990. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/96877/p307-chen/.

Chou:1990:MMG

[CG90b] Shang-Ching Chou and Xiao-Shan Gao. Methods for mechanical geometry formula deriving. In Watanabe and Nagata [WN90], pages 265–270. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95 .I57 1990. URL

http://www.acm.org:80/pubs/citations/proceedings/issac/
96877/p265-chou/.

Crouch:1992:ECI

[CG92] P. E. Crouch and R. L. Grossman. The explicit computation of integration algorithms and first integrals for ordinary differential equations with polynomial coefficients using trees. In Wang [Wan92b], pages 89–94. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN

QA76.95.I59 1992. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/143242/p89-crouch/. ACM

order number: 505920.

Cooperman:2001:SPC

[CG01] Gene Cooperman and Victor Grinberg. Scalable parallel coset enumeration using bulk definition. In Mourrain [Mou01], pages 77–84. ISBN 1-58113-417-7. LCCN QA76.95.I59 2001. ACM order number 505010.

Char:1986:CAU

[CGG^{+}86] B. W. Char, K. O. Geddes, G. H. Gonnet, B. J. Marshman, and P. J. Ponzo. Computer algebra in the undergraduate mathematics classroom. In Char [Cha86b], pages 135–140. ISBN 0-89791-199-7 (paperback). LCCN

QA155.7.E4 A281 1986. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/32439/p135-char/. ACM or-

der no. 505860.

Corless:1999:APD

[CGJ99] Robert M. Corless, Mark W. Giesbrecht, and David J. Jeffrey. Approximate polynomial decomposition. In Dooley [Doo99], pages 213–219. ISBN 1-58113-073-2. LCCN QA76.95

.I57 1999.  URL http://www.acm.org/pubs/contents/
proceedings/issac/309831/.

Chattopadhyay:2013:FBL

[CGK^{+}13] Arkadev Chattopadhyay, Bruno Grenet, Pascal Koiran, Natacha Portier, and Yann Strozecki. Factoring bivariate lacunary polynomials without heights. In Monagan et al. [MCG13], pages 141–148. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

Corless:2002:GNA

[CGKW02] Robert M. Corless, André Galligo, Ilias S. Kotsireas, and Stephen M. Watt. A geometric-numeric algorithm for absolute factorization of multivariate polynomials. In Mora [Mor02], pages 37–45. ISBN 1-58113-484-3. LCCN QA76.95. URL http:

//www.lifl.fr/ISSAC2002/.

Crouch:1991:CID

[CGL91] Peter Crouch, Robert Grossman, and Richard Larson. Computations involving differential operators and their actions on functions. In Watt [Wat91], pages 301–307. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/120694/p301-crouch/

.

Cheng:2009:RIB

[CGL09] Jin-San Cheng, Xiao-Shan Gao, and Jia Li. Root isolation for bivariate polynomial systems with local generic position method. In May [May09], pages 103–110. ISBN 1-60558-609-9. LCCN ????

Conti:2010:SBL

[CGR10] C. Conti, L. Gemignani, and L. Romani. Solving Bezout-like polynomial equations for the design of interpolatory subdivision schemes. In Watt [Wat10], pages 251–256. ISBN 1-4503-0150-9. LCCN QA76.95 .I59 2010.

Corless:1997:RSF

[CGT97] Robert M. Corless, Patrizia M. Gianni, and Barry M. Trager. A reordered Schur factorization method for zerodimensional polynomial systems with multiple roots. In Küchlin [Küc97], pages 133–140. ISBN 0-89791-875-4. LCCN

QA76.95. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/258726/p133-corless/.

Corless:1995:SVD

[CGTW95] Robert M. Corless, Patrizia M. Gianni, Barry M. Trager, and Stephen M. Watt. The singular value decomposition for polynomial systems. In Levelt [Lev95], pages 195–207. ISBN 0-89791-699-9. LCCN QA 76.95 I59 1995. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/220346/p195-

corless/. ACM order number: 505950.

Corless:2001:TFB

[CGvH^{+}01] Robert M. Corless, Mark W. Giesbrecht, Mark van Hoeij, Ilias S. Kotsireas, and Stephen M. Watt. Towards factoring bivariate approximate polynomials. In Mourrain [Mou01], pages 85–92. ISBN 1-58113-417-7. LCCN QA76.95.I59 2001. ACM order number 505010.

Cheng:2007:CNI

[CGY07] Jin-San Cheng, Xiao-Shan Gao, and Chee-Keng Yap. Complete numerical isolation of real zeros in zero-dimensional triangular systems. In Brown [Bro07], pages 92–99. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Chou:1993:AGT

[CGZ93] Shang-Ching Chou, Xiao-Shan Gao, and Jing-Zhong Zhang. Automated geometry theorem proving by vector calculation. In Bronstein [Bro93], pages 284–291. ISBN 0-89791-604-2. LCCN

QA 76.95 I59 1993. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/164081/p284-chou/. ACM

order number: 505930.

Chaffy:1986:HCM

[Cha86a] C. Chaffy. How to compute multivariate Padé approximants. In Char [Cha86b], pages 56–58. ISBN 0-89791-199-7 (paperback). LCCN QA155.7.E4 A281 1986. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/32439/p56-chaffy/. ACM

order no. 505860.

Char:1986:PSS

[Cha86b] Bruce W. Char, editor. Proceedings of the 1986 Symposium on Symbolic and Algebraic Computation: Symsac ’86, July 21–23, 1986, Waterloo, Ontario. ACM Press, New York, NY 10036, USA, 1986. ISBN 0-89791-199-7 (paperback). LCCN QA155.7.E4 A281 1986. ACM order no. 505860.

Char:1989:ARA

[Cha89] B. W. Char. Automatic reasoning about numerical stability of rational expressions. In Gonnet [Gon89], pages 234–241. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989.

US$29.00. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/74540/p234-char/. ACM order number:

505890. English and French.

Char:1990:PRS

[Cha90] B. W. Char. Progress report on a system for generalpurpose parallel symbolic algebraic computation. In Watanabe and Nagata [WN90], pages 96–103. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95

.I57 1990. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/96877/p96-char/.

Chaffy:1995:ACP

[Cha95] Claudine Chaffy. The analytic continuation process: From computer algebra to numerical analysis. In Levelt [Lev95], pages 216–222. ISBN 0-89791-699-9. LCCN QA 76.95

I59 1995. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/220346/p216-chaffy/. ACM order num-

ber: 505950.

Chapman:2003:EAA

[Cha03] Frederick W. Chapman. An elementary algorithm for the automatic derivation and proof of tensor product identities via computer algebra. In Senda [Sen03], pages 50–57. ISBN 1-58113-641-2. LCCN QA76.95. ACM order number 505030.

Chen:1989:CNF

[Che89] Guoting Chen. Computing the normal forms of matrices depending on parameters. In Gonnet [Gon89], pages 242–249. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989.

US$29.00. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/74540/p242-chen/. ACM order number:

505890. English and French.

Chen:1990:ACF

[Che90a] Guoting Chen. An algorithm for computing the formal solutions of differential systems in the neighborhood of an irregular singular point. In Watanabe and Nagata [WN90], pages 231–235. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95 .I57 1990. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/96877/p231-chen/.

Cherief:1990:AMP

[Che90b] F. Cherief. An algebraic model for the parallel interpretation of equationally defined functions (abstract). In Watanabe and Nagata [WN90], page 285. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95

.I57 1990. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/96877/p285-cherief/.

Cheze:2004:APF

[Chè04] Guillaume Chèze. Absolute polynomial factorization in two variables and the knapsack problem. In Gutierrez [Gut04], pages 87–94. ISBN 1-58113-827-X. LCCN QA76.95 .I57 2004.

Chou:1990:ARG

[Cho90] Shang-Ching Chou. Automated reasoning in geometries using the characteristic set method and Gröbner basis method. In Watanabe and Nagata [WN90], pages 255–260. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN

QA76.95 .I57 1990. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/96877/p255-chou/.

Cheng:2007:TSE

[CHT^{+}07] Howard Cheng, Guillaume Hanrot, Emmanuel Thomé, Paul Zimmermann, and Eugene Zima. Time-and space-efficient evaluation of some hypergeometric constants. In Brown [Bro07], pages 85–91. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Chistov:1997:PTA

[CIK97] Alexander Chistov, Gábor Ivanyos, and Marek Karpinski. Polynomial time algorithms for modules over finite dimensional algebras. In Küchlin [Küc97], pages 68–74. ISBN 0-89791-

875-4. LCCN QA76.95. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/258726/p68-chistov/.

Collins:1989:PRP

[CJ89a] G. E. Collins and J. R. Johnson. The probability of relative primality of Gaussian integers. In Gianni [Gia89], pages 252–258. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Collins:1989:QES

[CJ89b] G. E. Collins and J. R. Johnson. Quantifier elimination and the sign variation method for real root isolation. In Gonnet [Gon89], pages 264–271. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989.

US$29.00. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/74540/p264-collins/. ACM order num-

ber: 505890. English and French.

Corless:1997:SSL

[CJK97] Robert M. Corless, David J. Jeffrey, and Donald E. Knuth. A sequence of series for the Lambert W function. In Küchlin [Küc97], pages 197–204. ISBN 0-89791-875-4. LCCN

QA76.95. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/258726/p197-corless/.

Chen:2013:DEO

[CJKS13] Shaoshi Chen, Maximilian Jaroschek, Manuel Kauers, and Michael F. Singer. Desingularization explains order-degree curves for ore operators. In Monagan et al. [MCG13], pages 157–164. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

Char:1994:SEP

[CJSW94] B. Char, J. Johnson, D. Saunders, and A. P. Wack. Some experiments with parallel bignum arithmetic. In Hong [Hon94], pages 94–103. ISBN 981-02-2040-5. LCCN QA76.642.I58 1994.

Collins:1992:EAI

[CK92] George E. Collins and Werner Krandick. An efficient algorithm for infallible polynomial complex root isolation. In Wang [Wan92b], pages 189–194. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN QA76.95.I59 1992. URL http:

//www.acm.org:80/pubs/citations/proceedings/issac/143242/

p189-collins/. ACM order number: 505920.

Collins:1993:HMH

[CK93] George E. Collins and Werner Krandick. A hybrid method for high precision calculation of polynomial real roots. In Bronstein [Bro93], pages 47–52. ISBN 0-89791-604-2. LCCN QA 76.95

I59 1993. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/164081/p47-collins/. ACM order num-

ber: 505930.

Collins:1996:TMP

[CK96] George E. Collins and Werner Krandick. A tangent-secant method for polynomial complex root calculation. In Lakshman Y. N. [Lak96], pages 137–141. ISBN 0-89791-796-0. LCCN

QA 76.95 I59 1996. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/236869/p137-collins/.

Chtcherba:2000:CER

[CK00a] Arthur D. Chtcherba and Deepak Kapur. Conditions for exact resultants using the Dixon formulation. In Traverso [Tra00], pages 62–70. ISBN 1-58113-218-2. LCCN QA76.95.I59 2000.

URL http://www.acm.org/pubs/articles/proceedings/issac/
345542/p62-chtcherba/p62-chtcherba.pdf;   http://www.
acm.org/pubs/citations/proceedings/issac/345542/p62-

chtcherba/. ACM order number 505000.

Collins:2000:MFP

[CK00b] George E. Collins and Werner Krandick. Multiprecision floating point addition. In Traverso [Tra00], pages 71–77. ISBN 1-58113-218-2. LCCN QA76.95.I59 2000. URL http://www.

acm.org/pubs/articles/proceedings/issac/345542/p71-
collins/p71-collins.pdf;   http://www.acm.org/pubs/
citations/proceedings/issac/345542/p71-collins/. ACM

order number 505000.

Chtcherba:2002:EOD

[CK02] Arthur D. Chtcherba and Deepak Kapur. On the efficiency and optimality of Dixon-based resultant methods. In Mora [Mor02], pages 29–36. ISBN 1-58113-484-3. LCCN QA76.95. URL http:

//www.lifl.fr/ISSAC2002/.

Chtcherba:2004:SHR

[CK04] Arthur D. Chtcherba and Deepak Kapur. Support hull: relating the Cayley–Dixon resultant constructions to the support of a polynomial system. In Gutierrez [Gut04], pages 95–102. ISBN 1-58113-827-X. LCCN QA76.95 .I57 2004.

Chtcherba:2006:CDF

[CK06] Arthur D. Chtcherba and Deepak Kapur. Conditions for determinantal formula for resultant of a polynomial system. In Trager [Tra06], pages 55–62. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

Chen:2012:ODC

[CK12] Shaoshi Chen and Manuel Kauers. Order-degree curves for hypergeometric creative telescoping. In van der Hoeven and van Hoeij [vdHvH12], pages 122–129. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Comer:2012:SPI

[CKP12] Matthew T. Comer, Erich L. Kaltofen, and Clément Pernet. Sparse polynomial interpolation and Berlekamp/Massey algorithms that correct outlier errors in input values. In van der Hoeven and van Hoeij [vdHvH12], pages 138–145. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Chyzak:2009:NHS

[CKS09] Frédéric Chyzak, Manuel Kauers, and Bruno Salvy. A nonholonomic systems approach to special function identities. In May [May09], pages 111–118. ISBN 1-60558-609-9. LCCN ????

Chen:2012:TRA

[CKS12] Shaoshi Chen, Manuel Kauers, and Michael F. Singer. Telescopers for rational and algebraic functions via residues. In van der Hoeven and van Hoeij [vdHvH12], pages 130–137. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Canny:1989:SSN

[CKY89] J. F. Canny, E. Kaltofen, and L. Yagati. Solving systems of non-linear polynomial equations faster. In Gonnet [Gon89], pages 121–128. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989.

US$29.00. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/74540/p121-canny/. ACM order num-

ber: 505890. English and French.

Cabay:1989:FRA

[CL89] S. Cabay and G. Labahn. A fast, reliable algorithm for calculating Padé–Hermite forms. In Gonnet [Gon89], pages 95–100. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989.

US$29.00. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/74540/p95-cabay/. ACM order number:

505890. English and French.

Calmet:1996:DIS

[CL96] Jacques Calmet and Carla Limongelli, editors. Design and implementation of symbolic computation systems: International Symposium, DISCO ’96, Karlsruhe, Germany, September 18– 20, 1996: proceedings, volume 1128 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-61697-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.S88I576 1996.

Cheng:2001:CAF

[CL01] Howard Cheng and George Labahn. Computing all factorizations in Z_{N}[x]. In Mourrain [Mou01], pages 64–71. ISBN 1-58113-417-7. LCCN QA76.95.I59 2001. ACM order number 505010.

Cheng:2006:CPG

[CL06] Howard Cheng and George Labahn. On computing polynomial GCDs in alternate bases. In Trager [Tra06], pages 47–54. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

Cluzeau:2003:FDS

[Clu03] Thomas Cluzeau. Factorization of differential systems in characteristic p. In Senda [Sen03], pages 58–65. ISBN 1-58113-641-2. LCCN QA76.95. ACM order number 505030.

Covic:1995:SFC

[CM95] V. Covic and S. Markovic. Symbolic form computation of the complete dynamics of robotic systems with the closed chains structure. In Aityan [Ait95], pages 125–128. ISBN 0-9640398-9-3, 0-9640398-8-5. LCCN QA76.87 .I58 1995.

Cesari:1996:PFA

[CM96a] G. Cesari and R. Maeder. Parallel 3-primes FFT algorithm. In Calmet and Limongelli [CL96], pages 174–182. ISBN 3-540-61697-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.S88I576 1996.

Clark:1996:DOO

[CM96b] K. L. Clark and F. G. McCabe. Distributed and object oriented symbolic programming in April. In Briot et al. [BGY96], pages 104–124. ISBN 3-540-61487-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.64 .F7 1995.

Chen:2011:ACT

[CM11] Changbo Chen and Marc Moreno Maza. Algorithms for computing triangular decompositions of polynomial systems. In Schost and Emiris [SE11], pages 83–90. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Char:1989:DIC

[CMS89] B. W. Char, A. R. Macnaughton, and P. A. Strooper. Discovering inequality conditions in the analytical solutions of optimization problems. In Gianni [Gia89], pages 109–115. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Chen:2009:CCA

[CMXY09] Changbo Chen, Marc Moreno Maza, Bican Xia, and Lu Yang. Computing cylindrical algebraic decomposition via triangular decomposition. In May [May09], pages 95–102. ISBN 1-60558-609-9. LCCN ????

Cremanns:1994:CCP

[CO94] Robert Cremanns and Friedrich Otto. Constructing canonical presentations for subgroups of context-free groups in polynomial time-extended abstract. In ACM [ACM94], pages 147–153. ISBN 0-89791-638-7. LCCN QA76.95.I59 1994. URL http://

www.acm.org:80/pubs/citations/proceedings/issac/190347/
p147-cremanns/.

Cioffi:2001:CMG

[CO01] Francesca Cioffi and Ferruccio Orecchia. Computation of minimal generators of ideals of fat points. In Mourrain [Mou01], pages 72–76. ISBN 1-58113-417-7. LCCN QA76.95.I59 2001. ACM order number 505010.

Codognet:1990:EDU

[Cod90] P. Codognet. Equations, disequations and unsolvable subsets (abstract). In Watanabe and Nagata [WN90], page 289. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN

QA76.95 .I57 1990. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/96877/p289-codognet/.

Codutti:1992:NNL

[Cod92] M. Codutti. NODES: non linear ordinary differential equations solver. In Wang [Wan92b], pages 69–79. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN

QA76.95.I59 1992. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/143242/p69-codutti/. ACM

order number: 505920.

Cohn:2013:SES

[Coh13] Henry Cohn. Solving equations with size constraints for the solutions. In Monagan et al. [MCG13], pages 1–2. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

Colin:1997:RRP

[Col97] Antoine Colin. Relative resolvents and partition tables in Galois group computations. In Küchlin [Küc97], pages 78–84. ISBN 0-89791-875-4. LCCN QA76.95. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/258726/p78-colin/.

Cooperman:1986:SMC

[Coo86] Gene Cooperman. A semantic matcher for computer algebra. In Char [Cha86b], pages 132–134. ISBN 0-89791-199-7 (paperback). LCCN QA155.7.E4 A281 1986. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/32439/p132-

cooperman/. ACM order no. 505860.

Cook:1992:CGA

[Coo92] Grant O. Cook, Jr. Code generation in ALPAL using symbolic techniques. In Wang [Wan92b], pages 27–35. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN

QA76.95.I59 1992. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/143242/p27-cook/. ACM or-

der number: 505920.

Cooperman:1995:SMB

[Coo95a] G. Cooperman. STAR/MPI: Binding a parallel library to interactive symbolic algebra systems. In Levelt [Lev95], pages 126–132. ISBN 0-89791-699-9. LCCN QA 76.95 I59 1995. ACM order number: 505950.

Cooperman:1995:SBP

[Coo95b] Gene Cooperman. STAR/MPI: binding a parallel library to interactive symbolic algebra systems. In Levelt [Lev95], pages 126–132. ISBN 0-89791-699-9. LCCN QA 76.95

I59 1995. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/220346/p126-cooperman/. ACM order

number: 505950.

Corless:1994:SAC

[Cor94] Robert M. Corless. Sufficiency analysis for the calculus of variations. In ACM [ACM94], pages 197–204. ISBN 0-89791-638-7. LCCN QA76.95.I59 1994. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/190347/p197-
corless/.

Cormier:2001:LSL

[Cor01] Olivier Cormier. On Liouvillian solutions of linear differential equations of order 4 and 5. In Mourrain [Mou01], pages 93–100. ISBN 1-58113-417-7. LCCN QA76.95.I59 2001. ACM order number 505010.

Cox:2007:GBS

[Cox07] David A. Cox. Gröbner bases: a sampler of recent developments. In Brown [Bro07], pages 387–388. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Canny:1991:OCD

[CR91] John Canny and J. Maurice Rojas. An optimal condition for determining the exact number of roots of a polynomial system. In Watt [Wat91], pages 96–102. ISBN 0-89791-437-6.

LCCN QA 76.95 I59 1991. URL http://www.acm.org:80/
pubs/citations/proceedings/issac/120694/p96-canny/.

Char:1994:AIT

[CR94] Bruce W. Char and Mark F. Russo. Automatic identification of time scales in enzyme kinetics models. In ACM [ACM94], pages 74–83. ISBN 0-89791-638-7. LCCN

QA76.95.I59 1994. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/190347/p74-char/.

Caboara:1997:FIS

[CR97] Massimo Caboara and Lorenzo Robbiano. Families of ideals in statistics. In Küchlin [Küc97], pages 404–410. ISBN 0-89791-

875-4. LCCN QA76.95. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/258726/p404-caboara/.

Caboara:2001:FET

[CR01] Massimo Caboara and Lorenzo Robbiano. Families of estimable terms. In Mourrain [Mou01], pages 56–63. ISBN 1-58113-417-7. LCCN QA76.95.I59 2001. ACM order number 505010.

Cooperman:2003:MBD

[CR03] Gene Cooperman and Eric Robinson. Memory-based and diskbased algorithms for very high degree permutation groups. In Senda [Sen03], pages 66–73. ISBN 1-58113-641-2. LCCN QA76.95. ACM order number 505030.

Cannon:1997:CCC

[CS97] John Cannon and Bernd Souvignier. On the computation of conjugacy classes in permutation groups. In Küchlin [Küc97], pages 392–399. ISBN 0-89791-875-4. LCCN

QA76.95. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/258726/p392-cannon/.

Chen:2005:BBC

[CS05] Zhuliang Chen and Arne Storjohann. A BLAS based C library for exact linear algebra on integer matrices. In Kauers [Kau05], pages 92–99. ISBN 1-59593-095-7. LCCN ???? ACM Order Number 505050.

Cormier:2000:CGG

[CSU00] Olivier Cormier, Michael F. Singer, and Felix Ulmer. Computing the Galois group of a polynomial using linear differential equations. In Traverso [Tra00], pages 78–85. ISBN 1-58113-218-2. LCCN QA76.95.I59 2000. URL http://www.

acm.org/pubs/articles/proceedings/issac/345542/p78-
cormier/p78-cormier.pdf;   http://www.acm.org/pubs/
citations/proceedings/issac/345542/p78-cormier/. ACM

order number 505000.

Chen:2013:NVH

[CSV13] Jingwei Chen, Damien Stehlé, and Gilles Villard. A new view on HJLS and PSLQ: sums and projections of lattices. In Monagan et al. [MCG13], pages 149–156. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

Cohen:1991:OES

[CT91] Ian Cohen and Karl-Erik E. Thylwe. Obtaining exact steadystate responses in driven undamped oscillators. In Watt [Wat91], pages 319–320. ISBN 0-89791-437-6. LCCN QA 76.95

I59 1991. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/120694/p319-cohen/.

Cooperman:1996:NSP

[CT96] Gene Cooperman and Michael Tselman. New sequential and parallel algorithms for generating high dimension Hecke algebras using the condensation technique. In Lakshman Y. N. [Lak96], pages 155–160. ISBN 0-89791-796-0. LCCN QA 76.95

I59 1996. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/236869/p155-cooperman/.

Caboara:1998:EAI

[CT98] Massimo Caboara and Carlo Traverso. Efficient algorithms for ideal operations (extended abstract). In Gloor [Glo98], pages 147–152. ISBN 1-58113-002-3. LCCN QA155.7.E4 E88 1998.

URL http://www.acm.org/pubs/citations/proceedings/
issac/281508/p147-caboara/.

Cucker:2000:SPS

[Cuc00] Felipe Cucker. Solving polynomial systems: a complexity theory viewpoint. In Traverso [Tra00], page ?? ISBN 1-58113-218-2.

LCCN QA76.95.I59 2000. URL http://www.acm.org/pubs/
citations/proceedings/issac/345542/p-cucker/. ACM or-

der number 505000.

Cha:2009:LSI

[CvH09] Yongjae Cha and Mark van Hoeij. Liouvillian solutions of irreducible linear difference equations. In May [May09], pages 87–94. ISBN 1-60558-609-9. LCCN ????

Cha:2010:SRR

[CvHL10] Yongjae Cha, Mark van Hoeij, and Giles Levy. Solving recurrence relations using local invariants. In Watt [Wat10], pages 303–309. ISBN 1-4503-0150-9. LCCN QA76.95 .I59 2010.

Cheng:2000:AME

[CZ00] Howard Cheng and Eugene Zima. On accelerated methods to evaluate sums of products of rational numbers. In Traverso [Tra00], pages 54–61. ISBN 1-58113-218-2.

LCCN QA76.95.I59 2000. URL http://www.acm.org/pubs/
articles/proceedings/issac/345542/p54-cheng/p54-cheng.
pdf;   http://www.acm.org/pubs/citations/proceedings/
issac/345542/p54-cheng/. ACM order number 505000.

Czapor:1991:HSS

[Cza91] S. R. Czapor. A heuristic selection strategy for lexicographic Gröbner bases? In Watt [Wat91], pages 39–48. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/120694/p39-
czapor/.

Dahan:2009:SCL

[Dah09] Xavier Dahan. Size of coefficients of lexicographical Gröbner bases: the zero-dimensional, radical and bivariate case. In May [May09], pages 119–126. ISBN 1-60558-609-9. LCCN ????

Dalmas:1992:PFL

[Dal92] Stéphane Dalmas. A polymorphic functional language applied to symbolic computation. In Wang [Wan92b], pages 369–375. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN QA76.95.I59 1992. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/143242/p369-dalmas/

. ACM order number: 505920.

Datta:2003:UCA

[Dat03] Ruchira S. Datta. Using computer algebra to find Nash equilibria. In Senda [Sen03], pages 74–79. ISBN 1-58113-641-2. LCCN QA76.95. ACM order number 505030.

Davenport:1992:PTR

[Dav92] J. H. Davenport. Primality testing revisited. In Wang [Wan92b], pages 123–129. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN QA76.95.I59 1992. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/143242/p123-

davenport/. ACM order number: 505920.

Deng:2005:CBR

[DCS05] Jiansong Deng, Falai Chen, and Liyong Shen. Computing μbases of rational curves and surfaces using polynomial matrix factorization. In Kauers [Kau05], pages 132–139. ISBN 1-59593-095-7. LCCN ???? ACM Order Number 505050.

Deprit:1989:MPS

[DD89a] A. Deprit and E. Deprit. Massively parallel symbolic computation. In Gonnet [Gon89], pages 308–316. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989. US$29.00. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/74540/p308-

deprit/. ACM order number: 505890. English and French.

Dicrescenzo:1989:AEA

[DD89b] C. Dicrescenzo and D. Duval. Algebraic extensions and algebraic closure in Scratchpad II. In Gianni [Gia89], pages 440–446. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

DeFeo:2013:FAA

[DDS13] Luca De Feo, Javad Doliskani, and Eric Schost. Fast algorithms for l-adic towers over finite fields. In Monagan et al. [MCG13], pages 165–172. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

DAndrea:2001:HSR

[DE01] Carlos D’Andrea and Ioannis Z. Emiris. Hybrid sparse resultant matrices for bivariate systems. In Mourrain [Mou01], pages 24– 31. ISBN 1-58113-417-7. LCCN QA76.95.I59 2001. ACM order number 505010.

Dickenstein:2002:MRM

[DE02] Alicia Dickenstein and Ioannis Z. Emiris. Multihomogeneous resultant matrices. In Mora [Mor02], pages 46–54. ISBN 1-58113-484-3. LCCN QA76.95. URL http://www.lifl.fr/

ISSAC2002/.

Dickenstein:2013:CDR

[DEF13] Alicia Dickenstein, Ioannis Z. Emiris, and Vissarion Fisikopoulos. Combinatorics of 4-dimensional resultant polytopes. In Monagan et al. [MCG13], pages 173–180. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

Diochnos:2007:CRS

[DET07] Dimitrios I. Diochnos, Ioannis Z. Emiris, and Elias P. Tsigaridas. On the complexity of real solving bivariate systems. In Brown [Bro07], pages 127–134. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Devitt:1989:UCA

[Dev89] J. S. Devitt. Unleashing computer algebra on the mathematics curriculum. In Gonnet [Gon89], pages 218–227. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989. US$29.00. ACM order number: 505890. English and French.

Dewar:1989:IIS

[Dew89] M. C. Dewar. IRENA: an integrated symbolic and numerical computation environment. In Gonnet [Gon89], pages 171–179. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989. US$29.00. ACM order number: 505890. English and French.

Dewar:1992:UCA

[Dew92] Michael C. Dewar. Using computer algebra to select numerical algorithms. In Wang [Wan92b], pages 1–8. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN

QA76.95.I59 1992. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/143242/p1-dewar/. ACM or-

der number: 505920.

Dingle:1994:BCC

[DF94] Adam Dingle and Richard J. Fateman. Branch cuts in computer algebra. In ACM [ACM94], pages 250–257. ISBN 0-89791-638-7. LCCN QA76.95.I59 1994. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/190347/p250-dingle/

.

Doffou:1995:SCD

[DG95] M. J. Doffou and R. L. Grossman. The symbolic computation of differential invariants of polynomial vector field systems using trees. In Levelt [Lev95], pages 26–31. ISBN 0-89791-699-9.

LCCN QA 76.95 I59 1995. URL http://www.acm.org:80/
pubs/citations/proceedings/issac/220346/p26-doffou/.

ACM order number: 505950.

Graaf:1997:CFM

[dG97] W. A. de Graaf. Constructing faithful matrix representations of Lie algebras. In Küchlin [Küc97], pages 54–59. ISBN 0-89791-

875-4. LCCN QA76.95. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/258726/p54-de_graaf/.

Dumas:2002:FFL

[DGP02] Jean Guillaume Dumas, Thierry Gautier, and Clément Pernet. Finite field linear algebra subroutines. In Mora [Mor02], pages 63–74. ISBN 1-58113-484-3. LCCN QA76.95. URL http://

www.lifl.fr/ISSAC2002/.

Dumas:2004:FFF

[DGP04] Jean-Guillaume Dumas, Pascal Giorgi, and Clément Pernet. FFPACK: finite field linear algebra package. In Gutierrez [Gut04], pages 119–126. ISBN 1-58113-827-X. LCCN QA76.95 .I57 2004.

Dalmas:1994:DCA

[DGS94] S. Dalmas, M. Gaetano, and A. Sausse. Distributed computer algebra: the central control approach. In Hong [Hon94], pages 104–113. ISBN 981-02-2040-5. LCCN QA76.642.I58 1994.

Dolzmann:1998:APQ

[DGS98] Andreas Dolzmann, Oliver Gloor, and Thomas Sturm. Approaches to parallel quantifier elimination. In Gloor [Glo98], pages 88–95. ISBN 1-58113-002-3. LCCN QA155.7.E4

E88 1998. URL http://www.acm.org/pubs/citations/
proceedings/issac/281508/p88-dolzmann/.

Davenport:1991:SVA

[DGT91] J. H. Davenport, P. Gianni, and B. M. Trager. Scratchpad’s view of algebra. II. A categorical view of factorization. In Watt [Wat91], pages 32–38. ISBN 0-89791-437-6. LCCN QA 76.95

I59 1991. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/120694/p32-davenport/.

Dalmas:1997:OI

[DGW97] Stéphane Dalmas, Marc Gaëtano, and Stephen Watt. An Open-Math 1.0 implementation. In Küchlin [Küc97], pages 241–248. ISBN 0-89791-875-4. LCCN QA76.95. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/258726/p241-
dalmas/.

deGraaf:1999:CBF

[dGW99] W. A. de Graaf and J. Wisliceny. Constructing bases of finitely presented Lie algebras using Gröbner bases in free algebras. In Dooley [Doo99], pages 37–43. ISBN 1-58113-073-2.

LCCN QA76.95 .I57 1999. URL http://www.acm.org/pubs/
contents/proceedings/issac/309831/.

Din:2008:CGO

[Din08] Mohab Safey El Din. Computing the global optimum of a multivariate polynomial over the reals. In Jeffrey [Jef08], pages 71–78. ISBN 1-59593-904-0. LCCN ????

deJager:1991:SCZ

[dJ91] Bram de Jager. Symbolic calculation of zero dynamics for nonlinear control systems. In Watt [Wat91], pages 321–322. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/120694/p321-
de_jager/.

DeBosschere:1994:LCB

[DJ94] K. De Bosschere and J.-M. Jacquet. Local and conditional blackboard operations in log: Semantics, applicability, and implementation. In Hong [Hon94], pages 34–43. ISBN 981-02-2040-5. LCCN QA76.642.I58 1994.

Dimitrova:2007:GFM

[DJLS07] Elena S. Dimitrova, Abdul Salam Jarrah, Reinhard Laubenbacher, and Brandilyn Stigler. A Gröbner fan method for biochemical network modeling. In Brown [Bro07], pages 122–126. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Diaz:1995:CGC

[DK95] Angel Díaz and Erich Kaltofen. On computing greatest common divisors with polynomials given by black boxes for their evaluations. In Levelt [Lev95], pages 232–239. ISBN 0-89791-699-9. LCCN QA 76.95 I59 1995. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/220346/p232-

diaz/. ACM order number: 505950.

Diaz:1998:FSM

[DK98] Angel Díaz and Erich Kaltofen. FOXBOX: a system from manipulating symbolic objects in black box representation. In Gloor [Glo98], pages 30–37. ISBN 1-58113-002-3. LCCN

QA155.7.E4 E88 1998. URL http://www.acm.org/pubs/
citations/proceedings/issac/281508/p30-diaz/.

Dunstan:1998:LFM

[DKLM98] Martin Dunstan, Tom Kelsey, Steve Linton, and Ursula Martin. Lightweight formal methods for computer algebra systems. In Gloor [Glo98], pages 80–87. ISBN 1-58113-002-3. LCCN

QA155.7.E4 E88 1998. URL http://www.acm.org/pubs/
citations/proceedings/issac/281508/p80-dunstan/.

deKleine:2005:ANM

[dKMW05] Jennifer de Kleine, Michael Monagan, and Allan Wittkopf. Algorithms for the non-monic case of the sparse modular GCD algorithm. In Kauers [Kau05], pages 124–131. ISBN 1-59593-095-7. LCCN ???? ACM Order Number 505050.

Diaz:1991:DSD

[DKS^{+}91] A. Diaz, E. Kaltofen, K. Schmitz, T. Valente, M. Hitz, A. Lobo, and P. Smyth. DSC: a system for distributed symbolic computation. In Watt [Wat91], pages 323–332. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/120694/p323-diaz/.

Dora:2001:HC

[DMMRY01] Jean Della Dora, Aude Maignan, Mihaela Mirica-Ruse, and Sergio Yovine. Hybrid computation. In Mourrain [Mou01], pages 101–108. ISBN 1-58113-417-7. LCCN QA76.95.I59 2001. ACM order number 505010.

DelPozo-Prieto:1994:ISP

[DMN94] A. Del Pozo-Prieto and J. J. Moreno-Navarro. Independent subexpressions parallelism with delayed synchronization for functional logic languages. In Hong [Hon94], pages 316–325. ISBN 981-02-2040-5. LCCN QA76.642.I58 1994.

Daouda:2008:CTN

[DMR08] Diatta Niang Daouda, Bernard Mourrain, and Olivier Ruatta. On the computation of the topology of a non-reduced implicit space curve. In Jeffrey [Jef08], pages 47–54. ISBN 1-59593-904-0. LCCN ????

Dahan:2005:LTT

[DMS^{+}05] Xavier Dahan, Marc Moreno Maza, Eric Schost, Wenyuan Wu, and Yuzhen Xie. Lifting techniques for triangular decompositions. In Kauers [Kau05], pages 108–115. ISBN 1-59593-095-7. LCCN ???? ACM Order Number 505050.

Dooley:1998:CMC

[Doo98] Samuel S. Dooley. Coordinating mathematical content and presentation markup in interactive mathematical documents. In Gloor [Glo98], pages 54–61. ISBN 1-58113-002-3. LCCN

QA155.7.E4 E88 1998. URL http://www.acm.org/pubs/
citations/proceedings/issac/281508/p54-dooley/.

Dooley:1999:IJS

[Doo99] Sam Dooley, editor. ISSAC 99: July 29–31, 1999, Simon Fraser University, Vancouver, BC, Canada: proceedings of the 1999 International Symposium on Symbolic and Algebraic Computation. ACM Press, New York, NY 10036, USA, 1999. ISBN 1-58113-073-2. LCCN QA76.95 .I57 1999.

Dooley:2002:EMC

[Doo02] Samuel S. Dooley. Editing mathematical content and presentation markup in interactive mathematical documents. In Mora [Mor02], pages 55–62. ISBN 1-58113-484-3. LCCN QA76.95.

URL http://www.lifl.fr/ISSAC2002/.

Doye:1999:ACA

[Doy99] Nicolas J. Doye. Automated coercion for Axiom. In Dooley [Doo99], pages 229–235. ISBN 1-58113-073-2. LCCN QA76.95

.I57 1999.  URL http://www.acm.org/pubs/contents/
proceedings/issac/309831/.

Daberkow:1995:CRE

[DP95] M. Daberkow and M. Pohst. Computations with relative extensions of number fields with an application to the construction of Hilbert class fields. In Levelt [Lev95], pages 68–76. ISBN 0-89791-699-9. LCCN QA 76.95 I59 1995. URL http://www.

acm.org:80/pubs/citations/proceedings/issac/220346/

p68-daberkow/. ACM order number: 505950.

Dridi:2007:TNO

[DP07] Raouf Dridi and Michel Petitot. Towards a new ODE solver based on Cartan’s equivalence method. In Brown [Bro07], pages 135–142. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Dumas:2009:FMC

[DPS09] Jean-Guillaume Dumas, Clément Pernet, and B. David Saunders. On finding multiplicities of characteristic polynomial factors of black-box matrices. In May [May09], pages 135–142. ISBN 1-60558-609-9. LCCN ????

Dumas:2013:SCR

[DPS13] Jean-Guillaume Dumas, Clément Pernet, and Ziad Sultan. Simultaneous computation of the row and column rank profiles. In Monagan et al. [MCG13], pages 181–188. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

Dumas:2005:ECC

[DPW05] Jean-Guillaume Dumas, Clément Pernet, and Zhendong Wan. Efficient computation of the characteristic polynomial. In Kauers [Kau05], pages 140–147. ISBN 1-59593-095-7. LCCN ???? ACM Order Number 505050.

Davenport:1986:PSM

[DR86] J. H. Davenport and C. E. Roth. PowerMath: a system for the Macintosh. In Char [Cha86b], pages 13– 15. ISBN 0-89791-199-7 (paperback). LCCN QA155.7.E4

A281 1986. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/32439/p13-davenport/. ACM order no.

505860.

Dominguez:2001:MIC

[DR01] César Domínguez and Julio Rubio. Modeling inheritance as coercion in a symbolic computation system. In Mourrain [Mou01], pages 109–115. ISBN 1-58113-417-7. LCCN QA76.95.I59 2001. ACM order number 505010.

Dora:1997:ANA

[DRJ97] J. Della Dora and F. Richard-Jung. About the Newton algorithm for non-linear ordinary differential equations. In Küchlin [Küc97], pages 298–304. ISBN 0-89791-875-4. LCCN

QA76.95. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/258726/p298-dora/.

Denzinger:1994:RAP

[DS94] J. Denzinger and S. Schulz. Recording, analyzing and presenting distributed deduction processes. In Hong [Hon94], pages 114– 123. ISBN 981-02-2040-5. LCCN QA76.642.I58 1994.

Dolzmann:1997:GEP

[DS97] Andreas Dolzmann and Thomas Sturm. Guarded expressions in practice. In Küchlin [Küc97], pages 376–383. ISBN 0-89791-

875-4. LCCN QA76.95. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/258726/p376-dolzmann/.

Dolzmann:1999:ACS

[DS99] Andreas Dolzmann and Thomas Sturm. P -adic constraint solving. In Dooley [Doo99], pages 151–158. ISBN 1-58113-073-2.

LCCN QA76.95 .I57 1999. URL http://www.acm.org/pubs/
contents/proceedings/issac/309831/.

Din:2003:PVC

[DS03] Mohab Safey El Din and Éric Schost. Polar varieties and computation of one point in each connected component of a smooth real algebraic set. In Senda [Sen03], pages 224–231. ISBN 1-58113-641-2. LCCN QA76.95. ACM order number 505030.

Dahan:2004:SET

[DS04] Xavier Dahan and Éric Schost. Sharp estimates for triangular sets. In Gutierrez [Gut04], pages 103–110. ISBN 1-58113-827-X. LCCN QA76.95 .I57 2004.

Dolzmann:2004:EPO

[DSS04] Andreas Dolzmann, Andreas Seidl, and Thomas Sturm. Efficient projection orders for CAD. In Gutierrez [Gut04], pages 111–118. ISBN 1-58113-827-X. LCCN QA76.95 .I57 2004.

Dumas:2000:ISF

[DSV00] Jean-Guillaume Dumas, B. David Saunders, and Gilles Villard. Integer Smith form via the valence: experience with large sparse matrices from homology. In Traverso [Tra00], pages 95–105. ISBN 1-58113-218-2. LCCN QA76.95.I59 2000. URL http:/

/www.acm.org/pubs/articles/proceedings/issac/345542/
p95-dumas/p95-dumas.pdf;   http://www.acm.org/pubs/
citations/proceedings/issac/345542/p95-dumas/. ACM

order number 505000.

Dora:1986:FSL

[DT86] J. Della Dora and E. Tournier. Formal solutions of linear difference equations: method of Pincherle–Ramis. In Char [Cha86b], pages 192–196. ISBN 0-89791-199-7 (paperback). LCCN QA155.7.E4 A281 1986. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/32439/p192-

della_dora/. ACM order no. 505860.

Du:1994:ISA

[Du94] Hong Du. On the isomorphisms of smooth algebraic curves. In ACM [ACM94], pages 15–19. ISBN 0-89791-638-7. LCCN

QA76.95.I59 1994. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/190347/p15-du/.

Dumas:2008:QAT

[Dum08] Jean-Guillaume Dumas. Q-adic transform revisited. In Jeffrey [Jef08], pages 63–70. ISBN 1-59593-904-0. LCCN ????

Debeerst:2008:SDE

[DvHK08] Ruben Debeerst, Mark van Hoeij, and Wolfram Koepf. Solving differential equations in terms of Bessel functions. In Jeffrey [Jef08], pages 39–46. ISBN 1-59593-904-0. LCCN ????

Doleh:1990:SSI

[DW90] Y. Doleh and P. S. Wang. SUI: a system independent user interface for an integrated scientific computing environment. In Watanabe and Nagata [WN90], pages 88–95. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95

.I57 1990. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/96877/p88-doleh/.

Dolzmann:2000:LQE

[DW00] Andreas Dolzmann and Volker Weispfenning. Local quantifier elimination. In Traverso [Tra00], pages 86–94. ISBN 1-58113-218-2. LCCN QA76.95.I59 2000. URL http:/

/www.acm.org/pubs/articles/proceedings/issac/345542/
p86-dolzmann/p86-dolzmann.pdf;   http://www.acm.org/
pubs/citations/proceedings/issac/345542/p86-dolzmann/

. ACM order number 505000.

Dyer:1994:ASC

[Dye94] Charles C. Dyer. An application of symbolic computation in the physical sciences. In ACM [ACM94], pages 181–186. ISBN 0-89791-638-7. LCCN QA76.95.I59 1994. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/190347/p181-
dyer/.

Dayton:2005:CMS

[DZ05] Barry H. Dayton and Zhonggang Zeng. Computing the multiplicity structure in solving polynomial systems. In Kauers [Kau05], pages 116–123. ISBN 1-59593-095-7. LCCN ???? ACM Order Number 505050.

Eberly:1997:PEP

[Ebe97] Wayne Eberly. Processor-efficient parallel matrix inversion over abstract fields: two extensions. In Küchlin [Küc97], pages 38–

45. ISBN 0-89791-875-4. LCCN QA76.95. URL http://www.

acm.org:80/pubs/citations/proceedings/issac/266670/
p38-eberly/.

Eberly:2000:BBF

[Ebe00] Wayne Eberly. Black box Frobenius decompositions over small fields. In Traverso [Tra00], pages 106–113. ISBN 1-58113-218-2.

LCCN QA76.95.I59 2000. URL http://www.acm.org/pubs/
articles/proceedings/issac/345542/p106-eberly/p106-
eberly.pdf;   http://www.acm.org/pubs/citations/proceedings/
issac/345542/p106-eberly/. ACM order number 505000.

Eberly:2003:ETS

[Ebe03] Wayne Eberly. Early termination over small fields. In Senda [Sen03], pages 80–87. ISBN 1-58113-641-2. LCCN QA76.95. ACM order number 505030.

Eberly:2004:RKB

[Ebe04] Wayne Eberly. Reliable Krylov-based algorithms for matrix null space and rank. In Gutierrez [Gut04], pages 127–134. ISBN 1-58113-827-X. LCCN QA76.95 .I57 2004.

Eberly:2010:YAB

[Ebe10] Wayne Eberly. Yet another block Lanczos algorithm: how to simplify the computation and reduce reliance on preconditioners in the small field case. In Watt [Wat10], pages 289–296. ISBN 1-4503-0150-9. LCCN QA76.95 .I59 2010.

Emiris:1993:PMS

[EC93] Ioannis Emiris and John Canny. A practical method for the sparse resultant. In Bronstein [Bro93], pages 183–192. ISBN 0-89791-604-2. LCCN QA 76.95 I59 1993. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/164081/p183-

emiris/. ACM order number: 505930.

Edneral:1993:CGN

[Edn93] Victor F. Edneral. Computer generation of normalizing transformation for systems of nonlinear ODE. In Bronstein [Bro93], pages 14–19. ISBN 0-89791-604-2. LCCN QA 76.95

I59 1993. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/164081/p14-edneral/. ACM order num-

ber: 505930.

Edneral:1997:CEC

[Edn97] Victor F. Edneral. Computer evaluation of cyclicity in planar cubic system. In Küchlin [Küc97], pages 305–309. ISBN 0-89791-875-4. LCCN QA76.95. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/258726/p305-
edneral/.

Einwohner:1989:MPG

[EF89] T. H. Einwohner and R. J. Fateman. A MACSYMA package for the generation and manipulation of Chebyshev series. In Gonnet [Gon89], pages 180–185. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989. US$29.00. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/74540/p180-

einwohner/. ACM order number: 505890. English and French.

Einwohner:1995:STI

[EF95] T. H. Einwohner and Richard J. Fateman. Searching techniques for integral tables. In Levelt [Lev95], pages 133–139. ISBN 0-89791-699-9. LCCN QA 76.95 I59 1995. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/220346/p133-

einwohner/. ACM order number: 505950.

Eberly:1996:EDA

[EG96] W. Eberly and M. Giesbrecht. Efficient decomposition of associative algebras. In Lakshman Y. N. [Lak96], pages 170–178. ISBN 0-89791-796-0. LCCN QA 76.95

I59 1996. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/236869/p170-eberly/.

Elkadi:2007:STP

[EG07] Mohamed Elkadi and André Galligo. Systems of three polynomials with two separated variables. In Brown [Bro07], pages 159–166. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Eberly:2006:SSR

[EGG^{+}06] Wayne Eberly, Mark Giesbrecht, Pascal Giorgi, Arne Storjohann, and Gilles Villard. Solving sparse rational linear systems. In Trager [Tra06], pages 63–70. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

Eberly:2007:FIO

[EGG^{+}07] Wayne Eberly, Mark Giesbrecht, Pascal Giorgi, Arne Storjohann, and Gilles Villard. Faster inversion and other black box matrix computations using efficient block projections. In Brown [Bro07], pages 143–150. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Elkadi:2004:PSH

[EGL04] Mohamed Elkadi, André Galligo, and Thi Ha Lê. Parametrized surfaces in huge P^{3} of bidegree (1, 2). In Gutierrez [Gut04], pages 141–148. ISBN 1-58113-827-X. LCCN QA76.95 .I57 2004.

Elsheikh:2012:FCS

[EGNS12] Mustafa Elsheikh, Mark Giesbrecht, Andy Novocin, and B. David Saunders. Fast computation of Smith forms of sparse matrices over local rings. In van der Hoeven and van Hoeij [vdHvH12], pages 146–153. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Emiris:2010:RPE

[EGT10] Ioannis Z. Emiris, André Galligo, and Elias P. Tsigaridas. Random polynomials and expected complexity of bisection methods for real solving. In Watt [Wat10], pages 235–242. ISBN 1-4503-0150-9. LCCN QA76.95 .I59 2010.

Estevez:1995:CAE

[EK95] L. W. Estevez and N. D. Kehtarnavaz. Computer assisted enhancement of mammograms for detection of microcalcifications. In IEEE [IEE95], pages 16–23. ISBN 0-8186-7117-3. LCCN R858.A2 I155 1995. IEEE catalog number 95CH35813.

Eberly:1997:RLA

[EK97] Wayne Eberly and Erich Kaltofen. On randomized Lanczos algorithms. In Küchlin [Küc97], pages 176–183. ISBN 0-89791-

875-4. LCCN QA76.95. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/258726/p176-eberly/.

Erlingsson:1996:GGO

[EKM96] Úlfar Erlingsson, Erich Kaltofen, and David Musser. Generic Gram–Schmidt orthogonalization by exact division. In Lakshman Y. N. [Lak96], pages 275–282. ISBN 0-89791-796-0. LCCN

QA 76.95 I59 1996. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/236869/p275-erlingsson/.

Eigenwillig:2007:FEG

[EKW07] Arno Eigenwillig, Michael Kerber, and Nicola Wolpert. Fast and exact geometric analysis of real algebraic plane curves. In Brown [Bro07], pages 151–158. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

ElDin:2013:CPM

[El 13] Mohab Safey El Din. Critical point methods and effective real algebraic geometry: new results and trends. In Monagan et al. [MCG13], pages 5–6. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

Elkadi:1999:NAG

[EM99] Mohamed Elkadi and Bernard Mourrain. A new algorithm for the geometric decomposition of a variety. In Dooley [Doo99], pages 9–16. ISBN 1-58113-073-2. LCCN QA76.95

.I57 1999.  URL http://www.acm.org/pubs/contents/
proceedings/issac/309831/.

Emiris:2009:MRF

[EM09] Ioannis Z. Emiris and Angelos A. Mantzaflaris. Multihomogeneous resultant formulae for systems with scaled support. In May [May09], pages 143–150. ISBN 1-60558-609-9. LCCN ????

Emiris:2010:DBM

[EMT10] Ioannis Z. Emiris, Bernard Mourrain, and Elias P. Tsigaridas. The DMM bound: multivariate (aggregate) separation bounds. In Watt [Wat10], pages 243–250. ISBN 1-4503-0150-9. LCCN QA76.95 .I59 2010.

Encarnacion:1994:MAC

[Enc94] Mark J. Encarnación. On a modular algorithm for computing GCDs of polynomials over algebraic number fields. In ACM [ACM94], pages 58–65. ISBN 0-89791-638-7. LCCN

QA76.95.I59 1994. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/190347/p58-encarnacion/.

Encarnacion:1997:ANM

[Enc97a] Mark J. Encarnación. The average number of modular factors in Trager’s polynomial factorization algorithm. In Küchlin [Küc97], pages 278–281. ISBN 0-89791-875-4. LCCN

QA76.95. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/258726/p278-encarnacion/.

Encarnacion:1997:FPA

[Enc97b] Mark J. Encarnación. Factoring polynomials over algebraic number fields via norms. In Küchlin [Küc97], pages 265–270. ISBN 0-89791-875-4. LCCN QA76.95. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/258726/p265-
encarnacion/.

Emiris:1997:SSR

[EP97] Ioanis Z. Emiris and Victor Y. Pan. The structure of sparse resultant matrices. In Küchlin [Küc97], pages 189–196. ISBN 0-89791-875-4. LCCN QA76.95. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/258726/p189-emiris/

.

Egner:1998:SPR

[EP98] Sebastian Egner and Markus Püschel. Solving puzzles related to permutation groups. In Gloor [Glo98], pages 186–193. ISBN 1-58113-002-3. LCCN QA155.7.E4 E88 1998. URL http://

www.acm.org/pubs/citations/proceedings/issac/281508/
p186-egner/.

Eder:2011:SBA

[EP11] Christian Eder and John Edward Perry. Signature-based algorithms to compute Gröbner bases. In Schost and Emiris [SE11], pages 99–106. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Egner:1997:DPC

[EPB97] Sebastian Egner, Markus Püschel, and Thomas Beth. Decomposing a permutation into a conjugated tensor product. In Küchlin [Küc97], pages 101–108. ISBN 0-89791-875-4. LCCN

QA76.95. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/258726/p101-egner/.

Edelsbrunner:1989:TPS

[EPW89] H. Edelsbrunner, F. P. Preparata, and D. B. West. Tetrahedrizing point sets in three dimensions. In Gianni [Gia89], pages 315–331. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Emiris:1994:MBP

[ER94] Ioannis Z. Emiris and Ashutosh Rege. Monomial bases and polynomial system solving (extended abstract). In ACM [ACM94], pages 114–122. ISBN 0-89791-638-7. LCCN

QA76.95.I59 1994. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/190347/p114-emiris/.

Eder:2013:SRG

[ER13] Christian Eder and Bjarke Hammersholt Roune. Signature rewriting in Gröbner basis computation. In Monagan et al. [MCG13], pages 331–338. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

Eriksson:2004:TIH

[Eri04] Nicholas Eriksson. Toric ideals of homogeneous phylogenetic models. In Gutierrez [Gut04], pages 149–154. ISBN 1-58113-827-X. LCCN QA76.95 .I57 2004.

Emeliyanenko:2012:CSB

[ES12] Pavel Emeliyanenko and Michael Sagraloff. On the complexity of solving a bivariate polynomial system. In van der Hoeven and van Hoeij [vdHvH12], pages 154–161. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Eigenwillig:2006:ATR

[ESY06] Arno Eigenwillig, Vikram Sharma, and Chee K. Yap. Almost tight recursion tree bounds for the Descartes method. In Trager [Tra06], pages 71–78. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

Farcot:2006:SNA

[Far06] Etienne Farcot. Symbolic numeric analysis of attractors in randomly generated piecewise affine models of gene networks. In Trager [Tra06], pages 79–86. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

Fateman:1981:CAN

[Fat81] Richard J. Fateman. Computer algebra and numerical integration. In Wang [Wan81], pages 228–232. ISBN 0-89791-047-

8. LCCN QA155.7.E4 A28 1981. US$23.00. ACM order no. 505810.

Fateman:1989:LTR

[Fat89a] R. J. Fateman. Lookup tables, recurrences and complexity. In Gonnet [Gon89], pages 68–73. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989. US$29.00. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/74540/p68-fateman/.

ACM order number: 505890. English and French.

Fateman:1989:SSA

[Fat89b] R. J. Fateman. Series solutions of algebraic and differential equations: a comparison of linear and quadratic algebraic convergence. In Gonnet [Gon89], pages 11–16. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989. US$29.00. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/74540/p11-fateman/

. ACM order number: 505890. English and French.

Fateman:1990:ATD

[Fat90] R. J. Fateman. Advances and trends in the design and construction of algebraic manipulation systems. In Watanabe and Nagata [WN90], pages 60–67. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95

.I57 1990. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/96877/p60-fateman/.

Fateman:1991:CRL

[Fat91] Richard J. Fateman. Canonical representations in Lisp and applications to computer algebra systems. In Watt [Wat91], pages 360–369. ISBN 0-89791-437-6. LCCN QA 76.95

I59 1991. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/120694/p360-fateman/.

Fateman:1992:HPG

[Fat92] Richard Fateman. Honest plotting, global extrema, and interval arithmetic. In Wang [Wan92b], pages 216–223. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN

QA76.95.I59 1992. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/143242/p216-fateman/. ACM

order number: 505920.

Fateman:1996:SMS

[Fat96] Richard J. Fateman. Symbolic mathematics system evaluators (extended abstract). In Lakshman Y. N. [Lak96], pages 86–94. ISBN 0-89791-796-0. LCCN QA 76.95

I59 1996. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/236869/p86-fateman/.

Fateman:1997:NSS

[Fat97] Richard J. Fateman. Network servers for symbolic mathematics. In Küchlin [Küc97], pages 249–256. ISBN 0-89791-

875-4. LCCN QA76.95. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/258726/p249-fateman/.

Fateman:2003:HLP

[Fat03] Richard Fateman. High-level proofs of mathematical programs using automatic differentiation, simplification, and some common sense. In Senda [Sen03], pages 88–94. ISBN 1-58113-641-2. LCCN QA76.95. ACM order number 505030.

Faure:1990:MS

[Fau90] C. Faure. A Meta simplifier. In Watanabe and Nagata [WN90], page 290. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95 .I57 1990. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/96877/p290-
faure/.

Faugere:1994:PGB

[Fau94] J. C. Faugere. Parallelization of Gröbner basis. In Hong [Hon94], pages 124–132. ISBN 981-02-2040-5. LCCN QA76.642.I58 1994.

Faugere:2002:NEA

[Fau02] Jean Charles Faugère. A new efficient algorithm for computing Gröbner bases without reduction to zero (F_{5}). In Mora [Mor02], pages 75–83. ISBN 1-58113-484-3. LCCN QA76.95. URL http:

//www.lifl.fr/ISSAC2002/.

Faugere:2009:IBC

[Fau09] Jean-Charles Faugère. Interactions between computer algebra (Gröbner bases) and cryptology. In May [May09], pages 383– 384. ISBN 1-60558-609-9. LCCN ????

Fee:1990:CCC

[Fee90] G. J. Fee. Computation of Catalan’s constant using Ramanujan’s formula. In Watanabe and Nagata [WN90], pages 157–160. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95 .I57 1990. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/96877/p157-fee/.

Faugere:2010:CLR

[FES10] Jean-Charles Faugère, Mohab Safey El Din, and Pierre-Jean Spaenlehauer. Computing loci of rank defects of linear matrices using Gröbner bases and applications to cryptology. In Watt [Wat10], pages 257–264. ISBN 1-4503-0150-9. LCCN QA76.95 .I59 2010.

Faugere:2012:CPG

[FES12] Jean-Charles Faugère, Mohab Safey El Din, and Pierre-Jean Spaenlehauer. Critical points and Gröbner bases: the unmixed case. In van der Hoeven and van Hoeij [vdHvH12], pages 162– 169. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Faugere:2013:CCG

[FEV13] Jean-Charles Faugère, Mohab Safey El Din, and Thibaut Verron. On the complexity of computing Gröbner bases for quasihomogeneous systems. In Monagan et al. [MCG13], pages 189– 196. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

Fernandez-Ferreiros:2000:MDW

[FFdlAGM00] Pilar Fernandez-Ferreiros and Maria de los Angeles Gomez-Molleda. A method for deciding whether the Galois group is abelian. In Traverso [Tra00], pages 114–120. ISBN 1-58113-218-2. LCCN QA76.95.I59 2000. URL http://www.

acm.org/pubs/articles/proceedings/issac/345542/p114-
fernandez-ferreiros/p114-fernandez-ferreiros.pdf;
http://www.acm.org/pubs/citations/proceedings/issac/
345542/p114-fernandez-ferreiros/. ACM order number

505000.

Fernandez-Ferreiros:2002:PSR

[FFGMGV02] P. Fernandez-Ferreiros, M. A. Gomez-Molleda, and L. Gonzalez-Vega. Partial solvability by radicals. In Mora [Mor02], pages 84–91. ISBN 1-58113-484-3. LCCN QA76.95. URL http://

www.lifl.fr/ISSAC2002/.

Freire:1989:ASC

[FGPF89] E. Freire, E. Gamero, E. Ponce, and L. G. Franquelo. An algorithm for symbolic computation of center manifolds. In Gianni [Gia89], pages 218–230. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Fortuna:2002:CTR

[FGPT02] Elisabetta Fortuna, Patrizia Gianni, Paola Parenti, and Carlo Traverso. Computing the topology of real algebraic surfaces. In Mora [Mor02], pages 92–100. ISBN 1-58113-484-3. LCCN

QA76.95. URL http://www.lifl.fr/ISSAC2002/.

Fortuna:2001:CRP

[FGT01] Elisabetta Fortuna, Patrizia Gianni, and Barry Trager. Computation of the radical of polynomial ideals over fields of arbitrary characteristic. In Mourrain [Mou01], pages 116–120. ISBN 1-58113-417-7. LCCN QA76.95.I59 2001. ACM order number 505010.

Fang:1997:WCI

[FH97] Xin Gui Fang and George Havas. On the worst-case complexity of integer Gaussian elimination. In Küchlin [Küc97], pages 28– 31. ISBN 0-89791-875-4. LCCN QA76.95. URL http://www.

acm.org:80/pubs/citations/proceedings/issac/258726/
p28-fang/.

Freeman:1986:SMP

[FIK86] T. Freeman, G. Imirzian, and E. Kaltofen. A system for manipulating polynomials given by straight-line programs. In Char [Cha86b], pages 169–175. ISBN 0-89791-199-7 (paperback). LCCN QA155.7.E4 A281 1986. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/32439/p169-freeman/

. ACM order no. 505860.

Fitch:1989:CRB

[Fit89] J. Fitch. Can REDUCE be run in parallel? In Gonnet [Gon89], pages 155–162. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989.

US$29.00. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/74540/p155-fitch/. ACM order num-

ber: 505890. English and French.

Fitch:1990:DSR

[Fit90] J. Fitch. A delivery system for REDUCE. In Watanabe and Nagata [WN90], pages 76–81. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95

.I57 1990. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/96877/p76-fitch/.

Faradzev:1991:CPC

[FK91a] I. A. Faradzev and M. H. Klin. Computer package for computations with coherent configurations. In Watt [Wat91], pages 219–223. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991.

Faradzev:1991:CCC

[FK91b] I. A. Faradzev and M. H. Klin. For computations with coherent configurations. In Watt [Wat91], pages 219–223. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/120694/p219-
faradzev/.

Filatei:2006:ITF

[FLMS06] Akpodigha Filatei, Xin Li, Marc Moreno Maza, and Éric Schost. Implementation techniques for fast polynomial arithmetic in a high-level programming environment. In Trager [Tra06], pages 93–100. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

Fujise:1996:PDD

[FM96] Tetsuro Fujise and Hirokazu Murao. Parallel distinct degree factorization algorithm. In Lakshman Y. N. [Lak96], pages 18–25. ISBN 0-89791-796-0. LCCN QA 76.95

I59 1996. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/236869/p18-fujise/.

Foursov:2001:CAC

[FM01] Mikhail V. Foursov and Marc Moreno Maza. On computerassisted classification of coupled integrable equations. In Mourrain [Mou01], pages 129–136. ISBN 1-58113-417-7. LCCN QA76.95.I59 2001. ACM order number 505010.

Farzan:2006:SRF

[FM06] Arash Farzan and J. Ian Munro. Succinct representation of finite abelian groups. In Trager [Tra06], pages 87–92. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

Faugere:2011:FAC

[FM11] Jean-Charles Faugère and Chenqi Mou. Fast algorithm for change of ordering of zero-dimensional Gröbner bases with sparse multiplication matrices. In Schost and Emiris [SE11], pages 115–122. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Faugere:2008:CPT

[FMRD08] Jean-Charles Faugère, Guillaume Moroz, Fabrice Rouillier, and Mohab Safey El Din. Classification of the perspective-three-point problem, discriminant variety and real solving polynomial systems of inequalities. In Jeffrey [Jef08], pages 79–86. ISBN 1-59593-904-0. LCCN ????

Fujimura:1997:MSS

[FN97] Masayo Fujimura and Kiyoko Nishizawa. Moduli spaces and symmetry loci of polynomial maps. In Küchlin [Küc97], pages 342–348. ISBN 0-89791-875-4. LCCN QA76.95. URL http://

www.acm.org:80/pubs/citations/proceedings/issac/258726/
p342-fujimura/.

Fitch:1986:AIA

[FNM86] J. Fitch, A. Norman, and M. A. Moore. Alkahest III: automatic analysis of periodic weakly nonlinear ODEs. In Char [Cha86b], pages 34–38. ISBN 0-89791-199-7 (paperback). LCCN

QA155.7.E4 A281 1986. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/32439/p34-fitch/. ACM or-

der no. 505860.

Fortune:2001:PRF

[For01] Steven Fortune. Polynomial root finding using iterated eigenvalue computation. In Mourrain [Mou01], pages 121–128. ISBN 1-58113-417-7. LCCN QA76.95.I59 2001. ACM order number 505010.

FaugAre:2009:HOD

[FP09] Jean-Charles Faugère and Ludovic Perret. High order derivatives and decomposition of multivariate polynomials. In May [May09], pages 207–214. ISBN 1-60558-609-9. LCCN ????

Ferreira:1995:PAI

[FR95] Afonso Ferreira and Jose D. P. Rolim, editors. Parallel algorithms for irregularly structured problems: second international workshop, IRREGULAR 95, Lyon, France, September 4–6, 1995: proceedings, volume 980 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. CODEN LNCSD9. ISBN 3-540-60321-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.642 .I59 1995.

FaugAre:2009:SSP

[FR09] Jean-Charles Faugère and Sajjad Rahmany. Solving systems of polynomial equations with symmetries using SAGBI-Gröbner bases. In May [May09], pages 151–158. ISBN 1-60558-609-9. LCCN ????

Franova:1990:PIC

[Fra90] M. Franová. PRECOMAS. An implementation of constructive matching methodology. In Watanabe and Nagata [WN90], pages 16–23. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95 .I57 1990. URL http://

www.acm.org:80/pubs/citations/proceedings/issac/96877/
p16-franova/.

Fredet:2000:LDE

[Fre00] Anne Fredet. Linear differential equations, iterative logarithms and orderings on monomial differential extensions. In Traverso [Tra00], pages 121–128. ISBN 1-58113-218-2.

LCCN QA76.95.I59 2000. URL http://www.acm.org/pubs/
articles/proceedings/issac/345542/p121-fredet/p121-
fredet.pdf;   http://www.acm.org/pubs/citations/proceedings/
issac/345542/p121-fredet/. ACM order number 505000.

Fredet:2003:FLD

[Fre03] Anne Fredet. Factorization of linear differential operators in exponential extensions. In Senda [Sen03], pages 103–110. ISBN 1-58113-641-2. LCCN QA76.95. ACM order number 505030.

Frandsen:2004:RSE

[FS04] Gudmund S. Frandsen and Igor E. Shparlinski. On reducing a system of equations to a single equation. In Gutierrez [Gut04], pages 163–166. ISBN 1-58113-827-X. LCCN QA76.95 .I57 2004.

Feo:2009:FAA

[FS09] Luca De Feo and Éric Schost. Fast arithmetics in Artin–Schreier towers over finite fields. In May [May09], pages 127–134. ISBN 1-60558-609-9. LCCN ????

Faugere:2012:SPS

[FS12] Jean-Charles Faugère and Jules Svartz. Solving polynomial systems globally invariant under an action of the symmetric group and application to the equilibria of N vortices in the plane. In van der Hoeven and van Hoeij [vdHvH12], pages 170–178. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Faugere:2013:GBI

[FS13] Jean-Charles Faugere and Jules Svartz. Gröbner bases of ideals invariant under a commutative group: the non-modular case. In Monagan et al. [MCG13], pages 347–354. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

Feng:2004:RGS

[FsG04] Ruyong Feng and Xiao shan Gao. Rational general solutions of algebraic ordinary differential equations. In Gutierrez [Gut04], pages 155–162. ISBN 1-58113-827-X. LCCN QA76.95 .I57 2004.

Furukawa:1986:GBM

[FSK86] A. Furukawa, T. Sasaki, and H. Kobayashi. The Gröbner basis of a module over KUX1,…,Xne and polynomial solutions of a system of linear equations. In Char [Cha86b], pages 222– 224. ISBN 0-89791-199-7 (paperback). LCCN QA155.7.E4

A281 1986. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/32439/p222-furukawa/. ACM order no.

505860.

Fateman:2003:CCR

[FT03] Richard J. Fateman and Raymond Toy. Converting call-by-reference to call-by-value: Fortran and Lisp coexisting. In Senda [Sen03], pages 95–102. ISBN 1-58113-641-2. LCCN QA76.95. ACM order number 505030.

Fukuda:2008:EAS

[Fuk08] Komei Fukuda. Exact algorithms and software in optimization and polyhedral computation. In Jeffrey [Jef08], pages 333–334. ISBN 1-59593-904-0. LCCN ????

Fang:2011:DSO

[FvH11] Tingting Fang and Mark van Hoeij. 2-descent for second order linear differential equations. In Schost and Emiris [SE11], pages 107–114. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Faugere:2010:DGM

[FvzGP10] Jean-Charles Faugère, Joachim von zur Gathen, and Ludovic Perret. Decomposition of generic multivariate polynomials. In Watt [Wat10], pages 131–137. ISBN 1-4503-0150-9. LCCN QA76.95 .I59 2010.

Gates:1986:NCG

[Gat86] Barbara L. Gates. A numerical code generation facility for RE-DUCE. In Char [Cha86b], pages 94–99. ISBN 0-89791-199-7 (paperback). LCCN QA155.7.E4 A281 1986. URL http://www.

acm.org:80/pubs/citations/proceedings/issac/32439/p94-

gates/. ACM order no. 505860.

Gatemann:1990:SSP

[Gat90a] K. Gatemann. Symbolic solution polynomial equation systems with symmetry. In Watanabe and Nagata [WN90], pages 112–119. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95 .I57 1990. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/96877/p112-
gatemann/.

Gatermann:1990:SSP

[Gat90b] K. Gatermann. Symbolic solution of polynomial equation systems with symmetry. In Watanabe and Nagata [WN90], pages 112–119. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95 .I57 1990.

Gatermann:1991:MSS

[Gat91] Karin Gatermann. Mixed symbolic-numeric solution of symmetrical nonlinear systems. In Watt [Wat91], pages 431–432. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991. URL http://www.

acm.org:80/pubs/citations/proceedings/issac/120694/
p431-gatermann/.

Gao:1991:CPE

[GC91] Xiao-Shan Gao and Shang-Ching Chou. Computations with parametric equations. In Watt [Wat91], pages 122–127. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/120694/p122-
gao/.

Gao:1992:SPA

[GC92] Xiao-Shan Gao and Shang-Ching Chou. Solving parametric algebraic systems. In Wang [Wan92b], pages 335–341. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN

QA76.95.I59 1992. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/143242/p335-gao/. ACM or-

der number: 505920.

Godlevsky:1993:PPA

[GD93] A. B. Godlevsky and A. E. Doroshenko. Parallelizing programs with APS. In Bronstein [Bro93], pages 55–62. ISBN 0-89791-604-2. LCCN QA 76.95 I59 1993. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/164081/p55-

godlevsky/. ACM order number: 505930.

Greuet:2011:DRI

[GE11] Aurélien Greuet and Mohab Safey El Din. Deciding reachability of the infimum of a multivariate polynomial. In Schost and Emiris [SE11], pages 131–138. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Geddes:1986:NIS

[Ged86] K. O. Geddes. Numerical integration in a symbolic context. In Char [Cha86b], pages 185–191. ISBN 0-89791-199-7 (paperback). LCCN QA155.7.E4 A281 1986. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/32439/p185-

geddes/. ACM order no. 505860.

Gemignani:1997:GPB

[Gem97] Luca Gemignani. GCD of polynomials and Bézout matrices. In Küchlin [Küc97], pages 271–277. ISBN 0-89791-875-4. LCCN

QA76.95. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/258726/p271-gemignani/.

Gemignani:2001:GGI

[Gem01] Luca Gemignani. A generalized Graeffe’s iteration for evaluating polynomials and rational functions. In Mourrain [Mou01], pages 143–149. ISBN 1-58113-417-7. LCCN QA76.95.I59 2001. ACM order number 505010.

Gemignani:2007:SMM

[Gem07] Luca Gemignani. Structured matrix methods for polynomial root-finding. In Brown [Bro07], pages 175–180. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Gerhard:1998:HDS

[Ger98] Jürgen Gerhard. High degree solutions of low degree equations (extended abstract). In Gloor [Glo98], pages 284–289. ISBN 1-58113-002-3. LCCN QA155.7.E4 E88 1998. URL http://

www.acm.org/pubs/citations/proceedings/issac/281508/
p284-gerhard/.

Gerhard:2010:AFA

[Ger10] Jürgen Gerhard. Asymptotically fast algorithms for modern computer algebra. In Watt [Wat10], pages 9–10. ISBN 1-4503-0150-9. LCCN QA76.95 .I59 2010.

Guo:2010:GOP

[GEZ10] Feng Guo, Mohab Safey El Din, and Lihong Zhi. Global optimization of polynomials using generalized critical values and sums of squares. In Watt [Wat10], pages 107–114. ISBN 1-4503-0150-9. LCCN QA76.95 .I59 2010.

Guo:2013:CRS

[GEZ13] Qingdong Guo, Mohab Safey El Din, and Lihong Zhi. Computing rational solutions of linear matrix inequalities. In Monagan et al. [MCG13], pages 197–204. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

Geddes:1992:HSI

[GF92] K. O. Geddes and G. J. Fee. Hybrid symbolic-numeric integration in MAPLE. In Wang [Wan92b], pages 36–41. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN QA76.95.I59 1992. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/143242/p36-geddes/.

ACM order number: 505920.

Geddes:1989:NAC

[GG89] K. O. Geddes and G. H. Gonnet. A new algorithm for computing symbolic limits using hierarchical series. In Gianni [Gia89], pages 490–495. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Garcia:2012:RIA

[GG12] Maria Emilia Alonso Garcia and André Galligo. A root isolation algorithm for sparse univariate polynomials. In van der Hoeven and van Hoeij [vdHvH12], pages 35–42. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Geddes:1989:HMO

[GGS89] K. O. Geddes, G. H. Gonnet, and T. J. Smedley. Heuristic methods for operations with algebraic numbers. In Gianni [Gia89], pages 475–480. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Gerhard:2003:SDP

[GGSZ03] J. Gerhard, M. Giesbrecht, A. Storjohann, and E. V. Zima. Shiftless decomposition and polynomial-time rational summation. In Senda [Sen03], pages 119–126. ISBN 1-58113-641-2. LCCN QA76.95. ACM order number 505030.

Gao:2010:NIA

[GGV10] Shuhong Gao, Yinhua Guan, and Frank Volny IV. A new incremental algorithm for computing Gröbner bases. In Watt [Wat10], pages 13–19. ISBN 1-4503-0150-9. LCCN QA76.95 .I59 2010.

Green:2000:CER

[GHS00] Edward L. Green, Lenwood S. Heath, and Craig A. Struble. Constructing endomorphism rings via duals. In Traverso [Tra00], pages 129–136. ISBN 1-58113-218-2.

LCCN QA76.95.I59 2000. URL http://www.acm.org/pubs/
articles/proceedings/issac/345542/p129-green/p129-green.
pdf;   http://www.acm.org/pubs/citations/proceedings/
issac/345542/p129-green/. ACM order number 505000.

Gianni:1989:SAC

[Gia89] P. (Patrizia) Gianni, editor. Symbolic and algebraic computation: International Symposium ISSAC ’88, Rome, Italy, July 4– 8, 1988: proceedings, volume 358 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1989. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Giesbrecht:1994:FAR

[Gie94] Mark Giesbrecht. Fast algorithms for rational forms of integer matrices. In ACM [ACM94], pages 305–311. ISBN 0-89791-638-7. LCCN QA76.95.I59 1994. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/190347/p305-
giesbrecht/.

Giesbrecht:1995:FCS

[Gie95] Mark Giesbrecht. Fast computation of the Smith Normal Form of an integer matrix. In Levelt [Lev95], pages 110–118. ISBN 0-89791-699-9. LCCN QA 76.95 I59 1995. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/220346/p110-

giesbrecht/. ACM order number: 505950.

Giesbrecht:1997:EPS

[Gie97] Mark Giesbrecht. Efficient parallel solution of sparse systems of linear Diophantine equations. In Küchlin [Küc97], pages 1– 10. ISBN 0-89791-875-4. LCCN QA76.95. URL http://www.

acm.org:80/pubs/citations/proceedings/issac/266670/
p1-giesbrecht/.

Gil:1992:CJC

[Gil92] Isabelle Gil. Computation of the Jordan canonical form of a square matrix (using the Axiom programming language). In Wang [Wan92b], pages 138–145. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN

QA76.95.I59 1992. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/143242/p138-gil/. ACM or-

der number: 505920.

Giusti:1989:CRC

[Giu89] M. Giusti. On the Castelnuovo regularity for curves. In Gonnet [Gon89], pages 250–253. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989. US$29.00. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/74540/p250-giusti/.

ACM order number: 505890. English and French.

Giusti:2009:GAF

[Giu09] Marc Giusti. A Gröbner free alternative to solving and a geometric analogue to Cook’s thesis. In May [May09], pages 1–2. ISBN 1-60558-609-9. LCCN ????

Giorgi:2003:CPM

[GJV03] Pascal Giorgi, Claude-Pierre Jeannerod, and Gilles Villard. On the complexity of polynomial matrix computations. In Senda [Sen03], pages 135–142. ISBN 1-58113-641-2. LCCN QA76.95. ACM order number 505030.

Gladitz:1994:PIG

[GK94] K. Gladitz and H. Kuchen. Parallel implementation of the gamma-operation on bags. In Hong [Hon94], pages 154–163. ISBN 981-02-2040-5. LCCN QA76.642.I58 1994.

Gerhold:2005:PPS

[GK05] Stefan Gerhold and Manuel Kauers. A procedure for proving special function inequalities involving a discrete parameter. In Kauers [Kau05], pages 156–162. ISBN 1-59593-095-7. LCCN ???? ACM Order Number 505050.

Garoufalidis:2012:TQH

[GK12] Stavros Garoufalidis and Christoph Koutschan. Twisting qholonomic sequences by complex roots of unity. In van der Hoeven and van Hoeij [vdHvH12], pages 179–186. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Gao:2004:AFM

[GKM^{+}04] Shuhong Gao, Erich Kaltofen, John May, Zhengfeng Yang, and Lihong Zhi. Approximate factorization of multivariate polynomials via differential equations. In Gutierrez [Gut04], pages 167–174. ISBN 1-58113-827-X. LCCN QA76.95 .I57 2004.

Giesbrecht:2002:ACS

[GKsL02] Mark Giesbrecht, Erich Kaltofen, and Wen shin Lee. Algorithms for computing the sparsest shifts of polynomials via the Berlekamp/Massey algorithm. In Mora [Mor02], pages 101–108. ISBN 1-58113-484-3. LCCN QA76.95. URL http://www.lifl.

fr/ISSAC2002/.

Gray:1994:MPE

[GKW94] Simon Gray, Norbert Kajler, and Paul Wang. MP: a protocol for efficient exchange of mathematical expressions. In ACM [ACM94], pages 330–335. ISBN 0-89791-638-7. LCCN QA76.95.I59 1994. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/190347/p330-gray/.

Gebauer:1991:CCA

[GKWW91] R. Gebauer, M. Kalkbrener, B. Wall, and F. Winkler. CASA: a computer algebra package for constructive algebraic geometry. In Watt [Wat91], pages 403–410. ISBN 0-89791-437-6. LCCN

QA 76.95 I59 1991. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/120694/p403-gebauer/.

Gerdt:1990:SAS

[GKZ90] V. P. Gerdt, N. V. Khutornoy, and A. Yu. Zharkov. Solving algebraic systems which arise as necessary integrability conditions for polynomial-nonlinear evolution equations. In Watanabe and Nagata [WN90], page 299. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95

.I57 1990. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/96877/p299-gerdt/.

Gerdt:1991:LSC

[GKZ91] V. P. Gerdt, N. V. Khutornoy, and A. Yu. Zharkov. Lie– Bäcklund symmetries of coupled nonlinear Schrödinger equations. In Watt [Wat91], pages 313–314. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/120694/p313-gerdt/.

Gaudry:2007:GBI

[GKZ07] Pierrick Gaudry, Alexander Kruppa, and Paul Zimmermann. A gmp-based implementation of Schönhage–Strassen’s large integer multiplication algorithm. In Brown [Bro07], pages 167–174. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Guo:2012:CIH

[GKZ12] Feng Guo, Erich L. Kaltofen, and Lihong Zhi. Certificates of impossibility of Hilbert–Artin representations of a given degree for definite polynomials and functions. In van der Hoeven and van Hoeij [vdHvH12], pages 195–202. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Grossman:1989:LTE

[GL89] R. Grossman and R. G. Larson. Labeled trees and the efficient computation of derivations. In Gonnet [Gon89], pages 74–80. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989.

US$29.00. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/74540/p74-grossman/. ACM order num-

ber: 505890. English and French.

Grudtsin:1991:ISI

[GL91] S. N. Grudtsin and V. N. Larin. Integrated system IN-TERCOMP and computer language for physicists. In Watt [Wat91], pages 377–381. ISBN 0-89791-437-6. LCCN QA 76.95

I59 1991. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/120694/p377-grudtsin/.

Graebe:1994:PGF

[GL94] H.-G. Graebe and W. Lassner. A parallel Gröbner factorizer. In Hong [Hon94], pages 174–180. ISBN 981-02-2040-5. LCCN QA76.642.I58 1994.

Grigoriev:1995:ACS

[GL95] Dima Yu. Grigoriev and Lakshman Y. N. Algorithms for computing sparse shifts for multivariate polynomials. In Levelt [Lev95], pages 96–103. ISBN 0-89791-699-9. LCCN QA 76.95

I59 1995. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/220346/p96-grigoriev/. ACM order

number: 505950.

Guo:2011:ACS

[GL11] Leilei Guo and Feng Liu. An algorithm for computing settheoretic generators of an algebraic variety. In Schost and Emiris [SE11], pages 139–146. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Glauert:1997:OGR

[Gla97] John Glauert. Object graph rewriting: an experimental parallel implementation. In Küchlin [Küc97], pages 119–128. ISBN 0-89791-875-4. LCCN QA76.95. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/266670/p119-
glauert/.

Geddes:2004:DRN

[GLL04] Keith Geddes, Ha Le, and Ziming Li. Differential rational normal forms and a reduction algorithm for hyperexponential func. In Gutierrez [Gut04], pages 183–190. ISBN 1-58113-827-X. LCCN QA76.95 .I57 2004.

Gloor:1998:IPI

[Glo98] Oliver Gloor, editor. ISSAC 98: Proceedings of the 1998 International Symposium on Symbolic and Algebraic Computation, August 13–15, 1998, University of Rostock, Germany. ACM Press, New York, NY 10036, USA, 1998. ISBN 1-58113-002-3. LCCN QA155.7.E4 E88 1998.

Gaonzalez:1989:SS

[GLRR89a] L. Gaonzalez, H. Lombardi, T. Recio, and M.-F. Roy. Sturm–Habicht sequence. In Gonnet [Gon89], pages 136– 146. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989.

US$29.00. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/74540/p136-gaonzalez/. ACM order

number: 505890. English and French.

Gonzalez:1989:SS

[GLRR89b] L. Gonzalez, H. Lombardi, T. Recio, and M.-F. Roy. Sturm– Habicht sequence. In Gonnet [Gon89], pages 136–146. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989. US$29.00. ACM order number: 505890. English and French.

Giesbrecht:1998:CIS

[GLS98] M. Giesbrecht, A. Lobo, and B. D. Saunders. Certifying inconsistency of sparse linear systems. In Gloor [Glo98], pages 113– 119. ISBN 1-58113-002-3. LCCN QA155.7.E4 E88 1998. URL

http://www.acm.org/pubs/citations/proceedings/issac/
281508/p113-giesbrecht/.

Giesbrecht:2006:SNS

[GLsL06] Mark Giesbrecht, George Labahn, and Wen shin Lee. Symbolicnumeric sparse interpolation of multivariate polynomials. In Trager [Tra06], pages 116–123. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

Giusti:1989:ATP

[GLV89] M. Giusti, D. Lazard, and A. Valibouze. Algebraic transformations of polynomial equations, symmetric polynomials and elimination. In Gianni [Gia89], pages 309–314. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Gebauer:1986:BAS

[GM86] Rüdiger Gebauer and H. Michael Möller. Buchberger’s algorithm and staggered linear bases. In Char [Cha86b], pages 218–221. ISBN 0-89791-199-7 (paperback). LCCN QA155.7.E4

A281 1986. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/32439/p218-gebauer/. ACM order no.

505860.

Gaudry:2006:FAC

[GM06] P. Gaudry and F. Morain. Fast algorithms for computing the eigenvalue in the Schoof–Elkies–Atkin algorithm. In Trager [Tra06], pages 109–115. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

Granvilliers:2001:SIC

[GMB01] Laurent Granvilliers, Eric Monfroy, and Frédéric Benhamou. Symbolic-interval cooperation in constraint programming. In Mourrain [Mou01], pages 150–166. ISBN 1-58113-417-7. LCCN QA76.95.I59 2001. ACM order number 505010.

Giovini:1991:OSC

[GMN^{+}91] Alessandro Giovini, Teo Mora, Gianfranco Niesi, Lorenzo Robbiano, and Carlo Traverso. ‘one sugar cube, please’ or selection strategies in the Buchberger algorithm. In Watt [Wat91], pages 49–54. ISBN 0-89791-437-6. LCCN QA 76.95

I59 1991. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/120694/p49-giovini/.

Guergueb:1994:EAT

[GMR94] Ahmed Guergueb, Jean Mainguené, and Marie-Françoise Roy. Examples of automatic theorem proving in real geometry. In ACM [ACM94], pages 20–24. ISBN 0-89791-638-7. LCCN

QA76.95.I59 1994. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/190347/p20-guergueb/.

Ganzha:1990:ARS

[GMS90] V. G. Ganzha, S. V. Meleshko, and V. P. Shelest. Application of REDUCE system for analyzing consistency of systems of PDE’s. In Watanabe and Nagata [WN90], page 301. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95

.I57 1990. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/96877/p301-ganzha/.

Gianni:1989:DA

[GMT89] P. Gianni, V. Miller, and B. Trager. Decomposition of algebras. In Gianni [Gia89], pages 300–308. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Golden:1986:OAM

[Gol86] J. P. Golden. An operator algebra for Macsyma. In Char [Cha86b], pages 244–246. ISBN 0-89791-199-7 (paperback).

LCCN QA155.7.E4 A281 1986. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/32439/p244-golden/.

ACM order no. 505860.

Gonnet:1986:IOS

[Gon86a] G. H. Gonnet. An implementation of operators for symbolic algebra systems. In Char [Cha86b], pages 239– 243. ISBN 0-89791-199-7 (paperback). LCCN QA155.7.E4

A281 1986. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/32439/p239-gonnet/. ACM order no.

505860.

Gonnet:1986:NRR

[Gon86b] Gaston H. Gonnet. New results for random determination of equivalence of expressions. In Char [Cha86b], pages 127– 131. ISBN 0-89791-199-7 (paperback). LCCN QA155.7.E4

A281 1986. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/32439/p127-gonnet/. ACM order no.

505860.

Gonnet:1989:PAI

[Gon89] Gaston H. Gonnet, editor. Proceedings of the ACM-SIGSAM 1989 International Symposium on Symbolic and Algebraic Computation: ISSAC ’89 / July 17–19, 1989, Portland, Oregon. ACM Press, New York, NY 10036, USA, 1989. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989. US$29.00. ACM order number: 505890. English and French.

GonzalezVega:1991:STM

[Gon91a] L. Gonzalez Vega. A subresultant theory for multivariate polynomials. In Watt [Wat91], pages 79–85. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991.

GonzalezVega:1991:WRA

[Gon91b] L. Gonzalez Vega. Working with real algebraic plane curves in REDUCE: the GCUR package. In Watt [Wat91], pages 397– 402. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991.

Gupta:1997:EDD

[GP97] Gopal Gupta and Enrico Pontelli. Extended dynamic dependent And-parallelism in ACE. In Küchlin [Küc97], pages 68–79. ISBN 0-89791-875-4. LCCN QA76.95. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/266670/p68-
gupta/.

Galligo:2005:SBB

[GP05] André Galligo and Jean Pascal Pavone. Selfintersections of a Bézier bicubic surface. In Kauers [Kau05], pages 148–155. ISBN 1-59593-095-7. LCCN ???? ACM Order Number 505050.

Gaal:1991:RIF

[GPP91] I. Gaál, A. Pethö, and M. Pohst. On the resolution of index form equations. In Watt [Wat91], pages 185–186. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/120694/p185-gaal/.

Galligo:1989:GEC

[GPT89] André Galligo, Loïc Pottier, and Carlo Traverso. Greater easy common divisor and standard basis completion algorithms. In Gianni [Gia89], pages 162–176. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Gutierrez:1992:PIT

[GR92] Jaime Gutierrez and Tomas Recio. A practical implementation of two rational function decomposition algorithms. In Wang [Wan92b], pages 152–157. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN

QA76.95.I59 1992. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/143242/p152-gutierrez/. ACM

order number: 505920.

Gautier:1994:PSP

[GR94] T. Gautier and J.-L. Roch. PAC++ system and parallel algebraic numbers computation. In Hong [Hon94], pages 145–153. ISBN 981-02-2040-5. LCCN QA76.642.I58 1994.

Gautier:1997:NCG

[GR97] Thierry Gautier and Jean-Louis Roch. NC^{2} computation of gcd-free basis and alication to parallel algebraic numbers computation. In Küchlin [Küc97], pages 31–37. ISBN 0-89791-

875-4. LCCN QA76.95. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/266670/p31-gautier/.

Galligo:2001:SND

[GR01] André Galligo and David Rupprecht. Semi-numerical determination of irreducible branches of a reduced space curve. In Mourrain [Mou01], pages 137–142. ISBN 1-58113-417-7. LCCN QA76.95.I59 2001. ACM order number 505010.

Giesbrecht:2008:LPP

[GR08] Mark Giesbrecht and Daniel S. Roche. On lacunary polynomial perfect powers. In Jeffrey [Jef08], pages 103–110. ISBN 1-59593-904-0. LCCN ????

Gerdt:2010:CFD

[GR10] Vladimir P. Gerdt and Daniel Robertz. Consistency of finite difference approximations for linear PDE systems and its algorithmic verification. In Watt [Wat10], pages 53–59. ISBN 1-4503-0150-9. LCCN QA76.95 .I59 2010.

Giesbrecht:2011:DII

[GR11] Mark Giesbrecht and Daniel S. Roche. Diversification improves interpolation. In Schost and Emiris [SE11], pages 123–130. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Gupta:2000:FPA

[GRA00] Anshul Gupta, Pankaj Rohatgi, and Ramesh Agarwal. Fast practical algorithms for the Boolean-product-witness-matrix problem. In Traverso [Tra00], pages 146–152. ISBN 1-58113-218-2. LCCN QA76.95.I59 2000. URL

http://www.acm.org/pubs/articles/proceedings/issac/
345542/p146-gupta/p146-gupta.pdf;   http://www.acm.
org/pubs/citations/proceedings/issac/345542/p146-gupta/

. ACM order number 505000.

Grigorev:1989:CCC

[Gri89] D. Yu. Grigor’ev. Complexity of computing the characters and the genre of a system of exterior differential equations. In Gianni [Gia89], pages 534–543. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Grigoriev:1990:CIT

[Gri90a] D. Yu. Grigoriev. Complexity of irreducibility testing for a system of linear ordinary differential equations. In Watanabe and Nagata [WN90], pages 225–230. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95

.I57 1990. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/96877/p225-grigoriev/.

Grigoriev:1990:HTS

[Gri90b] D. Yu. Grigoriev. How to test in subexponential time whether two points can be connected by a curve in a semialgebraic set.

In Watanabe and Nagata [WN90], pages 104–105. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN

QA76.95 .I57 1990. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/96877/p104-grigoriev/.

Grigoriev:1996:TSP

[Gri96] D. Grigoriev. Testing shift-equivalence of polynomials using quantum machines. In Lakshman Y. N. [Lak96], pages 49–54. ISBN 0-89791-796-0. LCCN QA 76.95

I59 1996. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/236869/p49-grigoriev/.

Gutierrez:2001:UFT

[GRS01] Jamie Gutierrez, Rosario Rubio, and David Sevilla. Unirational fields of transcendence degree one and functional decomposition. In Mourrain [Mou01], pages 167–174. ISBN 1-58113-417-7. LCCN QA76.95.I59 2001. ACM order number 505010.

Gruntz:1993:NAC

[Gru93] Dominik Gruntz. A new algorithm for computing asymptotic series. In Bronstein [Bro93], pages 239–244. ISBN 0-89791-604-2. LCCN QA 76.95 I59 1993. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/164081/p239-gruntz/

. ACM order number: 505930.

Geddes:1989:RIM

[GS89] K. O. Geddes and L. Y. Stefanus. On the Risch–Norman integration method and its implementation in MAPLE. In Gonnet [Gon89], pages 212–217. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989. US$29.00. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/74540/p212-geddes/.

ACM order number: 505890. English and French.

Ganzha:1990:LAS

[GS90] V. G. Ganzha and M. Yu. Shashkov. Local approximation study of difference operators by means of REDUCE system. In Watanabe and Nagata [WN90], pages 185–192. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95

.I57 1990. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/96877/p185-ganzha/.

Grigoriev:1998:PAS

[GS98] D. Grigoriev and A. Slissenko. Polytime algorithm for the shortest path in a homotopy class amidst semi-algebraic obstacles in the plane. In Gloor [Glo98], pages 17–24. ISBN 1-58113-002-3. LCCN QA155.7.E4 E88 1998. URL http://

www.acm.org/pubs/citations/proceedings/issac/281508/
p17-grigoriev/.

Giusti:1999:SSO

[GS99] Marc Giusti and Éric Schost. Solving some overdetermined polynomial systems. In Dooley [Doo99], pages 1–8. ISBN 1-58113-073-2. LCCN QA76.95 .I57 1999. URL http://www.

acm.org/pubs/contents/proceedings/issac/309831/.

Grigoriev:2005:GLD

[GS05] Dima Grigoriev and Fritz Schwarz. Generalized Loewydecomposition of d-modules. In Kauers [Kau05], pages 163–170. ISBN 1-59593-095-7. LCCN ???? ACM Order Number 505050.

Guo:2006:ERB

[GS06] Li Guo and William Y. Sit. Enumeration of Rota–Baxter words. In Trager [Tra06], pages 124–131. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

Grigoriev:2008:LDT

[GS08] Dima Grigoriev and Fritz Schwarz. Loewy decomposition of third-order linear PDE’s in the plane. In Jeffrey [Jef08], pages 277–286. ISBN 1-59593-904-0. LCCN ????

Grigoriev:2010:AFN

[GS10] D. Grigoriev and F. Schwarz. Absolute factoring of nonholonomic ideals in the plane. In Watt [Wat10], pages 93–97. ISBN 1-4503-0150-9. LCCN QA76.95 .I59 2010.

Gupta:2011:CHF

[GS11] Somit Gupta and Arne Storjohann. Computing Hermite forms of polynomial matrices. In Schost and Emiris [SE11], pages 155–162. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Gonzalez-Sanchez:2009:AGB

[GSGVPN^{+}09] Jon Gonzalez-Sanchez, Laureano Gonzalez-Vega, Alejandro Piñera-Nicolas, Irene Polo-Blanco, Jorge Caravantes, and Ignacio F. Rua. Analyzing group based matrix multiplication algorithms. In May [May09], pages 159–166. ISBN 1-60558-609-9. LCCN ????

Ganzha:1991:SAD

[GSV91] V. G. Ganzha, B. Yu. Scobelev, and E. V. Vorozhtsov. Stability analysis of difference schemes by the catastrophe theory methods and by means of computer algebra. In Watt [Wat91], pages 427–428. ISBN 0-89791-437-6. LCCN QA 76.95

I59 1991. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/120694/p427-ganzha/.

Gleixner:2012:IAL

[GSW12] Ambros M. Gleixner, Daniel E. Steffy, and Kati Wolter. Improving the accuracy of linear programming solvers with iterative refinement. In van der Hoeven and van Hoeij [vdHvH12], pages 187–194. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Guo:2011:RPL

[GSZ11] Li Guo, William Y. Sit, and Ronghua Zhang. On Rota’s problem for linear operators in associative algebras. In Schost and Emiris [SE11], pages 147–154. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Glueck:1990:AMT

[GT90] R. Glueck and V. F. Turchin. Application of metasystem transition to function inversion and transformation. In Watanabe and Nagata [WN90], pages 286–287. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95

.I57 1990. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/96877/p286-glueck/.

Gonzalez:1994:MPE

[GT94a] A. Gonzalez and J. Tubella. The multipath parallel execution model for Prolog. In Hong [Hon94], pages 164–173. ISBN 981-02-2040-5. LCCN QA76.642.I58 1994.

Goriely:1994:HCM

[GT94b] Alain Goriely and Michael Tabor. How to compute the Melnikov vector? In ACM [ACM94], pages 205–210. ISBN 0-89791-638-7. LCCN QA76.95.I59 1994. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/190347/p205-
goriely/.

Gianni:1997:ICN

[GT97] Patrizia Gianni and Barry Trager. Integral closure of Noetherian rings. In Küchlin [Küc97], pages 212–216. ISBN 0-89791-

875-4. LCCN QA76.95. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/258726/p212-gianni/.

Graillat:2009:NAC

[GT09] Stef Graillat and Philippe Trébuchet. A new algorithm for computing certified numerical approximations of the roots of a zerodimensional system. In May [May09], pages 167–174. ISBN 1-60558-609-9. LCCN ????

Gutnik:1993:ACA

[Gut93] S. A. Gutnik. Application of computer algebra to investigation of the relative equilibria of a satellite. In Bronstein [Bro93], pages 63–64. ISBN 0-89791-604-2. LCCN QA 76.95

I59 1993. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/164081/p63-gutnik/. ACM order num-

ber: 505930.

Gutierrez:2004:IJU

[Gut04] Jaime Gutierrez, editor. ISAAC 2004: July 4–7, 2004, University of Cantabria, Santander, Spain: proceedings of the 2004 International Symposium on Symbolic and Algebraic Computation. ACM Press, New York, NY 10036, USA, 2004. ISBN 1-58113-827-X. LCCN QA76.95 .I57 2004.

Gonzalez-Vega:1991:STM

[GV91a] Laureano González-Vega. A subresultant theory for multivariate polynomials. In Watt [Wat91], pages 79–85. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/120694/p79-
gonzalez-vega/.

Gonzalez-Vega:1991:WRA

[GV91b] Laureano González-Vega. Working with real algebraic plane curves in REDUCE the GCUR package. In Watt [Wat91], pages 397–402. ISBN 0-89791-437-6. LCCN QA 76.95

I59 1991. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/120694/p397-gonzalez-vega/.

Ganzha:1993:PSM

[GV93] V. G. Ganzha and E. V. Vorozhtsov. A probabilistic symbolicnumerical method for the stability analyses of difference schemes for PDEs. In Bronstein [Bro93], pages 9–13. ISBN 0-89791-604-2. LCCN QA 76.95 I59 1993. URL http://www.

acm.org:80/pubs/citations/proceedings/issac/164081/

p9-ganzha/. ACM order number: 505930.

Grigoriev:2000:BNV

[GV00] Dima Grigoriev and Nicolai Vorobjov. Bounds on numers of vectors of multiplicities for polynomials which are easy to compute. In Traverso [Tra00], pages 137–146. ISBN 1-58113-218-2. LCCN QA76.95.I59 2000. URL

http://www.acm.org/pubs/articles/proceedings/issac/
345542/p137-grigoriev/p137-grigoriev.pdf;   http://
www.acm.org/pubs/citations/proceedings/issac/345542/

p137-grigoriev/. ACM order number 505000.

Ganzha:1994:SSI

[GVBv94] V. G. Ganzha, E. V. Vorozhtsov, J. Boers, and J. A. van Hulzen. Symbolic-numeric stability investigations of Jameson’s schemes for the thin-layer Navier–Stokes equations. In ACM [ACM94], pages 234–241. ISBN 0-89791-638-7. LCCN

QA76.95.I59 1994. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/190347/p234-ganzha/.

Gonzalez-Vega:1995:IPC

[GVT95] L. González-Vega and G. Trujillo. Implicitization of parametric curves and surfaces by using symmetric functions. In Levelt [Lev95], pages 180–186. ISBN 0-89791-699-9. LCCN QA 76.95

I59 1995. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/220346/p180-gonzalez-vega/. ACM or-

der number: 505950.

Ganzha:1992:NSA

[GVv92] V. G. Ganzha, E. V. Vorozhtsov, and J. A. van Hulzen. A new symbolic-numeric approach to stability analysis of difference schemes. In Wang [Wan92b], pages 9–15. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN

QA76.95.I59 1992. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/143242/p9-ganzha/. ACM

order number: 505920.

Galligo:1997:NAP

[GW97] André Galligo and Stephen Watt. A numerical absolute primality test for bivariate polynomials. In Küchlin [Küc97], pages 217–224. ISBN 0-89791-875-4. LCCN

QA76.95. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/258726/p217-galligo/.

Gao:2006:RSD

[GY06] Xiao-Shan Gao and Chun-Ming Yuan. Resolvent systems of difference polynomial ideals. In Trager [Tra06], pages 101–108. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

Gerdt:1990:CGN

[GZ90] V. P. Gerdt and A. Yu. Zharkov. Computer generation of necessary integrability conditions for polynomial nonlinear evolution systems. In Watanabe and Nagata [WN90], pages 250–254. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95 .I57 1990. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/96877/p250-gerdt/.

Geddes:2003:EFH

[GZ03a] Keith O. Geddes and Wei Wei Zheng. Exploiting fast hardware floating point in high precision computation. In Senda [Sen03], pages 111–118. ISBN 1-58113-641-2. LCCN QA76.95. ACM order number 505030.

Giesbrecht:2003:FDO

[GZ03b] Mark Giesbrecht and Yang Zhang. Factoring and decomposing Øre polynomials over F_{q}(t). In Senda [Sen03], pages 127– 134. ISBN 1-58113-641-2. LCCN QA76.95. ACM order number 505030.

Gao:2004:DDP

[GZ04] Xiao-Shan Gao and Mingbo Zhang. Decomposition of differential polynomials with constant coefficients. In Gutierrez [Gut04], pages 175–182. ISBN 1-58113-827-X. LCCN QA76.95 .I57 2004.

Gerdt:2008:PDA

[GZ08] Vladimir P. Gerdt and Mikhail V. Zinin. A Pommaret division algorithm for computing Gröbner bases in Boolean rings. In Jeffrey [Jef08], pages 95–102. ISBN 1-59593-904-0. LCCN ????

Hammond:1994:PFP

[Ham94] K. Hammond. Parallel functional programming: An introduction (invited tutorial). In Hong [Hon94], pages 181–193. ISBN 981-02-2040-5. LCCN QA76.642.I58 1994.

Hanke:2007:IPC

[Han07] Timo Hanke. The isomorphism problem for cyclic algebras and an application. In Brown [Bro07], pages 181–186. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Harris:1994:IRR

[Har94] Jason F. Harris. Inheritance of rewrite rule structures applied to symbolic computation. In ACM [ACM94], pages 318–323. ISBN 0-89791-638-7. LCCN QA76.95.I59 1994. URL http://www.

acm.org:80/pubs/citations/proceedings/issac/190347/
p318-harris/.

Harris:2000:ANM

[Har00] Jason Harris. Advanced notations in mathematica. In Traverso [Tra00], pages 153–160. ISBN 1-58113-218-2.

LCCN QA76.95.I59 2000. URL http://www.acm.org/pubs/
articles/proceedings/issac/345542/p153-harris/p153-
harris.pdf;   http://www.acm.org/pubs/citations/proceedings/
issac/345542/p153-harris/. ACM order number 505000.

Havas:1991:CES

[Hav91] George Havas. Coset enumeration strategies. In Watt [Wat91], pages 191–199. ISBN 0-89791-437-6. LCCN QA 76.95

I59 1991. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/120694/p191-havas/.

Huang:2005:SPS

[HC05] Fangjian Huang and Shengli Chen. Schur partition for symmetric ternary forms and readable proof to inequalities. In Kauers [Kau05], pages 185–192. ISBN 1-59593-095-7. LCCN ???? ACM Order Number 505050.

Hill:1994:VM

[HCB94] J. M. D. Hill, K. M. Clarke, and R. Bornat. The vectorisation monad. In Hong [Hon94], pages 204–213. ISBN 981-02-2040-5. LCCN QA76.642.I58 1994.

Halstead:1993:APS

[HCGW93] R. H. Halstead, T. Chikayama, R. Gabriel, and D. Waltz. Applications for parallel symbolic computation. In Halstead and Ito [HI93], page 417. ISBN 0-387-57396-8, 3-540-57396-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.P3785 1993.

Hur:2000:ERA

[HD00] Namhyun Hur and James H. Davenport. An exact real algebraic arithmetic with equality determination. In Traverso [Tra00], pages 169–174. ISBN 1-58113-218-2. LCCN QA76.95.I59 2000.

URL http://www.acm.org/pubs/articles/proceedings/issac/
345542/p169-hur/p169-hur.pdf;   http://www.acm.org/
pubs/citations/proceedings/issac/345542/p169-hur/. ACM

order number 505000.

Hong:2009:VRQ

[HD09] Hoon Hong and Mohab Safey El Din. Variant real quantifier elimination: algorithm and application. In May [May09], pages 183–190. ISBN 1-60558-609-9. LCCN ????

Hovinen:2005:RBL

[HE05] Bradford Hovinen and Wayne Eberly. A reliable block Lanczos algorithm over small finite fields. In Kauers [Kau05], pages 177– 184. ISBN 1-59593-095-7. LCCN ???? ACM Order Number 505050.

Hermenegildo:1997:WSC

[Her97] M. Hermenegildo. Workshop 19: Symbolic computation. In Lengauer et al. [LGG97], pages 1167–1168. ISBN 3-540-63440-1 (paperback). LCCN QA76.58.I5535 1997.

Halstead:1993:PSC

[HI93] Robert H. Halstead and Takayasu Ito, editors. Parallel symbolic computing: languages, systems, and applications: US/Japan workshop, Cambridge, MA, USA, October 14–17, 1992: proceedings, number 748 in Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 0-387-57396-8, 3-540-57396-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.P3785 1993.

Hietarinta:1991:SIP

[Hie91] Jarmo Hietarinta. Searching for integrable PDE’s by testing Hirota’s three-soliton condition. In Watt [Wat91], pages 295–300. ISBN 0-89791-437-6. LCCN QA 76.95

I59 1991. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/120694/p295-hietarinta/.

Hietarinta:1992:SCQ

[Hie92] Jarmo Hietarinta. Solving the constant quantum Yang–Baxter equation in 2 dimensions with massive use of factorizing Gröbner basis computations. In Wang [Wan92b], pages 350– 357. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN QA76.95.I59 1992. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/143242/p350-

hietarinta/. ACM order number: 505920.

Hitz:2005:CNS

[Hit05] Markus A. Hitz. On computing nearest singular Hankel matrices. In Kauers [Kau05], pages 171–176. ISBN 1-59593-095-7. LCCN ???? ACM Order Number 505050.

Hasegawa:1994:PMM

[HK94] R. Hasegawa and M. Koshimura. An AND parallelization method for MGTP and its evaluation. In Hong [Hon94], pages 194–203. ISBN 981-02-2040-5. LCCN QA76.642.I58 1994.

Hitz:1998:EAC

[HK98] Markus A. Hitz and Erich Kaltofen. Efficient algorithms for computing the nearest polynomial with constrained roots. In Gloor [Glo98], pages 236–243. ISBN 1-58113-002-3. LCCN

QA155.7.E4 E88 1998. URL http://www.acm.org/pubs/
citations/proceedings/issac/281508/p236-hitz/.

Houari:2004:ARC

[HK04] Hassan El Houari and M’hammed El Kahoui. Algorithms for recognizing coordinates in two variables over UFD’s. In Gutierrez [Gut04], pages 135–140. ISBN 1-58113-827-X. LCCN QA76.95 .I57 2004.

Hitz:1999:EAC

[HKL99] Markus A. Hitz, Erich Kaltofen, and Y. N. Lakshman. Efficient algorithms for computing the nearest polynomial with a real root and related problems. In Dooley [Doo99], pages 205–212. ISBN 1-58113-073-2. LCCN QA76.95 .I57 1999. URL http://

www.acm.org/pubs/contents/proceedings/issac/309831/.

HalAs:2009:SRD

[HKL^{+}09] Miroslav Halás, Ülle Kotta, Ziming Li, Huaifu Wang, and Chunming Yuan. Submersive rational difference systems and their accessibility. In May [May09], pages 175–182. ISBN 1-60558-609-9. LCCN ????

Hutton:2010:CRP

[HKZ10] Sharon Hutton, Erich L. Kaltofen, and Lihong Zhi. Computing the radius of positive semidefiniteness of a multivariate real polynomial via a dual of Seidenberg’s method. In Watt [Wat10], pages 227–234. ISBN 1-4503-0150-9. LCCN QA76.95 .I59 2010.

Hayden:1986:SBC

[HL86] Michael B. Hayden and Edmund A. Lamagna. Summation of binomial coefficients using hypergeometric functions. In Char [Cha86b], pages 77–81. ISBN 0-89791-199-7 (paperback). LCCN QA155.7.E4 A281 1986. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/32439/p77-hayden/. ACM

order no. 505860.

Hulpke:2003:TOS

[HL03] Alexander Hulpke and Steve Linton. Total ordering on subgroups and cosets. In Senda [Sen03], pages 156–160. ISBN 1-58113-641-2. LCCN QA76.95. ACM order number 505030.

Hubert:2012:RIS

[HL12] Evelyne Hubert and George Labahn. Rational invariants of scalings from Hermite normal forms. In van der Hoeven and van Hoeij [vdHvH12], pages 219–226. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Hadzikadic:1986:AKB

[HLY86] M. Hadzikadic, F. Lichtenberger, and D. Y. Y. Yun. An application of knowledge-base technology in education: a geometry theorem prover. In Char [Cha86b], pages 141– 147. ISBN 0-89791-199-7 (paperback). LCCN QA155.7.E4

A281 1986. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/32439/p141-hadzikadic/. ACM order

no. 505860.

Hanrot:2001:SRA

[HM01] G. Hanrot and F. Morain. Solvability by radicals from an algorithmic point of view. In Mourrain [Mou01], pages 175–182.

ISBN 1-58113-417-7. LCCN QA76.95.I59 2001. ACM order number 505010.

Hunt:2001:SLR

[HMS01] Harry B. Hunt, Madhav V. Marathe, and Richard E. Stearns. Strongly-local reductions and the complexity/efficient approximability of algebra and optimization on abstract algebraic structures. In Mourrain [Mou01], pages 183–191. ISBN 1-58113-417-7. LCCN QA76.95.I59 2001. ACM order number 505010.

Holt:2000:CWH

[Hol00] Derek F. Holt. Computation in word-hyperbolic groups. In Traverso [Tra00], page ?? ISBN 1-58113-218-2. LCCN

QA76.95.I59 2000. URL http://www.acm.org/pubs/citations/
proceedings/issac/345542/p-holt/. ACM order number

505000.

Hong:1990:IPO

[Hon90] Hooh Hong. An improvement of the projection operator in cylindrical algebraic decomposition. In Watanabe and Nagata [WN90], pages 261–264. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95

.I57 1990. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/96877/p261-hong/.

Hong:1992:SSF

[Hon92] Hoon Hong. Simple solution formula construction in cylindrical algebraic decomposition based quantifier elimination. In Wang [Wan92b], pages 177–188. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN QA76.95.I59 1992. URL http:

//www.acm.org:80/pubs/citations/proceedings/issac/143242/

p177-hong/. ACM order number: 505920.

Hong:1993:QEF

[Hon93] Hoon Hong. Quantifier elimination for formulas constrained by quadratic equations. In Bronstein [Bro93], pages 264–274. ISBN 0-89791-604-2. LCCN QA 76.95 I59 1993. URL http://www.

acm.org:80/pubs/citations/proceedings/issac/164081/

p264-hong/. ACM order number: 505930.

Hong:1994:FIS

[Hon94] Hoon Hong, editor. First International Symposium on Parallel Symbolic Computation, PASCO ’94, Hagenberg/Linz, Austria, September 26–28, 1994, volume 5 of Lecture notes series in computing. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1994. ISBN 981-02-2040-5. LCCN QA76.642.I58 1994.

Hong:1996:GBU

[Hon96] Hoon Hong. Gröbner basis under composition II. In Lakshman Y. N. [Lak96], pages 79–85. ISBN 0-89791-796-0. LCCN

QA 76.95 I59 1996. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/236869/p79-hong/.

Hentzel:1989:VNA

[HP89] I. R. Hentzel and D. J. Pokrass. Verification of non-identities in algebras. In Gianni [Gia89], pages 496–507. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Huang:1997:FRM

[HP97] Xiaohan Huang and Victor Y. Pan. Fast rectangular matrix multiplications and improving parallel matrix computations. In Küchlin [Küc97], pages 11–23. ISBN 0-89791-875-4. LCCN

QA76.95. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/266670/p11-huang/.

Hubert:2003:CPS

[HR03] E. Hubert and N. Le Roux. Computing power series solutions of a nonlinear PDE system. In Senda [Sen03], pages 148–155. ISBN 1-58113-641-2. LCCN QA76.95. ACM order number 505030.

Harvey:2010:PTF

[HR10] David Harvey and Daniel S. Roche. An in-place truncated Fourier transform and applications to polynomial multiplication. In Watt [Wat10], pages 325–329. ISBN 1-4503-0150-9. LCCN QA76.95 .I59 2010.

Hossain:2002:SIC

[HS02] Shahadat Hossain and Trond Steihaug. Sparsity issues in the computation of Jacobian matrices. In Mora [Mor02], pages 123– 130. ISBN 1-58113-484-3. LCCN QA76.95. URL http://www.

lifl.fr/ISSAC2002/.

Harrison:2006:RPD

[HS06] Michael Harrison and Josef Schicho. Rational parametrisation for degree 6 Del Pezzo surfaces using Lie algebras. In Trager [Tra06], pages 132–137. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

Henrion:2008:PGC

[HS08] Didier Henrion and Michael Sebek. Plane geometry and convexity of polynomial stability regions. In Jeffrey [Jef08], pages 111–116. ISBN 1-59593-904-0. LCCN ????

Hubert:1996:GSO

[Hub96] Evelyne Hubert. The general solution of an ordinary differential equation. In Lakshman Y. N. [Lak96], pages 189–195. ISBN 0-89791-796-0. LCCN QA 76.95 I59 1996. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/236869/p189-
hubert/.

Hubert:2004:ITD

[Hub04] Evelyne Hubert. Improvements to a triangulation-decomposition algorithm for ordinary differential systems in higher degree cases. In Gutierrez [Gut04], pages 191–198. ISBN 1-58113-827-X. LCCN QA76.95 .I57 2004.

Hubert:2010:AIT

[Hub10] Evelyne Hubert. Algebraic invariants and their differential algebras. In Watt [Wat10], pages 1–2. ISBN 1-4503-0150-9. LCCN QA76.95 .I59 2010.

Hulpke:1998:CNS

[Hul98] Alexander Hulpke. Computing normal subgroups. In Gloor [Glo98], pages 194–198. ISBN 1-58113-002-3. LCCN

QA155.7.E4 E88 1998. URL http://www.acm.org/pubs/
citations/proceedings/issac/281508/p194-hulpke/.

Hulpke:2013:CST

[Hul13] Alexander J. Hulpke. Calculation of the subgroups of a trivialfitting group. In Monagan et al. [MCG13], pages 205–210. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

Hart:2011:PPF

[HvHN11] William Hart, Mark van Hoeij, and Andrew Novocin. Practical polynomial factoring in polynomial time. In Schost and Emiris [SE11], pages 163–170. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Haridi:1997:ODD

[HVS97] Seif Haridi, Peter Van Roy, and Gert Smolka. An overview of the design of Distributed Oz. In Küchlin [Küc97], pages 176– 187. ISBN 0-89791-875-4. LCCN QA76.95. URL http://www.

acm.org:80/pubs/citations/proceedings/issac/266670/
p176-haridi/.

Hilali:1986:ACF

[HW86] A. Hilali and A. Wazner. Algorithm for computing formal invariants of linear differential systems. In Char [Cha86b], pages 197–201. ISBN 0-89791-199-7 (paperback). LCCN QA155.7.E4

A281 1986. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/32439/p197-hilali/. ACM order no.

505860.

Huang:2000:PMP

[HWSZ00] Yuzhen Huang, Wenda Wu, Hans J. Stetter, and Lihong Zhi. Pseudofactors of multivariate polynomials. In Traverso [Tra00], pages 161–168. ISBN 1-58113-218-2.

LCCN QA76.95.I59 2000. URL http://www.acm.org/pubs/
articles/proceedings/issac/345542/p161-huang/p161-huang.
pdf;   http://www.acm.org/pubs/citations/proceedings/
issac/345542/p161-huang/. ACM order number 505000.

Iglio:1998:SCC

[IA98] Pietro Iglio and Giuseppe Attardi. Software components for computer algebra. In Gloor [Glo98], pages 62–69. ISBN 1-58113-002-3. LCCN QA155.7.E4 E88 1998. URL http://www.

acm.org/pubs/citations/proceedings/issac/281508/p62-
iglio/.

Ida:2009:SAM

[Ida09] Tetsuo Ida. Symbolic and algebraic methods in computational origami: invited talk. In May [May09], pages 3–4. ISBN 1-60558-609-9. LCCN ????

IEEE:1995:PEI

[IEE95] IEEE, editor. Proceedings of the Eighth IEEE Symposium on Computer-Based Medical Systems / June 9–10, 1995, Lubbock, Texas. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-7117-3. LCCN R858.A2 I155 1995. IEEE catalog number 95CH35813.

Ilyin:1991:SST

[IK91] V. A. Ilyin and A. P. Kryukov. Symbolic simplification of tensor expressions using symmetries, dummy indices and identities. In Watt [Wat91], pages 224–228. ISBN 0-89791-437-6.

LCCN QA 76.95 I59 1991. URL http://www.acm.org:80/
pubs/citations/proceedings/issac/120694/p224-ilyin/.

Ishikawa:2012:ZHA

[IK12] Masao Ishikawa and Christoph Koutschan. Zeilberger’s holonomic ansatz for Pfaffians. In van der Hoeven and van Hoeij [vdHvH12], pages 227–233. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Ilyin:1991:PIF

[IKRT91] V. A. Ilyin, A. P. Kryukov, A. Ya. Rodionov, and A. Yu. Taranov. PC implementation of fast Dirac matrix trace calculations. In Watt [Wat91], pages 456–457. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/120694/p456-ilyin/.

Ivanyos:2009:SDP

[IKS09] Gábor Ivanyos, Marek Karpinski, and Nitin Saxena. Schemes for deterministic polynomial factoring. In May [May09], pages 191–198. ISBN 1-60558-609-9. LCCN ????

Ito:1993:MPA

[INPS93] T. Ito, R. Nikhil, J. Padget, and N. Suzuki. Massively parallel architectures and symbolic computation. In Halstead and Ito [HI93], pages 408–416. ISBN 0-387-57396-8, 3-540-57396-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.P3785 1993.

Ivanyos:2000:FRA

[Iva00] Gábor Ivanyos. Fast randomized algorithms for the structure of matrix algebras over finite fields (extended abstract). In Traverso [Tra00], pages 175–183. ISBN 1-58113-218-2. LCCN QA76.95.I59 2000. URL http://www.

acm.org/pubs/articles/proceedings/issac/345542/p175-
ivanyos/p175-ivanyos.pdf;   http://www.acm.org/pubs/
citations/proceedings/issac/345542/p175-ivanyos/. ACM

order number 505000.

Jacobs:1994:ANA

[Jac94] David P. Jacobs. The Albert nonassociative algebra system: a progress report. In ACM [ACM94], pages 41–44. ISBN 0-89791-638-7. LCCN QA76.95.I59 1994. URL http://www.

acm.org:80/pubs/citations/proceedings/issac/190347/
p41-jacobs/.

Jeannerod:2000:AEP

[Jea00] Claude-Pierre Jeannerod. An algorithm for the eigenvalue perturbation problem: reduction of a ?-matrix to a Lidskii matrix. In Traverso [Tra00], pages 184–191. ISBN 1-58113-218-2. LCCN QA76.95.I59 2000. URL http://

www.acm.org/pubs/citations/proceedings/issac/345542/

p184-jeannerod/. ACM order number 505000.

Jeannerod:2002:RFP

[Jea02] Claude Pierre Jeannerod. A reduced form for perturbed matrix polynomials. In Mora [Mor02], pages 131–137. ISBN 1-58113-484-3. LCCN QA76.95. URL http://www.lifl.fr/

ISSAC2002/.

Jebelean:1993:GBG

[Jeb93] T. Jebelean. A generalization of the binary GCD algorithm. In Bronstein [Bro93], pages 111–116. ISBN 0-89791-604-2. LCCN QA 76.95 I59 1993. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/164081/p111-

jebelean/. ACM order number: 505930.

Jebelean:1997:PID

[Jeb97] Tudor Jebelean. Practical integer division with Karatsuba complexity. In Küchlin [Küc97], pages 339–341. ISBN 0-89791-

875-4. LCCN QA76.95. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/258726/p339-jebelean/.

Jeffrey:1993:IOE

[Jef93] D. J. Jeffrey. Integration to obtain expressions valid on domains of maximum extent. In Bronstein [Bro93], pages 34–41. ISBN 0-89791-604-2. LCCN QA 76.95 I59 1993. URL http://www.

acm.org:80/pubs/citations/proceedings/issac/164081/

p34-jeffrey/. ACM order number: 505930.

Jeffrey:2008:PAM

[Jef08] David Jeffrey, editor. Proceedings of the 21st annual meeting of the International Symposium on Symbolic Computation, ISSAC 2008, July 20–23, 2008, Hagenberg, Austria. ACM Press, New York, NY 10036, USA, 2008. ISBN 1-59593-904-0. LCCN ????

Jenks:1976:SPA

[Jen76] Richard D. Jenks, editor. Symsac ’76: proceedings of the 1976 ACM Symposium on Symbolic and Algebraic Computation, August 10–12, 1976, Yorktown Heights, New York. ACM Press, New York, NY 10036, USA, 1976. LCCN QA155.7.E4 A15 1976. US$20.00.

Janovitz-Freireich:2006:ARI

[JFRS06] Itnuit Janovitz-Freireich, Lajos Rónyai, and Ágnes Szántó. Approximate radical of ideals with clusters of roots. In Trager [Tra06], pages 146–153. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

Janovitz-Freireich:2008:MMT

[JFSMR08] Itnuit Janovitz-Freireich, Agnes Szántó, Bernard Mourrain, and Lajos Ronyai. Moment matrices, trace matrices and the radical of ideals. In Jeffrey [Jef08], pages 125–132. ISBN 1-59593-904-0. LCCN ????

Johnson:1997:PRR

[JK97] J. R. Johnson and Werner Krandick. Polynomial real root isolation using approximate arithmetic. In Küchlin [Küc97], pages 225–232. ISBN 0-89791-875-4. LCCN

QA76.95. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/258726/p225-johnson/.

Johnson:2006:HPI

[JKL^{+}06] Jeremy R. Johnson, Werner Krandick, Kevin Lynch, David G. Richardson, and Anatole D. Ruslanov. High-performance implementations of the Descartes method. In Trager [Tra06], pages 154–161. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

Johansson:2013:FHS

[JKM13] Fredrik Johansson, Manuel Kauers, and Marc Mezzarobba. Finding hyperexponential solutions of linear ODEs by numerical evaluation. In Monagan et al. [MCG13], pages 211–218. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

Johnson:2005:AAC

[JKR05] Jeremy R. Johnson, Werner Krandick, and Anatole D. Ruslanov. Architecture-aware classical Taylor shift by 1. In Kauers [Kau05], pages 200–207. ISBN 1-59593-095-7. LCCN ???? ACM Order Number 505050.

Liu:1990:AFA

[jL90] Zhuo jun Liu. An algorithm for finding all isolated zeros of polynomial systems. In Watanabe and Nagata [WN90], page 300. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95 .I57 1990. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/96877/p300-liu/.

Jeffrey:1997:ISP

[JLVR97] D. J. Jeffrey, G. Labahn, M. Von Mohrenschildt, and A. D. Rich. Integration of the signum, piecewise and related functions. In Küchlin [Küc97], pages 324–330. ISBN 0-89791-

875-4. LCCN QA76.95. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/258726/p324-jeffrey/.

Javadi:2007:SMG

[JM07] Seyed Mohammad Mahdi Javadi and Michael Monagan. A sparse modular GCD algorithm for polynomials over algebraic function fields. In Brown [Bro07], pages 187–194. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Javadi:2009:FMP

[JM09] Seyed Mohammad Mahdi Javadi and Michael B. Monagan. On factorization of multivariate polynomials over algebraic number and function fields. In May [May09], pages 199–206. ISBN 1-60558-609-9. LCCN ????

Jeannerod:2010:CSG

[JM10] Claude-Pierre Jeannerod and Christophe Mouilleron. Computing specified generators of structured matrix inverses. In Watt [Wat10], pages 281–288. ISBN 1-4503-0150-9. LCCN QA76.95 .I59 2010.

Johnson:1992:RAN

[Joh92] J. R. Johnson. Real algebraic number computation using interval arithmetic. In Wang [Wan92b], pages 195– 205. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN QA76.95.I59 1992. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/143242/p195-

johnson/. ACM order number: 505920.

Jeannerod:1999:RAM

[JP99] C.-P. Jeannerod and E. Pflügel. A reduction algorithm for matrices depending on a parameter. In Dooley [Doo99], pages 121–128. ISBN 1-58113-073-2. LCCN QA76.95 .I57 1999. URL

http://www.acm.org/pubs/contents/proceedings/issac/
309831/.

Jeffrey:2005:ATA

[JPR05] D. J. Jeffrey, Pratibha, and K. B. Roach. Affine transformations of algebraic numbers. In Kauers [Kau05], pages 193–199. ISBN 1-59593-095-7. LCCN ???? ACM Order Number 505050.

Jeffrey:1998:RIP

[JR98] D. J. Jeffrey and A. D. Rich. Recursive integration of piecewisecontinuous functions. In Gloor [Glo98], pages 290–294. ISBN 1-58113-002-3. LCCN QA155.7.E4 E88 1998. URL http://

www.acm.org/pubs/citations/proceedings/issac/281508/
p290-jeffrey/.

Jenks:1994:HMA

[JT94] Richard D. Jenks and Barry M. Trager. How to make AX-IOM into a Scratchpad. In ACM [ACM94], pages 32–40. ISBN 0-89791-638-7. LCCN QA76.95.I59 1994. URL http://www.

acm.org:80/pubs/citations/proceedings/issac/190347/
p32-jenks/.

Juozapavicius:1989:SCW

[Juo89] A. Juozapavicius. Symbolic computation for Witt rings. In Gianni [Gia89], pages 271–273. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Jurkovic:1986:EES

[Jur86] N. Jurkovic. Edusym — educational symbolic manipulator on a microcomputer. In Char [Cha86b], pages 154– 156. ISBN 0-89791-199-7 (paperback). LCCN QA155.7.E4

A281 1986. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/32439/p154-jurkovic/. ACM order no.

505860.

Jurkovic:2001:DCS

[Jur01] Neven Jurkovic. Diagnosing and correcting student’s misconceptions in an educational computer algebra system. In Mourrain [Mou01], pages 195–200. ISBN 1-58113-417-7. LCCN QA76.95.I59 2001. ACM order number 505010.

Johnson:2007:GSD

[JX07] Jeremy Johnson and Xu Xu. Generating symmetric DFTs and equivariant FFT algorithms. In Brown [Bro07], pages 195–202. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Jinwang:2001:MPI

[JZXM01] Liu Jinwang, Liu Zhuojun, Liu Xiaoqi, and Wang Mingsheng. The membership problem for ideals of binomial skew polynomial rings. In Mourrain [Mou01], pages 192–195. ISBN 1-58113-417-7. LCCN QA76.95.I59 2001. ACM order number 505010.

Kaib:1994:FVG

[Kai94] M. Kaib. A fast variant of the Gaussian reduction algorithm. In Adleman and Huang [AH94], page 159. ISBN 0-387-58691-1 (New York), 3-540-58691-1 (Berlin). LCCN QA241.A43 1994.

Kajler:1992:CPE

[Kaj92] Norbert Kajler. CAS/PI: a portable and extensible interface for computer algebra systems. In Wang [Wan92b], pages 376– 386. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN QA76.95.I59 1992. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/143242/p376-

kajler/. ACM order number: 505920.

Kalkbrener:1990:SSB

[Kal90] M. Kalkbrener. Solving systems of bivariate algebraic equations by using primitive polynomial remainder sequences. In Watanabe and Nagata [WN90], page 295. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95

.I57 1990. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/96877/p295-kalkbrener/.

Kaltofen:1992:CDM

[Kal92] Erich Kaltofen. On computing determinants of matrices without divisions. In Wang [Wan92b], pages 342– 349. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN QA76.95.I59 1992. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/143242/p342-

kaltofen/. ACM order number: 505920.

Kalkbrener:1993:UBN

[Kal93] Michael Kalkbrener. An upper bound on the number of monomials in the Sylvester resultant. In Bronstein [Bro93], pages 161–163. ISBN 0-89791-604-2. LCCN QA 76.95

I59 1993. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/164081/p161-kalkbrener/. ACM order

number: 505930.

Kaltofen:1994:AFS

[Kal94] Erich Kaltofen. Asymptotically fast solution of Toeplitz-like singular linear systems. In ACM [ACM94], pages 297–304. ISBN 0-89791-638-7. LCCN QA76.95.I59 1994. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/190347/p297-
kaltofen/.

Kaltofen:1996:ISC

[Kal96] E. Kaltofen. ISSAC steering committee bylaws. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 30(1):31–33, March 1996. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Kaltofen:2002:OSV

[Kal02] Erich Kaltofen. An output-sensitive variant of the baby steps/ giant steps determinant algorithm. In Mora [Mor02], pages 138– 144. ISBN 1-58113-484-3. LCCN QA76.95. URL http://www.

lifl.fr/ISSAC2002/.

Kaltofen:2003:PFS

[Kal03] Erich Kaltofen. Polynomial factorization: a success story. In Senda [Sen03], pages 3–4. ISBN 1-58113-641-2. LCCN QA76.95. ACM order number 505030.

Kapur:1986:GTP

[Kap86] Deepak Kapur. Geometry theorem proving using Hilbert’s Nullstellensatz. In Char [Cha86b], pages 202–208. ISBN 0-89791-199-7 (paperback). LCCN QA155.7.E4 A281 1986. URL http:/

/www.acm.org:80/pubs/citations/proceedings/issac/32439/

p202-kapur/. ACM order no. 505860.

Kauers:2004:CPP

[Kau04] Manuel Kauers. Computer proofs for polynomial identities in arbitrary many variables. In Gutierrez [Gut04], pages 199–204. ISBN 1-58113-827-X. LCCN QA76.95 .I57 2004.

Kauers:2005:IJB

[Kau05] Manuel Kauers, editor. ISSAC ’05: July 24–27, 2005, Beijing, China: Proceedings of the 2005 International Symposium on Symbolic and Algebraic Computation. ACM Press, New York, NY 10036, USA, 2005. ISBN 1-59593-095-7. LCCN ???? ACM Order Number 505050.

Kauers:2008:IAF

[Kau08] Manuel Kauers. Integration of algebraic functions: a simple heuristic for finding the logarithmic part. In Jeffrey [Jef08], pages 133–140. ISBN 1-59593-904-0. LCCN ????

Kauers:2011:CT

[Kau11] Manuel Kauers. The concrete tetrahedron. In Schost and Emiris [SE11], pages 7–8. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Kunkle:2007:TSM

[KC07] Daniel Kunkle and Gene Cooperman. Twenty-six moves suffice for Rubik’s cube. In Brown [Bro07], pages 235–242. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Kunkle:2009:BTF

[KC09] Daniel Kunkle and Gene Cooperman. Biased tadpoles: a fast algorithm for centralizers in large matrix groups. In May [May09], pages 223–230. ISBN 1-60558-609-9. LCCN ????

Kemper:2008:AIT

[Kem08] Gregor Kemper. Algorithmic invariant theory. In Jeffrey [Jef08], pages 335–336. ISBN 1-59593-904-0. LCCN ????

Kesseler:1994:RGC

[Kes94] M. Kesseler. Reducing graph copying costs — time to wrap it up. In Hong [Hon94], pages 244–253. ISBN 981-02-2040-5. LCCN QA76.642.I58 1994.

Kornyak:1991:PSA

[KF91] V. V. Kornyak and W. I. Fushchich. A program for symmetry analysis of differential equations. In Watt [Wat91], pages 315–316. ISBN 0-89791-437-6. LCCN QA 76.95

I59 1991. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/120694/p315-kornyak/.

Kobayashi:1986:GBI

[KFS86] H. Kobayashi, A. Furukawa, and T. Sasaki. Gröbner bases of ideals of convergent power series. In Char [Cha86b], pages 225–227. ISBN 0-89791-199-7 (paperback). LCCN QA155.7.E4

A281 1986. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/32439/p225-kobayashi/. ACM order no.

505860.

Khanin:2001:DAC

[Kha01] Raya Khanin. Dimensional analysis in computer algebra. In Mourrain [Mou01], pages 201–208. ISBN 1-58113-417-7. LCCN QA76.95.I59 2001. ACM order number 505010.

Khetan:2002:DFC

[Khe02] Amit Khetan. Determinantal formula for the chow form of a toric surface. In Mora [Mor02], pages 145–150. ISBN 1-58113-484-3. LCCN QA76.95. URL http://www.lifl.fr/

ISSAC2002/.

Kirrinnis:1992:FCN

[Kir92] Peter Kirrinnis. Fast computation of numerical partial fraction decompositions and contour integrals of rational functions. In Wang [Wan92b], pages 16–26. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN

QA76.95.I59 1992. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/143242/p16-kirrinnis/. ACM

order number: 505920.

Kirrinnis:1998:FNI

[Kir98] Peter Kirrinnis. Fast numerical improvement of factors of polynomial and of partial fractions. In Gloor [Glo98], pages 260– 267. ISBN 1-58113-002-3. LCCN QA155.7.E4 E88 1998. URL

http://www.acm.org/pubs/citations/proceedings/issac/
281508/p260-kirrinnis/.

Krandick:1994:BEI

[KJ94] W. Krandick and T. Jebelean. Bidirectional exact integer division. In Hong [Hon94], pages 264–272. ISBN 981-02-2040-5. LCCN QA76.642.I58 1994.

Kirchner:1989:CER

[KK89] C. Kirchner and H. Kirchner. Constrained equational reasoning. In Gonnet [Gon89], pages 382–389. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989. US$29.00. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/74540/p382-

kirchner/. ACM order number: 505890. English and French.

Kleczka:1991:SCA

[KK91] W. Kleczka and E. Kreuzer. Systematic computer-aided analysis of dynamic systems. In Watt [Wat91], pages 429–430. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991. URL http://www.

acm.org:80/pubs/citations/proceedings/issac/120694/
p429-kleczka/.

Kaltofen:2005:CFB

[KK05] Erich Kaltofen and Pascal Koiran. On the complexity of factoring bivariate supersparse (lacunary) polynomials. In Kauers [Kau05], pages 208–215. ISBN 1-59593-095-7. LCCN ???? ACM Order Number 505050.

Kaltofen:2006:FSD

[KK06] Erich Kaltofen and Pascal Koiran. Finding small degree factors of multivariate supersparse (lacunary) polynomials over algebraic number fields. In Trager [Tra06], pages 162–168. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

Kaltofen:2008:EFT

[KK08] Erich Kaltofen and Pascal Koiran. Expressing a fraction of two determinants as a determinant. In Jeffrey [Jef08], pages 141– 146. ISBN 1-59593-904-0. LCCN ????

Kaltofen:1986:FPA

[KKS86] E. Kaltofen, M. Krishnamoorthy, and B. D. Saunders. Fast parallel algorithms for similarity of matrices. In Char [Cha86b], pages 65–70. ISBN 0-89791-199-7 (paperback). LCCN

QA155.7.E4 A281 1986. URL http://www.acm.org:80/
pubs/citations/proceedings/issac/32439/p65-kaltofen/

. ACM order no. 505860.

Kaltofen:1994:FHP

[KL94] Erich Kaltofen and Austin Lobo. Factoring high-degree polynomials by the black box Berlekamp algorithm. In ACM [ACM94], pages 90–98. ISBN 0-89791-638-7. LCCN

QA76.95.I59 1994. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/190347/p90-kaltofen/.

Kaltofen:1996:RPT

[KL96a] E. Kaltofen and A. Lobo. On rank properties of Toeplitz matrices over finite fields. In Lakshman Y. N. [Lak96], pages 241–249. ISBN 0-89791-796-0. LCCN QA 76.95

I59 1996. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/236869/p241-kaltofen/.

Karmarkar:1996:APG

[KL96b] N. Karmarkar and Lakshman Y. N. Approximate polynomial greatest common divisors and nearest singular polynomials. In Lakshman Y. N. [Lak96], pages 35–39. ISBN 0-89791-796-0. LCCN QA 76.95 I59 1996. URL http://www.

acm.org:80/pubs/citations/proceedings/issac/236869/
p35-karmarkar/.

Kaltofen:1990:MRS

[KLW90] E. Kaltofen, Lakshman Y. N., and J.-M. Wiley. Modular rational sparse multivariate polynomial interpolation. In Watanabe and Nagata [WN90], pages 135–139. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95

.I57 1990. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/96877/p135-kaltofen/.

Kaltofen:2008:ECG

[KLYZ08] Erich Kaltofen, Bin Li, Zhengfeng Yang, and Lihong Zhi. Exact certification of global optimality of approximate factorizations via rationalizing sums-of-squares with floating point scalars. In Jeffrey [Jef08], pages 155–164. ISBN 1-59593-904-0. LCCN ????

Kuhn:1989:MEC

[KM89] N. Kuhn and K. Madlener. A method for enumerating cosets of a group presented by a canonical system. In Gonnet [Gon89], pages 338–350. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989.

US$29.00. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/74540/p338-kuhn/. ACM order number:

505890. English and French.

Krishnan:1995:NAE

[KM95] Shankar Krishnan and Dinesh Manocha. Numeric-symbolic algorithms for evaluating one-dimensional algebraic sets. In Levelt [Lev95], pages 59–67. ISBN 0-89791-699-9. LCCN QA 76.95

I59 1995. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/220346/p59-krishnan/. ACM order

number: 505950.

Koenhagen:1996:OAC

[KM96a] Ulla Koenhagen and Ernst W. Mayr. An optimal algorithm for constructing the reduced Gröbner basis of binomial ideals. In Lakshman Y. N. [Lak96], pages 55–62. ISBN 0-89791-796-0. LCCN QA 76.95 I59 1996. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/236869/p55-
koppenhagen/.

Koppenhagen:1996:OAC

[KM96b] U. Koppenhagen and E. W. Mayr. An optimal algorithm for constructing the reduced Gröbner basis of binomial ideals. In Lakshman Y. N. [Lak96], pages 55–62. ISBN 0-89791-796-0. LCCN QA 76.95 I59 1996.

Kuhnle:1996:ESC

[KM96c] Klaus Kühnle and Ernst W. Mayr. Exponential space computation of Gröbner bases. In Lakshman Y. N. [Lak96], pages 63–71. ISBN 0-89791-796-0. LCCN QA 76.95

I59 1996. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/236869/p63-kuhnle/.

Kaltofen:1999:GMP

[KM99] Erich Kaltofen and Michael B. Monagan. On the genericity of the modular polynomial GCD algorithm. In Dooley [Doo99], pages 59–66. ISBN 1-58113-073-2. LCCN QA76.95

.I57 1999.  URL http://www.acm.org/pubs/contents/
proceedings/issac/309831/.

Kogan:2002:CCF

[KM02] Irina A. Kogan and Marc Moreno Maza. Computation of canonical forms for ternary cubics. In Mora [Mor02], pages 151–160. ISBN 1-58113-484-3. LCCN QA76.95. URL http://www.lifl.

fr/ISSAC2002/.

Kaltofen:2003:AIP

[KM03] Erich Kaltofen and John May. On approximate irreducibility of polynomials in several variables. In Senda [Sen03], pages 161– 168. ISBN 1-58113-641-2. LCCN QA76.95. ACM order number 505030.

Khodadad:2006:FRF

[KM06] Sara Khodadad and Michael Monagan. Fast rational function reconstruction. In Trager [Tra06], pages 184–190. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

Kavian:1996:MPR

[KMG96] Masoud Kavian, R. G. McLenaghan, and K. O. Geddes. MapleTensor: Progress report on a new system for performing indicial and component tensor calculations using symbolic computation. In Lakshman Y. N. [Lak96], pages 204–211. ISBN 0-89791-796-0. LCCN QA 76.95 I59 1996. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/236869/p204-
kavian/.

kavian:1997:AGA

[kMG97] M. kavian, R. G. McLenaghan, and K. O. Geddes. Application of genetic algorithms to the algebraic simplification of tensor polynomials. In Küchlin [Küc97], pages 93–100. ISBN 0-89791-

875-4. LCCN QA76.95. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/258726/p93-kavian/.

Kobayashi:1989:SSA

[KMH89] H. Kobayashi, S. Moritsugu, and R. W. Hogan. Solving systems of algebraic equations. In Gianni [Gia89], pages 139–149. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Kuhn:1990:TLC

[KMO90] N. Kuhn, K. Madlener, and F. Otto. A test for λconfluence for certain prefix rewriting systems with applications to the generalized word problem. In Watanabe and Nagata [WN90], pages 8–15. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95

.I57 1990. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/96877/p8-kuhn/.

Kuhn:1992:CPS

[KMO92] Norbert Kuhn, Klaus Madlener, and Friedrich Otto. Computing presentations for subgroups of context-free groups. In Wang [Wan92b], pages 240–250. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN QA76.95.I59 1992. URL http:

//www.acm.org:80/pubs/citations/proceedings/issac/143242/

p240-kuhn/. ACM order number: 505920.

Kaltofen:2005:GMM

[KMY05] Erich Kaltofen, Dmitriy Morozov, and George Yuhasz. Generic matrix multiplication and memory management in linBox. In Kauers [Kau05], pages 216–223. ISBN 1-59593-095-7. LCCN ???? ACM Order Number 505050.

Kislenkov:1998:MCR

[KMZ98] V. Kislenkov, V. Mitrofanov, and E. Zima. Multidimensional chains of recurrences. In Gloor [Glo98], pages 199–206. ISBN 1-58113-002-3. LCCN QA155.7.E4 E88 1998. URL http://

www.acm.org/pubs/citations/proceedings/issac/281508/
p199-kislenkov/.

Kislenkov:1999:HFC

[KMZ99] V. Kislenkov, V. Mitrofanov, and E. Zima. How fast can we compute products? In Dooley [Doo99], pages 75–82. ISBN 1-58113-073-2. LCCN QA76.95 .I57 1999. URL http://www.

acm.org/pubs/contents/proceedings/issac/309831/.

Kurata:2007:CDC

[KN07] Yosuke Kurata and Masayuki Noro. Computation of discrete comprehensive Gröbner bases using modular dynamic evaluation. In Brown [Bro07], pages 243–250. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Kaltofen:2011:SBB

[KN11] Erich L. Kaltofen and Michael Nehring. Supersparse black box rational function interpolation. In Schost and Emiris [SE11], pages 177–186. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Konno:1994:PMC

[KNKM94] K. Konno, M. Nagatsuka, N. Kobayashi, and S. Matsuoka. PARCS: An MPP-oriented CLP language. In Hong [Hon94], pages 254–263. ISBN 981-02-2040-5. LCCN QA76.642.I58 1994.

Knowles:1986:ILF

[Kno86] P. H. Knowles. Integration of Liouvillian functions with special functions. In Char [Cha86b], pages 179–184. ISBN 0-89791-199-7 (paperback). LCCN QA155.7.E4 A281 1986. URL http://

www.acm.org:80/pubs/citations/proceedings/issac/32439/

p179-knowles/. ACM order no. 505860.

Kaltofen:2011:QTC

[KNS11] Erich L. Kaltofen, Michael Nehring, and B. David Saunders. Quadratic-time certificates in linear algebra. In Schost and Emiris [SE11], pages 171–176. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Koepf:2003:PSB

[Koe03] Wolfram Koepf. Power series, Bieberbach conjecture and the de Branges and Weinstein functions. In Senda [Sen03], pages 169–175. ISBN 1-58113-641-2. LCCN QA76.95. ACM order number 505030.

Kohno:1990:RPT

[Koh90] M. Kohno. Reduction problems in the theory of differential equations. In Watanabe and Nagata [WN90], pages 244–249. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95 .I57 1990. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/96877/p244-kohno/.

Koiran:2012:UBR

[Koi12] Pascal Koiran. Upper bounds on real roots and lower bounds for the permanent. In van der Hoeven and van Hoeij [vdHvH12], page 8. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Kolyada:1990:SSC

[Kol90] S. V. Kolyada. Systems for symbolic computations in Boolean algebra. In Watanabe and Nagata [WN90], page 291. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN

QA76.95 .I57 1990. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/96877/p291-kolyada/.

Korelc:2001:HSM

[Kor01] Joze Korelc. Hybrid system for multi-language and multienvironment generation of numerical codes. In Mourrain [Mou01], pages 209–216. ISBN 1-58113-417-7. LCCN QA76.95.I59 2001. ACM order number 505010.

Kakas:1994:PAL

[KP94a] A. C. Kakas and G. A. Papadopoulos. Parallel abduction in logic programming. In Hong [Hon94], pages 214–224. ISBN 981-02-2040-5. LCCN QA76.642.I58 1994.

Kaltofen:1994:PST

[KP94b] E. Kaltofen and V. Pan. Parallel solution of Toeplitz and Toeplitz-like linear systems over fields of small positive characteristic. In Hong [Hon94], pages 225–233. ISBN 981-02-2040-5. LCCN QA76.642.I58 1994.

Kauers:2010:WCW

[KP10] Manuel Kauers and Veronika Pillwein. When can we detect that a P -finite sequence is positive? In Watt [Wat10], pages 195–201. ISBN 1-4503-0150-9. LCCN QA76.95 .I59 2010.

Khonji:2010:OSD

[KPR^{+}10] Majid Khonji, Clément Pernet, Jean-Louis Roch, Thomas Roche, and Thomas Stalinski. Output-sensitive decoding for redundant residue systems. In Watt [Wat10], pages 265–272. ISBN 1-4503-0150-9. LCCN QA76.95 .I59 2010.

Kryukov:1986:URC

[KR86] A. P. Kryukov and A. Y. Rodionov. Usage of REDUCE for computations of group-theoretical weight of Feynman diagrams in non-Abelian gauge theories. In Char [Cha86b], pages 91– 93. ISBN 0-89791-199-7 (paperback). LCCN QA155.7.E4

A281 1986. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/32439/p91-kryukov/. ACM order no.

505860.

Keady:1993:AIS

[KR93] G. Keady and M. G. Richardson. An application of IRENA to systems of nonlinear equations arising in equilibrium flows in networks. In Bronstein [Bro93], pages 311–320. ISBN 0-89791-604-2. LCCN QA 76.95 I59 1993. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/164081/p311-

keady/. ACM order number: 505930.

Kadyrsizova:2008:LRS

[KR08] Zhibek Kadyrsizova and Valery G. Romanovski. Linearizablity of 1:−3 resonant system with homogeneous cubic nonlinearities. In Jeffrey [Jef08], pages 255–260. ISBN 1-59593-904-0. LCCN ????

Kredel:1989:SDC

[Kre89] H. Kredel. Software development for computer algebra or from ALDES/SAC-2 to WEB/Modula-2. In Gianni [Gia89], pages 447–455. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Kryukov:1986:CRA

[KRL86] A. P. Kryukov, Y. Rodionov, and G. L. Litvinov. Construction of rational approximations by means of REDUCE. In Char [Cha86b], pages 31–33. ISBN 0-89791-199-7 (paperback). LCCN QA155.7.E4 A281 1986. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/32439/p31-kryukov/.

ACM order no. 505860.

Klimov:1993:SEN

[KRL93] D. M. Klimov, V. M. Rudenko, and V. V. Leonov. Symbolic evaluation in the nonlinear mechanical systems. In Bronstein [Bro93], pages 53–54. ISBN 0-89791-604-2. LCCN QA 76.95

I59 1993. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/164081/p53-klimov/. ACM order num-

ber: 505930.

Kaser:1994:HPR

[KRS94] O. Kaser, C. R. Ramakrishnan, and R. C. Sekar. A high performance runtime system for parallel evaluation of lazy languages. In Hong [Hon94], pages 234–243. ISBN 981-02-2040-5. LCCN QA76.642.I58 1994.

Kryukov:1986:DRE

[Kry86] A. P. Kryukov. Dialogue in REDUCE: experience and development. In Char [Cha86b], pages 107–109. ISBN 0-89791-199-7 (paperback). LCCN QA155.7.E4 A281 1986. URL http://www.

acm.org:80/pubs/citations/proceedings/issac/32439/p107-

kryukov/. ACM order no. 505860.

Kutzler:1986:AGT

[KS86] B. Kutzler and S. Stifter. Automated geometry theorem proving using Buchberger’s algorithm. In Char [Cha86b], pages 209–214. ISBN 0-89791-199-7 (paperback). LCCN QA155.7.E4

A281 1986. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/32439/p209-kutzler/. ACM order no.

505860.

Kapur:1995:CVM

[KS95] Deepak Kapur and Tushar Saxena. Comparison of various multivariate resultant formulations. In Levelt [Lev95], pages 187–194. ISBN 0-89791-699-9. LCCN QA 76.95

I59 1995. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/220346/p187-kapur/. ACM order num-

ber: 505950.

Kaltofen:1997:FPF

[KS97a] Erich Kaltofen and Victor Shoup. Fast polynomial factorization over high algebraic extensions of finite fields. In Küchlin [Küc97], pages 184–188. ISBN 0-89791-875-4. LCCN

QA76.95. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/258726/p184-kaltofen/.

Kapur:1997:EFD

[KS97b] Deepak Kapur and Tushar Saxena. Extraneous factors in the Dixon resultant formulation. In Küchlin [Küc97], pages 141– 148. ISBN 0-89791-875-4. LCCN QA76.95. URL http://www.

acm.org:80/pubs/citations/proceedings/issac/258726/
p141-kapur/.

Kauers:2006:AUS

[KS06] Manuel Kauers and Carsten Schneider. Application of unspecified sequences in symbolic summation. In Trager [Tra06], pages 177–183. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

Kauers:2007:SSR

[KS07] Manuel Kauers and Carsten Schneider. Symbolic summation with radical expressions. In Brown [Bro07], pages 219–226. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Kauers:2010:PDB

[KS10] Manuel Kauers and Carsten Schneider. Partial denominator bounds for partial linear difference equations. In Watt [Wat10], pages 211–218. ISBN 1-4503-0150-9. LCCN QA76.95 .I59 2010.

Kaminski:2011:UDC

[KS11a] Jeremy-Yrmeyahu Kaminski and Yann Sepulcre. Using discriminant curves to recover a surface of P^{4} from two generic linear projections. In Schost and Emiris [SE11], pages 187–192. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Kauers:2011:RDB

[KS11b] Manuel Kauers and Carsten Schneider. A refined denominator bounding algorithm for multivariate linear difference equations. In Schost and Emiris [SE11], pages 201–208. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Kerber:2011:ERR

[KS11c] Michael Kerber and Michael Sagraloff. Efficient real root approximation. In Schost and Emiris [SE11], pages 209–216. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Kawano:2013:QFT

[KS13] Yasuhito Kawano and Hiroshi Sekigawa. Quantum Fourier Transform over symmetric groups. In Monagan et al. [MCG13], pages 227–234. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

Khetan:2004:SRB

[KSG04] Amit Khetan, Ning Song, and Ron Goldman. Sylvesterresultants for bivariate polynomials with planar Newton polygons. In Gutierrez [Gut04], pages 205–212. ISBN 1-58113-827-X. LCCN QA76.95 .I57 2004.

Kato:1997:PIC

[KSI97] Shohei Kato, Hirohisa Seki, and Hidenori Itoh. A parallel implementation of cost-based abductive reasoning. In Küchlin [Küc97], pages 111–118. ISBN 0-89791-875-4. LCCN

QA76.95. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/266670/p111-kato/.

Kaltofen:2000:ETB

[KsLL00] Erich Kaltofen, Wen shin Lee, and Austin A. Lobo. Early termination in Ben-Or/Tiwari sparse interpolation and a hybrid of Zippel’s algorithm. In Traverso [Tra00], pages 192–201. ISBN 1-58113-218-2. LCCN QA76.95.I59 2000. URL http://www.

acm.org/pubs/articles/proceedings/issac/345542/p192-
kaltofen/p192-kaltofen.pdf;   http://www.acm.org/pubs/
citations/proceedings/issac/345542/p192-kaltofen/. ACM

order number 505000.

Khungurn:2007:MCP

[KSS07] Pramook Khungurn, Hiroshi Sekigawa, and Kiyoshi Shirayanagi. Minimum converging precision of the QR-factorization algorithm for real polynomial GCD. In Brown [Bro07], pages 227–234. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Kapur:2010:NAC

[KSW10] Deepak Kapur, Yao Sun, and Dingkang Wang. A new algorithm for computing comprehensive Gröbner systems. In Watt [Wat10], pages 29–36. ISBN 1-4503-0150-9. LCCN QA76.95 .I59 2010.

Kapur:2011:CCG

[KSW11] Deepak Kapur, Yao Sun, and Dingkang Wang. Computing comprehensive Gröbner systems and comprehensive Gröbner bases simultaneously. In Schost and Emiris [SE11], pages 193–200. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Kapur:1994:AGR

[KSY94] Deepak Kapur, Tushar Saxena, and Lu Yang. Algebraic and geometric reasoning using Dixon resultants. In ACM [ACM94], pages 99–107. ISBN 0-89791-638-7. LCCN

QA76.95.I59 1994. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/190347/p99-kapur/.

Kuchlin:1991:MCI

[Küc91] Wolfgang Küchlin. On the multi-threaded computation of integral polynomial greatest common divisors. In Watt [Wat91], pages 333–342. ISBN 0-89791-437-6. LCCN QA 76.95

I59 1991. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/120694/p333-kuchlin/.

Kuchlin:1997:PPS

[Küc97] Wolfgang W. Küchlin, editor. ISSAC 97: July 21–23, 1997, Maui, Hawaii, USA: proceedings of the 1997 International Symposium on Symbolic and Algebraic Computation. ACM Press, New York, NY 10036, USA, 1997. ISBN 0-89791-875-4.

LCCN QA76.95. URL http://www.acm.org/pubs/contents/
proceedings/issac/258726/.

Kutzler:1989:CAT

[Kut89] B. Kutzler. Careful algebraic translations of geometry theorems. In Gonnet [Gon89], pages 254–263. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989. US$29.00. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/74540/p254-kutzler/

. ACM order number: 505890. English and French.

Kunwar:2013:SOD

[KvH13] Vijay Jung Kunwar and Mark van Hoeij. Second order differential equations with hypergeometric solutions of degree three. In Monagan et al. [MCG13], pages 235–242. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

Kaltofen:1989:IVP

[KVY89] E. Kaltofen, T. Valente, and N. Yui. An improved Las Vegas primality test. In Gonnet [Gon89], pages 26–33. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989.

US$29.00. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/74540/p26-kaltofen/. ACM order num-

ber: 505890. English and French.

Kapur:1990:RPG

[KW90] D. Kapur and H. K. Wan. Refutational proofs of geometry theorems via characteristic set computation. In Watanabe and Nagata [WN90], pages 277–284. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95

.I57 1990. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/96877/p277-kapur/.

Kaltofen:1989:ISM

[KY89] E. Kaltofen and L. Yagati. Improved sparse multivariate polynomial interpolation algorithms. In Gianni [Gia89], pages 467– 474. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Kaltofen:2007:EAI

[KY07] Erich Kaltofen and Zhengfeng Yang. On exact and approximate interpolation of sparse rational functions. In Brown [Bro07], pages 203–210. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Kaltofen:2013:SMF

[KY13] Erich L. Kaltofen and Zhengfeng Yang. Sparse multivariate function recovery from values with noise and outlier errors. In Monagan et al. [MCG13], pages 219–226. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

Kanno:2007:POC

[KYAH07] Masaaki Kanno, Kazuhiro Yokoyama, Hirokazu Anai, and Shinji Hara. Parametric optimization in control using the sum of roots for parametric polynomial spectral factorization. In Brown [Bro07], pages 211–218. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Kanno:2008:SOA

[KYAH08] Masaaki Kanno, Kazuhiro Yokoyama, Hirokazu Anai, and Shinji Hara. Symbolic optimization of algebraic functions. In Jeffrey [Jef08], pages 147–154. ISBN 1-59593-904-0. LCCN ????

Kanno:2009:SAR

[KYAH09] Masaaki Kanno, Kazuhiro Yokoyama, Hirokazu Anai, and Shinji Hara. Solution of algebraic Riccati equations using the sum of roots. In May [May09], pages 215–222. ISBN 1-60558-609-9. LCCN ????

Kaltofen:2006:AGC

[KYZ06] Erich Kaltofen, Zhengfeng Yang, and Lihong Zhi. Approximate greatest common divisors of several polynomials with linearly constrained coefficients and singular polynomials. In Trager [Tra06], pages 169–176. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

Kaltofen:2006:HSN

[KZ06] Erich Kaltofen and Lihong Zhi. Hybrid symbolic-numeric computation. In Trager [Tra06], page 7. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

LakshmanYN:1996:IPI

[Lak96] Lakshman Y. N., editor. ISSAC ’96: Proceedings of the 1996 International Symposium on Symbolic and Algebraic Computation, July 24–26, 1996, Zurich, Switzerland. ACM Press, New York, NY 10036, USA, 1996. ISBN 0-89791-796-0. LCCN QA 76.95 I59 1996.

Lam:1997:MPP

[Lam97] Monica Lam. Maximizing performance on parallel machines (abstract). In Küchlin [Küc97], page 129. ISBN 0-89791-

875-4. LCCN QA76.95. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/266670/p129-lam/.

Langemyr:1991:ASA

[Lan91] Lars Langemyr. An analysis of the subresultant algorithm over an algebraic number field. In Watt [Wat91], pages 167–172. ISBN 0-89791-437-6. LCCN QA 76.95

I59 1991. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/120694/p167-langemyr/.

Laplagne:2006:ACR

[Lap06] Santiago Laplagne. An algorithm for the computation of the radical of an ideal. In Trager [Tra06], pages 191–195. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

Laubenbacher:2003:CAA

[Lau03] Reinhard Laubenbacher. A computer algebra approach to biological systems. In Senda [Sen03], pages 5–6. ISBN 1-58113-641-2. LCCN QA76.95. ACM order number 505030.

Lazard:2006:SKC

[Laz06] Daniel Lazard. Solving Kaltofen’s challenge on Zolotarev’s approximation problem. In Trager [Tra06], pages 196–203. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

Leung:1994:CSD

[LC94] H.-F. Leung and K. L. Clark. Constraint solving in distributed concurrent logic programming. In Hong [Hon94], pages 273–283. ISBN 981-02-2040-5. LCCN QA76.642.I58 1994.

Lakshman:1998:SCU

[LCJ98] Y. N. Lakshman, Bruce Char, and Jeremy Johnson. Software components using symbolic computation for problem solving environments. In Gloor [Glo98], pages 46–53. ISBN 1-58113-002-3. LCCN QA155.7.E4 E88 1998. URL http://www.

acm.org/pubs/citations/proceedings/issac/281508/p46-
lakshman/.

Li:2005:EAS

[LCW05] Biao Li, Yong Chen, and Qi Wang. Exact analytical solutions to the nonlinear Schrödinger equation model. In Kauers [Kau05], pages 224–230. ISBN 1-59593-095-7. LCCN ???? ACM Order Number 505050.

Liska:1990:FRP

[LD90] R. Liska and L. Drska. FIDE: a REDUCE package for automation of FInite difference method for solving pDE. In Watanabe and Nagata [WN90], pages 169–176. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95

.I57 1990. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/96877/p169-liska/.

Li:2011:APF

[LD11] Yue Li and Gabriel Dos Reis. An automatic parallelization framework for algebraic computation systems. In Schost and Emiris [SE11], pages 233–240. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Le:2002:SDS

[Le02] Ha Le. Simplification of definite sums of rational functions by creative symmetrizing method. In Mora [Mor02], pages 161–

167. ISBN 1-58113-484-3. LCCN QA76.95. URL http://www.

lifl.fr/ISSAC2002/.

Lecerf:2000:CED

[Lec00] Grégoire Lecerf. Computing an equidimensional decomposition of an algebraic variety by means of geometric resolutions. In Traverso [Tra00], pages 209–216. ISBN 1-58113-218-2.

LCCN QA76.95.I59 2000. URL http://www.acm.org/pubs/
articles/proceedings/issac/345542/p209-lecerf/p209-
lecerf.pdf;   http://www.acm.org/pubs/citations/proceedings/
issac/345542/p209-lecerf/. ACM order number 505000.

Lehobey:1997:RCR

[Leh97] Frédéric Lehobey. Resolvent computations by resultants without extraneous powers. In Küchlin [Küc97], pages 85–92. ISBN 0-89791-875-4. LCCN QA76.95. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/258726/p85-
lehobey/.

Lella:2012:EIA

[Lel12] Paolo Lella. An efficient implementation of the algorithm computing the Borel-fixed points of a Hilbert scheme. In van der Hoeven and van Hoeij [vdHvH12], pages 242–248. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Lenstra:2013:LS

[Len13] Hendrik Lenstra. Lattices with symmetry. In Monagan et al. [MCG13], pages 3–4. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

Leong:1986:IDU

[Leo86] B. L. Leong. Iris: design of an user interface program for symbolic algebra. In Char [Cha86b], pages 1– 6. ISBN 0-89791-199-7 (paperback). LCCN QA155.7.E4

A281 1986. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/32439/p1-leong/.  ACM order no.

505860.

Levelt:1995:IPI

[Lev95] A. H. M. Levelt, editor. ISSAC ’95: Proceedings of the 1995 International Symposium on Symbolic and Algebraic Computation: July 10–12, 1995, Montréal, Canada, ISSAC - PROCEEDINGS- 1995. ACM Press, New York, NY 10036, USA, 1995. ISBN 0-89791-699-9. LCCN QA 76.95 I59 1995. ACM order number: 505950.

Levandovskyy:2006:IIN

[Lev06] Viktor Levandovskyy. Intersection of ideals with noncommutative subalgebras. In Trager [Tra06], pages 212–219. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

Levin:2007:GBR

[Lev07] Alexander B. Levin. Gröbner bases with respect to several term orderings and multivariate dimension polynomials. In Brown [Bro07], pages 251–260. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Levandovskyy:2012:ECA

[Lev12] Viktor Levandovskyy. Elements of computer-algebraic analysis. In van der Hoeven and van Hoeij [vdHvH12], pages 9–10. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Levin:2013:MDD

[Lev13] Alexander B. Levin. Multivariate difference-differential dimension polynomials and new invariants of difference-differential field extensions. In Monagan et al. [MCG13], pages 267–274. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

Leykin:2008:NPD

[Ley08] Anton Leykin. Numerical primary decomposition. In Jeffrey [Jef08], pages 165–172. ISBN 1-59593-904-0. LCCN ????

Liao:1995:EHP

[LF95]  Hsin-Chao Liao and Richard J. Fateman.  Evaluation

of the heuristic polynomial GCD. In Levelt [Lev95], pages 240–247. ISBN 0-89791-699-9. LCCN QA 76.95

I59 1995. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/220346/p240-liao/. ACM order num-

ber: 505950.

Lengauer:1997:EPP

[LGG97] Christian Lengauer, Martin Griebl, and Sergei Gorlatch, editors. Euro-Par’97, parallel processing: third International Euro-Par Conference, Passau, Germany, August 26–29, 1997: proceedings, volume 1300 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 3-540-63440-1 (paperback). LCCN QA76.58.I5535 1997.

Li:2011:SDR

[LGY11] Wei Li, Xiao-Shan Gao, and Cum-Ming Yuan. Sparse differential resultant. In Schost and Emiris [SE11], pages 225–232. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Le:1999:CSC

[LH99] Ha Le and Chris Howlett. Client-server communication standards for mathematical computation. In Dooley [Doo99], pages 299–306. ISBN 1-58113-073-2. LCCN QA76.95 .I57 1999. URL

http://www.acm.org/pubs/contents/proceedings/issac/
309831/.

Li:2008:CBB

[LH08] Hongbo Li and Lei Huang. Complex brackets, balanced complex differences, and applications in symbolic geometric computing.

In Jeffrey [Jef08], pages 181–188. ISBN 1-59593-904-0. LCCN ????

LopezGarcia:1994:TGB

[LHD94] P. Lopez Garcia, M. Hermenegildo, and S. K. Debray. Towards granularity based control of parallelism in logic programs. In Hong [Hon94], pages 133–144. ISBN 981-02-2040-5. LCCN QA76.642.I58 1994.

Lamagna:1992:DUI

[LHJ92] Edmund A. Lamagna, Michael B. Hayden, and Catherine W. Johnson. The design of a user interface to a computer algebra system for introductory calculus. In Wang [Wan92b], pages 358–368. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN QA76.95.I59 1992. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/143242/p358-

lamagna/. ACM order number: 505920.

Linton:2010:ECS

[LHK^{+}10] S. Linton, K. Hammond, A. Konovalov, A. D. Al Zain, P. Trinder, P. Horn, and D. Roozemond. Easy composition of symbolic computation software: a new lingua franca for symbolic computation. In Watt [Wat10], pages 339–346. ISBN 1-4503-0150-9. LCCN QA76.95 .I59 2010.

Li:1998:STO

[Li98] Ziming Li. A subresultant theory for Øre polynomials with applications. In Gloor [Glo98], pages 132–139. ISBN 1-58113-002-3. LCCN QA155.7.E4 E88 1998. URL http://www.acm.org/

pubs/citations/proceedings/issac/281508/p132-li/.

Lisonek:2000:MIT

[LI00] Petr Lisonĕk and Robert B. Israel. Metric invariants of tetrahedra via polynomial elimination. In Traverso [Tra00], pages 217–219. ISBN 1-58113-218-2. LCCN QA76.95.I59 2000. URL

http://www.acm.org/pubs/articles/proceedings/issac/
345542/p217-lisonek/p217-lisonek.pdf;   http://www.
acm.org/pubs/citations/proceedings/issac/345542/p217-

lisonek/. ACM order number 505000.

Li:2004:SCH

[Li04] Hongbo Li. Symbolic computation in the homogeneous geometric model with Clifford algebra. In Gutierrez [Gut04], pages 221–228. ISBN 1-58113-827-X. LCCN QA76.95 .I57 2004.

Li:2007:RSG

[Li07] Hongbo Li. A recipe for symbolic geometric computing: long geometric product, BREEFS and Clifford factorization. In Brown [Bro07], pages 261–268. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Lichtblau:2005:HGF

[Lic05] Daniel Lichtblau. Half-GCD and fast rational recovery. In Kauers [Kau05], pages 231–236. ISBN 1-59593-095-7. LCCN ???? ACM Order Number 505050.

Linton:2004:FSI

[Lin04] Steve Linton. Finding the smallest image of a set. In Gutierrez [Gut04], pages 229–234. ISBN 1-58113-827-X. LCCN QA76.95 .I57 2004.

Liang:2008:CRC

[LJM08] Songxin Liang, David J. Jeffrey, and Marc Moreno Maza. The complete root classification of a parametric polynomial on an interval. In Jeffrey [Jef08], pages 189–196. ISBN 1-59593-904-0. LCCN ????

Letichevsky:1990:APA

[LK90] A. A. Letichevsky and J. V. Kapitonova. Algebraic programming in the APS system. In Watanabe and Nagata [WN90], pages 68–75. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95 .I57 1990. URL http://

www.acm.org:80/pubs/citations/proceedings/issac/96877/
p68-letichevsky/.

Li:1997:MGB

[LkGID97] Qiang Li, Yi ke Guo, Tetsuo Ida, and John Darlington. The minimised geometric Buchberger algorithm: an optimal algebraic algorithm for integer programming. In Küchlin [Küc97], pages 331–338. ISBN 0-89791-875-4. LCCN

QA76.95. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/258726/p331-li/.

Letichevsky:1991:APO

[LKK91] A. A. Letichevsky, J. V. Kapitonova, and S. V. Konozenko. Algebraic programs optimization. In Watt [Wat91], pages 370– 376. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991. URL

http://www.acm.org:80/pubs/citations/proceedings/issac/
120694/p370-letichevsky/.

Lin:1993:SRT

[LL93] Dongdai Lin and Zhuojun Liu. Some results on theorem proving in geometry over finite fields. In Bronstein [Bro93], pages 292–300. ISBN 0-89791-604-2. LCCN QA 76.95

I59 1993. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/164081/p292-lin/. ACM order number:

505930.

Labahn:2004:HSF

[LL04] George Labahn and Ziming Li. Hyperexponential solutions of finite-rank ideals in orthogonal Øre rings. In Gutierrez [Gut04], pages 213–220. ISBN 1-58113-827-X. LCCN QA76.95 .I57 2004.

Loera:2008:HNA

[LLMM08] J. A. De Loera, J. Lee, P. N. Malkin, and S. Margulies. Hilbert’s Nullstellensatz and an algorithm for proving combinatorial infeasibility. In Jeffrey [Jef08], pages 197–206. ISBN 1-59593-904-0. LCCN ????

Licciardi:1994:IHC

[LM94] Sandra Licciardi and Teo Mora. Implicitization of hypersurfaces and curves by the Primbasissatz and basis conversion. In ACM [ACM94], pages 191–196. ISBN 0-89791-638-7. LCCN QA76.95.I59 1994. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/190347/p191-
licciardi/.

Luks:2002:PTN

[LM02] Eugene M. Luks and Takunari Miyazaki. Polynomial-time normalizers for permutation groups with restricted composition factors. In Mora [Mor02], pages 176–183. ISBN 1-58113-484-3.

LCCN QA76.95. URL http://www.lifl.fr/ISSAC2002/.

Levandovskyy:2008:CMT

[LM08] Viktor Levandovskyy and Jorge Martin Morales. Computational D-module theory with singular, comparison with other systems and two new algorithms. In Jeffrey [Jef08], pages 173– 180. ISBN 1-59593-904-0. LCCN ????

Lamban:2013:CSM

[LMMRRR13] Laureano Lambán, Francisco J. Martín-Mateos, Julio Rubio, and José-Luis Ruiz-Reina. Certified symbolic manipulation: bivariate simplicial polynomials. In Monagan et al. [MCG13], pages 243–250. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

Li:2009:CMR

[LMP09] Xin Li, Marc Moreno Maza, and Wei Pan. Computations modulo regular chains. In May [May09], pages 239–246. ISBN 1-60558-609-9. LCCN ????

Lemaire:2008:WDE

[LMPX08] Francois Lemaire, Marc Moreno Maza, Wei Pan, and Yuzhen Xie. When does (T) equal Sat(T)? In Jeffrey [Jef08], pages 207–214. ISBN 1-59593-904-0. LCCN ????

Li:2007:FAT

[LMS07] Xin Li, Marc Moreno Maza, and Éric Schost. Fast arithmetic for triangular sets: from theory to practice. In Brown [Bro07], pages 269–276. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Lebreton:2013:CSB

[LMS13] Romain Lebreton, Esmaeil Mehrabi, and Eric Schost. On the complexity of solving bivariate systems: the case of non-singular solutions. In Monagan et al. [MCG13], pages 251–258. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

Li:1997:MAC

[LN97] Ziming Li and István Nemes. A modular algorithm for computing greatest common right divisors of Øre polynomials. In Küchlin [Küc97], pages 282–289. ISBN 0-89791-875-4. LCCN

QA76.95. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/258726/p282-li/.

Lamban:1999:SI

[LPR99] Laureano Lambán, Vico Pascual, and Julio Rubio. Specifying implementations. In Dooley [Doo99], pages 245–251. ISBN 1-58113-073-2. LCCN QA76.95 .I57 1999. URL http://www.

acm.org/pubs/contents/proceedings/issac/309831/.

Lisle:1995:ADS

[LRB95] I. G. Lisle, G. J. Reid, and A. Boulton. Algorithmic determination of structure of infinite Lie pseudogroups of symmetries of PDEs. In Levelt [Lev95], pages 1–6. ISBN 0-89791-699-9. LCCN

QA 76.95 I59 1995. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/220346/p1-lisle/. ACM or-

der number: 505950.

Luks:1994:CNP

[LRW94] Eugene M. Luks, Ferenc Rákóczi, and Charles R. B. Wright. Computing normalizers in permutation p-groups. In ACM [ACM94], pages 139–146. ISBN 0-89791-638-7. LCCN

QA76.95.I59 1994. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/190347/p139-luks/.

Llovet:1990:MAC

[LS90] J. Llovet and J. R. Sendra. A modular approach to the computation of the number of real roots. In Watanabe and Nagata [WN90], page 298. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95

.I57 1990. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/96877/p298-llovet/.

Lempken:1992:SPS

[LS92] W. Lempken and R. Staszewski. The structure of the PIMs of SL(3,4) in characteristic 2. In Wang [Wan92b], pages 233– 239. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN QA76.95.I59 1992. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/143242/p233-

lempken/. ACM order number: 505920.

LakshmanYN:1994:CSS

[LS94] Lakshman Y. N. and B. David Saunders. On computing sparse shifts for univariate polynomials. In ACM [ACM94], pages 108–113. ISBN 0-89791-638-7. LCCN

QA76.95.I59 1994. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/190347/p108-lakshman/.

Levandovskyy:2003:PCA

[LS03] Viktor Levandovskyy and Hans Schönemann. Plural: a computer algebra system for noncommutative polynomial algebras. In Senda [Sen03], pages 176–183. ISBN 1-58113-641-2. LCCN QA76.95. ACM order number 505030.

Lebreton:2012:AUD

[LS12] Romain Lebreton and Éric Schost. Algorithms for the universal decomposition algebra. In van der Hoeven and van Hoeij [vdHvH12], pages 234–241. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Levandovskyy:2013:ECG

[LSS13] Viktor Levandovskyy, Grischa Studzinski, and Benjamin Schnitzler. Enhanced computations of Gröbner bases in free algebras as a new application of the letterplace paradigm. In Monagan et al. [MCG13], pages 259–266. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

Li:2002:FZD

[LST02] Ziming Li, Fritz Schwarz, and Serguei P. Tsarev. Factoring zero-dimensional ideals of linear partial differential operators. In Mora [Mor02], pages 168–175. ISBN 1-58113-484-3. LCCN

QA76.95. URL http://www.lifl.fr/ISSAC2002/.

Landsmann:2000:SPP

[LSWH00] Günter Landsmann, Josef Schicho, Franz Winkler, and Erik Hillgarter. Symbolic parametrization of pipe and canal surfaces. In Traverso [Tra00], pages 202–208. ISBN 1-58113-218-2. LCCN QA76.95.I59 2000. URL http://

www.acm.org/pubs/citations/proceedings/issac/345542/

p202-landsmann/. ACM order number 505000.

Li:2006:RMD

[LSWZ06] Ziming Li, Michael F. Singer, Min Wu, and Dabin Zheng. A recursive method for determining the one-dimensional submodules of Laurent–Øre modules. In Trager [Tra06], pages 220–227. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

Li:2011:ARS

[LSYL11] Hongbo Li, Ruiyong Sun, Shoubin Yao, and Ge Li. Approximate rational solutions torational ODEs defined on discrete differentiable curves. In Schost and Emiris [SE11], pages 217–224. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Lemaire:2010:MSR

[LÜ10] François Lemaire and Asli Ürgüplü. A method for semirectifying algebraic and differential systems using scaling type Lie point symmetries with linear algebra. In Watt [Wat10], pages 85–92. ISBN 1-4503-0150-9. LCCN QA76.95 .I59 2010.

Lucks:1986:FIP

[Luc86] Michael Lucks. A fast implementation of polynomial factorization. In Char [Cha86b], pages 228–232. ISBN 0-89791-199-7 (paperback). LCCN QA155.7.E4 A281 1986. URL http://www.

acm.org:80/pubs/citations/proceedings/issac/32439/p228-

lucks/. ACM order no. 505860.

Liska:1991:ADS

[LuS91] Richard Liska and Michail Yu. u. Shashkov. Algorithms for difference schemes construction on non-orthogonal logically rectangular meshes. In Watt [Wat91], pages 419–426. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/120694/p419-
liska/.

Leff:1986:CSG

[LY86] L. Leff and D. Y. Y. Yun. Constructive solid geometry: a symbolic computation approach. In Char [Cha86b], pages 121–126. ISBN 0-89791-199-7 (paperback). LCCN

QA155.7.E4 A281 1986. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/32439/p121-leff/. ACM or-

der no. 505860.

Li:2013:SDR

[LYG13] Wei Li, Chun-Ming Yuan, and Xiao-Shan Gao. Sparse difference resultant. In Monagan et al. [MCG13], pages 275–282. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

Li:2010:BID

[LYZ10] Zijia Li, Zhengfeng Yang, and Lihong Zhi. Blind image deconvolution via fast approximate GCD. In Watt [Wat10], pages 155–162. ISBN 1-4503-0150-9. LCCN QA76.95 .I59 2010.

MacCallum:1989:ODE

[Mac89] M. A. H. MacCallum. An ordinary differential equation solver for REDUCE. In Gianni [Gia89], pages 196–205. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Macutan:1999:FSS

[Mac99] Y. O. Macutan. Formal solutions of scalar singularly-perturbed linear differential equations. In Dooley [Doo99], pages 113–120. ISBN 1-58113-073-2. LCCN QA76.95 .I57 1999. URL http://

www.acm.org/pubs/contents/proceedings/issac/309831/.

Maignan:1998:SOT

[Mai98] Aude Maignan. Solving one and two-dimensional exponential polynomial systems. In Gloor [Glo98], pages 215–221. ISBN 1-58113-002-3. LCCN QA155.7.E4 E88 1998. URL http://

www.acm.org/pubs/citations/proceedings/issac/281508/
p215-maignan/.

Manocha:1990:RCP

[Man90] D. Manocha. Regular curves and proper parametrizations. In Watanabe and Nagata [WN90], pages 271–276. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN

QA76.95 .I57 1990. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/96877/p271-manocha/.

Mandache:1994:GBA

[Man94a] Ana Maria Mandache. The Gröbner basis algorithm and subresultant theory. In ACM [ACM94], pages 123–128. ISBN 0-89791-638-7. LCCN QA76.95.I59 1994. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/190347/p123-
mandache/.

Manocha:1994:CSS

[Man94b] Dinesh Manocha. Computing selected solutions of polynomial equations. In ACM [ACM94], pages 1–8. ISBN 0-89791-638-7. LCCN QA76.95.I59 1994. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/190347/p1-manocha/.

Mansfield:2008:DAD

[Man08] Elizabeth L. Mansfield. Digital atlases and difference forms. In Jeffrey [Jef08], pages 3–4. ISBN 1-59593-904-0. LCCN ????

Marzinkewitsch:1991:OCA

[Mar91] Reiner Marzinkewitsch. Operating computer algebra systems by handprinted input. In Watt [Wat91], pages 411–413. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/120694/p411-
marzinkewitsch/.

Marje:1995:NLA

[Mar95] Prabhav Marje. A nearly linear algorithm for Sylow subgroups in small-base groups. In Levelt [Lev95], pages 270–277. ISBN 0-89791-699-9. LCCN QA 76.95 I59 1995. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/220346/p270-

marje/. ACM order number: 505950.

Mawata:1986:SDR

[Maw86] C. P. Mawata. A sparse distributed representation using prime numbers. In Char [Cha86b], pages 110–114. ISBN 0-89791-199-7 (paperback). LCCN QA155.7.E4 A281 1986. URL http://www.

acm.org:80/pubs/citations/proceedings/issac/32439/p110-

mawata/. ACM order no. 505860.

May:2009:PIS

[May09] John P. May, editor. Proceedings of the 2009 international symposium on Symbolic and algebraic computation, KIAS, Seoul, Korea, July 28–31, 2009. ACM Press, New York, NY 10036, USA, 2009. ISBN 1-60558-609-9. LCCN ????

Maza:2006:TDP

[Maz06] Marc Moreno Maza. Triangular decompositions of polynomial systems: from theory to practice. In Trager [Tra06], page 8. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

McCallum:2009:DVC

[MB09] Scott McCallum and Christopher W. Brown. On delineability of varieties in CAD-based quantifier elimination with two equational constraints. In May [May09], pages 71–78. ISBN 1-60558-609-9. LCCN ????

Mutrie:1992:AFE

[MBC92] Mark P. W. Mutrie, Richard H. Bartels, and Bruce W. Char. An approach for floating-point error analysis using computer algebra. In Wang [Wan92b], pages 284–293. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN QA76.95.I59 1992. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/143242/p284-mutrie/

. ACM order number: 505920.

Manocha:1991:ETM

[MC91] Dinesh Manocha and John Canny. Efficient techniques for multipolynomial resultant algorithms. In Watt [Wat91], pages 86–95. ISBN 0-89791-437-6. LCCN QA 76.95

I59 1991. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/120694/p86-manocha/.

Manocha:1992:MRL

[MC92] Dinesh Manocha and John F. Canny. Multipolynomial resultants and linear algebra. In Wang [Wan92b], pages 158– 167. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN QA76.95.I59 1992. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/143242/p158-

manocha/. ACM order number: 505920.

McCallum:1999:PCB

[McC99] Scott McCallum. On projection in CAD-based quantifier elimination with equational constraint. In Dooley [Doo99], pages 145–149. ISBN 1-58113-073-2. LCCN QA76.95 .I57 1999. URL

http://www.acm.org/pubs/contents/proceedings/issac/
309831/.

McCallum:2001:PEC

[McC01] Scott McCallum. On propagation of equational constraints in CAD-based quantifier elimination. In Mourrain [Mou01], pages 223–231. ISBN 1-58113-417-7. LCCN QA76.95.I59 2001. ACM order number 505010.

McCallum:2003:OIB

[McC03] Scott McCallum. On order-invariance of a binomial over a nullifying cell. In Senda [Sen03], pages 184–190. ISBN 1-58113-641-2. LCCN QA76.95. ACM order number 505030.

McCarron:2012:SHQ

[McC12] James McCarron. Small homogeneous quandles. In van der Hoeven and van Hoeij [vdHvH12], pages 257–264. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Monagan:2013:IPI

[MCG13] Michael Monagan, Gene Cooperman, and Mark Giesbrecht, editors. ISSAC 2013: Proceedings of the 2013 International Symposium on Symbolic and Algebraic Computation, June 26–29, 2013, Boston, MA, USA. ACM Press, New York, NY 10036, USA, 2013. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

Mezzarobba:2010:NPN

[Mez10] Marc Mezzarobba. NumGfun: a package for numerical and analytic computation with D-finite functions. In Watt [Wat10], pages 139–145. ISBN 1-4503-0150-9. LCCN QA76.95 .I59 2010.

Murao:1994:MAS

[MF94] H. Murao and T. Fujise. Modular algorithm for sparse multivariate polynomial interpolation and its parallel implementation. In Hong [Hon94], pages 304–315. ISBN 981-02-2040-5. LCCN QA76.642.I58 1994.

Murao:1997:TEI

[MF97] Hirokazu Murao and Tetsuro Fujise. Towards an efficient implementation of a fast algorithm for multipoint polynomial evaluation and its parallel processing. In Küchlin [Küc97], pages 24– 30. ISBN 0-89791-875-4. LCCN QA76.95. URL http://www.

acm.org:80/pubs/citations/proceedings/issac/266670/
p24-murao/.

Monagan:1994:SFA

[MG94] Michael B. Monagan and Gaston H. Gonnet. Signature functions for algebraic numbers. In ACM [ACM94], pages 291–296. ISBN 0-89791-638-7. LCCN QA76.95.I59 1994. URL http://

www.acm.org:80/pubs/citations/proceedings/issac/190347/
p291-monagan/.

Molenkamp:1991:IAA

[MGv91] J. H. J. Molenkamp, V. V. Goldman, and J. A. van Hulzen. An improved approach to automatic error cumulation control. In Watt [Wat91], pages 414–418. ISBN 0-89791-437-6. LCCN

QA 76.95 I59 1991. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/120694/p414-molenkamp/.

Majewski:1995:SEG

[MH95] Bohdan S. Majewski and George Havas. A solution to the extended gcd problem. In Levelt [Lev95], pages 248–253. ISBN 0-89791-699-9. LCCN QA 76.95 I59 1995. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/220346/p248-

majewski/. ACM order number: 505950.

Montelius:1997:EPS

[MH97] Johan Montelius and Seif Haridi. An evaluation of Penny: a system for fine grain implicit parallelism. In Küchlin [Küc97], pages 46–57. ISBN 0-89791-875-4. LCCN QA76.95. URL http://

www.acm.org:80/pubs/citations/proceedings/issac/266670/
p46-montelius/.

Mansfield:2001:TAW

[MH01] Elizabeth L. Mansfield and Peter E. Hydon. Towards approximations which preserve integrals. In Mourrain [Mou01], pages 217–222. ISBN 1-58113-417-7. LCCN QA76.95.I59 2001. ACM order number 505010.

Miller:1989:PGE

[Mil89] B. R. Miller. A program generator for efficient evaluation of Fourier series. In Gonnet [Gon89], pages 199– 206. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989.

US$29.00. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/74540/p199-miller/. ACM order num-

ber: 505890. English and French.

Milowski:2004:CII

[Mil04] R. Alexander Milowski. Computing irredundant irreducible decompositions of large scale monomial ideals. In Gutierrez [Gut04], pages 235–242. ISBN 1-58113-827-X. LCCN QA76.95 .I57 2004.

Miller:2011:CAE

[Mil11] Victor S. Miller. Computational aspects of elliptic curves and modular forms. In Schost and Emiris [SE11], pages 1–2. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Minkwitz:1995:COI

[Min95] Torsten Minkwitz. On the computation of ordinary irreducible representations of finite groups. In Levelt [Lev95], pages 278–284. ISBN 0-89791-699-9. LCCN QA 76.95

I59 1995. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/220346/p278-minkwitz/. ACM order

number: 505950.

Minimair:2003:FSR

[Min03] Manfred Minimair. Factoring sparse resultants of linearly combined polynomials. In Senda [Sen03], pages 207–214. ISBN 1-58113-641-2. LCCN QA76.95. ACM order number 505030.

Minimair:2006:RSC

[Min06] Manfred Minimair. Resultants of skewly composed polynomials. In Trager [Tra06], pages 228–233. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

Miyamoto:2000:CNP

[Miy00]  Izumi Miyamoto.  Computing normalizers of permuta-

tion groups efficiently using isomorphisms of association schemes. In Traverso [Tra00], pages 220–224. ISBN 1-58113-218-2. LCCN QA76.95.I59 2000. URL http://www.

acm.org/pubs/articles/proceedings/issac/345542/p220-
miyamoto/p220-miyamoto.pdf;   http://www.acm.org/pubs/
citations/proceedings/issac/345542/p220-miyamoto/. ACM

order number 505000.

Miyamoto:2006:IGN

[Miy06] Izumi Miyamoto. An improvement of GAP normalizer function for permutation groups. In Trager [Tra06], pages 234–238. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

Miyamoto:2007:CSM

[Miy07] Izumi Miyamoto. A computation of some multiply homogeneous superschemes from transitive permutation groups. In Brown [Bro07], pages 293–298. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Moritsugu:1999:MZS

[MK99] Schuichi Moritsugu and Kazuko Kuriyama. On multiple zeros of systems of algebraic equations. In Dooley [Doo99], pages 23–30. ISBN 1-58113-073-2. LCCN QA76.95 .I57 1999. URL http://

www.acm.org/pubs/contents/proceedings/issac/309831/.

Monagan:1997:TPM

[MM97] Michael B. Monagan and Gladys Monagan. A toolbox for program manipulation and efficient code generation with an application to a problem in computer vision. In Küchlin [Küc97], pages 257–264. ISBN 0-89791-875-4. LCCN

QA76.95. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/258726/p257-monagan/.

Mulholland:2001:ATP

[MM01] Jamie Mulholland and Michael Monagan. Algorithms for trigonometric polynomials. In Mourrain [Mou01], pages 245– 252. ISBN 1-58113-417-7. LCCN QA76.95.I59 2001. ACM order number 505010.

Mantzaflaris:2011:DCI

[MM11] Angelos Mantzaflaris and Bernard Mourrain. Deflation and certified isolation of singular zeros of polynomial systems. In Schost and Emiris [SE11], pages 249–256. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Marinari:1991:GBI

[MMM91] M. G. Marinari, H. M. Möller, and T. Mora. Gröbner bases of ideals given by dual bases. In Watt [Wat91], pages 55–63. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991. URL http://www.

acm.org:80/pubs/citations/proceedings/issac/120694/
p55-marinari/.

Marinari:1995:GDM

[MMM95] Maria Grazia Marinari, Teo Mora, and Hans Michael Möller. Gröbner duality and multiplicities in polynomial system solving. In Levelt [Lev95], pages 167–179. ISBN 0-89791-699-9. LCCN QA 76.95 I59 1995. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/220346/p167-

marinari/. ACM order number: 505950.

Mihailescu:2007:CES

[MMS07] P. Mihailescu, F. Morain, and É. Schost. Computing the eigenvalue in the Schoof–Elkies–Atkin algorithm using Abelian lifts. In Brown [Bro07], pages 285–292. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Moller:1992:GBC

[MMT92] H. Michael Möller, Teo Mora, and Carlo Traverso. Gröbner bases computation using syzygies. In Wang [Wan92b], pages 320–328. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN QA76.95.I59 1992. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/143242/p320-

moller/. ACM order number: 505920.

Monagan:1993:GAD

[MN93] Michael B. Monagan and Walter M. Neuenschwander. GRA-DIENT: algorithmic differentiation in Maple. In Bronstein [Bro93], pages 68–76. ISBN 0-89791-604-2. LCCN QA 76.95

I59 1993. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/164081/p68-monagan/. ACM order num-

ber: 505930.

Monagan:2004:MQR

[Mon04] Michael Monagan. Maximal quotient rational reconstruction: an almost optimal algorithm for rational reconstruction. In Gutierrez [Gut04], pages 243–249. ISBN 1-58113-827-X. LCCN QA76.95 .I57 2004.

Monagan:2005:PAC

[Mon05] Michael Monagan. Probabilistic algorithms for computing resultants. In Kauers [Kau05], pages 245–252. ISBN 1-59593-095-7. LCCN ???? ACM Order Number 505050.

Moody:2011:DPJ

[Moo11] Dustin Moody. Division polynomials for Jacobi quartic curves. In Schost and Emiris [SE11], pages 265–272. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Mora:1989:AAA

[Mor89a] T. Mora, editor. Applied Algebra, Algebraic Algorithms and Error-Correcting Codes. 6th International Conference, AAECC-6, Rome, Italy, July 4–8, 1988. Proceedings, volume 357 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1989. ISBN 3-540-51083-4. LCCN QA268 .A35 1988. Conference held jointly with ISSAC ’88.

Mora:1989:GBN

[Mor89b] T. Mora. Gröbner bases in noncommutative algebras. In Gianni [Gia89], pages 150–161. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Morain:1992:ENE

[Mor92] F. Morain. Easy numbers for the elliptic curve primality proving algorithm. In Wang [Wan92b], pages 263–268. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN QA76.95.I59 1992. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/143242/p263-morain/

. ACM order number: 505920.

Morje:1995:NLA

[Mor95] P. Morje. A nearly linear algorithm for Sylow subgroups in small-base groups. In Levelt [Lev95], pages 270–277. ISBN 0-89791-699-9. LCCN QA 76.95 I59 1995. ACM order number: 505950.

Mora:2002:IPI

[Mor02] Teo Mora, editor. ISSAC 2002: Proceedings of the 2002 International Symposium on Symbolic and Algebraic Computation, July 07–10, 2002, Université de Lille, Lille, France. ACM Press, New York, NY 10036, USA, 2002. ISBN 1-58113-484-3. LCCN

QA76.95. URL http://www.lifl.fr/ISSAC2002/.

Moroz:2006:CRP

[Mor06] Guillaume Moroz. Complexity of the resolution of parametric systems of polynomial equations and inequations. In Trager [Tra06], pages 246–253. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

Mourrain:1993:GPP

[Mou93] B. Mourrain. The 40 “generic” positions of a parallel robot. In Bronstein [Bro93], pages 173–182. ISBN 0-89791-604-2. LCCN QA 76.95 I59 1993. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/164081/p173-

mourrain/. ACM order number: 505930.

Mourrain:2001:IJU

[Mou01] Bernard Mourrain, editor. ISSAC 2001: July 22–25, 2001, University of Western Ontario, London, Ontario, Canada: proceedings of the 2001 International Symposium on Symbolic and Algebraic Computation. ACM Press, New York, NY 10036, USA, 2001. ISBN 1-58113-417-7. LCCN QA76.95.I59 2001. ACM order number 505010.

Mourrain:2005:GNF

[Mou05] Bernard Mourrain. Generalized normal forms and polynomial system solving. In Kauers [Kau05], pages 253–260. ISBN 1-59593-095-7. LCCN ???? ACM Order Number 505050.

Mishra:1990:ARA

[MP90] B. Mishra and P. Pedersen. Arithmetic with real algebraic numbers is in NC. In Watanabe and Nagata [WN90], pages 120–126. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley).

LCCN QA76.95 .I57 1990. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/96877/p120-mishra/.

Monagan:2006:RSM

[MP06] Michael Monagan and Roman Pearce. Rational simplification modulo a polynomial ideal. In Trager [Tra06], pages 239–245. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

Monagan:2009:PSP

[MP09] Michael Monagan and Roman Pearce. Parallel sparse polynomial multiplication using heaps. In May [May09], pages 263–

270. ISBN 1-60558-609-9. LCCN ????

Minh:1998:CMG

[MPV98] Hoang Ngoc Minh, Michel Petitot, and Joris Van Der Hoeven. Computation of the monodromy of generalized polylogarithms. In Gloor [Glo98], pages 276–283. ISBN 1-58113-002-

3. LCCN QA155.7.E4 E88 1998. URL http://www.acm.org/

pubs/citations/proceedings/issac/281508/p276-minh/.

Muller-Quade:1998:DLD

[MQR98] Jörn Müller-Quade and Martin Rötteler. Deciding linear disjointness of finitely generated fields. In Gloor [Glo98], pages 153–160. ISBN 1-58113-002-3. LCCN QA155.7.E4 E88 1998.

URL http://www.acm.org/pubs/citations/proceedings/
issac/281508/p153-muller-quade/.

Murray:1989:EPD

[MR89] N. V. Murray and E. Rosenthal. Employing path dissolution to shorten tableaux proofs. In Gonnet [Gon89], pages 373–381. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989.

US$29.00. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/74540/p373-murray/. ACM order num-

ber: 505890. English and French.

Murray:1990:RIT

[MR90] N. V. Murray and E. Rosenthal. Reexamining intractability of tableau methods. In Watanabe and Nagata [WN90], pages 52–59. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95 .I57 1990. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/96877/p52-murray/

.

Madlener:1993:CGB

[MR93] Klaus Madlener and Birgit Reinert. Computing Gröbner bases in monoid and group rings. In Bronstein [Bro93], pages 254–263. ISBN 0-89791-604-2. LCCN QA 76.95

I59 1993. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/164081/p254-madlener/. ACM order

number: 505930.

Marti:1994:CSS

[MR94] P. Marti and M. Rucher. A cooperative scheme for solving constraints over the reals. In Hong [Hon94], pages 284–293. ISBN 981-02-2040-5. LCCN QA76.642.I58 1994.

Mayr:2010:DBG

[MR10] Ernst W. Mayr and Stephan Ritscher. Degree bounds for Gröbner bases of low-dimensional polynomial ideals. In Watt [Wat10], pages 21–27. ISBN 1-4503-0150-9. LCCN QA76.95 .I59 2010.

Mayr:2011:SEG

[MR11] Ernst W. Mayr and Stephan Ritscher. Space-efficient Gröbner basis computation without degree bounds. In Schost and Emiris [SE11], pages 257–264. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Musser:1989:GP

[MS89] D. R. Musser and A. A. Stepanov. Generic programming. In Gianni [Gia89], pages 13–25. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Marinuzzi:1992:LNS

[MS92] Francesco Marinuzzi and Stefano Soliani. LISA: A new symbolic package for the definition, analysis and resolution of Markovian processes: symbolic and inductive techniques. In Wang [Wan92b], pages 303–311. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN

QA76.95.I59 1992. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/143242/p303-marinuzzi/. ACM

order number: 505920.

McKay:1997:FRA

[MS97] John McKay and Richard Stauduhar. Finding relations among the roots of an irreducible polynomial. In Küchlin [Küc97], pages 75–77. ISBN 0-89791-875-4. LCCN QA76.95. URL http:

//www.acm.org:80/pubs/citations/proceedings/issac/258726/
p75-mckay/.

Mulders:1998:M

[MS98] Thom Mulders and Arne Storjohann. The modulo N extended GCD problem for the polynomials. In Gloor [Glo98], pages 105–

112. ISBN 1-58113-002-3. LCCN QA155.7.E4 E88 1998. URL

http://www.acm.org/pubs/citations/proceedings/issac/
281508/p105-mulders/.

Mulders:1999:DLS

[MS99] Thom Mulders and Arne Storjohann. Diophantine linear system solving. In Dooley [Doo99], pages 181–188. ISBN 1-58113-073-2.

LCCN QA76.95 .I57 1999. URL http://www.acm.org/pubs/
contents/proceedings/issac/309831/.

Mulders:2000:RSS

[MS00] Thom Mulders and Arne Storjohann. Rational solutions of singular linear systems. In Traverso [Tra00], pages 242–249. ISBN 1-58113-218-2. LCCN QA76.95.I59 2000. URL http://www.

acm.org/pubs/articles/proceedings/issac/345542/p242-
mulders/p242-mulders.pdf;   http://www.acm.org/pubs/
citations/proceedings/issac/345542/p242-mulders/. ACM

order number 505000.

Matera:2002:DHF

[MS02] Guillermo Matera and Alexandre Sedoglavic. The differential Hilbert function of a differential rational mapping can be computed in polynomial time. In Mora [Mor02], pages 184–191. ISBN 1-58113-484-3. LCCN QA76.95. URL http://www.lifl.

fr/ISSAC2002/.

Mansfield:2003:ETD

[MS03a] E. L. Mansfield and A. Szanto. Elimination theory for differential difference polynomials. In Senda [Sen03], pages 191–198. ISBN 1-58113-641-2. LCCN QA76.95. ACM order number 505030.

Meunier:2003:EAG

[MS03b] Ludovic Meunier and Bruno Salvy. ESF: an automatically generated encyclopedia of special functions. In Senda [Sen03], pages 199–206. ISBN 1-58113-641-2. LCCN QA76.95. ACM order number 505030.

Mehlhorn:2009:IRR

[MS09] Kurt Mehlhorn and Michael Sagraloff. Isolating real roots of real polynomials. In May [May09], pages 247–254. ISBN 1-60558-609-9. LCCN ????

Morel:2009:HLU

[MSV09] Ivan Morel, Damien Stehlé, and Gilles Villard. H-LLL: using Householder inside LLL. In May [May09], pages 271–278. ISBN 1-60558-609-9. LCCN ????

May:2007:EMR

[MSW07] John P. May, David Saunders, and Zhendong Wan. Efficient matrix rank computation with application to the study of strongly regular graphs. In Brown [Bro07], pages 277–284. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Mehlhorn:2013:AFR

[MSW13] Kurt Mehlhorn, Michael Sagraloff, and Pengming Wang. From approximate factorization to root isolation. In Monagan et al. [MCG13], pages 283–290. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

Massey:1994:OCM

[MT94] B. Massey and E. Tick. Optimizing clause matching automata in committed-choice languages. In Hong [Hon94], pages 294– 303. ISBN 981-02-2040-5. LCCN QA76.642.I58 1994.

Mourrain:2000:SPC

[MT00] Bernard Mourrain and Philippe Trébuchet. Solving projective complete intersection faster. In Traverso [Tra00], pages 234–241. ISBN 1-58113-218-2. LCCN QA76.95.I59 2000. URL http:/

/www.acm.org/pubs/articles/proceedings/issac/345542/
p234-mourrain/p234-mourrain.pdf;   http://www.acm.org/
pubs/citations/proceedings/issac/345542/p234-mourrain/

. ACM order number 505000.

Mourrain:2012:BBR

[MT12] Bernard Mourrain and Philippe Trébuchet. Border basis representation of a general quotient algebra. In van der Hoeven and van Hoeij [vdHvH12], pages 265–272. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Muller:1998:MDM

[Mül98] Jürgen Müller. The 5-modular decomposition matrix of the sporadic simple Conway group Co_{3}. In Gloor [Glo98], pages 179– 185. ISBN 1-58113-002-3. LCCN QA155.7.E4 E88 1998. URL

http://www.acm.org/pubs/citations/proceedings/issac/
281508/p179-muller/.

Mazurik:1990:SCS

[MV90] S. I. Mazurik and E. V. Vorozhtsov. Symbolic-numerical computations in the stability analyses of difference schemes. In Watanabe and Nagata [WN90], pages 177–184. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95

.I57 1990. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/96877/p177-mazurik/.

Menezes:1989:SCA

[MvV89] A. J. Menezes, P. C. van Oorschot, and S. A. Vanstone. Some computational aspects of root finding in GF (q^{m}). In Gianni [Gia89], pages 259–270. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Man:1994:FPD

[MW94] Yiu-Kwong Man and Francis J. Wright. Fast polynomial dispersion computation and its application to indefinite summation. In ACM [ACM94], pages 175–180. ISBN 0-89791-638-7. LCCN QA76.95.I59 1994. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/190347/p175-man/.

Monagan:2000:DIB

[MW00] Michael B. Monagan and Allan D. Wittkopf. On the design and implementation of Brown’s algorithm over the integers and number fields. In Traverso [Tra00], pages 225–233. ISBN 1-58113-218-2. LCCN QA76.95.I59 2000. URL http://www.

acm.org/pubs/articles/proceedings/issac/345542/p225-
monagan/p225-monagan.pdf;   http://www.acm.org/pubs/
citations/proceedings/issac/345542/p225-monagan/. ACM

order number 505000.

Micciancio:2001:LSA

[MW01] Daniele Micciancio and Bogdan Warinschi. A linear space algorithm for computing the Hermite normal form. In Mourrain [Mou01], pages 231–236. ISBN 1-58113-417-7. LCCN QA76.95.I59 2001. ACM order number 505010.

Mao:2005:AWM

[MW05] Weibo Mao and Jinzhao Wu. Application of Wu’s method to symbolic model checking. In Kauers [Kau05], pages 237–244. ISBN 1-59593-095-7. LCCN ???? ACM Order Number 505050.

Mevissen:2009:SPS

[MYT09] Martin Mevissen, Kosuke Yokoyama, and Nobuki Takayama. Solutions of polynomial systems derived from the steady cavity flow problem. In May [May09], pages 255–262. ISBN 1-60558-609-9. LCCN ????

Mikhalev:1996:APE

[MZ96] Alexander A. Mikhalev and Andrej A. Zolotykh. Algorithms for primitive elements of free Lie algebras and Lie superalgebras. In Lakshman Y. N. [Lak96], pages 161–169. ISBN 0-89791-796-0. LCCN QA 76.95 I59 1996. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/236869/p161-
mikhalev/.

Mingsheng:2001:RGB

[MZ01] Wang Mingsheng and Liu Zhuojun. Remarks on Gröbner basis for ideals under composition. In Mourrain [Mou01], pages 237– 244. ISBN 1-58113-417-7. LCCN QA76.95.I59 2001. ACM order number 505010.

Ma:2011:MRG

[MZ11] Yue Ma and Lihong Zhi. The minimum-rank gram matrix completion via modified fixed point continuation method. In Schost and Emiris [SE11], pages 241–248. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Ma:2012:CRS

[MZ12] Yue Ma and Lihong Zhi. Computing real solutions of polynomial systems via low-rank moment matrix completion. In van der Hoeven and van Hoeij [vdHvH12], pages 249–256. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Nabeshima:2007:SAC

[Nab07] Katsusuke Nabeshima. A speed-up of the algorithm for computing comprehensive Gröbner systems. In Brown [Bro07], pages 299–306. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Nagasaka:2002:TCI

[Nag02] Kosaku Nagasaka. Towards certified irreducibility testing of bivariate approximate polynomials. In Mora [Mor02], pages 192–199. ISBN 1-58113-484-3. LCCN QA76.95. URL http://

www.lifl.fr/ISSAC2002/.

Nagasaka:2011:CSG

[Nag11] Kosaku Nagasaka. Computing a structured Gröbner basis approximately. In Schost and Emiris [SE11], pages 273–280. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Nam:1995:HSM

[Nam95] Trˆ àn Quoˆ ác Nam. A hybrid symbolic-numerical method for tracing surface-to-surface intersections. In Levelt [Lev95], pages 51–58. ISBN 0-89791-699-9. LCCN QA 76.95

I59 1995. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/220346/p51-nam/. ACM order number:

505950.

Norman:1996:MTA

[NF96] Arthur Norman and John Fitch. Memory tracing of algebraic calculations. In Lakshman Y. N. [Lak96], pages 113–119. ISBN 0-89791-796-0. LCCN QA 76.95 I59 1996. URL http://www.

acm.org:80/pubs/citations/proceedings/issac/236869/
p113-norman/.

Norman:1997:CPP

[NF97] Arthur Norman and John Fitch. CABAL: polynomial and power series algebra on a parallel computer. In Küchlin [Küc97], pages 196–203. ISBN 0-89791-875-4. LCCN

QA76.95. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/266670/p196-norman/.

Ng:1979:SAC

[Ng79] Edward W. Ng, editor. Symbolic and algebraic computation: EUROSAM ’79, an International Symposium on Symbolic and Algebraic Manipulation, Marseille, France, June 1979, volume 72 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNCSD9. ISBN 0-387-09519-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA155.7.E4 E88 1979.

Noda:1990:SHI

[NM90] Matu-Tarow T. Noda and E. Miyahiro. On the symbolic/ numeric hybrid integration. In Watanabe and Nagata [WN90], page 304. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95 .I57 1990. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/96877/p304-
noda/.

Noro:1997:CRF

[NM97] Masayuki Noro and John McKay. Computation of replicable functions on Risa/Asir. In Küchlin [Küc97], pages 130–138. ISBN 0-89791-875-4. LCCN QA76.95. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/266670/p130-
noro/.

Nocker:2000:SRP

[Nöc00] Michael Nöcker. Some remarks on parallel exponentiation (extended abstract). In Traverso [Tra00], pages 250–257. ISBN 1-58113-218-2. LCCN QA76.95.I59 2000. URL http://

www.acm.org/pubs/citations/proceedings/issac/345542/

p250-nocker/. ACM order number 505000.

Norman:1990:CBI

[Nor90] A. C. Norman. A critical-pair/completion based integration algorithm. In Watanabe and Nagata [WN90], pages 201–205. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95 .I57 1990. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/96877/p201-norman/.

Nordbeck:1998:CSB

[Nor98a] Patrik Nordbeck. Canonical subalgebraic bases in noncommutative polynomial rings. In Gloor [Glo98], pages 140– 146. ISBN 1-58113-002-3. LCCN QA155.7.E4 E88 1998. URL

http://www.acm.org/pubs/citations/proceedings/issac/
281508/p140-nordbeck/.

Norenberg:1998:CMA

[Nör98b] R. Nörenberg. Covering monomial algebras. In Gloor [Glo98], pages 161–164. ISBN 1-58113-002-3. LCCN QA155.7.E4

E88 1998. URL http://www.acm.org/pubs/citations/
proceedings/issac/281508/p161-norenberg/.

Norman:2000:FEJ

[Nor00] Arthur C. Norman. Further evaluation of Java for symbolic computation. In Traverso [Tra00], pages 258–265. ISBN 1-58113-218-2. LCCN QA76.95.I59 2000. URL http://www.

acm.org/pubs/articles/proceedings/issac/345542/p258-
norman/p258-norman.pdf;   http://www.acm.org/pubs/citations/
proceedings/issac/345542/p258-norman/. ACM order num-

ber 505000.

Noro:2006:MDE

[Nor06] Masayuki Noro. Modular dynamic evaluation. In Trager [Tra06], pages 262–268. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

Neunhoffer:2006:DSU

[NS06] Max Neunhöffer and Ákos Seress. A data structure for a uniform approach to computations with finite groups. In Trager [Tra06], pages 254–261. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

Noro:1992:RCA

[NT92] Masayuki Noro and Taku Takeshima. Risa/Asir — a computer algebra system. In Wang [Wan92b], pages 387–396. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN

QA76.95.I59 1992. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/143242/p387-noro/. ACM

order number: 505920.

Nussbaum:1997:RPP

[Nus97] Doron Nussbaum. Rectilinear p-piercing problems. In Küchlin [Küc97], pages 316–323. ISBN 0-89791-875-4. LCCN

QA76.95. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/258726/p316-nussbaum/.

Noro:2002:YAP

[NY02] Masayuki Noro and Kazuhiro Yokoyama. Yet another practical implementation of polynomial factorization over finite fields. In Mora [Mor02], pages 200–206. ISBN 1-58113-484-3. LCCN

QA76.95. URL http://www.lifl.fr/ISSAC2002/.

Oaku:1994:AFS

[Oak94] Toshinori Oaku. Algorithms for finding the structure of solutions of a system of linear partial differential equations.

In ACM [ACM94], pages 216–223. ISBN 0-89791-638-7. LCCN QA76.95.I59 1994. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/190347/p216-oaku/.

Oaku:2012:ACD

[Oak12] Toshinori Oaku. An algorithm to compute the differential equations for the logarithm of a polynomial. In van der Hoeven and van Hoeij [vdHvH12], pages 273–280. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Ohno:1997:IMC

[OIiM^{+}97] Kazuhiko Ohno, Masahiko Ikawa, Shin ichiro Mori, Hiroshi Nakashima, Shinji Tomita, and Masahiro Goshima. Improvement of message communication in concurrent logic language. In Küchlin [Küc97], pages 156–164. ISBN 0-89791-875-4. LCCN

QA76.95. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/266670/p156-ohno/.

Okada:1989:SNC

[Oka89] M. Okada. Strong normalizability for the combined system of the typed λ calculus and an arbitrary convergent term rewrite system. In Gonnet [Gon89], pages 357–363. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989. US$29.00. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/74540/p357-

okada/. ACM order number: 505890. English and French.

Okubo:1990:GTO

[Oku90] K. Okubo. Global theory of ordinary differential equations and formula manipulation. In Watanabe and Nagata [WN90], pages 193–200. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95 .I57 1990. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/96877/p193-
okubo/.

Ollivier:1989:IRM

[Oll89] F. Ollivier. Inversibility of rational mappings and structural identifiability in automatics. In Gonnet [Gon89], pages 43–54. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989.

US$29.00. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/74540/p43-ollivier/. ACM order num-

ber: 505890. English and French.

Orange:2009:CSS

[ORY09] Sébastien Orange, Guénaël Renault, and Kazuhiro Yokoyama. Computation schemes for splitting fields of polynomials. In May [May09], pages 279–286. ISBN 1-60558-609-9. LCCN ????

OHearn:1989:NTP

[OS89] P. O’Hearn and Z. Stachniak. Note on theorem proving strategies for resolution counterparts of nonclassical logics. In Gonnet [Gon89], pages 364–372. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989. US$29.00. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/74540/p364-o_hearn/

. ACM order number: 505890. English and French.

Oevel:1991:YES

[OS91] Walter Oevel and Klaus Strack. The Yang–Baxter equation and a systematic search for Poisson brackets on associative algebras. In Watt [Wat91], pages 229–236. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/120694/p229-oevel/.

Orden:2004:PNC

[OS04] David Orden and Francisco Santos. The polytope of noncrossing graphs on a planar point set. In Gutierrez [Gut04], pages 250–257. ISBN 1-58113-827-X. LCCN QA76.95 .I57 2004.

Oancea:2005:DEI

[OW05] Cosmin E. Oancea and Stephen M. Watt. Domains and expressions: an interface between two approaches to computer algebra. In Kauers [Kau05], pages 261–268. ISBN 1-59593-095-7. LCCN ???? ACM Order Number 505050.

Painter:1992:MES

[Pai92] Jeffrey F. Painter. The matrix editor for symbolic Jacobians in ALPAL. In Wang [Wan92b], pages 312–319. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN

QA76.95.I59 1992. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/143242/p312-painter/. ACM

order number: 505920.

Pan:1989:SCD

[Pan89] Victor Pan. On some computations with dense structured matrices. In Gonnet [Gon89], pages 34–42. ISBN 0-89791-325-6.

LCCN QA76.95.I59 1989. US$29.00. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/74540/p34-pan/

. ACM order number: 505890. English and French.

Pan:2000:MSP

[Pan00] Victor Y. Pan. Matrix structure, polynomial arithmetic, and erasure-resilient encoding/decoding. In Traverso [Tra00], pages 266–271. ISBN 1-58113-218-2. LCCN QA76.95.I59 2000. URL

http://www.acm.org/pubs/articles/proceedings/issac/
345542/p266-pan/p266-pan.pdf;   http://www.acm.org/
pubs/citations/proceedings/issac/345542/p266-pan/. ACM

order number 505000.

Pan:2001:UPN

[Pan01] Victor Y. Pan. Univariate polynomials: nearly optimal algorithms for factorization and rootfinding. In Mourrain [Mou01], pages 253–267. ISBN 1-58113-417-7. LCCN QA76.95.I59 2001. ACM order number 505010.

Parrilo:2004:SSP

[Par04] Pablo A. Parrilo. Sums of squares of polynomials and their applications. In Gutierrez [Gut04], page 1. ISBN 1-58113-827-X. LCCN QA76.95 .I57 2004.

Parrilo:2013:CAG

[Par13] Pablo A. Parrilo. Convex algebraic geometry and semidefinite optimization. In Monagan et al. [MCG13], pages 9–10. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

Paschel:2009:ASH

[Päs09] Markus Päschel. Automatic synthesis of high performance mathematical programs. In May [May09], pages 5–6. ISBN 1-60558-609-9. LCCN ????

Padget:1990:UPS

[PB90] J. Padget and A. Barnes. Univariate power series expansions in Reduce. In Watanabe and Nagata [WN90], pages 82–87. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95 .I57 1990. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/96877/p82-padget/.

Pak:1999:SGS

[PB99] Igor Pak and Sergey Bratus. On sampling generating sets of finite groups and product replacement algorithm: extended abstract. In Dooley [Doo99], pages 91–96. ISBN 1-58113-073-2.

LCCN QA76.95 .I57 1999. URL http://www.acm.org/pubs/
contents/proceedings/issac/309831/.

Perez-Diaz:2005:PDF

[PDS05] Sonia Pérez-Díaz and J. Rafael Sendra. Partial degree formulae for rational algebraic surfaces. In Kauers [Kau05], pages 301– 308. ISBN 1-59593-095-7. LCCN ???? ACM Order Number 505050.

Pecelli:1991:FMD

[Pec91] Giampiero Pecelli. Formal methods in delay-differential equations. In Watt [Wat91], pages 317–318. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/120694/p317-
pecelli/.

Petho:1991:AGB

[Pet91] Attila Pethö. Application of Gröbner bases to the resolution of systems of norm equations. In Watt [Wat91], pages 144–150. ISBN 0-89791-437-6. LCCN QA 76.95

I59 1991. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/120694/p144-petho/.

Petitjean:1994:ACS

[Pet94] Sylvain Petitjean. Automating the construction of stationary multiple-point classes. In ACM [ACM94], pages 9–14. ISBN 0-89791-638-7. LCCN QA76.95.I59 1994. URL http://www.

acm.org:80/pubs/citations/proceedings/issac/190347/
p9-petitjean/.

Pflugel:1997:ACE

[Pfl97] Eckhard Pflügel. An algorithm for computing exponential solutions of first order linear differential systems. In Küchlin [Küc97], pages 164–171. ISBN 0-89791-875-4. LCCN

QA76.95. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/258726/p164-pflugel/.

Pfluegel:2008:RDL

[Pfl08] Eckhard Pflüegel. A rational decomposition-lemma for systems of linear differential-algebraic equations. In Jeffrey [Jef08], pages 231–238. ISBN 1-59593-904-0. LCCN ????

Pontelli:1997:PSC

[PG97] E. Pontelli and G. Gupta. Parallel symbolic computation in ACE. In Baral et al. [BKLG97], pages 359–396. ISSN 1012-2443.

Pan:2013:TFA

[PHW13] Senshan Pan, Yupu Hu, and Baocang Wang. The termination of the F 5 algorithm revisited. In Monagan et al. [MCG13], pages 291–298. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

Pillwein:2013:TCP

[Pil13] Veronika Pillwein. Termination conditions for positivity proving procedures. In Monagan et al. [MCG13], pages 315–322. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

Peternell:2008:GSS

[PO08] Martin Peternell and Boris Odehnal. On generalized ln-surfaces in 4-space. In Jeffrey [Jef08], pages 223–230. ISBN 1-59593-904-0. LCCN ????

Poloskov:1998:CPP

[Pol98] Igor E. Poloskov. Compound program packages and a nonlinear random fluctuations analysis. In Gloor [Glo98], pages 70– 75. ISBN 1-58113-002-3. LCCN QA155.7.E4 E88 1998. URL

http://www.acm.org/pubs/citations/proceedings/issac/
281508/p70-poloskov/.

Porter:1989:DRA

[Por89] S. C. Porter. Dense representation of affine coordinate rings of curves with one point at infinity. In Gonnet [Gon89], pages 287–297. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989.

US$29.00. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/74540/p287-porter/. ACM order num-

ber: 505890. English and French.

Pospelov:2011:FFT

[Pos11] Alexey Pospelov. Fast Fourier Transforms over poor fields. In Schost and Emiris [SE11], pages 289–296. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Pottier:1996:EAD

[Pot96] Loïc Pottier. The Euclidean algorithm in dimension n. In Lakshman Y. N. [Lak96], pages 40–42. ISBN 0-89791-796-0. LCCN

QA 76.95 I59 1996. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/236869/p40-pottier/.

Poulisse:2006:CCA

[Pou06] Hennie Poulisse. Computational communicative algebra. In Trager [Tra06], pages 3–4. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

Pan:2011:RPM

[PQZ11] Victor Y. Pan, Guoliang Qian, and Ai-Long Zheng. Randomized preconditioning of the MBA algorithm. In Schost and Emiris [SE11], pages 281–288. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Powers:2006:QPT

[PR06] Victoria Powers and Bruce Reznick. A quantitative Pólya’s Theorem with corner zeros. In Trager [Tra06], pages 285–289. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

Poteaux:2008:GRP

[PR08] Adrien Poteaux and Marc Rybowicz. Good reduction of Puiseux series and complexity of the Newton–Puiseux algorithm over finite fields. In Jeffrey [Jef08], pages 239–246. ISBN 1-59593-904-0. LCCN ????

Petkovsek:1993:FAH

[PS93a] M. Petkovsek and B. Salvy. Finding all hypergeometric solutions of linear differential equations. In Bronstein [Bro93], pages 27–33. ISBN 0-89791-604-2. LCCN QA 76.95 I59 1993. ACM order number: 505930.

Petkovsek:1993:FAHb

[PS93b] Marko Petkovšek and Bruno Salvy. Finding all hypergeometric solutions of linear differential equations. In Bronstein [Bro93], pages 27–33. ISBN 0-89791-604-2. LCCN QA 76.95

I59 1993. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/164081/p27-petkovscaronek/. ACM or-

der number: 505930.

Pascal:2006:COB

[PS06] Cyril Pascal and Éric Schost. Change of order for bivariate triangular sets. In Trager [Tra06], pages 277–284. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

Pernet:2007:FAC

[PS07] Clément Pernet and Arne Storjohann. Faster algorithms for the characteristic polynomial. In Brown [Bro07], pages 307–314. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Pauderis:2012:DUC

[PS12] Colton Pauderis and Arne Storjohann. Deterministic unimodularity certification. In van der Hoeven and van Hoeij [vdHvH12], pages 281–288. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Pauderis:2013:CIS

[PS13] Colton Pauderis and Arne Storjohann. Computing the invariant structure of integer matrices: fast algorithms into practice. In Monagan et al. [MCG13], pages 307–314. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

Pan:2013:BCR

[PT13] Victor Y. Pan and Elias P. Tsigaridas. On the boolean complexity of real root refinement. In Monagan et al. [MCG13], pages 299–306. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

Popova:1998:SSN

[PU98] E. D. Popova and C. P. Ullrich. Simplification of symbolicnumerical interval expressions. In Gloor [Glo98], pages 207– 214. ISBN 1-58113-002-3. LCCN QA155.7.E4 E88 1998. URL

http://www.acm.org/pubs/citations/proceedings/issac/
281508/p207-popova/.

Purtilo:1986:ASI

[Pur86] J. Purtilo. Applications of a software interconnection system in mathematical problem solving environments. In Char [Cha86b], pages 16–23. ISBN 0-89791-199-7 (paperback). LCCN

QA155.7.E4 A281 1986. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/32439/p16-purtilo/. ACM

order no. 505860.

Purtilo:1989:MEO

[Pur89] J. M. Purtilo. Minion: an environment to organize mathematical problem solving. In Gonnet [Gon89], pages 147– 154. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989.

US$29.00. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/74540/p147-purtilo/. ACM order num-

ber: 505890. English and French.

Pericleous:2001:NCB

[PV01] Savvas Pericleous and Nicolai Vorobjov. New complexity bounds for cylindrical decompositions of sub-Pfaffian sets. In Mourrain [Mou01], pages 268–275. ISBN 1-58113-417-7. LCCN QA76.95.I59 2001. ACM order number 505010.

Pan:2002:AEA

[PW02] Victor Y. Pan and Xinmao Wang. Acceleration of Euclidean algorithm and extensions. In Mora [Mor02], pages 207–213. ISBN 1-58113-484-3. LCCN QA76.95. URL http://www.lifl.

fr/ISSAC2002/.

Pan:2006:UGA

[PW06] Wei Pan and Dongming Wang. Uniform Gröbner bases for ideals generated by polynomials with parametric exponents. In Trager [Tra06], pages 269–276. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

Pan:2010:RCP

[PZ10] Victor Y. Pan and Ai-Long Zheng. Real and complex polynomial root-finding with eigen-solving and preprocessing. In Watt [Wat10], pages 219–226. ISBN 1-4503-0150-9. LCCN QA76.95 .I59 2010.

Quere:1995:ARL

[QV95] M. P. Quéré and G. Villard. An algorithm for the reduction of linear DAE. In Levelt [Lev95], pages 223–231. ISBN 0-89791-699-9. LCCN QA 76.95 I59 1995. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/220346/p223-

quere/. ACM order number: 505950.

Raab:2013:IUF

[Raa13] Clemens G. Raab. Integration of unspecified functions and families of iterated integrals. In Monagan et al. [MCG13], pages 323–330. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

Rabinowitz:1989:CSS

[Rab89] S. Rabinowitz. On the computer solution of symmetric homogeneous triangle inequalities. In Gonnet [Gon89], pages 272–286. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989.

US$29.00. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/74540/p272-rabinowitz/. ACM order

number: 505890. English and French.

Rakoczi:1995:FRN

[Rák95] Ferenc Rákóczi. Fast recognition of the nilpotency of permutation groups. In Levelt [Lev95], pages 265–269. ISBN 0-89791-699-9. LCCN QA 76.95 I59 1995. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/220346/p265-

rakoczi/. ACM order number: 505950.

Reid:1991:RSD

[RB91] G. J. Reid and A. Boulton. Reduction of systems of differential equations to standard form and their integration using directed graphs. In Watt [Wat91], pages 308–312. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/120694/p308-reid/.

Roux:2006:RRC

[RB06] Nicolas Le Roux and Moulay Barkatou. Rank reduction of a class of Pfaffian systems in two variables. In Trager [Tra06], pages 204–211. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

Robinson:2006:PAD

[RC06] Eric Robinson and Gene Cooperman. A parallel architecture for disk-based computing over the Baby Monster and other large finite simple groups. In Trager [Tra06], pages 298–305. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

Rodriguez-Carbonell:2004:AGP

[RCK04] Enric Rodríguez-Carbonell and Deepak Kapur. Automatic generation of polynomial loop invariants: Algebraic foundations. In Gutierrez [Gut04], pages 266–273. ISBN 1-58113-827-X. LCCN QA76.95 .I57 2004.

Roque:1991:QRD

[Rd91] W. L. Roque and R. P. dos Santos. Qualitative reasoning, dimensional analysis and computer algebra. In Watt [Wat91], pages 460–461. ISBN 0-89791-437-6. LCCN QA 76.95

I59 1991. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/120694/p460-roque/.

Rondepierre:2005:ASN

[RD05] Aude Rondepierre and Jean-Guillaume Dumas. Algorithms for symbolic/numeric control of affine dynamical systems. In Kauers [Kau05], pages 277–284. ISBN 1-59593-095-7. LCCN ???? ACM Order Number 505050.

Reischert:1997:AFC

[Rei97] Daniel Reischert. Asymptotically fast computation of subresultants. In Küchlin [Küc97], pages 233–240. ISBN 0-89791-

875-4. LCCN QA76.95. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/258726/p233-reischert/.

Reinert:2000:SSL

[Rei00] Birgit Reinert. Solving systems of linear one-sided equations in integer monoid and group rings. In Traverso [Tra00], pages 281–287. ISBN 1-58113-218-2. LCCN QA76.95.I59 2000. URL

http://www.acm.org/pubs/articles/proceedings/issac/
345542/p281-reinert/p281-reinert.pdf;   http://www.
acm.org/pubs/citations/proceedings/issac/345542/p281-

reinert/. ACM order number 505000.

Renner:1991:NEE

[Ren91] Friedrich Renner. Nonlinear evolution equations and the Painleve analysis: a constructive approach with REDUCE. In Watt [Wat91], pages 289–294. ISBN 0-89791-437-6. LCCN

QA 76.95 I59 1991. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/120694/p289-renner/.

Renault:2006:CSF

[Ren06] Guénaël Renault. Computation of the splitting field of a dihedral polynomial. In Trager [Tra06], pages 290–297. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

Romero:2009:IBC

[RER09] Ana Romero, Graham Ellis, and Julio Rubio. Interoperating between computer algebra systems: computing homology of groups with kenzo and GAP. In May [May09], pages 303–310. ISBN 1-60558-609-9. LCCN ????

Richardson:1994:IPE

[RF94a] Dan Richardson and John Fitch. The identity problem for elementary functions and constants. In ACM [ACM94], pages 285–290. ISBN 0-89791-638-7. LCCN

QA76.95.I59 1994. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/190347/p285-richardson/.

RuizS:1994:AGG

[RF94b] O. E. Ruiz S. and P. M. Ferreira. Algebraic geometry and group theory in geometric constraint satisfaction. In ACM [ACM94], pages 224–233. ISBN 0-89791-638-7. LCCN QA76.95.I59 1994.

Roelofs:1991:IMO

[RG91] Marcel Roelofs and Peter K. H. Gragert. Implementation of multilinear operators in REDUCE and applications in mathematics. In Watt [Wat91], pages 390–396. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/120694/p390-
roelofs/.

Recio:1994:SIG

[RGL94] T. Recio and M. J. González-López. On the symbolic insimplification of the general 6R-manipulator kinematic equations. In ACM [ACM94], pages 354–358. ISBN 0-89791-638-7. LCCN QA76.95.I59 1994. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/190347/p354-recio/.

Rosales:2000:HCI

[RGSGG00] José Carlos Rosales, Pedro A. García-Sánchez, and Juan Ignacio García-García. How to check if a finitely generated commutative monoid is a principal ideal commutative monoid. In Traverso [Tra00], pages 288–291. ISBN 1-58113-218-2. LCCN QA76.95.I59 2000. URL http://www.

acm.org/pubs/articles/proceedings/issac/345542/p288-
rosales/p288-rosales.pdf;   http://www.acm.org/pubs/
citations/proceedings/issac/345542/p288-rosales/. ACM

order number 505000.

Richardson:1991:TCN

[Ric91] Daniel Richardson. Towards computing nonalgebraic cylindrical decompositions. In Watt [Wat91], pages 247–255. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/120694/p247-
richardson/.

Richardson:1992:ECP

[Ric92] Daniel Richardson. The elementary constant problem. In Wang [Wan92b], pages 108–116. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN QA76.95.I59 1992. URL http:

//www.acm.org:80/pubs/citations/proceedings/issac/143242/

p108-richardson/. ACM order number: 505920.

Richardson:1993:ZST

[Ric93] Daniel Richardson. A zero structure theorem for exponential polynomials. In Bronstein [Bro93], pages 144–151. ISBN 0-89791-604-2. LCCN QA 76.95 I59 1993. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/164081/p144-

richardson/. ACM order number: 505930.

Richardson:1995:SMR

[Ric95] Daniel Richardson. A simplified method of recognizing zero among elementary constants. In Levelt [Lev95], pages 104–109. ISBN 0-89791-699-9. LCCN QA 76.95 I59 1995. URL http://

www.acm.org:80/pubs/citations/proceedings/issac/220346/

p104-richardson/. ACM order number: 505950.

Richardson:1996:SES

[Ric96] Daniel Richardson. Solution of elementary systems of equations in a box in R^{n}. In Lakshman Y. N. [Lak96], pages 120–126. ISBN 0-89791-796-0. LCCN QA 76.95

I59 1996. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/236869/p120-richardson/.

Rioboo:1992:RAC

[Rio92] Renaud Rioboo. Real algebraic closure of an ordered field: implementation in Axiom. In Wang [Wan92b], pages 206–215. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN QA76.95.I59 1992. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/143242/p206-rioboo/

. ACM order number: 505920.

Rioboo:2002:TFR

[Rio02] Renaud Rioboo. Towards faster real algebraic numbers. In Mora [Mor02], pages 221–228. ISBN 1-58113-484-3. LCCN QA76.95.

URL http://www.lifl.fr/ISSAC2002/.

Ravenscroft:1989:SSG

[RL89] R. A. Ravenscroft, Jr. and E. A. Lamagna. Symbolic summation with generating functions. In Gonnet [Gon89], pages 228–233. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989. US$29.00. ACM order number: 505890. English and French.

Richardson:2002:SOF

[RL02] Daniel Richardson and Simon Langley. Some observations on familiar numbers. In Mora [Mor02], pages 214–220. ISBN 1-58113-484-3. LCCN QA76.95. URL http://www.lifl.fr/

ISSAC2002/.

Rudenko:1991:ACA

[RLBS91] V. M. Rudenko, V. V. Leonov, A. F. Bragazin, and I. P. Shmyglevsky. Application of computer algebra to the investigation of the orbital satellite motion. In Watt [Wat91], pages 450–451. ISBN 0-89791-437-6. LCCN QA 76.95

I59 1991. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/120694/p450-rudenko/.

Reid:1992:ADC

[RLBW92] G. J. Reid, I. G. Lisle, A. Boulton, and A. D. Wittkopf. Algorithmic determination of commutation relations for Lie symmetry algebras of PDEs. In Wang [Wan92b], pages 63– 68. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN QA76.95.I59 1992. URL http://www.acm.org:

80/pubs/citations/proceedings/issac/143242/p63-reid/

. ACM order number: 505920.

Renbao:1986:CAS

[RLZ86] Z. Renbao, X. Ling, and R. Zhaoyang. The computer algebra system CAS1 for the IBM-PC. In Char [Cha86b], pages 176–178. ISBN 0-89791-199-7 (paperback). LCCN QA155.7.E4

A281 1986. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/32439/p176-renbao/. ACM order no.

505860.

Robinson:2007:DBP

[RMC07] Eric Robinson, Jürgen Müller, and Gene Cooperman. A disk-based parallel implementation for direct condensation of large permutation modules. In Brown [Bro07], pages 315–322. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Reinert:1998:NNR

[RMM98] Birgit Reinert, Klaus Madlener, and Teo Mora. A note on Nielsen reduction and coset enumeration. In Gloor [Glo98], pages 171–178. ISBN 1-58113-002-3. LCCN QA155.7.E4

E88 1998. URL http://www.acm.org/pubs/citations/
proceedings/issac/281508/p171-reinert/.

Roach:1994:SNE

[Roa94] Kelly Roach. Symbolic-numeric nonlinear equation solving. In ACM [ACM94], pages 278–284. ISBN 0-89791-638-7. LCCN QA76.95.I59 1994. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/190347/p278-roach/.

Roach:1996:HFR

[Roa96] Kelly Roach. Hypergeometric function representations. In Lakshman Y. N. [Lak96], pages 301–308. ISBN 0-89791-796-0.

LCCN QA 76.95 I59 1996. URL http://www.acm.org:80/
pubs/citations/proceedings/issac/236869/p301-roach/.

Roach:1997:MGF

[Roa97] Kelly Roach. Meijer G functions representations. In Küchlin [Küc97], pages 205–211. ISBN 0-89791-875-4. LCCN

QA76.95. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/258726/p205-roach/.

Roche:2009:STE

[Roc09] Daniel S. Roche. Space- and time-efficient polynomial multiplication. In May [May09], pages 295–302. ISBN 1-60558-609-9. LCCN ????

Rolletschek:1989:SDC

[Rol89] H. Rolletschek. Shortest division chains in imaginary quadratic number fields. In Gianni [Gia89], pages 231–243. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Roune:2010:SAC

[Rou10] Bjarke Hammersholt Roune. A Slice algorithm for corners and Hilbert–Poincaré series of monomial ideals. In Watt [Wat10], pages 115–122. ISBN 1-4503-0150-9. LCCN QA76.95 .I59 2010.

Roy:2012:CDC

[Roy12] Marie-Françoise Roy. Complexity of deciding connectivity in real algebraic sets: recent results and future research directions. In van der Hoeven and van Hoeij [vdHvH12], pages 3–5. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Rust:1997:RPD

[RR97] C. J. Rust and G. J. Reid. Rankings of partial derivatives. In Küchlin [Küc97], pages 9–16. ISBN 0-89791-875-4. LCCN

QA76.95. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/258726/p9-rust/.

Rosenkranz:2008:IDP

[RR08] Markus Rosenkranz and Georg Regensburger. Integrodifferential polynomials and operators. In Jeffrey [Jef08], pages 261–268. ISBN 1-59593-904-0. LCCN ????

Regensburger:2009:SPA

[RRM09] Georg Regensburger, Markus Rosenkranz, and Johannes Middeke. A skew polynomial approach to integro-differential operators. In May [May09], pages 287–294. ISBN 1-60558-609-9. LCCN ????

Rust:1999:EUT

[RRW99] C. J. Rust, G. J. Reid, and A. D. Wittkopf. Existence and uniqueness theorems for formal power series solutions of analytic differential systems. In Dooley [Doo99], pages 105–112. ISBN 1-58113-073-2. LCCN QA76.95 .I57 1999. URL http://

www.acm.org/pubs/contents/proceedings/issac/309831/.

Reinhart:1996:DHD

[RS96] Georg M. Reinhart and William Sit. Differentially homogeneous differential polynomials. In Lakshman Y. N. [Lak96], pages 212–218. ISBN 0-89791-796-0. LCCN QA 76.95

I59 1996. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/236869/p212-reinhart/.

Romero:2012:PBT

[RS12a] Ana Romero and Francis Sergeraert. Programming before theorizing, a case study. In van der Hoeven and van Hoeij [vdHvH12], pages 289–296. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Roune:2012:PGB

[RS12b] Bjarke Hammersholt Roune and Michael Stillman. Practical Gröbner basis computation. In van der Hoeven and van Hoeij [vdHvH12], pages 203–210. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Roch:1989:CAM

[RSSRV89] J.-L. Roch, P. Senechaud, F. Siebert-Roch, and G. Villard. Computer algebra on MIMD machine. In Gianni [Gia89], pages 423–439. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Richardson:1996:AEE

[RSSV96] Daniel Richardson, Bruno Salvy, John Shackell, and Joris Van der Hoeven. Asymptotic expansions of exp–log functions. In Lakshman Y. N. [Lak96], pages 309–313. ISBN 0-89791-796-0. LCCN QA 76.95 I59 1996. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/236869/p309-
richardson/.

RochSiebert:1991:PFE

[RSV91a] F. Roch-Siebert and G. Villard. PAC: first experiments on a 128 transputers meganode. In Watt [Wat91], pages 343–351. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991.

Roch-Siebert:1991:PFE

[RSV91b] Françoise Roch-Siebert and Gilles Villard. PAC: first experiments on a 128 transputers méganode. In Watt [Wat91], pages 343–351. ISBN 0-89791-437-6. LCCN QA 76.95

I59 1991. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/120694/p343-roch-siebert/.

Recio:2004:HU

[RSV04] Tomas Recio, J. Rafael Sendra, and Carlos Villarino. From hypercircles to units. In Gutierrez [Gut04], pages 258–265. ISBN 1-58113-827-X. LCCN QA76.95 .I57 2004.

Rubio:2006:NIN

[RSV06] Rosario Rubio, J. Miguel Serradilla, and M. Pilar Vélez. A note on implicitization and normal parametrization of rational curves. In Trager [Tra06], pages 306–309. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

Reid:2003:CSN

[RTZ03] Greg Reid, Jianliang Tang, and Lihong Zhi. A complete symbolic-numeric linear method for camera pose determination. In Senda [Sen03], pages 215–223. ISBN 1-58113-641-2. LCCN QA76.95. ACM order number 505030.

Ruatta:2001:MWI

[Rua01] Olivier Ruatta. A multivariate Weierstrass iterative rootfinder. In Mourrain [Mou01], pages 276–283. ISBN 1-58113-417-7. LCCN QA76.95.I59 2001. ACM order number 505010.

Ruffo:2007:SLD

[Ruf07] James Ruffo. A straightening law for the Drinfel’d Lagrangian Grassmannian. In Brown [Bro07], pages 323–330. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Rump:2010:VMR

[Rum10] Siegfried M. Rump. Verification methods: rigorous results using floating-point arithmetic. In Watt [Wat10], pages 3–4. ISBN 1-4503-0150-9. LCCN QA76.95 .I59 2010.

Rupp:2010:SIC

[Rup10] Karl Rupp. Symbolic integration at compile time in finite element methods. In Watt [Wat10], pages 347–354. ISBN 1-4503-0150-9. LCCN QA76.95 .I59 2010.

Russo:1992:CSA

[Rus92] Mark F. Russo. A combined symbolic/numeric approach for the integration of stiff nonlinear systems of ODE’s. In Wang [Wan92b], pages 80–88. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN QA76.95.I59 1992. URL http:

//www.acm.org:80/pubs/citations/proceedings/issac/143242/

p80-russo/. ACM order number: 505920.

Roy:1993:AGA

[RV93] Marie-Françoise Roy and T. Van Effelterre. Aspect graphs of algebraic surfaces. In Bronstein [Bro93], pages 135–143. ISBN 0-89791-604-2. LCCN QA 76.95 I59 1993. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/164081/p135-

roy/. ACM order number: 505930.

Roy:1996:CCS

[RV96] Marie-Françoise Roy and Nicolai Vorobjov. Computing the complexification of a semi-algebraic set. In Lakshman Y. N. [Lak96], pages 26–34. ISBN 0-89791-796-0. LCCN QA 76.95

I59 1996. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/236869/p26-roy/.

Reid:2005:SNC

[RVWW05] Greg Reid, Jan Verschelde, Allan Wittkopf, and Wenyuan Wu. Symbolic-numeric completion of differential systems by homotopy continuation. In Kauers [Kau05], pages 269–276. ISBN 1-59593-095-7. LCCN ???? ACM Order Number 505050.

Rayes:1994:PGS

[RW94] Mohamed Omar Rayes and P. S. Wang. Parallel GCD for sparse multivariate polynomials on shared memory multiprocessors. In Hong [Hon94], pages 326–335. ISBN 981-02-2040-5. LCCN QA76.642.I58 1994.

Reid:2000:DMS

[RW00] Gregory J. Reid and Allan D. Wittkopf. Determination of maximal symmetry groups of classes of differential equations. In Traverso [Tra00], pages 272–280. ISBN 1-58113-218-2. LCCN QA76.95.I59 2000. URL http:/

/www.acm.org/pubs/articles/proceedings/issac/345542/
p272-reid/p272-reid.pdf;   http://www.acm.org/pubs/
citations/proceedings/issac/345542/p272-reid/. ACM

order number 505000.

Rayes:1994:PSM

[RWW94] Mohamed Omar Rayes, Paul S. Wang, and Kenneth Weber. Parallelization of the sparse modular GCD algorithm for multivariate polynomials on shared memory multiprocessors. In ACM [ACM94], pages 66–73. ISBN 0-89791-638-7. LCCN QA76.95.I59 1994. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/190347/p66-rayes/.

Renault:2008:MMA

[RY08] Guénaël Renault and Kazuhiro Yokoyama. Multi-modular algorithm for computing the splitting field of a polynomial. In Jeffrey [Jef08], pages 247–254. ISBN 1-59593-904-0. LCCN ????

Rybowicz:1991:ACI

[Ryb91] Marc Rybowicz. An algorithm for computing integral bases of an algebraic function field. In Watt [Wat91], pages 157–166. ISBN 0-89791-437-6. LCCN QA 76.95

I59 1991. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/120694/p157-rybowicz/.

Sagraloff:2012:WNM

[Sag12] Michael Sagraloff. When Newton meets Descartes: a simple and fast algorithm to isolate the real roots of a polynomial. In van der Hoeven and van Hoeij [vdHvH12], pages 297–304. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Salvy:2005:FAA

[Sal05] Bruno Salvy. D-finiteness: algorithms and applications. In Kauers [Kau05], pages 2–3. ISBN 1-59593-095-7. LCCN ???? ACM Order Number 505050.

Santas:1993:TSC

[San93] Phillip S. Santas. A type system for computer algebra (abstract). In Bronstein [Bro93], page 77. ISBN 0-89791-604-2.

LCCN QA 76.95 I59 1993. URL http://www.acm.org:80/
pubs/citations/proceedings/issac/164081/p77-santas/.

ACM order number: 505930.

Santos:2004:TPA

[San04] Francisco Santos. Triangulations of polytopes and algebraic geometry. In Gutierrez [Gut04], page 2. ISBN 1-58113-827-X. LCCN QA76.95 .I57 2004.

Sasaki:1986:SAE

[Sas86] Tateaki Sasaki. Simplification of algebraic expression by multiterm rewriting rules. In Char [Cha86b], pages 115– 120. ISBN 0-89791-199-7 (paperback). LCCN QA155.7.E4

A281 1986. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/32439/p115-sasaki/. ACM order no.

505860.

Sasaki:2001:AMP

[Sas01] Tateaki Sasaki. Approximate multivariate polynomial factorization based on zero-sum relations. In Mourrain [Mou01], pages 284–291. ISBN 1-58113-417-7. LCCN QA76.95.I59 2001. ACM order number 505010.

Sasaki:2003:SCC

[Sas03] Tateaki Sasaki. The subresultant and clusters of close roots. In Senda [Sen03], pages 232–239. ISBN 1-58113-641-2. LCCN QA76.95. ACM order number 505030.

Sato:1998:NTC

[Sat98] Yosuke Sato. A new type of canonical Gröbner bases in polynomial rings over Von Neumann regular rings. In Gloor [Glo98], pages 317–321. ISBN 1-58113-002-3. LCCN

QA155.7.E4 E88 1998. URL http://www.acm.org/pubs/
citations/proceedings/issac/281508/p317-sato/.

Saunders:2001:BBM

[Sau01] B. D. Saunders. Black box methods for least squares problems. In Mourrain [Mou01], pages 297–302. ISBN 1-58113-417-7. LCCN QA76.95.I59 2001. ACM order number 505010.

Sodan:1994:SAP

[SB94] A. Sodan and H. Bi. A semi-automatic approach for parallelizing symbolic processing programs. In Hong [Hon94], pages 363–372. ISBN 981-02-2040-5. LCCN QA76.642.I58 1994.

Shi:2010:CSR

[SC10] Xiaoran Shi and Falai Chen. Computing the singularities of rational space curves. In Watt [Wat10], pages 171–178. ISBN 1-4503-0150-9. LCCN QA76.95 .I59 2010.

LaScala:1994:AC

[Sca94a] R. La Scala. An algorithm for complexes. In ACM [ACM94], pages 264–268. ISBN 0-89791-638-7. LCCN QA76.95.I59 1994.

Scala:1994:AC

[Sca94b] Roberto La Scala. An algorithm for complexes. In ACM [ACM94], pages 264–268. ISBN 0-89791-638-7. LCCN

QA76.95.I59 1994. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/190347/p264-la_scala/.

Schwarz:1989:FAL

[Sch89] F. Schwarz. A factorization algorithm for linear ordinary differential equations. In Gonnet [Gon89], pages 17–25. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989.

US$29.00. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/74540/p17-schwarz/. ACM order num-

ber: 505890. English and French.

Schlegel:1991:DRS

[Sch91a] H. Schlegel. Determination of the root system of semisimple Lie algebras from the Dynkin diagram. In Watt [Wat91], pages 239–240. ISBN 0-89791-437-6. LCCN QA 76.95

I59 1991. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/120694/p239-schlegel/.

Schmitt:1991:EAA

[Sch91b] Joacheim Schmitt. An embedding algorithm for algebraic congruence function fields. In Watt [Wat91], pages 187–188. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991. URL http://www.

acm.org:80/pubs/citations/proceedings/issac/120694/
p187-schmitt/.

Schonhage:1991:FRC

[Sch91c] Arnold Schönhage. Fast reduction and composition of binary quadratic forms. In Watt [Wat91], pages 128–133. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/120694/p128-
schonhage/.

Schwarz:1991:ETP

[Sch91d] Fritz Schwarz. Existence theorems for polynomial first integrals. In Watt [Wat91], pages 256–264. ISBN 0-89791-437-6. LCCN

QA 76.95 I59 1991. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/120694/p256-schwarz/.

Schwarz:1992:RCA

[Sch92] Fritz Schwarz. Reduction and completion algorithms for partial differential equations. In Wang [Wan92b], pages 49–56. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN QA76.95.I59 1992. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/143242/p49-schwarz/

. ACM order number: 505920.

Schreiner:1994:PPI

[Sch94] W. Schreiner. A para-functional programming interface for a parallel computer algebra package. In Hong [Hon94], pages 346– 355. ISBN 981-02-2040-5. LCCN QA76.642.I58 1994.

Schwartz:1995:SOO

[Sch95a] F. Schwartz. Symmetries of 2nd and 3rd order ODEs. In Levelt [Lev95], pages 16–25. ISBN 0-89791-699-9. LCCN QA 76.95 I59 1995. ACM order number: 505950.

Schwarz:1995:SSS

[Sch95b] Fritz Schwarz. Symmetries of 2^{nd} and 3^{rd} order ODE’s. In Levelt [Lev95], pages 16–25. ISBN 0-89791-699-9. LCCN QA 76.95

I59 1995. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/220346/p16-schwarz/. ACM order num-

ber: 505950.

Schwarz:1996:JBO

[Sch96] Fritz Schwarz. Janet bases of 2nd order ordinary differential equations. In Lakshman Y. N. [Lak96], pages 179–188. ISBN 0-89791-796-0. LCCN QA 76.95 I59 1996. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/236869/p179-
schwarz/.

Schicho:1998:RPR

[Sch98] Josef Schicho. Rational parametrization of real algebraic surfaces. In Gloor [Glo98], pages 302–308. ISBN 1-58113-002-3. LCCN QA155.7.E4 E88 1998. URL http://

www.acm.org/pubs/citations/proceedings/issac/281508/
p302-schicho/.

Schicho:2000:PPS

[Sch00] Josef Schicho. Proper parametrization of surfaces with a rational pencil. In Traverso [Tra00], pages 292–300. ISBN 1-58113-218-2. LCCN QA76.95.I59 2000. URL http://www.

acm.org/pubs/articles/proceedings/issac/345542/p292-
schicho/p292-schicho.pdf;   http://www.acm.org/pubs/
citations/proceedings/issac/345542/p292-schicho/. ACM

order number 505000.

Schicho:2002:SSP

[Sch02a] Josef Schicho. Simplification of surface parametrizations. In Mora [Mor02], pages 229–237. ISBN 1-58113-484-3. LCCN

QA76.95. URL http://www.lifl.fr/ISSAC2002/.

Schost:2002:DBL

[Sch02b] Éric Schost. Degree bounds and lifting techniques for triangular sets. In Mora [Mor02], pages 238–245. ISBN 1-58113-484-3.

LCCN QA76.95. URL http://www.lifl.fr/ISSAC2002/.

Schneider:2004:SSS

[Sch04] Carsten Schneider. Symbolic summation with single-nested sum extensions. In Gutierrez [Gut04], pages 282–289. ISBN 1-58113-827-X. LCCN QA76.95 .I57 2004.

Schneider:2005:FTM

[Sch05a] Carsten Schneider. Finding telescopers with minimal depth for indefinite nested sum and product expressions. In Kauers [Kau05], pages 285–292. ISBN 1-59593-095-7. LCCN ???? ACM Order Number 505050.

Schost:2005:MPS

[Sch05b] Éric Schost. Multivariate power series multiplication. In Kauers [Kau05], pages 293–300. ISBN 1-59593-095-7. LCCN ???? ACM Order Number 505050.

Schwarz:2007:LDL

[Sch07] F. Schwarz. Loewy decomposition of linear differential equations. In Brown [Bro07], pages 389–390. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Schweighofer:2009:DCS

[Sch09] Markus Schweighofer. Describing convex semialgebraic sets by linear matrix inequalities. In May [May09], pages 385–386. ISBN 1-60558-609-9. LCCN ????

Scheiblechner:2012:ERC

[Sch12] Peter Scheiblechner. Effective de Rham cohomology: the hypersurface case. In van der Hoeven and van Hoeij [vdHvH12], pages 305–310. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Shavlik:1986:CUG

[SD86] Jude W. Shavlik and Gerald F. DeJong. Computer understanding and generalization of symbolic mathematical calculations: a case study in physics problem solving. In Char [Cha86b], pages 148–153. ISBN 0-89791-199-7 (paperback). LCCN QA155.7.E4

A281 1986. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/32439/p148-shavlik/. ACM order no.

505860.

Smith:2007:ADA

[SDJ07] Jacob Smith, Gabriel Dos Reis, and Jaakko Järvi. Algorithmic differentiation in Axiom. In Brown [Bro07], pages 347–354.

ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Slavici:2010:FML

[SDKC10] Vlad Slavici, Xin Dong, Daniel Kunkle, and Gene Cooperman. Fast multiplication of large permutations for disk, flash memory and RAM. In Watt [Wat10], pages 355–362. ISBN 1-4503-0150-9. LCCN QA76.95 .I59 2010.

Schost:2011:IPI

[SE11] Éric Schost and Ioannis Z. Emiris, editors. ISSAC 2011: Proceedings of the 2011 International Symposium on Symbolic and Algebraic Computation, June 7–11, 2011, San Jose, CA, USA. ACM Press, New York, NY 10036, USA, 2011. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Sedjelmaci:2001:PLE

[Sed01a] Sidi Mohammed Sedjelmaci. On a parallel Lehmer–Euclid GCD algorithm. In Mourrain [Mou01], pages 303–308. ISBN 1-58113-417-7. LCCN QA76.95.I59 2001. ACM order number 505010.

Sedoglavic:2001:PAT

[Sed01b] Alexandre Sedoglavic. A probabilistic algorithm to test local algebraic observability in polynomial time. In Mourrain [Mou01], pages 309–317. ISBN 1-58113-417-7. LCCN QA76.95.I59 2001. ACM order number 505010.

Sekigawa:2007:RFR

[Sek07] Hiroshi Sekigawa. On real factors of real interval polynomials. In Brown [Bro07], pages 331–338. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Sekigawa:2008:NPZ

[Sek08] Hiroshi Sekigawa. The nearest polynomial with a zero in a given domain from a geometrical viewpoint. In Jeffrey [Jef08], pages 287–294. ISBN 1-59593-904-0. LCCN ????

Senda:2003:IPI

[Sen03] J. Rafael Senda, editor. ISSAC 2003: Proceedings of the 2003 International Symposium on Symbolic and Algebraic Computation, August 3–6, 2003, Drexel University, Philadelphia, PA, USA. ACM Press, New York, NY 10036, USA, 2003. ISBN 1-58113-641-2. LCCN QA76.95. ACM order number 505030.

Sergeraert:1989:NRN

[Ser89] F. Sergeraert. From a noncomputability result to new interesting definitions and computability results. In Gianni [Gia89], pages 26–32. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Seress:2012:CCR

[Ser12] Ákos Seress. Construction of 2-closed M-representations. In van der Hoeven and van Hoeij [vdHvH12], pages 311–318. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Seymour:1986:CCM

[Sey86] Harlan R. Seymour. Conform: a conformal mapping system. In Char [Cha86b], pages 163–168. ISBN 0-89791-199-7 (paperback). LCCN QA155.7.E4 A281 1986. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/32439/p163-

seymour/. ACM order no. 505860.

Scott:1990:SAM

[SF90] T. C. Scott and G. J. Fee. Some applications of Maple symbolic computation to scientific and engineering problems. In Watanabe and Nagata [WN90], pages 302–303. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95

.I57 1990. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/96877/p302-scott/.

S:1994:AGG

[SF94] Oscar E. Ruiz S. and Placid M. Ferreira. Algebraic geometry and group theory in geometric constraint satisfaction. In ACM [ACM94], pages 224–233. ISBN 0-89791-638-7. LCCN QA76.95.I59 1994. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/190347/p224-ruiz_s.
/.

Salem:2004:FPP

[SGL04] Fatima Abu Salem, Shuhong Gao, and Alan G. B. Lauder. Factoring polynomials via polytopes. In Gutierrez [Gut04], pages 4–11. ISBN 1-58113-827-X. LCCN QA76.95 .I57 2004.

Shackell:1989:AEO

[Sha89a] J. Shackell. Asymptotic estimation of oscillating functions using an interval calculus. In Gianni [Gia89], pages 481–489. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Shackell:1989:DAF

[Sha89b] J. Shackell. A differential-equations approach to functional equivalence. In Gonnet [Gon89], pages 7–10. ISBN 0-89791-325-

6. LCCN QA76.95.I59 1989. US$29.00. ACM order number: 505890. English and French.

Shackle:1989:DAF

[Sha89c] J. Shackle. A differential-equations approach to functional equivalence. In Gonnet [Gon89], pages 7–10. ISBN 0-89791-325-

6. LCCN QA76.95.I59 1989. US$29.00. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/74540/p7-shackle/

. ACM order number: 505890. English and French.

Shackell:1993:NEH

[Sha93] John Shackell. Nested expansions and Hardy fields. In Bronstein [Bro93], pages 234–238. ISBN 0-89791-604-2. LCCN QA 76.95

I59 1993. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/164081/p234-shackell/. ACM order

number: 505930.

Shackell:1999:SPR

[Sha99] John Shackell. Star products and the representation of asymptotic growth. In Dooley [Doo99], pages 97–104. ISBN 1-58113-073-2. LCCN QA76.95 .I57 1999. URL http://www.acm.org/

pubs/contents/proceedings/issac/309831/.

Shaska:2003:DAG

[Sha03] Tanush Shaska. Determining the automorphism group of a hyperelliptic curve. In Senda [Sen03], pages 248–254. ISBN 1-58113-641-2. LCCN QA76.95. ACM order number 505030.

Sharma:2007:CRR

[Sha07] Vikram Sharma. Complexity of real root isolation using continued fractions. In Brown [Bro07], pages 339–346. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Shirayanagi:1990:IPF

[Shi90] K. Shirayanagi. On the isomorphism problem for finitedimensional binomial algebras. In Watanabe and Nagata [WN90], pages 106–111. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95 .I57 1990. URL

http://www.acm.org:80/pubs/citations/proceedings/issac/
96877/p106-shirayanagi/.

Shoup:1991:FDA

[Sho91] Victor Shoup. A fast deterministic algorithm for factoring polynomials over finite fields of small characteristic. In Watt [Wat91], pages 14–21. ISBN 0-89791-437-6. LCCN QA 76.95

I59 1991. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/120694/p14-shoup/.

Shoup:1999:ECM

[Sho99] Victor Shoup. Efficient computation of minimal polynomials in algebraic extensions of finite fields. In Dooley [Doo99], pages 53–58. ISBN 1-58113-073-2. LCCN QA76.95 .I57 1999. URL

http://www.acm.org/pubs/contents/proceedings/issac/
309831/.

Siegl:1993:PAS

[Sie93] K. Siegl. Parallelizing algorithms for symbolic computation using ∥Maple∥. In ACM [ACM93], pages 179–186. ISBN 0-89791-589-5. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.642.A27 1993.

Siegl:1994:PFT

[Sie94] K. Siegl. A parallel factorization tree Gröbner basis algorithm. In Hong [Hon94], pages 356–362. ISBN 981-02-2040-5. LCCN QA76.642.I58 1994.

Sims:1997:CSA

[Sim97] Charles C. Sims. Computing with subgroups of automorphism groups of finite groups. In Küchlin [Küc97], pages 400–403. ISBN 0-89791-875-4. LCCN QA76.95. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/258726/p400-
sims/.

Sims:1998:FMG

[Sim98] Charles C. Sims. Fast multiplication and growth in groups. In Gloor [Glo98], pages 165–170. ISBN 1-58113-002-3. LCCN

QA155.7.E4 E88 1998. URL http://www.acm.org/pubs/
citations/proceedings/issac/281508/p165-sims/.

Singer:1989:LFI

[Sin89] M. F. Singer. Liouvillian first integrals of differential equations. In Gianni [Gia89], pages 57–63. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Sincovec:1993:PSS

[Sin93] Richard F. Sincovec, editor. Proceedings of the Sixth SIAM Conference on Parallel Processing for Scientific Computing, Norfolk, VA, March, 1993. SIAM Press, Philadelphia, PA, USA, 1993. ISBN 0-89871-315-3. LCCN QA76.58.S55 1993. Two volumes.

Sit:1991:TPL

[Sit91] William Y. Sit. A theory for parametric linear systems. In Watt [Wat91], pages 112–121. ISBN 0-89791-437-6. LCCN QA 76.95

I59 1991. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/120694/p112-sit/.

Slavici:2012:EPM

[SKCL12] Vlad Slavici, Daniel Kunkle, Gene Cooperman, and Stephen Linton. An efficient programming model for memory-intensive recursive algorithms using parallel disks. In van der Hoeven and van Hoeij [vdHvH12], pages 327–334. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Sendra:1993:EAH

[SL93] Juan R. Sendra and Juan Llovet. Efficient algorithms for Hankel matrices over Z(x_{1}, …, x_{r}). In Bronstein [Bro93], pages 201–208. ISBN 0-89791-604-2. LCCN QA 76.95

I59 1993. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/164081/p201-sendra/. ACM order num-

ber: 505930.

Storjohann:1995:PRP

[SL95] Arne Storjohann and George Labahn. Preconditioning of rectangular polynomial matrices for efficient Hermite Normal Form computation. In Levelt [Lev95], pages 119–125. ISBN 0-89791-699-9. LCCN QA 76.95 I59 1995. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/220346/p119-

storjohann/. ACM order number: 505950.

Storjohann:1996:AFC

[SL96] Arne Storjohann and George Labahn. Asymptotically fast computation of the Hermite Normal Forms of integer matrices. In Lakshman Y. N. [Lak96], pages 259–266. ISBN 0-89791-796-0. LCCN QA 76.95 I59 1996. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/236869/p259-
storjohann/.

Saunders:1989:PIC

[SLA89] B. D. Saunders, H. R. Lee, and S. K. Abdali. A parallel implementation of the cylindrical algebraic decomposition algorithm. In Gonnet [Gon89], pages 298–307. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989. US$29.00. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/74540/p298-

saunders/. ACM order number: 505890. English and French.

Shemyakova:2008:MFL

[SM08] Ekaterina Shemyakova and Elizabeth L. Mansfield. Moving frames for Laplace invariants. In Jeffrey [Jef08], pages 295–302. ISBN 1-59593-904-0. LCCN ????

Smedley:1989:NMA

[Sme89] T. J. Smedley. A new modular algorithm for computation of algebraic number polynomial gcds. In Gonnet [Gon89], pages 91–94. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989.

US$29.00. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/74540/p91-smedley/. ACM order num-

ber: 505890. English and French.

Smedley:1990:DAD

[Sme90] T. J. Smedley. Detecting algebraic dependencies between unnested radicals (abstract). In Watanabe and Nagata [WN90], pages 292–293. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95 .I57 1990. URL http://

www.acm.org:80/pubs/citations/proceedings/issac/96877/
p292-smedley/.

Soiffer:1995:MTM

[Soi95] Neil Soiffer. Mathematical typesetting in Mathematica. In Levelt [Lev95], pages 140–149. ISBN 0-89791-699-9. LCCN QA 76.95 I59 1995. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/220346/p140-

soiffer/. ACM order number: 505950.

Sommeling:1994:CCI

[Som94] Ron Sommeling. Characteristic classes for irregular singularities. In ACM [ACM94], pages 163–168. ISBN 0-89791-638-7. LCCN QA76.95.I59 1994. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/190347/p163-
sommeling/.

Sorenson:1995:ALE

[Sor95] Jonathan Sorenson. An analysis of Lehmer’s Euclidean GCD algorithm. In Levelt [Lev95], pages 254–258. ISBN 0-89791-699-9. LCCN QA 76.95 I59 1995. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/220346/p254-

sorenson/. ACM order number: 505950.

SchulzePillot:1991:ACG

[SP91a] R. Schulze-Pillot. An algorithm for computing genera of ternary and quaternary quadratic forms. In Watt [Wat91], pages 134– 143. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991.

Schulze-Pillot:1991:ACG

[SP91b] Rainer Schulze-Pillot. An algorithm for computing genera of ternary and quaternary quadratic forms. In Watt [Wat91], pages 134–143. ISBN 0-89791-437-6. LCCN QA 76.95

I59 1991. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/120694/p134-schulze-pillot/.

Siebert-Roch:1989:PAH

[SR89] F. Siebert-Roch. Parallel algorithms for Hermite normal form of an integer matrix. In Gonnet [Gon89], pages 317–321. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989.

US$29.00. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/74540/p317-siebert-roch/. ACM order

number: 505890. English and French.

Saenz:1994:SMP

[SRHW94] F. Saenz, J. J. Ruz, W. Hans, and S. Winkler. A stackbased machine for parallel execution of Babel programs. In Hong [Hon94], pages 336–345. ISBN 981-02-2040-5. LCCN QA76.642.I58 1994.

Smith:1986:MUI

[SS86] C. J. Smith and N. Soiffer. MathScribe: a user interface for computer algebra systems. In Char [Cha86b], pages 7– 12. ISBN 0-89791-199-7 (paperback). LCCN QA155.7.E4

A281 1986. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/32439/p7-smith/.  ACM order no.

505860.

Surguladze:1991:APC

[SS91] Levan R. Surguladze and Mark A. Samuel. Algebraic perturbative calculations in high energy physics. Methods, algorithms, computer programs and physical applications. In Watt [Wat91], pages 439–447. ISBN 0-89791-437-6. LCCN QA 76.95

I59 1991. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/120694/p439-surguladze/.

Salvy:1992:AEF

[SS92] Bruno Salvy and John Shackell. Asymptotic expansions of functional inverses. In Wang [Wan92b], pages 130–137. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN QA76.95.I59 1992. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/143242/p130-salvy/.

ACM order number: 505920.

Shevchenko:1993:SRP

[SS93] Ivan I. Shevchenko and Andrej G. Sokolsky. Studies of regular precessions of a symmetric satellite by means of computer algebra. In Bronstein [Bro93], pages 65–67. ISBN 0-89791-604-2. LCCN QA 76.95 I59 1993. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/164081/p65-

shevchenko/. ACM order number: 505930.

Schonert:1994:FBI

[SS94] Martin Schönert and Ákos Seress. Finding blocks of imprimitivity in small-base groups in nearly linear time. In ACM [ACM94], pages 154–157. ISBN 0-89791-638-7. LCCN

QA76.95.I59 1994. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/190347/p154-schonert/.

Semmler:1995:NUH

[SS95] Klaus-Dieter Semmler and Mika Seppälä. Numerical uniformization of hyperelliptic curves. In Levelt [Lev95], pages 208–215. ISBN 0-89791-699-9. LCCN QA 76.95

I59 1995. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/220346/p208-semmler/. ACM order

number: 505950.

Sato:2001:DCG

[SS01] Yosuke Sato and Akira Suzuki. Discrete comprehensive Gröbner bases. In Mourrain [Mou01], pages 292–296. ISBN 1-58113-417-7. LCCN QA76.95.I59 2001. ACM order number 505010.

Suzuki:2002:AAC

[SS02] Akira Suzuki and Yosuke Sato. An alternative approach to comprehensive Gröbner bases. In Mora [Mor02], pages 255–261. ISBN 1-58113-484-3. LCCN QA76.95. URL http://www.lifl.

fr/ISSAC2002/.

Seidl:2003:GPO

[SS03] Andreas Seidl and Thomas Sturm. A generic projection operator for partial cylindrical algebraic decomposition. In Senda [Sen03], pages 240–247. ISBN 1-58113-641-2. LCCN QA76.95. ACM order number 505030.

Sekigawa:2006:LRM

[SS06a] Hiroshi Sekigawa and Kiyoshi Shirayanagi. Locating real multiple zeros of a real interval polynomial. In Trager [Tra06], pages 310–317. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

Sexton:2006:AMS

[SS06b] Alan Sexton and Volker Sorge. Abstract matrices in symbolic computation. In Trager [Tra06], pages 318–325. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

Suzuki:2006:SAC

[SS06c] Akira Suzuki and Yosuke Sato. A simple algorithm to compute comprehensive Gröbner bases using Gröbner bases. In Trager [Tra06], pages 326–331. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

Sato:2009:CIR

[SS09] Yosuke Sato and Akira Suzuki. Computation of inverses in residue class rings of parametric polynomial ideals. In May [May09], pages 311–316. ISBN 1-60558-609-9. LCCN ????

Sarkar:2011:NRR

[SS11] Soumojit Sarkar and Arne Storjohann. Normalization of row reduced matrices. In Schost and Emiris [SE11], pages 297–304. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Sexton:2009:CAM

[SSW09] Alan P. Sexton, Volker Sorge, and Stephen M. Watt. Computing with abstract matrix structures. In May [May09], pages 325– 332. ISBN 1-60558-609-9. LCCN ????

Stetter:1998:SSP

[ST98] Hans J. Stetter and Günther H. Thallinger. Singular systems of polynomials. In Gloor [Glo98], pages 9–16. ISBN 1-58113-002-3. LCCN QA155.7.E4 E88 1998. URL http://www.acm.org/

pubs/citations/proceedings/issac/281508/p9-stetter/.

Strzebonski:2011:URR

[ST11a] Adam Strzebonski and Elias Tsigaridas. Univariate real root isolation in an extension field. In Schost and Emiris [SE11], pages 321–328. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Sturm:2011:VSU

[ST11b] Thomas Sturm and Ashish Tiwari. Verification and synthesis using real quantifier elimination. In Schost and Emiris [SE11], pages 329–336. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Strzebonski:2012:URR

[ST12] Adam Strzeboński and Elias P. Tsigaridas. Univariate real root isolation in multiple extension fields. In van der Hoeven and van Hoeij [vdHvH12], pages 343–350. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Stachniak:1990:RPS

[Sta90] Z. Stachniak. Resolution proof systems with weak transformation rules. In Watanabe and Nagata [WN90], pages 38–43. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN

QA76.95 .I57 1990. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/96877/p38-stachniak/.

Stetter:1996:AZC

[Ste96] Hans J. Stetter. Analysis of zero clusters in multivariate polynomial systems. In Lakshman Y. N. [Lak96], pages 127–136. ISBN 0-89791-796-0. LCCN QA 76.95 I59 1996. URL http://www.

acm.org:80/pubs/citations/proceedings/issac/236869/
p127-stetter/.

Stetter:1997:SPS

[Ste97] Hans J. Stetter. Stabilization of polynomial systems solving with Gröbner bases. In Küchlin [Küc97], pages 117–124. ISBN 0-89791-875-4. LCCN QA76.95. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/258726/p117-
stetter/.

Stein:2008:CWC

[Ste08] William A. Stein. Can we create a viable free open source alternative to Magma, Maple, Mathematica and Matlab? In Jeffrey [Jef08], pages 5–6. ISBN 1-59593-904-0. LCCN ????

Steffy:2013:ELI

[Ste13] Daniel E. Steffy. Exact linear and integer programming: tutorial abstract. In Monagan et al. [MCG13], pages 11–12. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

Stifter:1989:GRM

[Sti89] S. Stifter. A generalization of the Roider method to solve the robot collision problem in 3D. In Gianni [Gia89], pages 332–343. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Stifter:1993:GTP

[Sti93] Sabine Stifter. Geometry theorem proving in vector spaces by means of Gröbner bases. In Bronstein [Bro93], pages 301–310. ISBN 0-89791-604-2. LCCN QA 76.95 I59 1993. URL http://

www.acm.org:80/pubs/citations/proceedings/issac/164081/

p301-stifter/. ACM order number: 505930.

Stiller:1997:SCO

[Sti97] Peter F. Stiller. Symbolic computation of object/image equations. In Küchlin [Küc97], pages 359–364. ISBN 0-89791-

875-4. LCCN QA76.95. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/258726/p359-stiller/.

Sperber:1997:DPE

[STK97] Michael Sperber, Peter Thiemann, and Hervert Klaeren. Distributed partial evaluation. In Küchlin [Küc97], pages 80–87. ISBN 0-89791-875-4. LCCN QA76.95. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/266670/p80-
sperber/.

Storjohann:1996:NOA

[Sto96] Arne Storjohann. Near optimal algorithms for computing Smith normal forms of integer matrices. In Lakshman Y. N. [Lak96], pages 267–274. ISBN 0-89791-796-0. LCCN QA 76.95

I59 1996. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/236869/p267-storjohann/.

Storjohann:1997:SEG

[Sto97] Arne Storjohann. A solution to the extended GCD problem with applications. In Küchlin [Küc97], pages 109–116. ISBN 0-89791-

875-4. LCCN QA76.95. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/258726/p109-storjohann/.

Storjohann:1998:I

[Sto98] Arne Storjohann. An O(n^{3}) algorithm for the Frobenius normal form. In Gloor [Glo98], pages 101–105. ISBN 1-58113-002-3. LCCN QA155.7.E4 E88 1998. URL http://

www.acm.org/pubs/citations/proceedings/issac/281508/
p101-storjohann/.

Storjohann:2002:HOL

[Sto02] Arne Storjohann. High-order lifting. In Mora [Mor02], pages 246–254. ISBN 1-58113-484-3. LCCN QA76.95. URL http://

www.lifl.fr/ISSAC2002/.

Storjohann:2009:IMR

[Sto09] Arne Storjohann. Integer matrix rank certification. In May [May09], pages 333–340. ISBN 1-60558-609-9. LCCN ????

Strzebonski:2008:RRI

[Str08] Adam Strzebonski. Real root isolation for exp-log functions. In Jeffrey [Jef08], pages 303–314. ISBN 1-59593-904-0. LCCN ????

Strzebonski:2009:RRI

[Str09] Adam Strzebonski. Real root isolation for tame elementary functions. In May [May09], pages 341–350. ISBN 1-60558-609-9. LCCN ????

Strzebonski:2010:CSS

[Str10] Adam Strzeboński. Computation with semialgebraic sets represented by cylindrical algebraic formulas. In Watt [Wat10], pages 61–68. ISBN 1-4503-0150-9. LCCN QA76.95 .I59 2010.

Strassen:2012:ASM

[Str12a] Volker Strassen. Asymptotic spectrum and matrix multiplication. In van der Hoeven and van Hoeij [vdHvH12], pages 6–7. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Strzebonski:2012:SPS

[Str12b] Adam Strzeboński. Solving polynomial systems over semialgebraic sets represented by cylindrical algebraic formulas. In van der Hoeven and van Hoeij [vdHvH12], pages 335–342. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Singer:1992:LST

[SU92] Michael F. Singer and Felix Ulmer. Liouvillian solutions of third order linear differential equations: new bounds and necessary conditions. In Wang [Wan92b], pages 57–62. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN

QA76.95.I59 1992. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/143242/p57-singer/. ACM

order number: 505920.

Sudan:2008:AAC

[Sud08] Madhu Sudan. Algebraic algorithms and coding theory. In Jeffrey [Jef08], page 337. ISBN 1-59593-904-0. LCCN ????

Sullivant:2012:AS

[Sul12] Seth Sullivant. Algebraic statistics. In van der Hoeven and van Hoeij [vdHvH12], page 11. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Storjohann:2005:CRS

[SV05] Arne Storjohann and Gilles Villard. Computing the rank and a small nullspace basis of a polynomial matrix. In Kauers [Kau05], pages 309–316. ISBN 1-59593-095-7. LCCN ???? ACM Order Number 505050.

Sottile:2010:SSP

[SVV10] Frank Sottile, Ravi Vakil, and Jan Verschelde. Solving Schubert problems with Littlewood–Richardson homotopies. In Watt [Wat10], pages 179–186. ISBN 1-4503-0150-9. LCCN QA76.95 .I59 2010.

Sharma:1989:SDA

[SW89] N. Sharma and P. S. Wang. Symbolic derivation and automatic generation of parallel routines for finite element analysis. In Gianni [Gia89], pages 33–56. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Sendra:2001:CDR

[SW01] J. Rafael Sendra and Franz Winkler. Computation of the degree of rational maps between curves. In Mourrain [Mou01], pages 317–322. ISBN 1-58113-417-7. LCCN QA76.95.I59 2001. ACM order number 505010.

Saunders:2004:SNF

[SW04] David Saunders and Zhendong Wan. Smith Normal Form of dense integer matrices fast algorithms into practice. In Gutierrez [Gut04], pages 274–281. ISBN 1-58113-827-X. LCCN QA76.95 .I57 2004.

Sevilla:2010:PIR

[SW10] David Sevilla and Daniel Wachsmuth. Polynomial integration on regions defined by a triangle and a conic. In Watt [Wat10], pages 163–170. ISBN 1-4503-0150-9. LCCN QA76.95 .I59 2010.

Sun:2011:GCS

[SW11] Yao Sun and Dingkang Wang. A generalized criterion for signature related Gröbner basis algorithms. In Schost and Emiris [SE11], pages 337–344. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Sun:2012:SBA

[SWMZ12] Yao Sun, Dingkang Wang, Xiaodong Ma, and Yang Zhang. A signature-based algorithm for computing Gröbner bases in solvable polynomial algebras. In van der Hoeven and van Hoeij [vdHvH12], pages 351–358. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Saunders:2011:NSE

[SWY11] B. David Saunders, David Harlan Wood, and Bryan S. Youse. Numeric-symbolic exact rational linear system solver. In Schost and Emiris [SE11], pages 305–312. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

She:2011:AAA

[SXZ11] Zhikun She, Bai Xue, and Zhiming Zheng. Algebraic analysis on asymptotic stability of continuous dynamical systems. In Schost and Emiris [SE11], pages 313–320. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Sasaki:1998:ACE

[SY98] Tateaki Sasaki and Satoshi Yamaguchi. An analysis of cancellation error in multivariate Hensel construction with floatingpoint number arithmetic. In Gloor [Glo98], pages 1–8. ISBN 1-58113-002-3. LCCN QA155.7.E4 E88 1998. URL http://

www.acm.org/pubs/citations/proceedings/issac/281508/
p1-sasaki/.

Saunders:2009:LMS

[SY09] B. David Saunders and Bryan S. Youse. Large matrix, small rank. In May [May09], pages 317–324. ISBN 1-60558-609-9. LCCN ????

Sagraloff:2011:SEE

[SY11] Michael Sagraloff and Chee K. Yap. A simple but exact and efficient algorithm for complex root isolation. In Schost and Emiris [SE11], pages 353–360. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Sharma:2012:NOT

[SY12] Vikram Sharma and Chee K. Yap. Near optimal tree size bounds on a simple real root isolation algorithm. In van der Hoeven and van Hoeij [vdHvH12], pages 319–326. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Stein:1991:ADR

[SZ91] Andreas Stein and Horst Günter Zimmer. An algorithm for determining the regulator and the fundamental unit of a hyperelliptic congruence function field. In Watt [Wat91], pages 183–184. ISBN 0-89791-437-6. LCCN QA 76.95

I59 1991. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/120694/p183-stein/.

Sturm:2010:PQS

[SZ10] Thomas Sturm and Christoph Zengler. Parametric quantified SAT solving. In Watt [Wat10], pages 77–84. ISBN 1-4503-0150-9. LCCN QA76.95 .I59 2010.

Szanto:1997:CWR

[Szá97] Ágnes Szántó. Complexity of the Wu–Ritt decomposition. In Küchlin [Küc97], pages 139–149. ISBN 0-89791-875-4. LCCN

QA76.95. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/266670/p139-szanto/.

Szanto:2011:HSN

[Sza11] Agnes Szanto. Hybrid symbolic-numeric methods for the solution of polynomial systems: tutorial overview. In Schost and Emiris [SE11], pages 9–10. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Takayama:1990:ACI

[Tak90a] N. Takayama. An algorithm of constructing the integral of a module — an infinite dimensional analog of Gröbner basis. In Watanabe and Nagata [WN90], pages 206–211. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN

QA76.95 .I57 1990. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/96877/p206-takayama/.

Takayama:1990:GBI

[Tak90b] N. Takayama. Gröbner basis, integration and transcendental functions. In Watanabe and Nagata [WN90], pages 152– 156. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95 .I57 1990. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/96877/p152-
takayama/.

Takesue:1994:PSP

[Tak94] M. Takesue. Parallel symbolic processing with the distributed lists. In Hong [Hon94], pages 373–381. ISBN 981-02-2040-5. LCCN QA76.642.I58 1994.

Tao:1990:SAM

[Tao90] Qingsheng Tao. Symbolic and algebraic manipulation for formulae of interpolation and quadrature. In Watanabe and Nagata [WN90], page 306. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95

.I57 1990. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/96877/p306-tao/.

Todd:1989:SAP

[TC89] P. H. Todd and G. W. Cherry. Symbolic analysis of planar drawings. In Gianni [Gia89], pages 344–355. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Teo:1997:DEP

[TCT97] Yong Meng Teo, Wei-Ngan Chin, and Soon Huat Tan. Deriving efficient parallel programs for complex recurrences. In Küchlin [Küc97], pages 101–110. ISBN 0-89791-875-4. LCCN

QA76.95. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/266670/p101-teo/.

Traverso:1989:EGB

[TD89] C. Traverso and L. Donati. Experimenting the Gröbner basis algorithm with the A1PI system. In Gonnet [Gon89], pages 192–198. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989.

US$29.00. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/74540/p192-traverso/. ACM order

number: 505890. English and French.

Tan:1990:OTS

[TD90] H. Q. Tan and X. Dong. Optimization techniques for symbolic equation solver in engineering applications. In Watanabe and Nagata [WN90], page 305. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95

.I57 1990. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/96877/p305-tan/.

Teitelbaum:1989:CCR

[Tei89] J. Teitelbaum. On the computational complexity of the resolution of plane curve singularities. In Gianni [Gia89], pages 285–292. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Terui:2009:IMC

[Ter09] Akira Terui. An iterative method for calculating approximate GCD of univariate polynomials. In May [May09], pages 351– 358. ISBN 1-60558-609-9. LCCN ????

Thomas:1996:SCI

[Tho96] G. Thomas. Symbolic computation of the index of quasilinear differential-algebraic equations. In Lakshman Y. N.

[Lak96], pages 196–203. ISBN 0-89791-796-0. LCCN QA 76.95

I59 1996. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/236869/p196-thomas/.

Thome:2001:FCL

[Tho01] Emmanuel Thomé. Fast computation of linear generators for matrix sequences and application to the block Wiedemann algorithm. In Mourrain [Mou01], pages 323–331. ISBN 1-58113-417-7. LCCN QA76.95.I59 2001. ACM order number 505010.

Thuemmel:1996:CCT

[Thu96a] A. Thuemmel. Computing character tables of p-groups. In Lakshman Y. N. [Lak96], pages 150–154. ISBN 0-89791-796-0. LCCN QA 76.95 I59 1996.

Thummel:1996:CCT

[Thü96b] Andreas Thümmel. Computing character tables of p-groups. In Lakshman Y. N. [Lak96], pages 150–154. ISBN 0-89791-796-0. LCCN QA 76.95 I59 1996. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/236869/p150-
thummel/.

Tiwari:2010:TRV

[Tiw10] Ashish Tiwari. Theory of reals for verification and synthesis of hybrid dynamical systems. In Watt [Wat10], pages 5–6. ISBN 1-4503-0150-9. LCCN QA76.95 .I59 2010.

Tong:1994:IDC

[TL94] B.-M. Tong and H.-F. Leung. Implementation of a data-parallel concurrent constraint programming system. In Hong [Hon94], pages 382–393. ISBN 981-02-2040-5. LCCN QA76.642.I58 1994.

Todd:1994:SSA

[TMH94] Philip H. Todd, Robin J. Y. McLeod, and Marcia Harris. A system for the symbolic analysis of problems in engineering mechanics. In ACM [ACM94], pages 84–89. ISBN 0-89791-638-7. LCCN QA76.95.I59 1994. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/190347/p84-todd/.

Tournier:2005:ADS

[Tou05] Laurent Tournier. Approximation of dynamical systems using s-systems theory: application to biological systems. In Kauers [Kau05], pages 317–324. ISBN 1-59593-095-7. LCCN ???? ACM Order Number 505050.

Traverso:1989:GTA

[Tra89] C. Traverso. Gröbner trace algorithms. In Gianni [Gia89], pages 125–138. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Tran:1998:PCG

[Tra98] Q.-N. Tran. Parallel computation and Gröbner bases: An application for converting bases with the Gröbner walk. In Buchberger and Winkler [BW98], pages 519–531. ISBN 0-521-63298-6. LCCN QA251.3.G76 1998.

Traverso:2000:IAU

[Tra00] Carlo Traverso, editor. ISSAC 2000: 7–9 August 2000, University of St. Andrews, Scotland: proceedings of the 2000 International Symposium on Symbolic and Algebraic Computation. ACM Press, New York, NY 10036, USA, 2000. ISBN 1-58113-218-2. LCCN QA76.95.I59 2000. URL http://www.acm.

org/pubs/contents/proceedings/issac/345542/. ACM or-

der number 505000.

Trager:2006:PIS

[Tra06] Barry Trager, editor. Proceedings of the 2006 International Symposium on Symbolic and Algebraic Computation, Genoa, Italy July 09–12, 2006. ACM Press, New York, NY 10036, USA, 2006. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

Tsarev:1996:ACE

[Tsa96] S. P. Tsarev. An algorithm for complete enumeration of all factorizations of a linear ordinary differential operator. In Lakshman Y. N. [Lak96], pages 226–231. ISBN 0-89791-796-0. LCCN

QA 76.95 I59 1996. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/236869/p226-tsarev/.

Tsarev:1997:SMI

[Tsa97] S. P. Tsarev. Symbolic manipulation of integrodifferential expressions and factorization of linear ordinary differential operators over transcendental extensions of a differential field. In Küchlin [Küc97], pages 310–315. ISBN 0-89791-875-4. LCCN

QA76.95. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/258726/p310-tsarev/.

Tsarev:1999:FNO

[Tsa99] S. P. Tsarev. On factorization of nonlinear ordinary differential equations. In Dooley [Doo99], pages 159–164. ISBN 1-58113-073-2. LCCN QA76.95 .I57 1999. URL http://www.acm.org/

pubs/contents/proceedings/issac/309831/.

Tsarev:2005:GLT

[Tsa05] Sergey P. Tsarev. Generalized Laplace transformations and integration of hyperbolic systems of linear partial differential equations. In Kauers [Kau05], pages 325–331. ISBN 1-59593-095-7. LCCN ???? ACM Order Number 505050.

Tsarev:2010:TFP

[Tsa10] S. P. Tsarev. Transformation and factorization of partial differential systems: applications to stochastic systems. In Watt [Wat10], pages 11–12. ISBN 1-4503-0150-9. LCCN QA76.95 .I59 2010.

Trenkov:1991:ARS

[TSD91] I. Trenkov, M. Spiridonova, and M. Daskalova. An application of the REDUCE system for solving a mathematical geodesy problem. In Watt [Wat91], pages 448–449. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/120694/p448-
trenkov/.

Turner:2005:PSB

[Tur05] William J. Turner. Preconditioners for singular black box matrices. In Kauers [Kau05], pages 332–339. ISBN 1-59593-095-7. LCCN ???? ACM Order Number 505050.

Turner:2006:BWR

[Tur06] William J. Turner. A block Wiedemann rank algorithm. In Trager [Tra06], pages 332–339. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

Trevisan:1991:PFU

[TW91] Vilmar Trevisan and Paul Wang. Practical factorization of univariate polynomials over finite fields. In Watt [Wat91], pages 22–31. ISBN 0-89791-437-6. LCCN QA 76.95

I59 1991. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/120694/p22-trevisan/.

Wu:2005:FKT

[tW05] Wen tsun Wu. On a finite kernel theorem for polynomial-type optimization problems and some of its applications. In Kauers [Kau05], page 4. ISBN 1-59593-095-7. LCCN ???? ACM Order Number 505050.

Traverso:2002:NSS

[TZ02] Carlo Traverso and Alberto Zanoni. Numerical stability and stabilization of Gröbner basis computation. In Mora [Mor02], pages 262–269. ISBN 1-58113-484-3. LCCN QA76.95. URL

http://www.lifl.fr/ISSAC2002/.

Ulmer:1990:LSH

[UC90] F. Ulmer and J. Calmet. On Liouvillian solutions of homogeneous linear differential equations. In Watanabe and Nagata [WN90], pages 236–243. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95

.I57 1990. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/96877/p236-ulmer/.

Grigoriev:1991:ASR

[uGK91] Dima Yu. u. Grigoriev and Marek Karpinski. Algorithms for sparse rational interpolation. In Watt [Wat91], pages 7–13. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991. URL http://

www.acm.org:80/pubs/citations/proceedings/issac/120694/
p7-grigoriev/.

Grigoriev:1992:ESP

[uGKO92] Dima Y. u. Grigoriev, Marek Karpinski, and Andrew M. Odlyzko. Existence of short proofs for nondivisibility of sparse polynomials under the extended Riemann hypothesis. In Wang [Wan92b], pages 117–122. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN

QA76.95.I59 1992. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/143242/p117-grigoriev/. ACM

order number: 505920.

Umans:2006:GTA

[Uma06] Christopher Umans. Group-theoretic algorithms for matrix multiplication. In Trager [Tra06], page 5. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

Valibouze:1989:RSF

[Val89] A. Valibouze. Resolvents and symmetric functions. In Gonnet [Gon89], pages 390–399. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989. US$29.00. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/74540/p390-

valibouze/. ACM order number: 505890. English and French.

Vallier:1993:ACN

[Val93] L. Vallier. An algorithm for the computation of normal forms and invariant manifolds. In Bronstein [Bro93], pages 225–233. ISBN 0-89791-604-2. LCCN QA 76.95

I59 1993. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/164081/p225-vallier/. ACM order

number: 505930.

vanHoeij:1994:CPR

[van94] Mark van Hoeij. Computing parameterizations of rational algebraic curves. In ACM [ACM94], pages 187–190. ISBN 0-89791-638-7. LCCN QA76.95.I59 1994. URL http://www.acm.org:

80/pubs/citations/proceedings/issac/190347/p187-van_
hoeij/.

vanHoeij:1995:ACW

[van95] Mark van Hoeij. An algorithm for computing the Weierstrass normal form. In Levelt [Lev95], pages 90–95. ISBN 0-89791-699-9. LCCN QA 76.95 I59 1995. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/220346/p90-

van_hoeij/. ACM order number: 505950.

vanHoeij:1996:RSM

[van96] Mark van Hoeij. Rational solutions of the mixed differential equation and its application to factorization of differential operators. In Lakshman Y. N. [Lak96], pages 219–225. ISBN 0-89791-796-0. LCCN QA 76.95 I59 1996. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/236869/p219-
van_hoeij/.

vanderHoeven:1997:LMF

[van97] Joris van der Hoeven. Lazy multiplication of formal power series. In Küchlin [Küc97], pages 17–20. ISBN 0-89791-

875-4. LCCN QA76.95. URL http://www.acm.org:80/
pubs/citations/proceedings/issac/258726/p17-van_der_
hoeven/.

vanHoeij:1998:RSL

[van98] Mark van Hoeij. Rational solutions of linear difference equations. In Gloor [Glo98], pages 120–123. ISBN 1-58113-002-3. LCCN QA155.7.E4 E88 1998. URL http://

www.acm.org/pubs/citations/proceedings/issac/281508/
p120-van_hoeij/.

Vinette:1989:USC

[VC89] F. Vinette and J. Cizek. The use of symbolic computation in solving some nonrelativistic quantum mechanical problems. In Gianni [Gia89], pages 85–95. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Hoeven:2002:NZT

[vdH02] Joris van der Hoeven. A new zero-test for formal power series. In Mora [Mor02], pages 117–122. ISBN 1-58113-484-3. LCCN

QA76.95. URL http://www.lifl.fr/ISSAC2002/.

Hoeven:2003:RMU

[vdH03] Joris van der Hoeven. Relaxed multiplication using the middle product. In Senda [Sen03], pages 143–147. ISBN 1-58113-641-2. LCCN QA76.95. ACM order number 505030.

vanderHoeven:2004:TFT

[vdH04] Joris van der Hoeven. The truncated Fourier transform and applications. In Gutierrez [Gut04], pages 290–296. ISBN 1-58113-827-X. LCCN QA76.95 .I57 2004.

vanderHoeven:2006:ERN

[vdH06] Joris van der Hoeven. Effective real numbers in Mmxlib. In Trager [Tra06], pages 138–145. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

vanderHoeven:2012:CMB

[vdHL12] Joris van der Hoeven and Grégoire Lecerf. On the complexity of multivariate blockwise polynomial multiplication. In van der Hoeven and van Hoeij [vdHvH12], pages 211–218. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

vanderHoeven:2013:IMC

[vdHL13] Joris van der Hoeven and Grégoire Lecerf. Interfacing MATHEMAGIX with C++. In Monagan et al. [MCG13], pages 363–370. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

vanderHoeven:2013:SFT

[vdHLS13] Joris van der Hoeven, Romain Lebreton, and Éric Schost. Structured FFT and TFT: symmetric and lattice polynomials. In Monagan et al. [MCG13], pages 355–362. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

vanderHoeven:2012:IPI

[vdHvH12] Joris van der Hoeven and Mark van Hoeij, editors. ISSAC 2012: Proceedings of the 2012 International Symposium on Symbolic and Algebraic Computation, July 22–25, 2012, Grenoble, France. ACM Press, New York, NY 10036, USA, 2012. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

vandeWoestijne:2005:DES

[vdW05] Christiaan van de Woestijne. Deterministic equation solving over finite fields. In Kauers [Kau05], pages 348–353. ISBN 1-59593-095-7. LCCN ???? ACM Order Number 505050.

vandeWoestijne:2006:SPD

[vdW06] Christiaan van de Woestijne. Surface parametrisation without diagonalisation. In Trager [Tra06], pages 340–344. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

Verschelde:2004:NAG

[Ver04] Jan Verschelde. Numerical algebraic geometry and symbolic computation. In Gutierrez [Gut04], page 3. ISBN 1-58113-827-X. LCCN QA76.95 .I57 2004.

Viklund:1992:OLS

[VF92] Lars Viklund and Peter Fritzson. An object-oriented language for symbolic computation — applied to machine element analysis. In Wang [Wan92b], pages 397–405. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN

QA76.95.I59 1992. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/143242/p397-viklund/. ACM

order number: 505920.

VonzurGathen:1996:AFP

[vG96] Joachim von zur Gathen and Jürgen Gerhard. Arithmetic and factorization of polynomial over F_{2} (extended abstract). In Lakshman Y. N. [Lak96], pages 1–9. ISBN 0-89791-796-0.

LCCN QA 76.95 I59 1996. URL http://www.acm.org:80/
pubs/citations/proceedings/issac/236869/p1-von_zur_
gathen/.

vanderPut:1993:RRK

[vH93] Marius van der Put and Peter A. Hendriks. A rationality result for Kovacic’s algorithm. In Bronstein [Bro93], pages 4–8. ISBN 0-89791-604-2. LCCN QA 76.95 I59 1993. URL http://www.

acm.org:80/pubs/citations/proceedings/issac/164081/

p4-van_der_put/. ACM order number: 505930.

vonzurGathen:1996:FMP

[vH96] Joachim von zur Gathen and Silke Hartlieb. Factoring modular polynomials (extended abstract). In Lakshman Y. N. [Lak96], pages 10–17. ISBN 0-89791-796-0. LCCN QA 76.95

I59 1996. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/236869/p10-von_zur_gathen/.

vanHoeij:2007:STO

[vH07] Mark van Hoeij. Solving third order linear differential equations in terms of second order equations. In Brown [Bro07], pages 355–360. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

vanHoeij:2013:CFU

[vH13] Mark van Hoeij. The complexity of factoring univariate polynomials over the rationals: tutorial abstract. In Monagan et al. [MCG13], pages 13–14. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

vanHulzen:1989:COP

[vHGv89] J. A. van Hulzen, B. J. A. Hulshof, B. L. Gates, and M. C. van Heerwaarden. A code optimization package for REDUCE. In Gonnet [Gon89], pages 163–170. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989. US$29.00. URL http://www.acm.org:

80/pubs/citations/proceedings/issac/74540/p163-van_

hulzen/. ACM order number: 505890. English and French.

vanHoeij:2011:GS

[vHKN11] Mark van Hoeij, Jürgen Klüners, and Andrew Novocin. Generating subfields. In Schost and Emiris [SE11], pages 345–352. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

vanHoeij:2010:LSI

[vHL10] Mark van Hoeij and Giles Levy. Liouvillian solutions of irreducible second order linear difference equations. In Watt [Wat10], pages 297–301. ISBN 1-4503-0150-9. LCCN QA76.95 .I59 2010.

Hoeij:2002:MGA

[vHM02] Mark van Hoeij and Michael Monagan. A modular GCD algorithm over number fields presented with multiple extensions. In Mora [Mor02], pages 109–116. ISBN 1-58113-484-3. LCCN

QA76.95. URL http://www.lifl.fr/ISSAC2002/.

vanHoeij:2004:APG

[vHM04] Mark van Hoeij and Michael Monagan. Algorithms for polynomial GCD computation over algebraic function fields. In Gutierrez [Gut04], pages 297–304. ISBN 1-58113-827-X. LCCN QA76.95 .I57 2004.

vanHoeij:2005:SSO

[vHW05] M. van Hoeij and J.-A. Weil. Solving second order linear differential equations with Klein’s theorem. In Kauers [Kau05], pages 340–347. ISBN 1-59593-095-7. LCCN ???? ACM Order Number 505050.

vanHoeij:2010:FAB

[vHY10] Mark van Hoeij and Quan Yuan. Finding all Bessel type solutions for linear differential equations with rational function coefficients. In Watt [Wat10], pages 37–44. ISBN 1-4503-0150-

9. LCCN QA76.95 .I59 2010.

Villard:1992:PLB

[Vil92] Gilles Villard. Parallel lattice basis reduction. In Wang [Wan92b], pages 269–277. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN QA76.95.I59 1992. URL http:

//www.acm.org:80/pubs/citations/proceedings/issac/143242/

p269-villard/. ACM order number: 505920.

Villard:1993:CSN

[Vil93] Gilles Villard. Computation of the Smith normal form of polynomial matrices. In Bronstein [Bro93], pages 209–217. ISBN 0-89791-604-2. LCCN QA 76.95 I59 1993. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/164081/p209-

villard/. ACM order number: 505930.

Villard:1994:FPC

[Vil94] Gilles Villard. Fast parallel computation of the Smith normal form of polynomial matrices. In ACM [ACM94], pages 312–317. ISBN 0-89791-638-7. LCCN QA76.95.I59 1994. URL http://

www.acm.org:80/pubs/citations/proceedings/issac/190347/
p312-villard/.

Villard:1996:CPH

[Vil96] G. Villard. Computing Popov and Hermite forms of polynomial matrices. In Lakshman Y. N. [Lak96], pages 250–258. ISBN 0-89791-796-0. LCCN QA 76.95 I59 1996. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/236869/p250-
villard/.

Villard:1997:FAC

[Vil97] G. Villard. Further analysis of Coppersmith’s block Wiedemann algorithm for the solution of sparse linear systems. In Küchlin [Küc97], pages 32–39. ISBN 0-89791-875-4. LCCN

QA76.95. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/258726/p32-villard/.

Villard:2007:CFL

[Vil07a] Gilles Villard. Certification of the QR factor R and of lattice basis reducedness. In Brown [Bro07], pages 361–368. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Villard:2007:SRP

[Vil07b] Gilles Villard. Some recent progress in exact linear algebra and related questions. In Brown [Bro07], pages 391–392. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Villard:2011:RPL

[Vil11] Gilles Villard. Recent progress in linear algebra and lattice basis reduction. In Schost and Emiris [SE11], pages 3–4. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Vinette:1991:FSC

[Vin91] F. Vinette. Features of symbolic computation exploited in the calculation of lower energy bounds of cyclic polyene models. In Watt [Wat91], pages 458–459. ISBN 0-89791-437-6. LCCN

QA 76.95 I59 1991. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/120694/p458-vinette/.

Villard:1999:AIA

[VM99] Dominique Villard and Michael B. Monagan. ADrien: an implementation of automatic differentiation in Maple. In Dooley [Doo99], pages 221–228. ISBN 1-58113-073-2.

LCCN QA76.95 .I57 1999. URL http://www.acm.org/pubs/
contents/proceedings/issac/309831/.

Volcheck:1993:NSS

[Vol93a] E. J. Volcheck. Noether’s S-transformation simplifies curve singularities rationally: a local analysis. In Bronstein [Bro93], pages 164–172. ISBN 0-89791-604-2. LCCN QA 76.95 I59 1993. ACM order number: 505930.

Volcheck:1993:NTS

[Vol93b] Emil J. Volcheck. Noether’s S-transformation simplifies curve singularities rationally: a local analysis. In Bronstein [Bro93], pages 164–172. ISBN 0-89791-604-2. LCCN QA 76.95

I59 1993. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/164081/p164-volcheck/. ACM order

number: 505930.

Volcheck:1997:CDP

[Vol97] Emil Volcheck. On computing the dual of a plane algebraic curve. In Küchlin [Küc97], pages 356–358. ISBN 0-89791-

875-4. LCCN QA76.95. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/258726/p356-volcheck/.

Vollmer:2003:NHB

[Vol03] Ulrich Vollmer. A note on the Hermite basis computation of large integer matrices. In Senda [Sen03], pages 255–257. ISBN 1-58113-641-2. LCCN QA76.95. ACM order number 505030.

vonzurGathen:1990:PFF

[von90] J. von zur Gathen. Polynomials over finite fields with large images. In Watanabe and Nagata [WN90], pages 140–144. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95 .I57 1990. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/96877/p140-von_zur_
gathen/.

Vakhidov:1998:ACA

[VT98] Akmal A. Vakhidov and Irina V. Tupikova. Application of computer algebra methods to the construction of an asteroid motion theory based on Lie transforms. In Gloor [Glo98], pages 76– 79. ISBN 1-58113-002-3. LCCN QA155.7.E4 E88 1998. URL

http://www.acm.org/pubs/citations/proceedings/issac/
281508/p76-vakhidov/.

V:1997:MRP

[VZ97] Eugene V. and Zima. Mixed representation of polynomials oriented towards fast parallel shift. In Küchlin [Küc97], pages 150– 155. ISBN 0-89791-875-4. LCCN QA76.95. URL http://www.

acm.org:80/pubs/citations/proceedings/issac/266670/
p150-zima/.

vonzurGathen:2001:ITF

[vzG01] Joachim von zur Gathen. Irreducible trinomials over finite fields. In Mourrain [Mou01], pages 332–336. ISBN 1-58113-417-7. LCCN QA76.95.I59 2001. ACM order number 505010.

vonzurGathen:2006:WWW

[vzG06] Joachim von zur Gathen. Who was who in polynomial factorization: 1. In Trager [Tra06], page 2. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

vonzurGathen:2007:CRS

[vzG07] Joachim von zur Gathen. Counting reducible and singular bivariate polynomials. In Brown [Bro07], pages 369–376. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

vonzurGathen:2009:NDU

[vzG09] Joachim von zur Gathen. The number of decomposable univariate polynomials. extended abstract. In May [May09], pages 359–366. ISBN 1-60558-609-9. LCCN ????

vonzurGathen:1997:FAT

[vzGG97] Joachim von zur Gathen and Jürgen Gerhard. Fast algorithms for Taylor shifts and certain difference equations. In Küchlin [Küc97], pages 40–47. ISBN 0-89791-875-4. LCCN

QA76.95. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/258726/p40-von_zur_gathen/.

vonzurGathen:2010:CCP

[vzGGZ10] Joachim von zur Gathen, Mark Giesbrecht, and Konstantin Ziegler. Composition collisions and projective polynomials: statement of results. In Watt [Wat10], pages 123–130. ISBN 1-4503-0150-9. LCCN QA76.95 .I59 2010.

vonzurGathen:1999:CSP

[vzGN99] Joachim von zur Gathen and Michael Nöcker. Computing special powers in finite fields: extended abstract. In Dooley [Doo99], pages 83–90. ISBN 1-58113-073-2. LCCN QA76.95

.I57 1999.  URL http://www.acm.org/pubs/contents/
proceedings/issac/309831/.

Wang:1981:SPA

[Wan81] Paul S. Wang, editor. SYMSAC ’81: proceedings of the 1981 ACM Symposium on Symbolic and Algebraic Computation, Snowbird, Utah, August 5–7, 1981. ACM Press, New York, NY 10036, USA, 1981. ISBN 0-89791-047-8. LCCN QA155.7.E4 A28 1981. US$23.00. ACM order no. 505810.

Wang:1990:PUP

[Wan90]  P. S. Wang.  Parallel univariate polynomial factoriza-

tion on shared-memory multiprocessors. In Watanabe and Nagata [WN90], pages 145–151. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95

.I57 1990. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/96877/p145-wang/.

Wang:1991:TMI

[Wan91] Dongming Wang. A toolkit for manipulating indefinite summations with application to neural networks. In Watt [Wat91], pages 462–463. ISBN 0-89791-437-6. LCCN QA 76.95

I59 1991. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/120694/p462-wang/.

Wang:1992:PUA

[Wan92a] Paul S. Wang. Parallel univariate p-adic lifting on sharedmemory multiprocessors. In Proceedings of ISSAC ’92. International Symposium on Symbolic and Algebraic Computation [Wan92b], pages 168–176. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN QA76.95.I59 1992. URL http:

//www.acm.org:80/pubs/citations/proceedings/issac/143242/

p168-wang/. ACM order number: 505920.

Wang:1992:PII

[Wan92b] Paul S. Wang, editor. Proceedings of ISSAC ’92. International Symposium on Symbolic and Algebraic Computation. ACM Press, New York, NY 10036, USA, 1992. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN QA76.95.I59 1992. ACM order number: 505920.

Wang:1994:PPO

[Wan94] P. S. Wang. Parallel polynomial operations: A progress report (invited tutorial). In Hong [Hon94], pages 394–404. ISBN 981-02-2040-5. LCCN QA76.642.I58 1994.

Wang:1997:TPD

[Wan97] Paul S. Wang. Tools for parallel/distributed mathematical computation. In Küchlin [Küc97], pages 188–195. ISBN 0-89791-

875-4. LCCN QA76.95. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/266670/p188-wang/.

Wang:1999:DPI

[Wan99] Paul S. Wang. Design and protocol for Internet accessible mathematical computation. In Dooley [Doo99], pages 291–298. ISBN 1-58113-073-2. LCCN QA76.95 .I57 1999. URL http://www.

acm.org/pubs/contents/proceedings/issac/309831/.

Watt:1989:FPM

[Wat89] S. M. Watt. A fixed point method for power series computation. In Gianni [Gia89], pages 206–217. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Watt:1991:IPI

[Wat91] Stephen M. Watt, editor. ISSAC ’91: proceedings of the 1991 International Symposium on Symbolic and Algebraic Computation, July 15–17, 1991, Bonn, Germany. ACM Press, New York, NY 10036, USA, 1991. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991.

Watt:2010:IPI

[Wat10] Stephen M. Watt, editor. ISSAC 2010: Proceedings of the 2010 International Symposium on Symbolic and Algebraic Computation, July 25–28, 2010, Munich, Germany. ACM Press, New York, NY 10036, USA, 2010. ISBN 1-4503-0150-9. LCCN QA76.95 .I59 2010.

Wavrik:1999:CTE

[Wav99] John J. Wavrik. Commutativity theorems: examples in search of algorithms. In Dooley [Doo99], pages 31–36. ISBN 1-58113-073-2. LCCN QA76.95 .I57 1999. URL http://www.acm.org/

pubs/contents/proceedings/issac/309831/.

Willis:1993:CSP

[WB93] T. J. Willis and E. A. Bogucz. Coupling symbolic processing with parallel numeric computation. In Sincovec [Sin93], pages 788–792. ISBN 0-89871-315-3. LCCN QA76.58.S55 1993. Two volumes.

Watt:1994:FRA

[WBD^{+}94] Stephen M. Watt, Peter A. Broadbery, Samuel S. Dooley, Pietro Iglio, Scott C. Morrison, Jonathan M. Steinbach, and Robert S. Sutor. A first report on the Ahash compiler. In ACM [ACM94], pages 25–31. ISBN 0-89791-638-7. LCCN QA76.95.I59 1994. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/190347/p25-watt/.

Wang:2008:IPQ

[WCD08] Xuhui Wang, Falai Chen, and Jiansong Deng. Implicitization and parametrization of quadratic surfaces with one simple base point. In Jeffrey [Jef08], pages 31–38. ISBN 1-59593-904-0. LCCN ????

Weber:1994:ATI

[Web94a] Andreas Weber. Algorithms for type inference with coercions. In ACM [ACM94], pages 324–329. ISBN 0-89791-638-7. LCCN QA76.95.I59 1994. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/190347/p324-weber/.

Weber:1994:PIA

[Web94b] K. Weber. Parallel implementation of the accelerated integer GCD algorithm. In Hong [Hon94], pages 405–411. ISBN 981-02-2040-5. LCCN QA76.642.I58 1994.

Weispfenning:1989:EDP

[Wei89] V. Weispfenning. Efficient decision procedures for locally finite theories. II. In Gianni [Gia89], pages 262–273. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. For pt. I, see AECC-3, Grenoble, Springer LNCS, vol. 229.

Weispfenning:1992:FGB

[Wei92a] V. Weispfenning. Finite Gröbner bases in non-Noetherian skew polynomial rings. In Wang [Wan92b], pages 329– 334. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN QA76.95.I59 1992. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/143242/p329-

weispfenning/. ACM order number: 505920.

Weisss:1992:HDP

[Wei92b] Jürgen Weisß. Homogeneous decomposition of polynomials. In Wang [Wan92b], pages 146–151. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN

QA76.95.I59 1992. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/143242/p146-weiszlig/. ACM

order number: 505920.

Weispfenning:1993:DTP

[Wei93a] V. Weispfenning. Differential term-orders. In Bronstein [Bro93], pages 245–253. ISBN 0-89791-604-2. LCCN QA 76.95 I59 1993. ACM order number: 505930.

Weispfenning:1993:DT

[Wei93b] Volker Weispfenning. Differential term-orders. In Bronstein [Bro93], pages 245–253. ISBN 0-89791-604-2. LCCN QA 76.95

I59 1993. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/164081/p245-weispfenning/. ACM or-

der number: 505930.

Weil:1994:USS

[Wei94a] Jacques-Arthur Weil. The use of the special semi-groups for solving differential equations. In ACM [ACM94], pages 341–347. ISBN 0-89791-638-7. LCCN QA76.95.I59 1994. URL http://

www.acm.org:80/pubs/citations/proceedings/issac/190347/
p341-weil/.

Weispfenning:1994:QER

[Wei94b] Volker Weispfenning. Quantifier elimination for real algebrathe cubic case. In ACM [ACM94], pages 258–263. ISBN 0-89791-638-7. LCCN QA76.95.I59 1994. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/190347/p258-
weispfenning/.

Weispfenning:1997:CUE

[Wei97] Wolker Weispfenning. Complexity and uniformity of elimination in Presburger arithmetic. In Küchlin [Küc97], pages 48–53. ISBN 0-89791-875-4. LCCN QA76.95. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/258726/p48-
weispfenning/.

Weispfenning:1999:MRI

[Wei99] Volker Weispfenning. Mixed real-integer linear quantifier elimination. In Dooley [Doo99], pages 129–136. ISBN 1-58113-073-2.

LCCN QA76.95 .I57 1999. URL http://www.acm.org/pubs/
contents/proceedings/issac/309831/.

Weispfenning:2002:CCG

[Wei02] Volker Weispfenning. Canonical comprehensive Gröbner bases. In Mora [Mor02], pages 270–276. ISBN 1-58113-484-3. LCCN

QA76.95. URL http://www.lifl.fr/ISSAC2002/.

Wallack:1998:MMM

[WEM98] Aaron Wallack, Ioannis Z. Emiris, and Dinesh Manocha. MARS: a MAPLE/MATLAB/C resultant-based solver. In Gloor [Glo98], pages 244–251. ISBN 1-58113-002-3. LCCN

QA155.7.E4 E88 1998. URL http://www.acm.org/pubs/
citations/proceedings/issac/281508/p244-wallack/.

Weibel:1991:AP

[WG91] Trudy Weibel and Gaston H. Gonnet. An algebra of properties. In Watt [Wat91], pages 352–359. ISBN 0-89791-437-6. LCCN

QA 76.95 I59 1991. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/120694/p352-weibel/.

Wang:2001:IAP

[WGK^{+}01] P. Wang, S. Gray, N. Kajler, D. Lin, W. Liao, and X. Zou. IAMC architecture and prototyping: a progress report. In Mourrain [Mou01], pages 337–344. ISBN 1-58113-417-7. LCCN QA76.95.I59 2001. ACM order number 505010.

Wikstroem:1994:DPE

[Wik94] C. Wikstroem. Distributed programming in Erlang. In Hong [Hon94], pages 412–421. ISBN 981-02-2040-5. LCCN QA76.642.I58 1994.

Winkler:1989:GDA

[Win89a] F. Winkler. A geometrical decision algorithm based on the Gröbner bases algorithm. In Gianni [Gia89], pages 356–363. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Winkler:1989:KPB

[Win89b]  F. Winkler.  Knuth–Bendix procedure and Buchberger

algorithm — a synthesis. In Gonnet [Gon89], pages 55–67. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989.

US$29.00. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/74540/p55-winkler/. ACM order num-

ber: 505890. English and French.

Wissmann:1989:ART

[Wis89] D. Wissmann. Applying rewriting techniques to groups with power-commutation-presentations. In Gianni [Gia89], pages 378–389. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Jinzhao:1993:RPG

[WJLL93] Wu-Jinzhao and Li-Lian. The regular problem and Green equivalences for special monoids. In Bronstein [Bro93], pages 78–85. ISBN 0-89791-604-2. LCCN QA 76.95 I59 1993. ACM order number: 505930.

Wang:2003:WTW

[WKZZ03] Paul S. Wang, Norbert Kajler, Yi Zhou, and Xiao Zou. WME: towards a Web for Mathematics Education. In Senda [Sen03], pages 258–265. ISBN 1-58113-641-2. LCCN QA76.95. ACM order number 505030.

Wu:1993:RPG

[WL93] Jinzhao Wu and Lian Li. The regular problem and green equivalences for special monoids. In Bronstein [Bro93], pages 78–85. ISBN 0-89791-604-2. LCCN QA 76.95

I59 1993. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/164081/p78-wu/. ACM order number:

505930.

White:1989:CF

[WM89] N. L. White and T. McMillan. Cayley factorization. In Gianni [Gia89], pages 521–533. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Watanabe:1990:IPI

[WN90] Shunro Watanabe and Morio Nagata, editors. ISSAC ’90: proceedings of the International Symposium on Symbolic and Algebraic Computation: August 20–24, 1990, Tokyo, Japan. ACM Press and Addison-Wesley, New York, NY 10036, USA and Reading, MA, USA, 1990. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95 .I57 1990.

Wolf:1995:PAS

[Wol95] Thomas Wolf. Programs for applying symmetries of PDEs. In Levelt [Lev95], pages 7–15. ISBN 0-89791-699-9. LCCN

QA 76.95 I59 1995. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/220346/p7-wolf/. ACM or-

der number: 505950.

Wolfram:2013:CAY

[Wol13] Stephen Wolfram. Computer algebra: a 32-year update. In Monagan et al. [MCG13], pages 7–8. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

Wu:2006:ANA

[WR06] Wenyuan Wu and Greg Reid. Application of numerical algebraic geometry and numerical linear algebra to PDE. In Trager [Tra06], pages 345–352. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

Wu:2007:SNC

[WR07] Wenyuan Wu and Greg Reid. Symbolic-numeric computation of implicit Riquier Bases for PDE. In Brown [Bro07], pages 377– 386. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Wu:2013:FPR

[WR13] Wenyuan Wu and Greg Reid. Finding points on real solution components and applications to differential polynomial systems. In Monagan et al. [MCG13], pages 339–346. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

Wu:1993:ACU

[Wu93] Hongzhong Wu. On the assignment complexity of uniform trees. In Bronstein [Bro93], pages 95–104. ISBN 0-89791-604-2. LCCN

QA 76.95 I59 1993. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/164081/p95-wu/. ACM order

number: 505930.

Weerawarana:1989:GPC

[WW89] S. Weerawarana and P. S. Wang. GENCRAY: a portable code generator for Cray Fortran. In Gonnet [Gon89], pages 186–191. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989.

US$29.00. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/74540/p186-weerawarana/. ACM order

number: 505890. English and French.

Wang:2005:SAB

[WX05] Dongming Wang and Bican Xia. Stability analysis of biological systems with real solution classification. In Kauers [Kau05], pages 354–361. ISBN 1-59593-095-7. LCCN ???? ACM Order Number 505050.

Wu:2008:CMS

[WZ08] Xiaoli Wu and Lihong Zhi. Computing the multiplicity structure from geometric involutive form. In Jeffrey [Jef08], pages 325–332. ISBN 1-59593-904-0. LCCN ????

Yamasaki:1990:DLP

[Yam90] S. Yamasaki. Dataflow for logic program as substitution manipulator. In Watanabe and Nagata [WN90], pages 44–51. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN

QA76.95 .I57 1990. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/96877/p44-yamasaki/.

Yap:2009:ENC

[Yap09] Chee K. Yap. Exact numerical computation in algebra and geometry. In May [May09], pages 387–388. ISBN 1-60558-609-9. LCCN ????

Ye:1992:SLI

[YC92] Honglin Ye and Robert M. Corless. Solving linear integral equations in Maple. In Wang [Wan92b], pages 95–102. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN

QA76.95.I59 1992. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/143242/p95-ye/. ACM order

number: 505920.

Yang:2004:EME

[YF04] Michael Yang and Richard Fateman. Extracting mathematical expressions from PostScript documents. In Gutierrez [Gut04], pages 305–311. ISBN 1-58113-827-X. LCCN QA76.95 .I57

2004. URL http://www.ocf.berkeley.edu/~mlyang/papers/
MichaelYangPsmath.pdf.

Yakubovich:1991:EIS

[YL91] S. B. Yakubovich and Yu. F. Luchko. The evaluation of integrals and series with respect to indices (parameters) of hypergeometric functions. In Watt [Wat91], pages 271–280. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/120694/p271-
yakubovich/.

Yokoyama:1995:FRU

[YLN95] Kazuhiro Yokoyama, Ziming Li, and István Nemes. Finding roots of unity among quotients of the roots of an integral polynomial. In Levelt [Lev95], pages 85–89. ISBN 0-89791-699-9. LCCN QA 76.95 I59 1995. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/220346/p85-

yokoyama/. ACM order number: 505950.

Yokoyama:1990:DSP

[YNT90a] K. Yokoyama, M. Noro, and T. Takeshima. On determining the solvability of polynomials. In Watanabe and Nagata [WN90], pages 127–134. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95

.I57 1990. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/96877/p127-yokoyama/.

Yokoyama:1990:FMP

[YNT90b] Kazuhiro Yokoyama, Masayuki Noro, and Taku Takeshima. On factoring multi-variate polynomials over algebraically closed fields (abstract). In Watanabe and Nagata [WN90], page 297. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95 .I57 1990. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/96877/p297-
yokoyama/.

Yokoyama:2004:SAE

[Yok04] Kazuhiro Yokoyama. On systems of algebraic equations with parametric exponents. In Gutierrez [Gut04], pages 312–319. ISBN 1-58113-827-X. LCCN QA76.95 .I57 2004.

Yokoyama:1993:HCE

[YT93] Kazuhiro Yokoyama and Taku Takeshima. On Hensel construction of eigenvalues and eigenvectors of matrices with polynomial entries. In Bronstein [Bro93], pages 218–224. ISBN 0-89791-604-2. LCCN QA 76.95 I59 1993. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/164081/p218-

yokoyama/. ACM order number: 505930.

Yun:1986:FCF

[YZ86] D. Y. Y. Yun and C. N. Zhang. A fast carry-free algorithm and hardware design for extended integer GCD computation. In Char [Cha86b], pages 82–84. ISBN 0-89791-199-7 (paperback). LCCN QA155.7.E4 A281 1986. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/32439/p82-yun/. ACM

order no. 505860.

Yang:2005:OPM

[YZ05] Lu Yang and Zhenbing Zeng. An open problem on metric invariants of tetrahedra. In Kauers [Kau05], pages 362–364. ISBN 1-59593-095-7. LCCN ???? ACM Order Number 505050.

Yang:2013:VEB

[YZZ13] Zhengfeng Yang, Lihong Zhi, and Yijun Zhu. Verified error bounds for real solutions of positive-dimensional polynomial systems. In Monagan et al. [MCG13], pages 371–378. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

Zanoni:2010:ITC

[Zan10] Alberto Zanoni. Iterative Toom–Cook methods for very unbalanced long integer multiplication. In Watt [Wat10], pages 319–323. ISBN 1-4503-0150-9. LCCN QA76.95 .I59 2010.

Zhang:1994:CSD

[ZB94] H. Zhang and M. P. Bonacina. Cumulating search in a distributed computing environment: A case study in parallel satisfiability. In Hong [Hon94], pages 422–431. ISBN 981-02-2040-5. LCCN QA76.642.I58 1994.

Zeng:2004:AGI

[ZD04] Zhonggang Zeng and Barry H. Dayton. The approximate GCD of inexact polynomials. In Gutierrez [Gut04], pages 320–327. ISBN 1-58113-827-X. LCCN QA76.95 .I57 2004.

Zeng:2003:MCM

[Zen03] Zhonggang Zeng. A method computing multiple roots of inexact polynomials. In Senda [Sen03], pages 266–272. ISBN 1-58113-641-2. LCCN QA76.95. ACM order number 505030.

Zeng:2009:AIF

[Zen09] Zhonggang Zeng. The approximate irreducible factorization of a univariate polynomial: revisited. In May [May09], pages 367– 374. ISBN 1-60558-609-9. LCCN ????

Zhang:2000:RCC

[ZG00] Ming Zhang and Ron Goldman. Rectangular corner cutting and Sylvester-resultants. In Traverso [Tra00], pages 301– 308. ISBN 1-58113-218-2. LCCN QA76.95.I59 2000. URL

http://www.acm.org/pubs/articles/proceedings/issac/
345542/p301-zhang/p301-zhang.pdf;   http://www.acm.
org/pubs/citations/proceedings/issac/345542/p301-zhang/

. ACM order number 505000.

Zharkov:1993:ASF

[Zha93] Alexey Y. Zharkov. On algebraic solutions of first order Riccatti equation. In Bronstein [Bro93], pages 1–3. ISBN 0-89791-604-2.

LCCN QA 76.95 I59 1993. URL http://www.acm.org:80/
pubs/citations/proceedings/issac/164081/p1-zharkov/.

ACM order number: 505930.

Ziel:1991:RFD

[Zie91] Richard Ziel. Rational function decomposition. In Watt [Wat91], pages 1–6. ISBN 0-89791-437-6. LCCN QA 76.95

I59 1991. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/120694/p1-zippel/.

Zima:1993:NCO

[Zim93] E. V. Zima. Numeric code optimization in computer algebra systems and recurrent relations technique. In Bronstein [Bro93], pages 42–46. ISBN 0-89791-604-2. LCCN QA 76.95

I59 1993. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/164081/p42-zima/. ACM order number:

505930.

Zima:1995:SOT

[Zim95] Eugene V. Zima. Simplification and optimization transformations of chains of recurrences. In Levelt [Lev95], pages 42–50. ISBN 0-89791-699-9. LCCN QA 76.95

I59 1995. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/220346/p42-zima/. ACM order number:

505950.

Zimmer:1997:SCA

[Zim97] Horst G. Zimmer. SIMATH — a computer algebra system for number theoretic applications. In Küchlin [Küc97], pages 365– 375. ISBN 0-89791-875-4. LCCN QA76.95. URL http://www.

acm.org:80/pubs/citations/proceedings/issac/258726/
p365-zimmer/.

Zima:2001:CPC

[Zim01] Eugene V. Zima. On computational properties of chains of recurrences. In Mourrain [Mou01], page 345. ISBN 1-58113-417-7. LCCN QA76.95.I59 2001. ACM order number 505010.

Zippel:1991:RFD

[Zip91] R. Zippel. Rational function decomposition. In Watt [Wat91], pages 1–6. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991.

Zhou:2009:ECO

[ZL09] Wei Zhou and George Labahn. Efficient computation of order bases. In May [May09], pages 375–382. ISBN 1-60558-609-9. LCCN ????

Zhang:2012:FDO

[ZL12] Mingbo Zhang and Yong Luo. Factorization of differential operators with ordinary differential polynomial coefficients. In van der Hoeven and van Hoeij [vdHvH12], pages 359–365. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Zhou:2013:CCB

[ZL13] Wei Zhou and George Labahn. Computing column bases of polynomial matrices. In Monagan et al. [MCG13], pages 379– 386. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

Zhou:2012:CMN

[ZLS12] Wei Zhou, George Labahn, and Arne Storjohann. Computing minimal nullspace bases. In van der Hoeven and van Hoeij [vdHvH12], pages 366–373. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Zobnin:2005:AOF

[Zob05] Aleksey Zobnin. Admissible orderings and finiteness criteria for differential standard bases. In Kauers [Kau05], pages 365–372. ISBN 1-59593-095-7. LCCN ???? ACM Order Number 505050.

Zolotykh:1991:PCS

[Zol91] A. A. Zolotykh. A package for computations in simple Lie algebra representations. In Watt [Wat91], pages 237–238. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991. URL http://www.acm.

org:80/pubs/citations/proceedings/issac/120694/p237-
zolotykh/.

Zolotykh:1997:TPD

[Zol97] Andrij A. Zolotykh. Tensor product decomposition and other algorithms for representations of large simple Lie algebras. In Küchlin [Küc97], pages 60–67. ISBN 0-89791-875-4. LCCN

QA76.95. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/258726/p60-zolotykh/.

Zilic:1999:FMP

[ZR99] Zeljko Zilic and Katarzyna Radecka. On feasible multivariate polynomial interpolations over arbitrary fields. In Dooley [Doo99], pages 67–74. ISBN 1-58113-073-2. LCCN QA76.95

.I57 1999.  URL http://www.acm.org/pubs/contents/
proceedings/issac/309831/.

Zhao:1996:MPM

[ZSST96] Yanjie Zhao, Tetsuya Sakurai, Hiroshi Sugiura, and Tatsuo Torii. A methodology of parsing mathematical notation for mathematical computation. In Lakshman Y. N. [Lak96], pages 292–300. ISBN 0-89791-796-0. LCCN QA 76.95

I59 1996. URL http://www.acm.org:80/pubs/citations/
proceedings/issac/236869/p292-zhao/.

Zhou:2006:GAB

[ZW06] Meng Zhou and Franz Winkler. Gröbner bases in differencedifferential modules. In Trager [Tra06], pages 353–360. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.